A Bibliography of Papers in Lecture Notes in Computer Science (2002) (Part 1 of 4)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

29 July 2017 Version 1.13

Title word cross-reference

{0, 1} [581]. 2 [1771, 1676, 918, 1096, 161, 1711, 1353]. 26 [895]. 3 [918, 1327, 911, 1817, 88, 425, 192, 1607, 1096, 782, 1820, 1711, 915, 900, 1806, 1080, 27, 1766]. 4 [898]. 5/2 [776]. 6 [900]. 9 [782]. ^{∗} [752]. ^{2} [1310]. ^{3} [1037]. ^{TM} [1314]. 12 [1688]. _{6} [1688]. _{YX} [571]. α [1407]. β [506]. d [1131]. h [1843]. k [487]. κ [1638]. L [488]. λ [742]. O(ln ρ/ ln ln ρ) [422]. μ [1394, 1314]. n [1655, 1090]. ω [757]. P [900, 502]. pq [1205]. q [1452]. r [1843]. r^{th} [480]. t [1658]. × [1125]. ZZ^{n} [894].

- [1843]. -Approximation [422]. -Ary [1131, 1658, 1452]. -Calculus [1394].

-Coloring [488]. -D [1817, 192, 1353]. -definable [742]. -Dimensional [1327]. -Histograms [1205]. -Level [1676]. -Metric [581]. -pseudosnakes [487]. -Queen [1090]. -SAT [425]. -Simple [900]. -skeletons [506]. -Subiteration [900]. -Surfaces [895].

11th [1503].

4th [1519, 1514].

6th [1510].

7 [1281]. 7th [1515].

= [1793, 32].

A/D [624]. Abbreviations [203].

Abduction [336]. Abductive [1165]. able [519]. Absence [299]. Absorbed [926]. Abstract [733, 1364, 750, 1385, 1524, 214, 748, 1017, 916, 1386, 1430, 737, 193]. Abstraction [616, 301, 1116, 1384, 1387, 1407, 21, 1371, 892, 168, 293, 751]. Abstractions [333, 644, 646, 16]. Acceleration [1622, 427]. Acceptance [749]. Accepting [1214]. Access [604, 1081, 886, 711, 575, 964, 1799, 552]. Accessibility [705]. Accessing [1571]. Accidents [1255]. Accountability [597]. Accumulation [1208]. Accumulative [1009]. Accuracy [1771, 1098, 1792, 1467, 1244]. Accuracy-Based [1467]. Accurate [1809]. ACO [258, 263, 272]. Acquaintance [815]. Acquired [1096]. Acquisition [332]. Acquisitions [329]. ACS2 [1476]. Action [1237]. Actions [823, 524]. Active [1098, 1590, 601, 1252, 1236, 342, 184, 1428]. Activities [1152]. Activity [1285, 1101]. Actuators [645]. Acyclic [53, 309, 771]. Ad [1797, 1309, 602]. Ad-hoc [602]. Adaptability [1490]. Adaptable [143]. Adaptation [877, 1425]. Adapting [1298, 1204]. Adaptive [135, 132, 136, 1487, 1254, 884, 1604, 1079, 314, 142, 1468, 1830, 1492, 1164, 1256, 1599, 697, 1507, 1589, 159, 532, 143, 134, 1161]. Adaptivity [795]. Addition [1332, 757]. Address [588]. Addressing [879, 161]. Adjunct [244]. Adjustment [1278]. ADL [17, 1034]. Admissibility [214]. Admitting [965]. Adopting [1445]. Advanced [1575, 1511, 1549, 406, 1430]. Advances [1325, 907]. Advice [1205]. Aesthetic [74]. Affecting [1098]. Affine [150, 639, 152]. after [1276]. AG [1805]. AG-IVE [1805]. against [1328, 601, 1432, 608, 607, 606]. AGD [103, 187]. Age [1290]. Ageing [1732]. Agent [794, 604, 7, 1297, 795, 1491, 1490, 1239, 8, 797, 1158, 3, 1153, 1237, 4, 802, 17, 1496, 1488, 804, 633, 1492, 1673, 9, 1479, 809, 1780, 1063, 13, 1228, 1501, 1315, 1497, 18, 1494, 14, 16, 1805, 10, 798, 818, 1240, 12, 1557, 1466, 801, 1757, 1305, 803, 1470, 1310, 1610, 808, 1311, 812, 813, 814, 815, 1224, 817, 821, 1152, 19]. Agent-Based [7, 17, 1315, 1497]. Agent-Enhanced [1158]. Agent-Oriented [1063, 13, 14]. Agents [10, 1487, 1482, 799, 6, 1238, 1489, 807, 1102, 1493, 605, 1480, 791, 849, 1478]. Agentspace [794]. Aggregate [559]. Aggregation [1706, 1102, 557]. Agile [1438]. Aging [1448]. Agreement [1395]. AH3 [131]. AHP [1782]. Aiming [1159]. Aircraft [286]. Airline [286]. Airship [33]. Algebra [1118, 1029, 955, 832, 213, 835]. Algebra-Coalgebra [955]. Algebraic [612, 834, 779, 712, 714, 1713, 958, 444]. Algebras [829, 1124]. Algorithm [1176, 242, 1485, 474, 1655, 1577, 924, 1838, 1212, 1679, 1204, 1767, 165, 1104, 57, 180, 1640, 172, 169, 1082, 187, 1165, 368, 315, 425, 181, 1563, 1613, 374, 164, 266, 1762, 1610, 691, 1644, 483, 422, 900, 345, 545, 927, 811, 494, 1214, 1312, 1143, 485, 558, 1684, 171, 273, 1568, 193]. Algorithmic [1476, 1552, 350]. Algorithms [516, 514, 866, 473, 96, 170, 1128, 1141, 507, 1132, 1685, 496, 1656, 911, 1557, 167, 285, 1483, 308, 1680, 1134, 265, 1687, 1816, 183, 1072, 745, 1210, 1213, 1795, 1832, 471, 547, 1650, 1230, 466, 1215, 243, 546, 1831, 249, 235, 1701, 1155, 1150, 1696, 1318, 1068, 326, 152, 1080, 717, 469]. Alias [1591]. Aliasing [168]. ALICE [1785]. Alignment [473, 424]. Allele [243]. Allocated [982]. Allocation [1797, 307, 1640, 1560, 995, 824]. Alloy [291]. Alpha [636]. Alpha-Beta [636]. Alphabets [776]. Alternating [1291]. Alternating-Time [1291]. Alternation [1394]. Alternation-Free [1394]. Alternative [261, 1076]. Alzheimer [1275]. Ambient [520]. Ambients [1023, 750]. among [253]. Amplitudes [435]. AMR [1694]. Analyser [708]. Analyses [1018]. Analysing [733, 1355]. Analysis [1177, 616, 618, 1389, 105, 1525, 695, 145, 1676, 15, 1577, 8, 536, 740, 240, 1679, 991, 461, 1439, 1024, 627, 1706, 1540, 1305, 371, 1591, 1661, 1748, 472, 985, 353, 1310, 372, 355, 1396, 311, 251, 806, 1613, 855, 356, 374, 358, 1689, 1543, 1213, 809, 290, 984, 742, 994, 1020, 584, 1197, 1754, 726, 1594, 907, 18, 297, 585, 340, 862, 1010, 867, 1056, 1598]. Analytic [1536]. Analytical [694]. Analyzing [1544, 368, 736]. Anaphora [1217]. Anastomosing [1734]. Ancestor [508, 490]. And/Or [1033]. Angular [83]. Animated [184, 925]. Animating [167]. Animation [166, 172, 169, 851, 164, 1375, 171, 1348]. Animations [614]. Anisotropic [1789, 1766]. Annealing [1656, 1211, 1659]. Annotate [1270]. Answer [554]. Answering [1740]. Answers [441]. Ant [287]. Anthropomorphic [848]. Anticipatory [1464]. Anydevice [1382]. Anytime [1176, 1165, 1382]. Anywhere [1382]. API [95]. Appearance [23]. Appearance-Based [23]. Appendix [408, 202, 396]. Applet [1027]. Applicability [317]. Application [1045, 95, 1667, 1580, 1577, 920, 1679, 274, 1090, 1103, 139, 850, 804, 542, 354, 355, 417, 265, 1688, 1813, 1034, 1715, 1083, 1146, 1211, 337, 1664, 1133, 1665, 1578, 377, 1570, 1600, 352, 561, 1562]. Applications [369, 436, 1421, 1181, 1233, 838, 1184, 1756, 1668, 1499, 1447, 1636, 1648, 758, 189, 1574, 499, 1305, 1279, 1409, 713, 161, 1795, 1406, 1046, 1783, 1341, 155, 1701, 348, 1799, 992, 478, 445, 1291]. Applied [1232, 1807, 1489, 1246, 1479, 863, 282, 241]. Applying [1496, 115, 351, 1148, 1532, 119, 708]. Approach [242, 135, 1654, 1498, 11, 1125, 1828, 1026, 1036, 52, 539, 1763, 628, 1550, 81, 1642, 1169, 1489, 263, 671, 1396, 146, 893, 250, 294, 1747, 1236, 1044, 923, 1110, 1016, 1646, 984, 1220, 1274, 635, 521, 1001, 1754, 1207, 1174, 1702, 1696, 645, 1071, 1474, 143, 883, 1330]. Approaches [1517, 1437, 260, 944, 271, 1438]. Appropriate [1209]. Approximants [974]. Approximate [484, 417, 712, 437, 530, 550, 339, 791, 1262]. Approximating [752]. Approximation [1234, 1708, 422, 1089, 717]. Approximations [428, 1827]. Arbitrary [897]. Architecting [660]. Architectural [1033, 1295, 654, 1514]. Architecture [669, 1444, 797, 877, 1158, 1303, 676, 1439, 1822, 1670, 542, 633, 1639, 1122, 653, 668, 1595, 1418, 821, 1582, 159, 873, 282, 1568, 876, 883, 665, 1314]. Architecture-Based [877]. Architecture-Centric [1444]. Architectures [675, 676, 678, 680, 872, 250, 1621, 155, 1618, 18]. Area [886, 915, 881]. Argument [1715, 221]. Argumentation [1487, 1483]. Argumentation-Based [1487]. Argumentative [1165]. Arguments [1426]. Arms [253]. Array [1192, 1191, 1001]. Arrays [492, 744]. Arrival [1528]. Art [264]. ART2 [139]. Artifacts [1425]. Artificial [1757, 1103, 1230, 1088, 1834]. Artist [124]. Ary [1131, 1641, 1658, 1452]. Ascent [258]. Ashkin [1692]. ASIC [1663]. Aspects [795, 1756, 837, 805, 1552, 27, 350, 681]. Assembly [622, 786]. Asserting [1017]. Assertions [1183, 885, 383, 390]. Assess [1245, 1354]. Assessment [1334, 1467, 1335]. Asset [241]. Assets [659, 1443, 116, 377]. Assignment [1614, 54, 1090, 365]. Assistance [365]. Assistant [1164]. Assistants [220]. Assisted [1248, 1806]. Association [526, 374, 1085, 349, 1053, 1058, 1057, 1054]. Associative [1299, 1658]. Asymptotic [422]. Asynchronous [254, 295]. ATAM [680]. ATL [752]. Atmosphere [1777].

Atmospheric [1811]. AToM [1037]. ATSP [428]. Attack [607]. Attackers [1328]. Attacks [593, 601, 608]. Attention [1367, 859, 1366]. Attitudes [799]. Attractor [271]. Attribute [300, 664, 668, 989]. Audio [119]. Audiovisual [1081]. Augmentation [135, 266]. Augmentative [1076]. Augmented [1150]. Auld [118]. Auroral [356]. Authentication [603, 602, 592]. Author [20, 978, 1002, 1025, 1051, 1106, 1126, 1156, 1189, 1201, 1258, 1288, 1324, 1343, 1383, 1408, 1446, 1463, 1477, 1502, 1522, 1548, 1581, 1633, 1735, 1845, 41, 112, 144, 163, 195, 222, 255, 289, 325, 378, 464, 517, 576, 610, 647, 686, 710, 719, 755, 790, 826, 836, 870, 891, 930]. Authoring [1074, 1041, 134]. Authorisation [604]. Auto [1246, 373]. Auto-regressive [1246]. Auto-tuning [373]. Autocalibration [284]. Automata [298, 756, 1812, 769, 504, 773, 739, 295, 801, 777, 961, 1815, 1813, 476, 1094, 775, 1155, 1702, 642, 448, 765, 646, 1814, 749]. Automata-Based [801, 1702]. Automata-Theoretic [739]. Automated [796, 302, 1660, 1234, 1406, 684, 1725, 55, 18]. Automatic [843, 695, 247, 1284, 97, 250, 1044, 1777, 351, 1020, 177, 1214, 93, 747, 1823]. Automating [1038, 1506]. Automaton [629]. Automorphisms [1637]. Automotive [680, 871]. Autonomous [33, 1238, 1489, 809, 39, 43]. Avatar [853]. Average [1798, 449]. Avoiding [1587, 1147]. AVOW [1379]. AVS [1835]. Aware [575, 875]. Awareness [1254]. Axial [1253]. Axiomatising [459]. Axiomatizable [772]. Axiomatizing [458]. Axioms [212].

B} [1525, 303]. Background [721]. Backwarding [1146]. Bad [221]. Balanced [418]. Balancing [1488, 1567, 1798, 1612, 1603, 1561, 1645].

BAM [697]. BAN [793]. Bandwidth [1459, 1618]. bank [998]. Bar [1470]. Base [102]. Based [516, 1442, 96, 135, 1119, 7, 1402, 1653, 666, 11, 1299, 904, 1042, 140, 1300, 1239, 136, 1487, 928, 1151, 1285, 877, 1111, 1384, 1544, 28, 858, 911, 1336, 1113, 1807, 3, 1237, 844, 1537, 801, 734, 1346, 1234, 17, 1306, 1414, 1489, 1367, 1559, 263, 985, 1249, 1050, 147, 689, 1819, 1470, 1246, 1467, 312, 609, 1823, 1613, 807, 1114, 374, 1762, 633, 1281, 1803, 563, 1096, 1484, 23, 1731, 1420, 900, 1044, 567, 1406, 1738, 1639, 1646, 1663, 176, 1579, 887, 635, 636, 1107, 584, 1791, 707, 1034, 815, 1215, 1312, 703, 564, 1532, 1331, 220]. Based [683, 1083, 1263, 697, 1147, 1315, 1497, 1645, 1699, 1253, 338, 558, 1824, 1155, 1702, 1266, 1573, 1264, 1066, 643, 40, 1794, 1265, 548, 1095, 690, 1568, 1099, 288, 869, 1080, 717, 1321, 367, 134, 284, 120, 552, 599, 1436, 1356, 825, 1805]. Basic [1765, 1619, 454]. Basics [398]. Basin [275]. Basins [640]. Basis [976]. Batch [1528]. Bayes [1186]. Bayesian [698, 1091, 1078, 1173, 1244, 1265, 690]. BDD [1111, 745]. BDD-Based [1111]. BDDs [326]. Be [441, 1159, 1709]. BE/FE [1709]. Because [122]. Become [1386]. Behavior [1781, 1815, 634, 1199, 813, 863, 197]. Behavior-Oriented [813]. Behaviour [1298, 1470, 805, 1479, 809, 819]. Behavioural [804, 2]. Behaviours [1102]. Belief [1370]. Beliefs [1178]. Benchmark [1584, 1616, 1627, 946, 1600]. Benchmarking [1651]. Benchmarks [1628, 1620, 1617, 1615]. Bends [82]. Benefit [607]. Benefits [662]. Beowulf [1694]. Beowulf-Class [1694]. Best [227, 333, 345, 936, 346, 708]. Best-Match [708]. Best-Solution [936]. Beta [481, 636]. Beta-Expansions [481]. Better [1032]. between [735, 785, 1575, 232, 475, 1337, 383, 390, 553].

Beverages [1834]. Beyond [305, 1550]. BFS [486]. BGK [1771]. Bi [443, 1206]. Bi-directional [1206]. Bi-Immunity [443]. Biasing [1464]. Bibliography [1387, 409]. Bicolored [903]. Biconnectivity [266]. Bidomain [1789]. Bilateral [242]. Bilingual [706]. Bills [1286]. Binary [926, 1130, 590, 1558, 1748, 264, 1263, 283]. Biological [848, 1732]. Biologists [1127]. Biology [1492]. Biometrics [1249]. BioPath [94]. Biposets [773]. Biprocessor [1643]. Birge [1755]. Bisimilarity [454]. Bisimulation [734, 633, 977]. Bit [985]. Bits [511]. Black [1160, 316, 1597]. Black-Box [1160]. Blind [600]. Bloat [1147]. Block [1675, 1597, 494, 1696, 477, 288]. Bluetooth [889, 888]. Bluetooth-Enabled [889]. Boltzmann [1771, 1807]. Book [566, 122, 76]. Boolean [227, 440, 511]. Boosting [360, 1148, 272, 348, 350]. Borrowing [971]. Boru [278]. Bottleneck [1369, 1396]. Bottom [358]. Bound [1485, 59, 626, 414, 1318, 445]. Boundary [1728, 1712]. Bounded [456, 969, 76]. Bounds [416, 1679, 1327, 434, 961, 1820, 1400, 1001, 513]. Box [1160, 316]. Boxes [665]. Brain [340]. Branch [1485, 66, 884, 1604, 1594, 1318, 1593]. Branching [1743, 1011, 445]. Bridging [553]. Brightness [1370]. Bringing [4]. Broadcast [1299, 595, 1278, 552, 553]. Broadcast-Based [552]. Broadening [125]. Brute [250]. Brute-Force [250]. BTexact [1181]. Bucket [234]. Buffer [588, 1593]. Builders [95]. Building [980, 706, 1443, 288, 568]. Builds [474]. Buoy [1262]. Business [1460, 536, 1518, 368, 1780, 1480]. Busy [1545]. Bytecode [1389, 733, 988].

C [431, 993, 1613, 223, 994, 1069, 1704]. Cable [99]. Cache [754, 996, 1607, 1609, 1625, 1593].

Cache-Related [1609]. Cached [884]. Caches [1605, 158]. Caching [1546]. Caesar [97]. Calculating [831]. Calculi [958, 1123]. Calculus [830, 1394, 1039, 309, 723, 973, 1019, 954]. Call [983, 1162, 189, 1004, 1762]. Call-by-Value [1004]. Cam [160]. Cam-E-leon [160]. Camellia [579]. Camera [160, 865]. Can [441]. Cancellation [1634]. Cannot [1185]. CAP [561]. Capacity [1820]. Caption [1284, 1285, 1282]. Capture [640, 847]. Car [269, 1143, 1202]. Car-Like [1202]. Card [1396]. Carlo [1128, 1692, 1749, 1570]. Cars [871]. CASC [1338]. Case [1235, 1240, 1358, 1234, 871, 1815, 1777, 562, 1432, 407, 1215, 1531, 1225, 1315, 1253, 1359, 157, 1745, 1802]. Case-Based [1215]. Casebase [338]. Cases [1423, 1044, 1066]. Casl [1108]. Cast [1089]. Castings [1712]. Categorical [973]. Categoricity [212]. Categorising [702]. Categorization [701, 703, 1273]. Category [499]. Causal [1247, 816, 1056]. Causal-Knowledge [816]. Causality [1375, 1055]. CBIR [1271]. cc [1614]. cc-NUMA [1614]. CCA2 [582]. CCM [1035]. CDEG [1351]. Cell [1629, 1329, 1339, 1328, 1326, 1330]. Cellphones [1403]. Cellular [1812, 504, 1815, 1813, 1155, 1702, 1814]. Census [1340]. Center [1792]. Centric [1444, 1290]. CFB [585]. CFD [1810]. Chains [1525, 1698, 1161]. Challenge [245, 1721]. Challenges [520, 1188, 1480]. Chances [334]. Changing [218]. Channel [1294, 595, 1278, 608]. Chaos [669]. Chapter [829, 830, 834, 827, 832, 831, 833, 835, 828]. Characterising [497]. Characteristics [1691, 578]. Characterization [1344, 58, 960, 1210, 1545, 775, 912]. Chart [146]. Charts [495, 453]. Chauffeur [636]. Cheap [1196]. Cheating [594]. Check [198, 199, 1001, 716]. Checker [1535].

Checking [301, 1389, 306, 733, 1402, 1394, 1111, 746, 1387, 1113, 461, 1401, 307, 462, 314, 1404, 1392, 1456, 753, 1388, 961, 2, 294, 966, 967, 1393, 968, 309, 1003, 293, 816, 1206, 738, 737, 1030, 1049, 470]. Checkpointing [1543]. Chemical [1499, 1806]. Children [332]. Chinese [845, 846]. Choice [747]. Chomsky [785]. Chop [966]. Chorus [235]. Church [784]. CIL [994]. Circal [1029]. Circuits [1104, 1215]. Circular [973]. Circulation [1811, 1624, 1695]. CIS [392]. CISC [250]. Clairvoyant [297]. Class [1829, 7, 97, 632, 812, 1694, 1840, 513]. Class-Constrained [513]. Classes [439, 462, 419]. Classical [264]. Classification [1162, 1184, 1793, 227, 709, 1086, 546, 1174, 1426, 347, 1704]. Classifier [1471, 1464, 1465, 1466, 1470, 1467, 1469, 1474]. Classifiers [232, 1093]. Clausal [341]. Clean [1200]. Client [543]. Clique [515, 267]. Clocks [810]. Close [906]. Close-Open [906]. Closed [1284, 1285, 896]. Closing [738]. Cloud [1602]. CLP [1120, 1227]. CLPS [303]. CLPS-B [303]. Cluster [1629, 1635, 1551, 1559, 1567, 265, 1764, 1612, 1560, 1621, 1674, 1665, 1554, 1814, 1556, 381, 1573]. Clustered [56, 1605]. Clustering [229, 1092, 702, 1565, 1083, 1082]. Clusters [1585, 1575, 1666, 1784, 1574, 1671, 1657, 1590, 1562, 1639, 1646, 1612, 1694, 1652, 1638, 1568]. Co [628, 958, 1821]. Co-algebraic [958]. Co-generation [628]. Co-Operative [1821]. Coalgebra [955]. Coalgebraic [833]. Coalgebras [829]. Coalition [1485, 815, 1501]. Coalitions [818]. Coarse [1261, 1572, 1625]. Coarse-Grain [1572]. Coarsest [318]. Coastal [1624]. Cobra [564]. Code [1296, 1629, 1457, 493, 993, 613, 250, 1777, 1650, 1003, 1313, 236, 1000]. Codes [590, 761, 1820, 716]. Coding [1276, 1690, 860]. Coefficient [1707]. Coercion [235]. Coevolution [1151, 253, 1155]. Cognitive [1804, 724]. Coherence [754, 4, 1589]. Coherent [923]. Coke [1808]. CoLaS [1302]. Colimit [1306]. Colimit-Based [1306]. Collaboration [1575]. Collaborative [1233, 140, 141, 1553, 1238, 139, 115, 1101]. Collapsing [768]. Collection [207, 1840]. Collections [1220]. Collective [1815, 819, 822]. Collision [1206]. Collusion [590, 1452]. Collusion-Secure [590]. Colony [287]. Color [1261, 419, 1263, 1266]. Colored [419]. Coloring [488, 769, 418, 1140, 417, 268]. Colors [418]. Colour [1209]. Coloured [1425]. Columnwise [1675]. Combination [885, 697]. Combinational [1215]. Combinatorial [899, 711, 259]. Combinators [941, 940]. Combined [1761]. Combining [1117, 733, 1118, 741, 689, 1114, 228, 343, 1725, 690, 1115]. Combustion [355]. Commerce [597, 544, 1555]. Commercial [1417]. Commitments [800]. Commodity [1590]. Communicating [787, 1241, 871, 1069]. Communication [1301, 787, 880, 1466, 1641, 1307, 1355, 434, 1567, 1076, 1796, 1644, 422, 969, 1650, 1381, 878, 1618, 788, 412, 1573, 159, 1345]. Communications [711, 1832]. Communicative [1237]. Community [1737]. Commutative [975]. Compact [970, 1475]. Compaction [62]. Comparative [1471, 689, 1088, 1198]. Compare [1260]. Comparing [1401, 254, 264, 448]. Comparison [1194, 623, 261, 1249, 1687, 1803, 1782, 1230, 1739, 494, 1717, 869]. Comparisons [1144, 1619]. Compatible [949]. Competitive [34]. Compilation [987, 154, 1191]. Compiler [1585, 981, 1630, 1198]. Compilers [1627]. Compiling [993]. Complementary [1329, 143]. Complete [1641, 460, 452, 828, 726]. Completeness [301, 48, 49, 443, 974].

Completeness-Driven [49]. Complex [1781, 1577, 1369, 529, 1687, 1768, 38, 18, 365, 1365, 1105]. Complexity [432, 1119, 420, 424, 439, 1368, 1661, 444, 542, 1229, 434, 34, 419, 446, 778, 1020, 763]. Compliant [1494]. Component [1442, 1415, 666, 1510, 36, 1028, 1034, 651, 1697, 120, 1436]. Component-Based [1442, 666, 1034, 120, 1436]. Component-Oriented [1510, 36]. Components [1035, 1298, 804, 198, 1412]. Composed [923, 1509]. Composing [644]. Composite [1112]. Composition [1294, 1026, 526, 1541, 1410, 1323]. Compositional [1027, 954, 740, 1352, 322, 158]. Comprehensible [376]. Comprehension [1346, 1354]. Comprehensive [673, 1001, 184, 1594]. Compressible [1828]. Compressing [1008, 344]. Compression [890, 343, 982]. Compressors [1253]. Computable [215, 283]. Computation [595, 1483, 264, 630, 1145, 691, 1796, 1137, 1692, 435, 972, 364, 1089, 718]. Computational [1787, 1119, 1756, 622, 1631, 1838, 1810, 240, 211, 356, 1722, 332, 440, 1808, 1555, 1786, 725, 1753, 271, 359, 1105, 468]. Computations [1614, 1017, 1691, 915, 1016, 1652, 427, 1754, 1705]. Computer [89, 1127, 837, 1248, 1344, 465, 1584, 1707, 1673, 842, 1804, 1603, 1806, 715, 1665]. Computer-Assisted [1806]. Computerized [1351]. Computers [1586, 1682]. Computing [127, 617, 604, 1184, 1666, 1784, 886, 1828, 1553, 1822, 1549, 1670, 1551, 871, 1823, 1688, 125, 430, 1311, 909, 1778, 545, 1750, 1622, 360, 1674, 1825, 130, 1786, 1582, 511, 129, 1602, 1556, 120, 128]. Concealment [1276]. Concept [950, 794, 1162, 1667, 879, 1836, 697]. Concepts [951, 115, 1413, 1171, 329, 1405, 1478].

Conceptual [1380, 1738, 1220]. Conceptualization [1046]. Conceptualizations [1345]. Concerning [781]. Concerns [1511]. Conclusion [50, 201, 945, 1059]. Conclusions [729, 394]. Concrete [585]. Concurrency [917]. Concurrent [1293, 1007, 952, 1660, 319, 176, 1412, 1724]. Condition [1029]. Condition/Event [1029]. Conditions [1712, 280, 661, 749]. Conducting [1355]. Configurable [1639, 1646]. Configuration [12]. Confirmation [588]. Conflict [964]. Conflicts [388, 380]. Confluence [1110]. Conformance [1669]. Congested [371]. Congestion [1587, 1763]. Congruential [784]. Conjugate [1563]. Conjunction [311]. Conjunctive [523]. Connected [906, 1603]. Connecting [571]. Connection [723]. Connectionist [1703]. Connections [830, 1033, 1556]. Conquer [545]. Consequence [1154]. Considering [651]. Consistency [1566, 145, 1655]. Consistent [30]. Constant [485]. Constant-Space [485]. Constants [1729]. Constrained [306, 482, 1212, 625, 1775, 1352, 1213, 640, 513, 1150, 925, 171]. Constraint [1007, 666, 952, 303, 1118, 1113, 1121, 264, 1123, 1107, 1222, 1724, 941, 932]. Constraint-Based [666, 1113, 1107]. Constraint-Lambda [1123]. Constraints [473, 612, 432, 433, 1042, 296, 180, 1120, 1819, 1662, 446, 995, 637, 1068, 1069, 1112]. Constructing [369, 801, 1707, 181, 338, 1805]. Construction [983, 981, 1657, 290, 1178, 38, 1376, 385, 384, 386, 391, 1286]. Constructions [594]. Constructive [209, 213, 212, 216, 1717]. Constructs [1619]. Consumers [1168]. Consumption [998, 1593]. Containment [545]. Content [1081, 140, 1459, 211, 689, 1281, 1738, 887, 564, 1263, 1452, 598, 1265, 690, 561]. Content-Based [689, 1281, 887, 564, 1263, 1265, 690]. Contest [111]. Context [135, 7, 1042, 680, 874, 780, 875, 995, 771, 974, 40, 538, 792]. Context-Aware [875]. Context-Based [135, 1042]. Context-Free [780, 974]. Context-Sensitive [771]. Contexts [1364, 526]. Contextual [118]. Contingency [1327]. Continuations [1012]. Continuous [620, 117, 1141, 844, 531, 1723, 1150, 532]. Contour [1837]. Contract [1486]. Contradictory [1782]. Contribute [1376]. Control [615, 866, 604, 153, 621, 1419, 1182, 1160, 226, 1498, 247, 623, 1023, 28, 457, 624, 1325, 625, 33, 189, 888, 628, 801, 1500, 848, 868, 1246, 865, 964, 634, 1084, 1711, 1663, 684, 1341, 637, 638, 639, 1214, 199, 1225, 1094, 1146, 1253, 1753, 641, 1150, 643, 645, 1495, 869, 876, 241, 867]. Controller [629]. Controlling [1547]. Convection [1726]. Convection-Diffusion [1726]. Convergence [1700]. Conversational [849]. Conversations [4, 1486]. Conversion [475, 1199]. Convex [507]. Cookies [587]. CoolBOT [36]. Cooperating [1482]. Cooperation [818, 129, 1291]. Cooperative [1151, 1240, 1722, 1256, 1152, 1319, 385]. Coordinate [54, 83]. Coordinates [608]. Coordinating [1296, 1304, 1318, 823]. Coordination [1292, 1294, 1299, 1302, 1303, 1305, 1307, 1047, 1308, 1315, 1317, 1290]. Copies [511]. Copy [1462]. Copyright [1461, 1462, 591]. Coq [212]. Coral [1743]. Core [1666, 1736]. CoreMedia(R) [561]. Corporate [232]. Correcting [273]. Correction [1776, 1216]. Correctness [952, 1194, 1195]. Correlation [262, 586]. Correspondence [735]. Corresponding [1350]. Cost [1307, 1096, 1197, 1592]. Costs [1796, 1609]. Cotree [58]. Counter [961]. Counterexample [1384]. Countermeasure [609, 607]. Counting [1227]. Coupled [1809, 1817, 1770, 1777, 544]. Coupling [542, 1712]. Courses [134]. Coverage [312, 1014]. Covering [262]. Covers [783]. CPlanT [815, 1501]. CPS [956]. Cracking [1808]. Crafty [1450]. Craters [354]. Crates [1834]. Crawler [278]. Creation [117]. Crew [1072]. Criteria [1305, 1388, 1662, 44]. Criterions [333]. Critical [58, 60, 885, 338]. CRL [1314]. CrocoCosmos [90]. Cross [604, 258, 707, 1381, 1198]. Cross-Domain [604]. Cross-Entropy [258]. Cross-Language [707]. Cross-linguistic [1381]. Cross-Platform [1198]. Crossing [59, 60, 61]. Crossing-Critical [60]. Crossover [1104, 243]. Cryptanalysis [580, 1459, 593, 579, 581, 607, 606]. Cryptanalytic [578]. Cryptographic [436, 1397, 736]. Cryptography [471, 577]. Crypton [580]. Cryptosystem [582]. Cryptosystems [583, 584]. CSP [1721]. CTI [1050]. CTR [585]. CTR-CFB [585]. CTR-OFB [585]. CTW [343]. Cubic [481]. Cultural [1212, 1138]. Current [1760, 156]. Curry [211]. Curse [1754]. Curve [608]. Curves [479, 894]. Customer [1162]. Customization [672]. Cut [66, 525]. Cut-and-Pick [525]. Cycle [1598]. Cycles [430].

D [1771, 918, 911, 624, 1817, 88, 898, 192, 1607, 1096, 161, 1820, 1711, 915, 900, 1806, 1080, 27, 1353, 1766]. D.I.K.E. [387]. DAGs [416]. Damages [1834]. DAML [1293, 519]. DAML-S [1293]. DaParT [1664]. Data [540, 572, 1357, 331, 1781, 1471, 1162, 141, 1180, 1333, 880, 1636, 1472, 539, 169, 537, 1542, 1328, 357, 79, 1680, 1249, 333, 353, 278, 1015, 358, 556, 1087, 1796, 1795, 714, 1168, 279, 218, 342, 562, 1778, 1171, 984, 1274, 361, 48, 1287, 334, 343, 546, 1664, 339, 543, 1794, 548, 349, 384, 386, 391, 544, 552, 1596, 1834, 982, 557]. Data-Dependent [1542]. Data-Flow [1778, 984, 1274]. Data-Intensive [1795]. Data-Parallel [1664]. Database [92, 551, 570, 542, 1252, 1226, 1269, 1273, 1275, 560]. Database-Supported [570]. Databases [1421, 331, 1167, 1325, 688, 1234, 531, 1274, 1064, 530, 546, 543, 533, 1272, 366, 392, 1058, 1057, 1055]. Dataflow [145]. Datasets [1092]. Dataspaces [1300]. Dates [489]. DBdraw [92]. DCT [1277]. DCT-Domain [1277]. DDG [1674]. Deadline [1500]. Deaf [851]. Deal [1177, 1754]. Debug [177]. Debugging [319, 194, 176, 1366]. Decennial [1340]. Decentralised [194]. Decentralized [586]. Decidability [295]. Decidable [71]. Decision [1117, 229, 739, 1751, 760, 1114, 440, 228, 1168, 345, 337, 1066, 348, 1163]. Decoding [581]. Decompiling [988]. Decomposable [462, 1068]. Decomposition [516, 1677, 1767, 897, 1770, 1687, 1683, 1769]. Decompositions [770]. Decouple [1430]. Decryption [582]. Dedicated [1643]. Deduced [357]. Deduction [1124, 1725]. Deductive [1100]. Defeasible [1067]. Defeat [1483]. Defense [588]. Defensive [735]. Definability [1125, 463]. definable [742]. Defined [404]. Defining [919]. Definition [1372, 1097, 1418, 660, 344]. Deforestation [1152]. Deformable [925]. Deformations [1248]. Degree [76]. Delaunay [1770]. Delay [422]. Delayed [1730]. Delays [1641]. Demand [983]. Demonic [957]. Demonstration [569, 1050, 1406, 202, 1405]. Demonstrator [566]. Dense [293]. Dense-Time [293]. Density [1142, 626, 227, 716]. Density-Classification [227]. Deontic [1237]. Dependence [1606, 152]. Dependencies [527, 547, 1430]. Dependency [1024, 1019]. Dependent [207, 296, 1730, 475, 1542, 1642, 1816, 503]. Deposition [1808]. Derivation [1532, 18, 388]. Derivations [330]. Derivative [970]. Deriving [251].

Describing [439, 1373]. Description [1476, 1303, 676, 1261, 1280]. Descriptions [1663, 510]. Descriptive [1426]. Descriptors [1281]. Design [1048, 619, 1045, 664, 861, 666, 132, 15, 603, 160, 3, 1380, 1741, 803, 1310, 146, 556, 5, 1186, 584, 885, 148, 668, 1143, 1844, 1824, 281, 1753, 324, 1289, 159, 1552, 989, 1802, 1080, 1290, 882, 553, 16, 149]. Designing [17, 1382, 527, 359]. Designs [711, 1030, 1049]. Desires [799]. Detecting [66, 875, 1220]. Detection [515, 1284, 1285, 1209, 278, 964, 586, 1739, 351, 280, 1844]. Determine [486, 972]. Determining [13]. determinism [1197]. Deterministic [1135, 774, 483, 1701]. Detour [430]. Developers [189]. Developing [1786, 1664, 134]. Development [1177, 1108, 1422, 803, 1047, 1187, 125, 979, 971, 841, 1063, 1034, 651, 683, 1824, 1411, 1506, 660, 120, 16, 1436, 1598]. Developmental [1492]. Developments [1493, 156]. Deviations [1220]. Device [1844]. Devices [671, 889]. DFS [58]. Diagnosis [1257, 1079, 2, 1173, 635]. Diagnostic [1257]. Diagram [429, 1367, 1377, 1361, 1374, 1371]. Diagram-Based [1367]. Diagrammatic [1357, 1364, 1344, 1358, 1351, 1348]. Diagrams [7, 1362, 1369, 1379, 1380, 1347, 1349, 97, 1350, 1352, 1375, 177, 55, 1354]. Dialects [1302]. Dialogue [1075, 841]. Dialogues [1481]. Diameter [486, 909]. Dictionaries [706]. Dictionary [344]. Dietary [1172]. Diff [1157]. Diff-Serv [1157]. Difference [1817, 365]. Differences [480]. Different [260, 1779, 1650, 248, 1652, 1701, 126]. Differential [617, 612, 1728, 580, 1730, 579, 1710, 1681]. Differentiated [1762]. Differentiation [1284]. Diffusion [1726, 1798, 243]. Digital [1461, 904, 1460, 924, 1459, 1453, 575, 916, 1613, 1454, 927, 1283, 560]. Digitisation [293]. Digitized [915]. Dihomotopy [472]. Dilemma [1779]. DIM [1579]. DIME} [1219]. Dimension [504, 483, 705, 339]. Dimensional [1564, 1829, 1629, 1327, 1815, 358, 1692, 1716, 1831, 76, 1654, 424, 1830, 339]. Dimensionality [1754]. Dimensions [77, 897, 67]. Diophantine [1662]. Direct [1371]. Directed [53, 1393, 68]. Directing [142]. directional [1206]. Disambiguation [990]. Disassembly [814]. disciplinary [1802]. Discipline [1523]. Disclosure [1334, 1341, 1331, 1337, 1335]. Disconnected [81]. Discouraging [1448]. Discover [368]. Discovering [375, 361]. Discovery [1347, 227, 373, 228, 1287, 328, 334, 1215, 370, 376, 363, 364, 359, 344, 340, 366, 349, 367]. Discrete [919, 621, 921, 928, 583, 920, 627, 479, 1774, 892, 929, 1841, 1150, 907, 642, 1697, 912]. Discrete-Time [627, 642]. Discretization [918]. Discussion [677, 670]. Disease [1275]. Disjoint [514]. Disjointness [434]. Disjunctive [643]. Disks [429]. DisMedJava [1580]. Dispersion [1133]. Displacement [1677]. Displaying [902]. Disruption [286]. Dissipation [996, 997]. Distance [226, 262, 1096, 1083, 913, 1260]. Distances [578]. Distortion [78]. Distributed [540, 1586, 1485, 1623, 153, 666, 1575, 1504, 1491, 1498, 1299, 1655, 1668, 1571, 1685, 1580, 1797, 675, 296, 1558, 1254, 1648, 6, 1574, 254, 1500, 1483, 1811, 1488, 1516, 1520, 1610, 1796, 1731, 1311, 875, 1579, 1825, 1714, 1312, 878, 1801, 939, 1725, 821, 1570, 1569, 1799, 1814, 1552, 1821, 299, 602, 824]. Distribution [1134, 1761, 1672, 1526, 578, 352, 1474]. Distributions [699]. Diversity [226, 238, 671, 1129, 670]. Divide [545]. Divide-and-Conquer [545]. Divisible [1647]. Do [1344]. Doctor [1386].

Document [695, 696, 698, 705, 690, 704]. Documentation [659]. Documents [1100, 529, 702, 528, 706]. Does [1344]. Domain [1421, 604, 1422, 1419, 482, 1363, 1240, 1767, 1817, 709, 1120, 1414, 871, 1770, 1087, 1277, 667, 1216, 1769, 681]. Domain-Specific [1363]. Domains [171, 1105, 126]. Dominance [433]. Domination [516]. Domination-like [516]. Domino [903, 574]. Down [348]. Downconversion [1277]. DPA [609]. Drawer [104]. Drawing [74, 96, 92, 111, 91, 65, 83, 108, 69, 187, 67, 73, 68, 70, 1381, 76]. Drawings [75, 63, 62, 77, 82, 1353]. Drawn [1354]. DrawNET [1540]. Drift [1162]. Driven [983, 664, 1235, 1409, 1165, 1278, 49, 46, 1046, 550]. DRM [1458]. Drug [228]. DSC [1646]. DSM [1566]. DSM-PM [1566]. DSP [162, 668]. Dual [1624, 1149, 717]. Dual-Level [1624]. Dualization [421]. Due [489]. Duplication [422, 225]. Durational [967]. During [1172, 1366, 651]. Dynamic [1632, 1378, 1838, 624, 1237, 1196, 1250, 1546, 987, 1456, 415, 1813, 535, 1400, 1538, 1561, 1645, 412, 1411, 1376, 1578, 282, 1556, 27, 568, 574]. Dynamical [1247, 618]. Dynamically [1298, 1639, 1646, 1568, 982]. Dynamics [1829, 1142, 1729, 1810, 12, 1775, 1249, 355, 2, 806, 1650, 1791, 1837, 1696].

E-Commerce [544, 1555]. e-Market [1759]. e-Science [1736, 1583]. e-Services [563]. e-XML [537]. Eagerness [1196]. Early [11, 1412]. Early-Reply [1412]. Earth [1600]. EAs [276]. EASEA [1144]. Easing [672]. Easy [423, 418, 963]. ECA [1252]. Eclipse [1788]. Ecology [1745]. Economics [1434, 652]. Ecosystems [194, 1152]. Eddy [1807]. Eden [1190]. Edge [514, 503, 1844]. Edge-Disjoint [514]. Edges [65]. editing [116, 377]. Editors [1372, 1371]. EDOC [1035]. Education [179]. Educational [1079]. Effect [1523]. Effective [226, 999].

Effects [1362, 959, 1761, 1606, 1136]. Efficiency [1552, 553]. Efficient [540, 331, 673, 1809, 529, 526, 1559, 534, 1590, 962, 600, 589, 528, 1711, 342, 547, 816, 148, 1451, 1754, 1595, 1844, 159, 532, 1272, 544, 883, 1112]. Efficiently [360, 283]. Effort [240]. EHL [1788]. Elastic [1709]. Elasticity [1677]. Elastostatics [1716]. Election [960]. Electric [1474]. Electricity [1379]. Electromagnetic [1629]. Electronic [501, 597, 1480]. Element [1828, 1706, 1687, 970, 1663, 1827, 1622, 1826, 1837]. Elementary [757, 1836, 1310]. Elements [1663]. Eliminating [788]. Elimination [1675, 1116, 219, 1001]. Elitism [1701]. Ellipse [1788]. Elliptic [479, 608]. Elongated [913]. Embedded [321, 657, 189, 1508, 1307, 613, 500, 146, 1122, 148, 611, 658, 1556, 882, 156]. Embeddings [78, 76]. Embodiment [1490]. Emergence [1470]. Emergencies [1479]. Emotion [1178]. Emotional [1075]. Empirical [132, 1329, 1687, 1230, 727, 1342, 1428]. Enabled [1742, 162, 889]. Enabling [1160, 1546]. Encapsulation [168]. Encoding [746, 697, 235]. Encounters [864]. Encryption [411]. End [1143]. Endogenous [1295]. Endoscopic [1689]. Energy [998, 1609, 1595, 1593, 159, 553]. Energy-Efficiency [553]. Engine [1363, 1339, 1256, 687, 1569, 548]. Engineering [1421, 1442, 657, 1038, 1444, 1517, 1810, 1036, 1453, 187, 115, 1183, 649, 174, 1413, 667, 173, 1043, 13, 1824, 681]. Engines [1122, 935]. English [1217]. Enhanced [1158, 692]. Enhancement [999]. Enhancing [1415, 587]. Enriching [1039]. Enrichment [117]. Enterprise [1822, 115]. EnterTheGrid [1741]. Entries [1327]. Entropy [258, 767]. Enumeration [730]. Enumerative [455]. Envelope [1599]. Environment [604, 1553, 539, 1537, 1822, 1047, 1192, 1355, 1373, 805, 1803, 1689, 1406, 176, 684, 979, 842, 1203, 1825, 133, 1395, 1034, 1801, 1824, 129, 1802, 1821]. Environmental [1153, 1768]. Environments [1238, 1823, 22, 839, 120, 552]. EO [1144]. EPIC [1608]. Epistemic [1226, 1291]. EPL [1705]. EPL-Julia [1705]. Equality [612, 434]. Equally [418]. Equation [620, 1730]. Equational [758, 246, 1725]. Equations [1728, 1116, 451, 327, 1727, 1816, 1710, 1827, 1789, 1713, 1818, 1681, 1715, 151, 367]. Equivalence [1533, 961, 737]. Equivalence-Checking [961, 737]. Era [872]. Erdos [467]. Error [1276, 1679, 1090, 276, 320, 1709, 426, 273, 1843]. Error-Correcting [273]. Errors [1016, 1207]. Escape [743]. Estimated [842]. Estimates [1709]. Estimating [1648, 1796, 554]. Estimation [1729, 259, 1707, 631, 1134, 1613, 915, 1714, 1526, 1831, 555]. Etch [1766]. Ethernet [1764]. Ethical [837]. Euclidean [1351]. Euler [1829, 1349, 1350]. Evaluating [983, 687]. Evaluation [1628, 1329, 1537, 689, 1649, 1764, 287, 1217, 727, 1627, 1749, 550, 558, 708, 1617, 1080]. Evaluations [949, 426]. Event [1300, 1254, 586, 1738, 1841, 878, 1697]. Event-Based [1300, 1738]. Events [165, 1241]. Everlasting [411]. Evidence [1245, 690, 693]. Evidences [1342]. Evidential [37]. Evoking [861]. Evolution [232, 1759, 1304, 234, 1139, 1085, 654, 1514, 248, 1070, 249, 270, 283, 561]. Evolutionary [1787, 618, 242, 1726, 1685, 274, 1153, 260, 1223, 264, 1145, 1137, 279, 1138, 1149, 1143, 235, 1133, 1150, 271, 1136, 284, 1705]. Evolvable [281]. Evolving [232, 1130, 229, 1815, 1145, 1155, 1152, 1179, 282]. Exact [496, 712, 708, 152]. Exact-Match [708]. Example [1235, 1353]. Example-Driven [1235]. Examples [21, 337]. Exception [987, 363]. Exceptions [1012]. Exchange [596, 188, 825]. Exchanges [1555]. Exchanging [109]. Excitation [1789]. Exclusion [1655, 1640, 326]. Executable [1030, 1049]. Executing [208, 1663]. Execution [1623, 1776, 296, 1648, 175, 1412, 1071, 1251]. Executions [536]. Exercise [1421]. Exercises [833]. Existential [451, 1005]. Exogenous [1295]. Exons [236]. Expanding [1599]. Expansion [694]. Expansions [481]. Expectation [15]. Expectation-Oriented [15]. Experience [1437, 680, 1326]. Experiment [1636, 748, 1785, 701]. Experimental [1389, 244, 268]. Experiments [1355, 1806, 391, 388, 390, 389]. Expert [807]. Explanation [193]. ExPlanTech [1497]. Explicit [497, 226, 1401, 974, 1365, 299]. Explicitly [1679, 1670]. Exploitation [1624, 1497]. Exploiting [262, 1120, 629, 1313, 40]. Exploration [89, 1464, 1534, 148, 1207]. Explorations [1473]. Exploratory [180]. Explorer [1788, 938]. Exploring [308, 1792, 1733]. Explosion [461]. Expressing [1574]. Expression [227, 562, 674]. Expressions [452]. Expressive [1359]. expressiveness [1012]. EXPSPACE [460]. EXPSPACE-Complete [460]. Extended [1067, 214, 115, 374, 5, 1646, 1386, 1317, 1105]. Extending [1131, 1407, 1328, 1391, 1430, 1322]. Extensible [1033, 676, 1018]. Extension [1200, 478]. Extensions [1628, 1295, 1329, 1040, 1455]. External [457, 1376]. Extracting [1230, 1365]. Extraction [1833, 526, 26, 358, 1097, 1269, 687, 598, 1267, 349, 383, 390, 389, 382, 381, 380]. Extrapolation [1031]. eXtreme [1179]. Eye [191].

F [570]. Face [859]. Face-to-Face [859].

Facilitate [673]. Facilities [1691, 1430, 1322]. Facility [482]. Facing [1159]. Factor [1166]. Factories [1430]. Factorization [1559]. Factors [1098, 786]. Failures [1543, 1022]. Fair [317, 596]. Families [785, 960, 654, 679]. Family [676, 671, 649, 1698, 651, 660, 681, 665]. Farol [1470]. Fast [54, 582, 1282, 62, 1644, 39, 584, 608, 200, 558, 1262, 1764]. Fat [65]. Fate [320]. Fault [1800, 239, 197, 1159, 1554, 606]. Fault-Behavior [197]. Faults [1257]. Faulty [511]. FCA [697]. FCA-BAM [697]. FD [1120, 1227]. FDBS [542]. FDTD [1577]. Feasibility [924]. Feature [354, 1413, 1208, 1269, 1083, 1509, 1267, 1231, 1262]. Feature-Based [1083]. Features [675, 689, 907, 1264, 126]. Feedback [625, 1578, 1265, 693]. Feedforward [869]. Feeding [35]. FEM [1819, 1652]. FEM-Based [1819]. Few [75]. Fibre [1245]. Fidelity [181, 1280]. Field [1132, 531, 1831, 1174, 149]. Fields [1756, 899, 130, 478]. Fiffo [500]. Figures [204]. File [1571, 1334, 1651]. Film [1731]. Filter [581, 887, 281]. Filtering [1128, 141, 37, 709, 139, 1267]. Filters [990, 906]. Final [965]. Finding [57, 503, 362, 346, 341, 224]. FINDS [1245]. Fine [879, 1261, 311]. Fine-Grain [311]. Fine-Grained [879]. Finger [846]. Finger-Spelling [846]. Fingerprinting [590, 1269, 1452]. Finite [1130, 1828, 1023, 1817, 1706, 1687, 893, 1016, 1622, 477, 448, 470]. Finite-Control [1023]. FIPA [1494]. Fire [1818]. First [316, 956, 1400, 121, 367]. First-Order [956]. Fit [361]. Fitness [262, 251, 282, 283]. Fitting [1527]. Fixed [830, 1131, 1182, 52, 758, 966]. Fixed-Length [1131]. Fixed-Parameter [52]. Flexibility [1550, 147]. Flexibility/ Cost [147]. Flexibility/Cost-Tradeoffs [147]. Flexible [1398, 1676, 798, 1488, 357, 605, 1228]. Floating [1017]. Floating-Point [1017]. Flood [1790, 1686]. Floor [80]. Floor-Planning [80]. Flow [1771, 509, 1141, 1828, 275, 261, 189, 734, 169, 1643, 1778, 984, 1274, 1693, 1826, 280, 1773, 1253, 1018, 1659]. Flows [1329, 1807, 1691, 1719]. Fluid [1810]. fMRI [340]. Focus [1366]. FOHM [119]. Follow [863]. Font [1284]. Foraging [1153]. Foraminiferal [1744]. Forbidden [786]. Force [622, 250]. Force-Guided [622]. Forensic [1245]. Forest [1818]. Forests [421]. Form [995, 608, 1089]. Formal [1048, 735, 797, 1040, 783, 1486, 1041, 804, 1396, 1102, 304, 1505, 1046, 1043, 1314, 822]. Formal-Driven [1046]. Formalism [1529, 1541, 1187, 1037]. Formalization [1429]. Formalizing [1297, 216]. Formally [217]. Formation [815]. Formed [1076, 1612]. Forms [1080]. Formulas [1354]. Formulation [958]. Formulations [643]. ForSpec [310]. Fortran [1629, 1626]. Fortresses [1431]. Forward [740]. Forwarding [989]. Fossil [1246]. Foundation [916]. Foundations [795, 1107, 359]. Fourier [718]. Foxy [1450]. FPGA [1663, 1683]. FPGA-Based [1663]. FPGAs [223]. Fractal [1690]. Fractional [1767, 417, 1710]. Fragment [460, 786]. Fragmentation [1651]. Fragments [759]. Framework [838, 622, 899, 1239, 1429, 36, 1302, 1028, 1268, 1237, 1496, 1559, 192, 1114, 556, 1484, 1173, 1263, 736, 1318, 660, 1320, 27, 1321]. Frameworks [3]. Frameworks-Based [3]. Framing [1177]. Free [1394, 780, 320, 772, 984, 1077, 747, 974]. French [857]. Frequency [699]. Frequent [420, 1175, 1592]. Front [1143]. Frontal [1809]. Froth [1775]. Fujitsu [1628]. Full [293]. Fully [1809, 750, 1603]. Fully-Coupled [1809]. Function [1544, 1229, 1096, 1213, 1150, 40, 282, 1089].

Function-Based [40]. Functional

[300, 1608, 1344, 1524, 831, 1613, 304, 527, 1193, 399, 400, 997, 157]. Functional-Level [1613]. Functioning [1758]. Functions [1125, 227, 1836, 712, 1144, 742, 403, 511, 283, 1843]. Fundamental [213, 427, 1716, 1718, 284]. Further [425, 450]. Fusion [1259, 570, 1082, 821]. Future [190, 290, 393, 1059, 1626]. Fuzzy [1247, 1151, 1363, 818, 229, 1763, 252, 1079, 1216, 1094, 1157, 1699, 1703, 1246]. Fuzzy-Logic [1216]. Fuzzy-RED [1157].

GA [1794, 1080]. GA-Based [1794]. Gabor [1267]. Gabriel [506]. Gains [630]. GALE [1471]. GALib [1144]. Galois [830]. Game [1347, 426, 402]. Games [614, 442, 739, 1224, 1289, 737]. Gap [553]. Gas [1829]. Gateway [1786]. Gaussian [1116]. Geist [1539]. Gender [1757]. Gene [227, 562]. General [1828, 1811, 1527, 1145, 605, 1682, 1695, 523]. Generalisation [282]. Generalised [953]. Generalization [1117, 220]. Generalized [990, 493, 758, 462, 479, 1680]. Generated [774, 968]. Generating [420, 1349, 99]. Generation [1421, 1074, 1441, 153, 132, 247, 323, 1104, 1234, 1031, 312, 1572, 1823, 1770, 1539, 1658, 1711, 1044, 1777, 1663, 1768, 863, 1217, 1214, 661, 1356, 1766, 628]. Generative [1513, 1441, 1447, 1435, 869]. Generator [1246]. Generators [581]. Generic [1200, 1566, 170, 1472, 1028, 961, 21, 1286]. Genetic [696, 226, 1132, 245, 247, 240, 1204, 1104, 275, 229, 238, 285, 254, 308, 1169, 252, 244, 277, 246, 239, 251, 265, 1610, 1698, 1072, 1210, 1213, 1832, 267, 223, 1070, 1148, 1215, 243, 1147, 280, 1146, 1831, 249, 1700, 1701, 236, 225, 1080, 241]. Genre [709]. Genres [1301]. GeoCast [879]. Geodesic [920]. GeoFEM [1621]. Geomagnetic [351]. Geometric [170, 1449, 30, 555]. Geometries [1775]. Geometry [904, 920, 916, 32, 265, 1351, 35, 533, 468, 917]. GeoWin [170]. Gestural [866]. Gesture [838, 840, 841, 852, 869, 849]. Gestures [861, 864, 859, 865, 851, 842, 852, 839, 860, 862, 867]. ’Get [1185]. GHQUAR [1339]. Giant [1377]. Gigabit [1396, 1694]. Giganet [1764]. Gilgamesh [1582]. GILK [1538]. Given [474, 1327]. Givens [1683]. GLIDE [106]. Global [996, 357, 1096, 634, 1760, 1723, 1700]. Goldreich [436]. Good [83]. GP [230]. GP2002 [1435]. Gradient [258, 1563]. Grain [1572, 311, 1625]. Grained [879]. Grammar [787, 244, 1532, 1147, 1356]. Grammar-Based [1532, 1147, 1356]. Grammars [300, 779, 780, 989]. Grammatical [232, 234, 248, 249]. Grand [1432]. Granular [1810]. Granularities [557]. Granularity [1649]. Graph [74, 983, 1108, 96, 95, 111, 474, 91, 451, 88, 108, 189, 1346, 81, 69, 187, 1140, 53, 896, 417, 230, 419, 713, 73, 984, 68, 70, 1842, 1532, 86, 715, 273, 188, 76]. Graph-Drawing [111]. Graph-Free [984]. Graph-Theoretical [715]. Graphical [861, 1355, 1345, 704, 126]. Graphics [1008, 1804]. Graphing [1792]. GraphML [110]. Graphs [96, 56, 506, 488, 72, 902, 1678, 486, 83, 58, 960, 372, 60, 67, 102, 84, 419, 422, 456, 1220, 971, 61, 185, 71, 268, 82, 93, 109, 178]. Graphviz [108]. Grasping [25]. Gravitation [810]. Grayscale [923]. Greedy [1642]. Greek [856, 1076]. Grid [1667, 1549, 1830, 1316, 1695, 1824, 1736, 1583]. Grid-Adaptive [1830]. Grids [1787, 1784, 1631, 77, 1737, 487, 1621, 1693, 913]. Ground [456, 968]. Group [1765, 1254, 823]. Grouping [365]. Groups [138, 1270]. Growing [771]. Growth [1731, 1743, 236]. Gsm [603]. GSOS [458]. GSPN [1534, 1530]. Guaranteed [620]. Guarantees [563]. Guarded [460]. Guards [1457]. GUI [1409]. GUI-Driven [1409]. Guided [622, 1384, 1090, 244, 1578, 273, 1000]. Guidelines [1413]. Guiding [1367]. GXL [109, 188].

Haar [362]. Hacking [588]. Halting [216]. Hamilton [620]. Hand [868]. Handel [223]. Handel-C [223]. Handling [572, 1121, 671, 1172, 1494]. Hard [61, 454, 224]. Hardware [223, 281, 282, 606]. Harmony [1240]. Hashing [1449]. Having [361]. Haystacks [224]. Hazard [1426]. Head [856, 1098]. Header [890]. Heart [1789, 563]. Heating [1817]. Heavy [1523]. Height [476]. Heilbronn [483]. Helps [368]. Hermes [89]. Heterogeneous [1357, 138, 354, 1373, 634, 1673, 1691, 1171, 971, 148, 1561, 1645, 1360]. Heuristic [1489, 1404, 1330]. Heuristics [63, 1329, 261, 1112]. Hidden [1077, 478]. Hierarchical [702, 495, 1280, 1698, 633, 70, 1545, 639, 1837, 1752, 1330]. Hierarchical-Element [1837]. Hierarchies [774]. Hierarchy [785, 776, 1191, 1596]. High [1564, 886, 1553, 1459, 1584, 1680, 1785, 993, 358, 1786, 1595, 280, 1405, 1095, 1705, 1629, 1626]. High-Bandwidth [1459]. High-Level [993]. High-Performance [886, 1584, 1705]. High-Resolution [1095]. Higher [208, 750, 504, 327, 963, 1010]. Higher-Order [750, 327, 963, 1010]. Highway [1812]. Higman [221]. Hilbert [329]. Hints [68]. Histograms [1205]. History [1604, 368, 73, 1137, 1626]. Hitachi [1617]. HiTaS [1330]. HiWEP [1626]. Hoc [1797, 1309, 602]. HOL [399, 1108]. Hol-Casl [1108]. Holes [124]. Holographic [505]. Holon [1496]. Holonic [1498, 1499, 1482, 1500]. Holons [1493]. Homicidal [636]. Homogeneous [1662]. Homonymies [380]. Hopfield [1090]. Horizontal [54]. HPC2002 [1584]. HPF [1632, 1628, 1631, 1630, 1627]. HPF/JA [1628]. Hull [507]. Human [1746, 349, 1080]. Humanitarian [1501].

Husserl [329]. Hybrid [614, 615, 616, 1585, 619, 620, 621, 1184, 623, 840, 1336, 624, 628, 1257, 1223, 631, 632, 294, 1621, 635, 1618, 1094, 640, 641, 611, 642, 643, 644, 646, 1105]. Hybrid-Parallel [1585]. Hydrodynamics [1599]. Hygro [1809]. Hygro-Thermo-Mechanical [1809]. Hyper [411, 1279]. Hyper-Encryption [411]. Hyper-Links [1279]. Hypercube [1333]. Hyperdatabases [518]. Hypergraphs [902]. Hyperheuristics [256]. Hypermedia [114, 135, 136, 1079, 115, 142, 125, 692, 118, 1046, 119, 143]. HyperQueries [568]. Hyperspace [142]. Hypertext [132, 143, 126]. Hypertextbooks [184]. Hypertexts [121]. Hyponymies [389, 382]. Hypotheses [360]. Hypothesis [341]. Hysteresis [645].

I-cache [1593]. I/O [1631]. IA [1670]. IA-64 [1670]. IBM [1620]. ICCG [1597]. ICGO [392]. Iconic [839]. Iconicity [854]. ICSR7 [1433, 1436]. ICT [1802]. Ideal [329]. Ideas [1380, 1376]. Identification [1285, 1250, 863, 1077]. Identifiers [203]. Identity [1249]. Idescat [1342]. IEC [1579]. II [69]. Illustrator [1107]. ILOG [86]. ILP [1166]. Image [904, 1580, 902, 28, 911, 688, 285, 1261, 26, 1281, 1689, 1690, 923, 1449, 1579, 122, 362, 1263, 1273, 281, 1266, 1264, 1265, 159, 283]. Image-Based [904, 28]. ImageGrouper [1270]. Imagery [354]. Images [923, 1270, 913, 907, 925, 1260, 340]. Imaging [1095]. Immortality [1137]. Immune [594, 1243, 1702, 606]. Immune-Like [1702]. Immunity [443]. Immunological [1747]. Impact [132, 1417, 1153, 1591, 1607]. Impacts [1493]. Imperative [150, 1005]. Implementation [321, 136, 210, 1508, 1496, 1811, 803, 223, 1826, 1683, 1494]. Implementing [1566, 1029, 1795, 1312, 942, 943].

Implication [1699]. Implication-Based [1699]. Implications [1461, 1460, 1344]. Implicit [629, 693]. Implict [1365]. Importance [655, 1134]. Impossible [580]. Imprecise [1168]. Improve [282]. Improved [516, 496, 580, 59, 425, 1456, 434, 691, 811, 494, 1589]. Improvement [583, 1103]. Improving [292, 746, 1604, 1588, 1014, 1078, 1426, 288, 1231]. Impulse [640]. Impulses [1730]. In-Place [1006, 507, 744]. Inapproximability [512]. Incidents [280]. Inclusion [547]. Inclusions [617]. Incomplete [787]. Incompressible [1727, 1826]. Inconsistent [1782]. Incorporated [1794]. Incorporating [1075, 1150]. Incremental [924, 1465, 541]. IND-CCA2 [582]. Independence [1173]. Independent [844, 235]. Index [20, 978, 1002, 1025, 1051, 1106, 1126, 1156, 1189, 1201, 1258, 1288, 1324, 1343, 1383, 1408, 1446, 1463, 1477, 1502, 1522, 1548, 1581, 1633, 1735, 1845, 41, 112, 144, 163, 195, 222, 255, 289, 325, 378, 464, 517, 576, 610, 647, 686, 710, 719, 755, 790, 826, 836, 870, 891, 930, 397, 410, 732, 948, 1061, 484, 101, 548]. Index-Based [548]. Indexed [688]. Indexing [534, 531, 1738, 707, 1263, 1262]. Indirect [1594]. Individual [1745]. Individual-Oriented [1745]. Indoor [1097, 1207, 1073]. Induced [57]. Induction [1751, 250, 406]. Inductive [369, 499, 353, 332, 1225, 335, 221]. Inductively [404]. Industrial [793, 1181, 1188, 1402, 1437, 1807, 104, 1257, 1103, 31, 1808, 1495]. Industries [1499]. Inequality [612, 428]. Inertial [838]. Inference [1325, 499, 748, 1079, 1256, 1703, 690, 1065]. Inferences [1367]. Inferential [1359]. Inferential-Expressive [1359]. Inferring [1175]. InfiniBand [1588]. Infinite [1660, 968, 765]. Influence [292, 998, 1390]. Influences [1177, 299]. Informal [1041, 1382]. Informality [1187]. Informatic [970]. Information [1177, 1357, 114, 127, 138, 1587, 1461, 292, 1259, 1369, 787, 570, 1268, 1742, 709, 734, 1368, 1488, 142, 374, 194, 692, 43, 44, 1804, 636, 565, 679, 518, 1337, 377, 538, 1365, 1099, 385, 574]. Information-Based [636]. Informational [873]. Infrared [276]. Infrastructure [518]. Infrastructures [16]. Infrequent [420]. Initial [652]. INM [1811]. Innovative [1178]. Input [626, 638, 511]. Insight [79]. INSPIRE [133]. Inspired [1492]. Instabilities [1253]. installment [1647]. Instruction [1670, 133]. Instrumentation [1538]. Intake [1172]. Integer [77, 760, 362]. Integer-Type [362]. Integers [1115]. Integral [1496]. Integrated [1406, 1593, 1030]. Integrating [1048, 1108, 141, 1111, 1301, 1334, 1792, 1409, 542, 659, 43, 521, 1418, 126, 1598]. Integration [114, 127, 951, 1364, 537, 1783, 335, 29, 873, 1073]. Integrators [1839]. Intellectual [116, 377]. Intelligence [520, 1103, 1804, 822, 1105]. Intelligent [1081, 1181, 1100, 1182, 1240, 801, 1164, 1251, 1478, 133]. Intensional [143, 1010]. Intensity [279]. Intensive [1795]. Intentional [805]. Intentions [601]. Inter [1556]. Inter-Processor [1556]. Interaction [1242, 24, 1237, 859, 804, 865, 5, 31, 9, 1509, 1323]. Interactional [863]. Interactions [1247, 1027, 1757, 1312, 819]. Interactive [1481, 1398, 170, 88, 180, 1750, 979, 863, 708, 377, 1805, 665]. Interesting [165, 375, 363]. Interface [861, 1668, 1524, 1396, 842, 177, 1565, 704, 126]. Interfaces [1382, 1827, 841]. Interference [1832]. Intermediate [994]. International [1434]. Internet [872, 466, 1062, 376, 1554, 134, 792]. Interpretable [459, 368]. Interpretation [733, 1385, 748, 745, 1199, 1804, 839]. Interpreter [980, 1017]. Interrupt [1634]. Intersection [72, 452, 1773]. Intersections [1227]. Interval [1116]. Intervals [446, 1717]. Intestinal [1746]. Intractability [431]. Intrinsic [522]. Introducing [685, 1042, 933, 658]. Introduction [131, 827, 190, 179, 537, 164, 648, 173, 185, 42, 196, 113, 720, 931, 716, 765, 123, 379, 1052]. Intrusion [586]. Intuitions [329]. Invariant [74, 1250, 1267]. Invariants [193]. Inventions [225]. Inverse [1708]. Invertible [922]. Investigating [1759]. Investigation [1472, 1122, 248, 268]. Investment [1418]. Invited [1626]. IO [1194]. IP [1544, 1528]. IPM [1755]. IPMP [1455]. IPSI [566]. IQ [822]. IR [700, 697]. IRIS [1788]. Irregular [1614, 1666, 1784, 1000]. ISO [1579]. Isomorphism [419, 273]. Issues [656, 1188, 650, 649, 1555, 881]. Iterated [267, 268, 477]. Iterative [1706, 154, 1449, 1621, 1753]. Itineraries [605]. IVE [1805].

J.CP [1222]. Jacobi [620]. Japanese

[1627]. Java [950, 96, 733, 1575, 1576, 1668, 302, 186, 1574, 162, 1404, 1505, 176, 988, 999, 177, 175, 1840, 1001, 1366]. Java-Enabled [162]. JavaCard [735]. Javacards [603]. JavaSpaces [1314]. JAVAVIS [177]. JaVis [176]. JDI [177]. Jeliot [166]. Jerks [351]. JGraph [96]. Jini [1672]. JINT [1672]. Job [298, 1801, 1647]. Job-Shop [298]. Jobs [296]. Joins [545, 530, 555]. Journey [277]. JSL [855]. Judgements [1782]. Judgments [330]. Juggling [867]. Julia [1705]. Jumps [1032]. Jury [501]. JViews [86].

K-Ary [1641]. KCDSA [599]. KECho [878]. Kernel [1538, 878, 1833]. Kernels [1620, 1565, 640]. Key [582, 583, 584, 1451, 577, 1478, 1048]. Keynote [872, 871]. Keys [270]. Keystroke [1249]. Keyword [1218]. Kit [803]. KMP [485]. Know [565]. Know-ME [565].

Knowledge [1424, 1833, 101, 1346, 1392, 1087, 1175, 373, 1230, 1287, 329, 816, 1224, 370, 376, 821, 359, 366, 383, 390, 694, 1291]. Knowledge-Based [1346]. Known [593]. Known-Plaintext [593]. Koza [240]. Kripke [214, 753, 967]. Kripke-Style [214]. Krylov [1789].

Lab [1806]. Labeled [309]. Labeling [63, 415, 860]. Labelled [1124]. Labelling [1216]. Lagrangian [1150]. Lambda [1123]. LAN [1567]. LAN-WAN [1567]. Lanczos [1679]. Landmark [30]. Landmarks [1097, 1203]. Landscapes [259, 1203]. Langford [1721]. Language [794, 1074, 1293, 310, 1585, 843, 837, 1363, 858, 1303, 676, 12, 6, 758, 844, 1661, 850, 519, 181, 1076, 332, 778, 979, 994, 707, 854, 992, 847, 989, 823, 188, 1356, 856, 845, 846, 857]. Languages [455, 785, 953, 1372, 766, 1036, 1512, 783, 774, 1669, 1307, 1005, 761, 806, 5, 450, 476, 1008, 771, 784, 767, 510]. Laplacian [1645]. Large [331, 105, 1784, 1810, 845, 1765, 1807, 57, 307, 1092, 1706, 308, 1087, 422, 487, 1274, 530, 1681, 1526, 30, 611, 1615, 1602, 19, 1057]. Large-Eddy [1807]. Large-Scale [1784, 1810, 307, 1706, 611, 1602]. Latin [1333]. Lattice [1771, 1807, 338]. Lattice-Based [338]. Lattice-Boltzmann [1807]. Lattices [742, 828]. Law [1461, 621, 367]. Laws [625]. Layer [52, 53, 889, 1554]. Layered [698, 1411]. Layers [75]. Layout [81, 97, 86, 93, 1356]. Lazy [1206, 1202]. LCS [1473]. Lead [1680]. Leaky [118]. Leap [1377]. Learnability [463, 1374, 1138]. Learners [1225, 441]. Learning [369, 1233, 1471, 1131, 1464, 1151, 1833, 1235, 1232, 1154, 1347, 180, 368, 1229, 808, 1084, 1236, 39, 342, 1230, 811, 1088, 1215, 1094, 184, 1806, 819, 38, 336, 337, 1474, 1163]. Learnt [865]. Least [1084, 1826]. Least-Squares [1084, 1826]. Left [757].

Legacy [659, 654]. Lemma [713, 221]. Length [473, 1131, 1604, 243, 1166]. Lengths [446]. leon [160]. Letter [766, 776]. Level [508, 1676, 1524, 884, 1334, 776, 1624, 1785, 993, 1613, 1565, 1405, 149]. Levin [436, 1751]. LHC [1785]. LHS [1336]. LHS-Based [1336]. Libraries [575, 1823, 1778, 1840, 1411]. Library [95, 1145, 373, 1600, 1705]. Life [1276, 1733]. Lifecycle [19]. Lifelike [849]. Lifting [362]. Light [670, 602, 933, 942]. Light-Weight [670, 602]. Lightweight [1630]. LIGO [1636]. Like [1202, 1702, 1159, 516]. Limbs [848]. Limit [215]. Limited [1141]. Linda [1317, 1322]. Line [1442, 1415, 657, 924, 902, 655, 1437, 678, 77, 680, 659, 662, 1607, 1413, 345, 1443, 684, 652, 1440, 661, 1474, 917, 1353, 592, 514, 198, 199]. Line-Graphs [902]. Linear [1119, 621, 226, 924, 626, 1706, 1250, 475, 1229, 630, 230, 1662, 995, 1228, 243, 485, 1682, 1684, 447, 1150]. Linear-Graph [230]. Linear-Time [485]. Linearity [977]. Lines [685, 650, 928, 1414, 1445, 674, 658]. linguistic [1381]. Link [135, 117, 613, 118, 690, 1161]. Linked [1596]. Linking [119]. Links [1279]. Linux [1538]. LISA [979]. List [204, 205]. Lists [440]. Literary [364]. Lithosphere [1696]. Live [167]. Liveness [754]. Load [1488, 1798, 1672, 1673, 1612, 1560, 1603, 1561, 1645]. Loads [1576]. Local [473, 1587, 1142, 695, 1394, 316, 1248, 915, 1400, 267, 309, 268, 269, 819, 881, 1744, 976]. Locality [1592]. Localization [1419, 1282, 23]. Locally [462]. Location [482, 1714, 1313, 1227, 881, 574]. Log [525]. Logarithm [583, 479]. LOGCFL [452]. Logic [1242, 1067, 310, 208, 759, 1237, 627, 312, 332, 966, 1216, 1226, 1228, 1215, 1225, 335, 447, 1291, 341]. Logical [211, 1317, 510, 340]. Logics [793, 1119, 955, 965, 1224]. Long [1221]. Longest [416]. Longest-in-System [416].

Longitudinal [1781]. Look [121, 393]. Loop [153, 1776, 424, 150, 998, 999, 1578]. Loops [624]. Loosely [544]. Loosely-Coupled [544]. LOSEST [431]. Loss [1337, 1474]. LOTOS [981]. Low [78, 181, 280, 1592, 716, 876]. Low-Cost [1592]. Low-Density [716]. Low-Distortion [78]. Lower [59, 626, 414, 961, 445]. LQN [1532]. LR [990]. LTF [1704]. LTF-C [1704]. LTL [746]. LU [1559]. Lukasiewicz [767]. Lukasiewicz-Languages [767]. Lunar [354].

M [1564, 797]. M-Agent [797]. M-Tree [1564]. Machine [1576, 1550, 1496, 1093, 368, 278, 782, 250, 1215, 1749, 1195]. Machine-Holon [1496]. Machines [735, 1130, 1092, 490, 1638, 448]. Macintosh [1635]. Macro [1177, 1012]. Macro-expressiveness [1012]. Magic [467]. Magnetohydrodynamic [1830]. Magnification [1690]. Maintaining [238]. Maintenance [1441, 539, 541]. Making [1066]. MAMSolver [1536]. Man [31]. Man-Robot [31]. Manage [1458]. Management [1424, 1461, 1460, 1239, 798, 6, 1453, 286, 562, 1451, 1275, 1159, 560, 574]. Managements [241]. Manager [1108, 101, 802]. Managing [1120, 1086, 573, 1179]. manifold [897]. Manipulation [1371, 1112]. Manipulations [1695]. Manipulators [1204]. Manufacturing [798, 1482, 1500, 1489]. Many [441]. Map [911]. Mapping [1177, 1035, 165, 1034, 744]. Mappings [339]. Maps [624, 139, 1531, 30, 1073]. Marginal [1327]. Market [1759, 568]. Markets [1760]. Markov [1533, 1132, 1698, 1526, 1077, 1161]. Markovian [1528]. Marriage [512]. MAS [820, 1495]. Masked [1337]. Masking [239, 609, 1337]. Mass [672]. Masses [1635]. Massive [515]. Match [708]. Matching [484, 495, 962, 528, 1208, 1267, 1322]. Material [1810, 1494, 1286]. Materialization [377]. Mathematical [329, 1565, 1354]. Mathematics [215]. Matlab [1833]. Matrices [926, 1680, 760]. Matrix [1688, 427, 1536, 284]. Matrogenic [488]. Matter [805]. Max [270, 496]. MAX-SAT [496]. Maximal [420, 493, 730, 485]. Maximum [430, 267]. Maxwell [1816]. Maxwellian [352]. May [607]. Maya [1108]. McNaughton [785]. ME [565]. Mean [1344, 1643, 630]. Meaning [1377]. Means [840, 671, 246, 183, 351, 225]. Measure [499, 1244, 822]. Measurement [838, 1576, 1649, 1129]. Measures [1700]. Measuring [678, 1133]. Mechanical [1809, 1712]. Mechanics [1706]. Mechanism [1603, 1317]. Mechanisms [1134, 1650, 1760]. Mechanized [1022]. Media [117, 1817, 1691, 148, 116, 377]. Mediated [865]. Mediators [1099]. Medical [1580, 285, 1690, 667, 1172, 366, 681]. MEDIMAGE [1275]. Meet [1045]. Melody [687]. Membership [495, 452]. Meme [116, 377]. Memetic [266, 273]. Memory [1586, 1623, 1655, 588, 1021, 1811, 1192, 1590, 998, 1747, 1639, 1616, 1573, 1582, 1568, 1758, 1647, 1596, 178]. Memory/ Message [1573]. Mercury [993]. Merge [1586]. Merging [1400]. Mesh [1838, 1770, 1768, 1599, 910, 1766]. Meshes [1687]. Mesoscopic [1747]. Message [495, 563, 453, 1684, 8]. Message-Based [563]. Message/UML [8]. Messages [810]. Messy [1204]. Meta [369, 1045, 1036, 1037]. Meta-Learning [369]. Meta-modeling [1045]. Meta-Modelling [1036, 1037]. Metacomputing [1783]. Metadata [1742]. Metadata-Enabled [1742]. Metamodel [1420]. Metamodel-Based [1420]. Metamodeling [1507]. Method [369, 1771, 1117, 664, 1776, 1676, 1653, 1828, 1091, 1748, 1708, 609, 1597, 961, 1563, 1709, 1798, 1782, 1710, 929, 287, 1789, 1692, 343, 608, 1755, 1716, 1202, 1837, 1717, 578, 708, 365, 1265, 1018]. Methodological [854]. Methodologies [149]. Methodology [1544, 1707, 146, 1210, 1723, 1211]. Methods [1048, 612, 834, 1634, 1726, 1116, 1447, 1534, 1104, 1817, 1439, 258, 734, 1249, 264, 1727, 1803, 1213, 1084, 99, 1449, 45, 1536, 641, 715, 1769, 1506, 1720, 1843]. METIOREW [140]. Metocean [1756]. Metric [1564, 484, 1368, 581, 1260, 1262, 1073]. Metrics [509]. Metrology [904]. MFS [1718]. Micro [1333, 200]. Microaggregation [1336]. Microcoded [156]. Microdata [1332, 1336, 1337, 1335]. Microflora [1746]. MiCRoM [1260]. Microorganisms [1070]. Microprocessor [872, 883]. Microscopic [1772]. Microwave [1817]. Migration [661]. Mind [191, 805, 1582]. Minerva [1589]. Minimal [420, 783, 626, 1466, 898, 221]. Minimal-Bad-Sequence [221]. Minimization [61, 1744]. Minimize [1643]. Minimizes [493]. Minimizing [274, 958, 1609, 1832]. Mining [572, 331, 1471, 1180, 526, 1472, 1249, 354, 1087, 1796, 525, 342, 547, 1255, 364, 1794, 524]. Mining-Based [1249]. Minkowski [922]. Misses [1820]. Missing [613, 441]. Missions [290]. Mixed [1241, 1665]. Mixin [1004]. Mixture [700, 352]. MLP [1232]. MMSC [1800]. Mn [1688]. Mobie [796]. Mobile [604, 1296, 1297, 1515, 796, 1023, 750, 1024, 872, 575, 874, 535, 1832, 39, 605, 1207, 1724, 16, 1073, 792]. Mobility [1537, 1313]. Möbius [1524]. Mod [234]. Modal [1125, 753, 309, 633]. Mode [631, 161]. Model [919, 301, 1389, 306, 733, 1402, 1394, 232, 1299, 1571, 698, 136, 1111, 746, 750, 818, 1387, 1113, 461, 1401, 307, 1466, 801, 1540, 1346, 462, 702, 1811, 214, 1774, 314, 1404, 1050, 1031, 1470, 1246, 1310, 1392, 102, 753, 1388, 21, 1722, 332, 1658, 1535, 294, 966, 967, 1420, 1252, 1393, 968, 1732, 1738, 1777, 309, 1743, 1622, 1789, 1692, 351, 1621, 635, 293, 1138, 815, 816, 1331, 697, 1831, 1243, 29, 652, 1638, 738, 1752, 822, 1697, 690, 352, 1416, 602, 1030, 1049]. Model-Based [1050, 635, 815]. Model-Checking [462, 968, 309]. Modeling [1247, 1398, 1781, 1141, 11, 1218, 1028, 1742, 628, 1414, 1624, 357, 1790, 1686, 1746, 1528, 5, 1813, 634, 1492, 1413, 714, 1747, 1840, 1773, 1440, 1360, 1744, 1416, 299, 1045]. Modelling [321, 1525, 673, 1160, 1132, 675, 1810, 1036, 275, 33, 1529, 291, 1373, 805, 1102, 1698, 1164, 1712, 1225, 14, 1037]. Models [614, 1424, 1756, 1534, 1530, 254, 1757, 848, 1425, 1551, 613, 700, 1733, 1750, 1043, 1841, 1652, 1532, 1531, 197, 1618, 1526, 1507, 38, 336, 1695, 412, 1473, 1077, 1696, 925, 1745, 1320, 19, 1321]. Moderated [1031]. Modes [585]. Modified [1472]. Modular [991, 1021, 1390, 1110, 584, 989, 1018]. Modularity [1109]. Module [1390, 86]. Modules [1004, 1323]. Modulo [502]. Molecular [1650, 365]. MOMBASA [1537]. Moments [1266, 912]. Monadic [949, 745]. Monadically [772]. Monads [1109, 972]. Monitoring [1667, 1668, 1673, 1251, 1835]. Monitors [313]. Monoids [449, 477]. Monotonic [901]. Monotonicity [742]. Monte [1128, 1692, 1749, 1570]. Montgomery [608]. Montgomery-Form [608]. Monthly [351]. Morphisms [781]. Morphological [908]. Mortem [167]. Most [503]. Motion [1285, 69, 32, 1202, 349]. Motivated [1026, 1479]. Motivational [799]. Motive [219]. Motorways [277, 280]. Movement [1345]. Movements [856]. Moving [22, 533]. MPEG [1284, 1280, 1455, 1281]. MPEG-4 [1455]. MPEG-7 [1281, 1280]. MPI [1665, 1620, 1600]. MPSA [1211].

MRMSolve [1526]. mRNA [432]. MSC [953, 450, 969]. MT [1217]. Mu [1039, 309]. Mu-Calculus [1039, 309]. Muller [749]. Multi [10, 1654, 1490, 921, 798, 1259, 818, 1240, 424, 12, 1557, 1466, 1751, 801, 1757, 1082, 1305, 802, 596, 803, 1708, 1065, 1470, 1310, 998, 601, 1562, 1830, 1610, 808, 633, 1492, 1037, 1644, 1311, 812, 813, 814, 815, 1501, 1224, 817, 339, 821, 1665, 1573, 1152, 1802, 19, 1647]. Multi-Agent [1490, 1492, 1501, 10, 798, 818, 1240, 12, 1557, 1466, 801, 1757, 1305, 803, 1470, 1310, 1610, 808, 1311, 812, 813, 814, 815, 1224, 817, 821, 1152, 19]. Multi-application [1562]. Multi-bank [998]. Multi-cluster [1573]. Multi-clustering [1082]. Multi-dimensional [1654, 424, 1830, 339]. Multi-disciplinary [1802]. Multi-formalism [1037]. Multi-inference [1065]. Multi-installment [1647]. Multi-modal [633]. Multi-neurules [1065]. Multi-objective [1802]. Multi-party [596]. Multi-phase [1708]. Multi-processor [1644]. Multi-scale [921]. Multi-sensor [1259]. Multi-signature [601]. Multi-trees [1751]. Multiagent [1242]. MultiCAD [1080]. MultiCAD-GA [1080]. Multicast [1537, 589]. Multicast-Based [1537]. Multigrid [1769]. Multilevel [1706]. Multilingual [1074]. Multimedia [1398, 1160, 1504, 161, 155, 1275, 1272]. Multiobjective [1143]. Multiparadigm [1512]. Multiparty [1312]. Multiple [711, 884, 1588, 32, 1610, 22, 1646, 271, 1264, 511, 532, 557]. Multiple-Features [1264]. Multiple-Robot [32]. Multiples [1339]. Multiplexer [1465]. Multiplication [608]. Multiplicatively [475]. Multiprocessing [1590]. Multiprocessor [1572]. Multiprocessors [1614, 1811, 1560]. Multirelational [1235]. Multistart [267]. Multithreaded [1566, 302]. Multithreading [950]. Multivariate [1756, 361, 352]. Multiversion [529]. Multiway [530]. Music [1460, 867]. Musical [1240]. Mutation [1135, 237]. Mutations [288]. Mutual [1655, 1493, 326]. Myrinet [1764].

N [570, 239, 772]. N-Free [772]. N-Version [239]. Name [958]. Nanoscience [265]. Naphtha [1808]. NAS [1628, 1620]. Nash [221]. Nash-Williams [221]. Natural [1074, 1706, 218, 842, 841]. Navier [1727, 1827]. Navigating [565]. Navigation [614, 1362, 1205, 34, 1097, 1073]. Navigator [374]. Near [276]. Nearest [1262]. Nearest-Neighbor [1262]. Nearly [345]. Needles [224]. Negation [856]. Negative [480, 366, 1054]. Negotiation [1487, 1158, 1489]. Negotiation-Based [1489]. Negotiations [242]. Neighbor [337, 1262]. Neighborhood [1653, 902, 271]. Neighbourhoods [908]. Nested [150]. Nesting [447]. Net [1486, 315, 761, 1044, 1841]. NetCoach [134]. Netherlands [1341]. Nets [1029, 1529, 1425, 1541]. Network [1587, 1239, 698, 1329, 1158, 1091, 1544, 182, 1232, 1765, 1763, 1546, 1093, 355, 1396, 1673, 1088, 1603, 1244, 270, 881, 690, 869, 1290, 1704, 882, 792]. Network-Centric [1290]. Networked [671]. Networking [1694]. Networks [1247, 89, 105, 1422, 1182, 1797, 880, 150, 1090, 1588, 371, 139, 1309, 1229, 1562, 1528, 503, 415, 228, 1078, 875, 1230, 1157, 412, 881, 1734, 1089, 1474, 602, 1834, 824]. Neumann [1676]. Neural [795, 1232, 1229, 355, 228, 1230, 1088, 1089, 1704, 1834, 1833]. Neuro [1246, 1699]. Neuro-Fuzzy [1699, 1246]. Neuron [1729]. Neurophysiology [1083]. neurules [1065]. Neutrality [224]. Neville [1675]. News [1280]. Newton [1789]. Next [1584, 1516, 1520]. Ni [1688]. Niching [1128, 696]. NMS [855]. No [489, 235]. Nodes [1567]. Noise [1162, 1332, 703].

Nominal [361]. Non [669, 480, 1204, 1767, 1104, 759, 259, 897, 898, 1815, 1197, 1150, 862]. Non-determinism [1197]. Non-linear [1150]. Non-manifold [897]. Non-obvious [862]. Non-optimal [669]. Non-overlapping [1767]. Non-parametric [259]. Non-redundant [1204]. Non-regular [759]. Non-simple [898]. Non-trivial [1815]. Non-uniform [1104]. Nondedicated [1671]. Nonlinear [1729, 1726, 1685, 1116, 625, 581, 1713]. Nonmonotonic [335]. Normalisation [497]. Normalization [699, 914, 1718]. Norms [799, 741]. Notation [10, 855]. Notations [1358]. Note [769, 73, 975]. Notions [443, 972]. Novel [1369]. NP [61, 443]. NP-Completeness [443]. NP-Hard [61]. NT [981]. NUMA [1614, 1638]. Number [478, 606]. Numbers [481, 209, 59, 418, 218, 762]. Numeration [475]. Numeric [1336, 361]. Numerical [1611, 612, 1729, 1726, 1686, 1761, 373, 1710, 1712, 1818, 1717, 1744, 1834].

O [1631]. Object [1048, 1503, 1167, 1296, 1504, 1517, 1515, 1518, 1757, 1279, 1508, 278, 1516, 1520, 1514, 177, 591, 1507, 1222, 1267, 347, 1506, 381, 1030, 1049]. Object-Models [1507]. Object-Oriented [1048, 1503, 1167, 1296, 1517, 1518, 1508, 1514, 591, 1222, 1506, 1030, 1049]. Object/ Component [1504]. ObjectCheck [1049]. Objective [140, 1315, 1802]. Objects [905, 919, 1575, 1540, 1546, 534, 1352, 1145, 22, 535, 714, 915, 168, 38, 533]. Obligations [799, 800]. Observability [951, 955]. Observation [1321]. Observational [1039]. Observed [356]. Observer [1043]. Observers [619]. Obstacle [1769]. Obtained [1834]. obvious [862]. Occluded [905]. Ocean [1624, 358, 1777]. Ocean-Atmosphere [1777]. OCL [1039]. OCR [1232]. ODEs [1836, 1792].

Odometric [1207]. Odour [1557]. OFB [585]. Off [1359, 592]. Off-Line [592]. Offensive [735]. OHS7 [113]. Oil [1774, 519]. OLAP [540]. Omnidirectional [23]. OMP2001 [1615]. On-Line [345, 1474, 514, 198, 199]. One [766, 956, 1779, 1377, 961, 1815, 61, 270, 607]. One-Counter [961]. One-Letter [766]. One-Max [270]. One-Pass [956]. Online [1460, 70, 986, 513]. onto [1586]. Ontologies [1241]. Ontology [1233, 519, 1099]. Ontology-Based [1099]. OntoWebber [573]. OO [1512, 1519]. Open [114, 1417, 906, 108, 115, 1454, 118, 119, 468, 665]. Open-Close [906]. Open-Source [1417]. OpenCoLaS [1302]. Opening [1369]. OpenMP} [1665, 1623, 1624, 1830, 1620, 1616, 1621, 1625, 1619, 1573, 1617]. OpenMP/MPI} [1573, 1621]. OpenTella [1612]. Operating [1671, 1519]. Operation [1247, 115, 585]. Operational [1308, 1050]. Operations [1364, 1765, 908, 1501]. Operative [1821]. Operators [1701, 1317]. Opportunities [1188, 1480]. Optical [1765]. Optimal [621, 1558, 1641, 700, 360, 638, 1714, 1150, 645, 1227, 669]. Optimisation [696, 256, 1802]. Optimistic [596]. Optimization [1787, 1632, 834, 1685, 1212, 1232, 985, 1819, 265, 1139, 1072, 1609, 1213, 287, 1625, 49, 1215, 269, 1723, 1149, 1337, 1794, 824]. Optimizations [1192, 998]. Optimizing [997, 543]. Option [1803]. Orchestration [1292]. ORDB [567]. ORDB-Based [567]. Order [208, 750, 949, 956, 759, 1013, 327, 963, 1710, 1393, 1312, 1010]. Order-Based [1312]. Ordered [491, 828, 775, 641]. Ordering [1748, 1597]. Orderly [80]. Orders [918]. Ordinals [757]. Ordinary [1728, 1710]. Organisation [806]. Organisations [812]. Organising [688]. Organization [2, 1272]. Organizational [1301]. Organizations [16]. Organize [1270]. Organizing [139]. Oriented [1048, 1503, 657, 1167, 1296, 1510, 15, 1517, 140, 36, 8, 1518, 1508, 1015, 1236, 1778, 1514, 1743, 1063, 13, 1841, 813, 591, 817, 1222, 820, 1427, 1745, 1506, 1030, 1049, 14]. Origins [1733, 1138]. Orography [1761]. Orthogonal [75, 63, 62, 82, 76]. Other [130, 944]. Out-of-Core [1666]. Out-of-Vocabulary [1218]. Outlook [201]. Overapproximations [620]. Overfitting [1146]. Overflow [588]. Overlap [474, 1739]. Overlapping [1693, 1767]. Overlappings [382]. Overview [1387, 1325, 1670, 1157, 1273, 1095, 847, 188, 149]. Ownership [168]. Oz [933, 942, 938].

P [1122, 789]. P-Systems [789]. Package [1245, 1682]. Packet [371]. Packets [875]. Packing [1654, 81, 513, 492]. Page [551]. Page-Replacement [551]. PageRank [691]. Pages [375]. Pairing [479, 1072]. Pairwise [1782]. Pajek [105]. Panel [1520]. Paper [1347]. Paraconsistency [1167]. Paradigm [1194, 1308, 13, 1427, 1744]. Parallel [1564, 1628, 1585, 1586, 1776, 1677, 1654, 1653, 1394, 1141, 1666, 1631, 1577, 1828, 1838, 1558, 787, 1656, 1635, 773, 260, 285, 254, 1601, 1669, 1542, 1670, 1811, 1169, 1661, 1671, 1657, 1721, 32, 1790, 1686, 1009, 1597, 1563, 1688, 1658, 1689, 373, 1649, 1795, 1691, 1639, 1663, 1399, 1606, 706, 1622, 1789, 1692, 1621, 1825, 1826, 1694, 1652, 1618, 788, 1699, 1682, 1695, 1837, 1664, 1155, 1725, 454, 1318, 1799, 1089, 1602, 1720, 231, 1659]. Parallelisation [1557, 315]. Parallelism [1574, 1624, 1830, 449]. Parallelization [1572, 1650, 1723]. Parallelize [1211]. Parallelizing [1586, 1198]. Parameter [52, 1714, 1831]. Parameterized [754, 305, 1113, 307, 431, 439]. Parameters [498, 1214]. Parametric [461, 1794, 544, 259]. Parametrized [428]. Pareto [699]. Parity [716]. Parity-Check [716]. pARMS [1682]. Parsers [990]. Part [35]. Partial [1393, 1136]. Partially [491, 688, 775]. Partially-Ordered [775]. Particle [1629, 1838, 37]. Particle-in-Cell [1629]. Particle-Mesh [1838]. Particles [356]. Particular [812]. Partition [1653, 318]. Partitioned [28, 1765, 1603]. Partitioning [1838, 1678, 1644, 1842, 152]. Partitions [480]. party [596]. Pass [956]. Passage [1533, 1221, 1740]. Passage-Time [1533]. Passing [958, 1573, 1684]. Passive [1765]. Past [1066]. Patented [225]. Path [1204, 60, 183, 1206, 1073]. Path-Width [60]. Paths [514, 920, 880, 503, 430]. Patients [1275]. Pattern [549, 1669, 495, 807, 1174, 550, 364, 1231]. Patterns [1104, 1741, 368, 1175, 764, 337, 820, 346, 363, 344, 383, 390, 524]. PaViS [1681]. PC [1629]. PCs [1792, 1814]. PDE [1607]. PDES [1745]. Peer [1612, 121]. Peer-Reviewed [121]. Peer-to-Peer [1612]. Penalty [1213, 1150]. Component [1504]. Correcting [712]. Cost-Tradeoffs [147]. CPU64 [157]. Event [1029]. FE [1709]. HMM [844]. Inductive [337]. JA [1628]. Message [1573]. MPI} [1573, 1621]. Or [1033]. SPIN [1398]. Subscribe [887]. UML [8]. Pending [1608]. Penna [1732]. PEP [324]. PEPA [1529]. Perception [33, 29]. Perceptual [1449, 1452]. Perfect [669]. Perform [1815]. Performance [416, 232, 886, 1629, 1553, 1584, 261, 260, 1537, 1540, 1542, 1529, 1680, 689, 542, 1467, 1764, 1650, 158, 1599, 1532, 1531, 1786, 1619, 1595, 1589, 1627, 1198, 708, 1617, 1615, 1626, 1705, 1598]. Performances [1481]. Performer [862]. Periodic [781, 1278]. Perl [850]. Permutations [762]. Persistence [1733]. Persistent [1458]. Personalized [796, 1271, 192, 133]. Perspective [1188, 1460]. Perspectives [166, 190]. Perturbation [1720]. Perturbative [1335]. Pervasive [877, 792]. Petaflops [1582]. Petaflops-Scale [1582]. Peterson [1655]. Petri [1541, 1044, 1425, 315, 761, 1841]. Petroleum [489]. Phase [617, 465, 1775, 1527, 1691, 1692, 1708]. Phase-Type [1527]. PhD [1503]. Phenomena [1712]. Pheromone [287]. Pheromones [258, 811]. PhFit [1527]. Philips [665]. Philips-Open [665]. Philosophical [328]. Phones [796]. Phonetics [847]. Phrases [701]. Phylogenetic [1657]. Physical [556, 1715, 873, 607]. Pick [525]. PicSOM [1281]. Pieces [927]. Piecewise [639, 1089]. Piecewise-Affine [639]. Pigeonhole [763]. Pipeline [117, 489]. Pipelined [223]. Piracy [1448]. Pirates [1450]. Pisot [481]. Pitfalls [988]. Place [1006, 507, 744]. Places [568]. Plaintext [593]. Planar [507, 83, 57, 430, 559]. Planarity [102]. Planarization [56, 52]. Plane [924, 429, 893, 82]. Planes [927]. Planner [1206]. Planning [1485, 1204, 32, 80, 46, 47, 1501, 1497, 1316, 1206, 1202, 1418, 1495, 1251, 1073]. Plans [99, 38, 1071]. Plant [104, 1246]. Plants [253, 628]. Plasma [357, 356, 352]. Platform [1566, 1559, 663, 147, 874, 667, 1621, 1783, 683, 817, 1198, 561]. Platform-Based [147]. Platforms [946]. Playing [1347, 1289]. Plenty [520]. Plug [841]. Plug-In [841]. PM [1566]. PN [1310]. Point [830, 1017, 966, 714, 909, 328, 854, 1416]. Pointer [490]. Pointing [842]. Points [758, 1779, 900, 559, 1227]. Polar [83, 84]. Policies [539, 1319]. Policy [6, 964]. Pollutant [1761]. Pollution [1843]. Polygon [64]. Polygonal [617, 482]. Polygonalization [919]. Polygons [922]. Polynomial [483, 1197, 1645, 558]. Polynomial-Cost [1197]. Polynomials [361]. Polyomino [81]. Polyp [1743]. Polyphemus [89]. Pomsets [772]. Pools [116]. Populated [22]. Population [263, 1129, 653, 1342]. Populations [1610].

Porous [1817, 1691]. Portal [1741, 1786]. Portfolio [1139]. Portraits [617]. Position [1362, 235]. Positive [451, 366]. Positivity [1730]. Possibilistic [1176]. Possible [960]. Post [167, 1216, 1337, 225]. Post-2000 [225]. Post-Correction [1216]. Post-Masking [1337]. Poster [1521]. Postures [868]. Power [766, 996, 486, 628, 1246, 777, 1613, 440, 1839, 997, 1474, 876]. PPM [343]. PPRO [569]. Practical [1445, 597, 577, 1440]. Practically [346]. Practice [1385, 1387, 1183, 151]. Practices [1417]. Praxical [864]. Precedence [1640]. Precise [1035, 1036, 987]. Precision [1017, 1016]. Predicate [616]. Predict [345]. Prediction [1611, 1184, 1130, 275, 276, 1542, 1467, 277, 1560, 1594, 1161]. Predictive [1246]. Predictor [1593]. Predictors [884, 1604, 1592]. Preemption [1609]. Preemptive [298, 417, 1609]. Preferences [522, 808]. Prefetch [1595]. Prefetching [1000, 1596]. Pregnancy [1172]. Preliminary [1135, 1472]. Preorder [751]. Prepaid [796]. Preprocessing [272]. Prescribed [72, 62]. Presence [1162]. Presentations [1398]. Presenting [1362, 181, 401]. Preserving [1533]. Pricing [1803]. Primal [717]. Primal-Dual [717]. Primitive [349]. Primitives [664, 1266]. Principle [438, 367]. Principles [207, 1387, 763]. Priority [1762]. Priority-Scheduling [1762]. PRISM [1535, 294]. Privacy [1453, 437, 1454]. Private [595]. Probabilistic [583, 1091, 957, 631, 425, 1535, 294, 1171, 1202, 1071, 747]. Probabilities [1748]. Probe [1095]. Probe-Based [1095]. Problem [508, 1654, 1141, 924, 1558, 769, 1465, 461, 257, 1090, 261, 211, 1763, 1822, 479, 1642, 318, 1367, 1721, 1470, 216, 1798, 556, 1803, 1782, 483, 422, 267, 1395, 452, 1403, 1211, 270, 1824, 271, 1066, 273, 478, 1494, 1821, 1659]. Problems [516, 1728, 1809, 834, 1676, 1677, 1726, 623, 1369, 431, 1706, 262, 1489, 1708, 512, 264, 1722, 1803, 988, 1768, 1826, 1064, 490, 269, 946, 1150, 1769, 468, 1720, 1704]. ProC [1525]. ProC/B} [1525]. Procedure [1050, 852]. Procedures [1117, 1114]. Process [1247, 1525, 656, 1422, 795, 666, 536, 1029, 1499, 682, 150, 1086, 1246, 1528, 1740, 1287, 814, 38, 1253, 55, 660]. Processes [1533, 1299, 1655, 1810, 295, 1257, 1103, 670, 974, 1753, 1133, 454, 359, 1436]. Processing [540, 866, 1580, 858, 1671, 32, 1597, 1762, 1579, 530, 1699, 543, 559, 532, 282, 1647]. Processor [1608, 1613, 1663, 1560, 197, 200, 198, 199, 155, 1582, 1556, 1644]. Processor-in-Memory [1582]. Processors [1605, 1562, 202, 997, 882, 156]. PROD [1530]. Produces [253, 1367]. Product [1442, 1415, 657, 685, 677, 650, 655, 1437, 676, 678, 957, 680, 1414, 671, 659, 662, 649, 1445, 1413, 654, 1443, 684, 653, 651, 683, 679, 674, 652, 1440, 658, 661, 660, 681, 665]. Product-Line [1442]. Production [852, 1497, 611, 824]. Products [666, 451, 421, 489, 477]. Profile [695, 357, 279, 1532, 1000]. Profiling [986]. Profit [524]. Program [166, 991, 832, 1736, 1583, 1606, 177, 1366, 1600]. Programmable [146, 1596]. Programmers [1193]. Programming [1200, 1585, 1513, 1007, 1510, 226, 245, 247, 1300, 36, 1118, 240, 1447, 1512, 275, 1304, 229, 254, 227, 1669, 1306, 1169, 252, 1551, 244, 1183, 277, 246, 239, 251, 1435, 332, 223, 1778, 1008, 979, 1621, 1825, 399, 400, 1652, 1842, 1228, 1148, 243, 1225, 1618, 1147, 1222, 280, 1146, 1724, 335, 932, 1427, 1665, 1573, 236, 225, 1179, 1552, 823, 1356]. Programs [1067, 1632, 306, 952, 247, 1194, 302, 150, 186, 238, 1021, 1542, 831, 1661, 1404, 1032, 1649, 1505, 1663, 176, 994, 175, 1755, 445]. Progress [110]. Progressive [558]. Prohibition [235]. Project [1424, 1338, 1087, 1219]. Projection [1727].

Projections [926, 903, 910]. Projective [608]. ProjectLeader [666]. Prolog [1069]. PROMELA [1398, 1391]. PROMELA/ SPIN [1398]. Promotion [996]. Pronominal [1217]. Proof [213, 1122, 220, 763, 1166, 1018]. Proofs [734, 444, 217, 973]. Propagation [1176, 1557, 1601, 1016]. Properties [733, 1297, 1125, 259, 313, 251, 304, 1397, 343, 725, 724, 728, 974, 1719]. Property [442, 310, 906, 743]. Property-Specification [310]. ProPolyne [558]. Propositional [1119]. Propositions [217]. Prostate [1248]. Protect [1339, 1328]. Protecting [1457, 1454, 605, 1342]. Protection [1332, 1459, 1336, 1333, 591, 1331, 1335]. Protein [432]. Protocol [754, 603, 182, 1486, 596, 5, 1612, 407, 1395, 1589, 1323]. Protocols [793, 1566, 1307, 804, 501, 597, 9, 1397, 816, 736, 889]. Prototype [217, 1019]. Prototyping [256, 1046, 885]. Provably [600, 584]. Prove [741]. Providers [142]. Providing [128]. Proving [952, 1194, 961, 1110, 1193, 1348, 1195]. Provision [1309]. Provisioning [1158]. Proxy [525]. Pruning [345]. PSEs [1788]. pseudosnakes [487]. PSPACE [454]. PSPACE-Hard [454]. Psychopathology [1758]. Public [582, 583, 22, 584, 577]. Public-Key [582, 584, 577]. Publish [887]. Publish/Subscribe [887]. Publishable [121]. Pulsatile [1771]. Purchase [368]. Pure [498, 433, 490]. Purpose [1145]. Pursuit [363]. Pushdown [777, 963]. Puzzle [245]. PVS [304]. PWA [627]. Pyramid [899]. Pyramids [892]. Pyrolysis [1808].

Q [811]. Q-Learning [811]. Qi [1755]. QoS [1160, 1544]. QR [1683]. QR-Decomposition [1683]. Quadrangulations [896]. Qualification [1064]. Qualitative [618, 1119, 1250, 634, 722]. Quality [572, 664, 1504, 1544, 1439, 1103, 1086, 663, 44, 46, 45, 668, 881]. Quality-Driven [46]. Quantification [546]. Quantifying [662]. Quantitative [1517, 862]. Quantization [626]. Quantized [626, 638]. Quantum [436, 756, 438, 434, 437, 435, 718, 469]. Quasi [515, 1691]. Quasi-Characteristics [1691]. Quasi-Clique [515]. Queen [1090]. Queries [528, 535, 1221, 550, 558, 708, 532, 1262, 523, 554]. Query [540, 696, 1259, 529, 1363, 1165, 49, 46, 47, 521, 1206, 441, 708, 1099, 694]. Query-Driven [1165]. Querying [1357, 1100, 522, 1274, 548]. Question [1740]. Questions [760]. Quicksort [273]. Quotient [770].

Race [253]. Radiosity [929]. Railway [684]. Ramification [1064]. Random [1132, 480, 491, 426, 270]. Random-Error [426]. Randomized [777, 608, 427]. Randomness [595]. Range [26, 558]. Range-Sum [558]. Rank [1336]. Ranking [45, 548]. Rapid [256]. Rascal [802]. Rate [1276]. Ratio [506, 430, 1236]. Ratio-Oriented [1236]. Rational [766]. Rationales [1499]. Rationals [213]. RBF [1091]. RCC [726, 725, 727, 730, 728]. RCC-8 [726, 725, 727, 730, 728]. Re [116, 377]. Re-editing [116, 377]. Reachability [616, 951, 955, 311]. Reacting [1241]. Reaction [1726]. Reactions [1808]. Reactive [1300, 500, 1102, 1094, 736, 751]. Real [1498, 296, 845, 209, 840, 1040, 801, 807, 1609, 1804, 1791, 334, 1547, 324, 27]. Real-Time [1498, 296, 845, 840, 1040, 801, 807, 1547, 324, 27]. Realistic [1812]. Realities [1178]. Reals [212]. Reason [519]. Reason-able [519]. Reasoning [1119, 954, 1344, 37, 1121, 1165, 597, 1215, 1412, 727, 722, 791, 40, 976]. Receiver [1278]. Receiver-Driven [1278]. Receptive [899]. Recharge [796]. Recognition [843, 838, 924, 845, 840, 844, 1093, 846, 807, 927, 1231, 1834]. Recognition-Based [807]. Recognizable [772, 453]. Recognizers [840]. Recognizing [71]. Recommendation [138, 141]. Recommending [140]. Recompilation [1632]. Recomputation [937]. Reconciling [955]. Reconfigurability [1556]. Reconfigurable [145, 160, 157, 1568, 883, 156]. Reconfiguration [639]. Reconfiguring [1498]. Reconstruction [905, 926, 903, 1096, 279, 925, 1719, 231]. Record [1334]. Recover [200]. Recovery [1774, 1311, 1674]. Recrystallization [1813]. Recurrence [151, 1684]. Recursion [1606, 406]. Recursive [306, 908]. Red [1597, 1157]. Red-Black [1597]. Redefining [1836]. Redistributing [116, 377]. Reduce [996, 1593]. Reduced [596]. Reducing [327, 1207]. Reduction [632, 1388, 1393, 703, 339, 744, 1474, 1030]. Redundancy [493]. Redundant [1204, 1644]. Reeb [896]. Reengineering [1441]. Reentrant [950]. Refactoring [1429]. Reference [1418]. References [206, 395, 731, 947, 51, 1060]. Refinement [301, 1833, 1259, 1384, 743, 1599, 521, 642]. Reflection [1550]. Reflections [124]. Reflective [1303]. Reformulation [695]. Refutable [337]. Refutable/Inductive [337]. Region [356, 723, 1264]. Region-Based [1264]. Regions [1261, 1350]. Register [995]. Registers [996]. Registration [285]. Regression [252, 1707, 340, 233]. regressive [1246]. Regular [455, 953, 1678, 759, 1031, 450, 784, 452, 510]. Regularity [713, 454]. Regularization [1091]. Regulated [789]. Reinforced [365]. Reinforcement [1084]. Rejection [1218]. Related [431, 1609, 839]. Relating [1067]. Relation [338]. Relational [1118, 1226, 546]. Relations [957, 1640, 403, 724, 730]. Relationship [785, 232]. Relationships [130]. Relative [301, 287, 392]. Relaxation [549, 1829]. Release [1566]. Relevance [848, 1265, 548, 693]. Reliabilities [503]. Reliability [657, 1748, 1749, 1554]. Reliability-Oriented [657]. Relief [1501]. Remote [888, 133, 1146, 1799]. Rendering [1569]. Rendezvous [889]. Replacement [551]. Replanning [1484]. Reply [1412]. Report [656, 111, 677, 110, 682, 680, 649, 670]. Reports [137]. Repositories [369]. Repository [1100, 567, 384, 391, 1428]. Represent [338]. Representation [1369, 985, 1467, 1203, 697, 722, 1366, 910]. Representational [1177, 728]. Representations [1491, 260, 629, 84, 892, 857, 762, 1359, 1376, 690]. Representing [676, 1380, 1]. Requirement [12]. Requirements [11, 1655, 596, 1041, 2, 304, 1479, 809, 1043, 1179, 1647]. Rescheduling [286]. Research [837, 1465]. Researchers [1433]. Resection [1248]. Residue [606]. Resistance [1456]. Resolution [83, 1721, 964, 1277, 1095]. Resolving [1602]. Resource [306, 1797, 307, 802, 1484, 637, 824]. Resource-Based [1484]. Resource-Constrained [306]. Resources [1571, 1644, 184]. Respect [1819]. Response [563, 1547, 553]. Responsiveness [1075]. Restarted [1679]. Restarts [1136]. Restricted [77, 435, 1366, 445]. Restrictions [661]. Restructuring [541]. Result [781]. Results [1135, 28, 252, 1707, 512, 244, 1198, 1766]. Retarded [1715]. Retrieval [696, 698, 1268, 1261, 1742, 689, 1281, 1221, 1740, 1738, 707, 564, 705, 1266, 339, 1264, 1265, 704, 693]. Retrieving [1353]. Retrospective [708]. Reusability [1415, 1429]. Reusable [1423, 1443]. Reuse [1424, 1441, 1434, 1425, 1420, 567, 1418, 1427, 1426, 1428, 1438]. Reusing [1417, 1409]. Reveillès [928]. Reverend [1186]. Reversals [494]. Reverse [174]. Reversible [476]. Review [1088, 1105]. Reviewed [121]. Reviewing [680]. Revisited [7, 1033, 780, 47]. Revisiting [137]. Revocation [589]. Reward [1526]. Rewrite [1114]. Rewriting [456, 968, 788, 789, 523]. Reynolds [1818]. Rhythm [1282]. Ridgelet [928]. Rights [1461, 1460, 1453, 560]. Rijndael [580]. Rings [514]. RISC [883]. Risk [1334, 1748, 1337, 1335]. RNA [1733]. Road [1132, 769, 1255]. Roadmap [1202]. Roadmapping [653]. Robot [32, 1097, 31, 23, 39, 814, 1073]. Robotic [622, 33]. Robotics [36, 27]. Robots [247, 863, 1207, 1202]. Robust [551, 914, 890, 1449, 637, 29, 1071]. ROC [374]. Role [1087, 1178, 1077, 825]. Roles [800, 1101]. Rollback [1543, 200]. Root [630, 778]. Root-Mean-Square [630]. Roots [502]. ROSPER [1122]. Rosser [784]. Rotation [1718]. Rough [1234, 366]. Roundoff [1016]. Routine [225]. Routing [257, 1588, 414, 1603, 288]. Royal [1132]. RSA [606]. RTSP [119]. rube [192]. Rule [1151, 1114, 234, 1253, 340, 1053, 1056, 1054]. Rule-Based [1151, 1253]. Rules [1793, 526, 227, 1121, 214, 368, 234, 1085, 1252, 1216, 402, 863, 363, 349, 1058, 1057]. Rulesets [1475]. Run [160]. Run-Time [160]. Running [96]. Runtime [1199].

S [1293, 431]. Safety [754, 313, 885]. Safety-Critical [885]. Salesman [1211, 271]. Sample [1229]. Sampling [1333, 1700]. SASEPA [1640]. SAT [1142, 496, 1111, 746, 1384, 425]. SAT-Based [1111, 1384]. Satellite [356]. Satisfaction [264]. Satisfiability [1402, 1662]. Saturation [1370]. SavingsAnts [257]. SC3 [123]. Scalable [1280, 1795, 376]. Scalar [608]. Scale [1784, 1810, 307, 1706, 1087, 1732, 1844, 611, 1582, 1602, 921]. Scale-Space [1844]. Scan [995]. Scannerless [990]. Scatter [1140]. Scattered [279]. Scattering [1709]. Scenario [1424]. Scenarios [1531]. Schedulability [295]. Schedule [553]. Scheduler [1800]. Scheduling [298, 423, 1369, 256, 261, 260, 1640, 1641, 1591, 1642, 1643, 1489, 417, 311, 1562, 1762, 1644, 422, 1646, 1760, 1625, 151, 1316, 1801, 1155, 1702, 297, 1578, 1799, 1568]. Schema [541, 550, 271, 737, 717, 561]. Schema-Driven [550]. Schema-Restructuring [541]. Schemas [92]. Schemata [1606]. Schematic [99]. Scheme [1678, 601, 1280, 1191, 1278, 1691, 1599, 1645, 235]. Schemes [1829, 415, 589, 1105, 599]. Schwarz [1789]. Science [837, 465, 1379, 1736, 1583]. Scientific [1784, 1183, 356, 1778, 328, 543, 367]. Scientists [1127]. SCIRun [1788]. Scope [655]. Scoping [1440]. SCR [304]. SCR-Style [304]. Screen [859]. Scripts [322]. SDC [1340]. SDL [1391, 738]. Search [1421, 1564, 432, 1142, 1653, 316, 1363, 1558, 1751, 1140, 327, 1795, 426, 1400, 267, 1270, 636, 248, 268, 1700, 936, 939, 944, 938, 935, 934, 1133, 1725, 1794, 1136, 548, 544]. Searching [491, 688, 326]. Second [949, 759, 1013, 365]. Second-Order [949, 759, 1013]. Secret [426, 594]. Section [985]. Secure [604, 1027, 590, 601, 600, 1454, 589, 584, 608, 1452]. Security [793, 733, 411, 734, 587, 1397, 407, 577, 585]. Seek [122]. Segmentation [911, 1261, 893, 923, 1216]. Segmented [1260]. Segments [72, 917]. Seismic [1601, 1719]. Seismograph [358]. Selecting [539, 1209, 333]. Selection [1131, 1154, 1134, 703, 651, 694]. Selections [555]. Selective [426]. Selectivity [555]. Self [551, 1483, 139, 1456, 1468]. Self-Adaptive [1468]. Self-Checking [1456]. Self-Organizing [139]. Self-Stabilizing [1483]. Self-tuning [551]. Semantic [1081, 536, 857, 976]. Semantical [949]. Semantics [1293, 300, 571, 1306, 1308, 987, 804, 1374, 1352, 1373, 965, 1016, 973, 1022, 1323]. Semi [1533, 444, 1823]. Semi-algebraic [444]. Semi-automatic [1823]. Semi-Markov [1533]. Semidefinite [1842]. Semiotised [864]. Semistructured [370]. Sense [1782]. Sensing [1146, 1095]. Sensitive [1075, 1333, 1750, 771]. Sensitivity [1754]. Sensor [875, 1259]. Sensors [1714]. Sentences [1076, 857, 344]. Separates [443]. Separation [1511, 895, 417]. Sequence [495, 453, 494, 177, 221]. Sequences [905, 455, 1130, 26, 1171, 347]. Sequencing [269]. Sequential [1104, 1173, 1412, 789, 1069]. Serial [1595]. Serially [1096]. Series [773, 353, 358, 351, 1083, 646]. Series-Parallel [773]. Serv [1157]. Server [91, 118, 1545]. Service [1523, 1504, 1158, 1254, 864, 1309, 874, 31, 679, 1427, 1430]. Services [1293, 563, 1454, 1786, 878, 1554, 129, 128, 792]. Serving [692]. Servo [28]. Servoing [27]. Session [656, 682, 649, 1521]. Sessions [670]. Set [474, 262, 909, 545, 665]. Set-Top [665]. Sets [620, 420, 491, 898, 1223, 772, 453, 404, 403, 828, 1337, 366, 1115]. Setting [1004]. Several [1328, 1231]. SFB [567]. SFB-501-Reuse-Repository [567]. Shakedown [1819]. Shannon [493]. Shape [1685, 910, 839, 1267, 1843]. Shape-Related [839]. Shapes [914]. Shaping [1276]. Shared [1623, 1491, 1655, 1300, 1021, 595, 1192, 1639, 1616, 1313, 1573, 1568]. Shared-Memory [1021]. Sharing [594]. Sharpe [1236]. Shells [1744]. Shop [298, 1141, 261, 1659]. Short [73, 765]. Shostak [1117]. Shot [1285]. Should [1127]. Shuffle [770]. Side [608, 543]. Sided [61]. Sign [856, 843, 837, 845, 858, 844, 850, 846, 857, 854, 847]. Signal [1613, 358, 362, 282]. Signals [274, 504, 199]. Signature [600, 601]. Signatures [598]. Signcryption [599]. Signer [844, 601]. Signer-Independent [844]. Signing [851, 853]. SignSynth [850]. SignWriting [858]. SignWriting-Based [858]. Silico [1746]. Silicon [1731]. Silicon-Based [1731]. Similar [463]. Similarities [887, 388, 381]. Similarity [1564, 432, 1234, 364]. Similarity-Based [1234]. Simple [54, 1017, 368, 1102, 1762, 900, 1226, 898]. Simplicity [895]. Simplification [406]. Simplified [508, 927]. Simulate [1745]. Simulated [1481, 1656, 1211, 1659]. Simulation [317, 1812, 1248, 1828, 1544, 1540, 1757, 1601, 318, 1470, 806, 1731, 1841, 1840, 1772, 1749, 1801, 1752, 1152, 1697, 751, 1814, 1835, 1834, 1805, 1766]. Simulation-Based [1544]. Simulations [1807, 1707, 1688, 1761, 1830, 1813, 1693, 1599, 813, 1837, 1570]. Simulator [890, 1602, 1600]. Simultaneity [1375]. Simultaneous [1640]. Since [447]. Single [633, 1206]. Single-Agent [633]. Single-Query [1206]. Singular [1728, 1720, 1843]. SIP6 [1762]. Sites [573, 1161]. Situation [575]. Size [1284, 1765, 62, 419, 744, 431]. Sized [1190]. Sizes [554]. Skeleton [1009]. Skeletons [1653, 1192, 1190, 506]. Sketch [689]. Skin [1209]. SLA [1158]. Slices [1096]. Slicing [1032]. Slide [593]. Slopes [72]. Small [1377, 658, 1018, 1058]. Small-Step [1018]. Smart [802, 875, 889, 645]. SmartKom [861, 860]. SML [1198]. Smooth [1827]. SMP [1585, 1620, 1646, 1621, 1589, 1638, 1665]. SMP-Machines [1638]. SMS [1078]. Soap [1775]. Social [1237, 1470, 1]. Societies [3]. Society [1757]. Sockets [1022]. Soft [1188, 1007, 1184]. Software [301, 1038, 1623, 1441, 1422, 1364, 1444, 650, 1298, 1517, 1417, 1457, 877, 1387, 1439, 179, 1304, 1401, 1537, 1822, 1414, 1425, 803, 613, 192, 1388, 291, 1187, 1448, 1823, 304, 174, 672, 1445, 1609, 1413, 1170, 1390, 290, 1399, 1122, 173, 1622, 1480, 185, 1063, 13, 1070, 591, 1565, 1531, 668, 674, 1431, 611, 1427, 1289, 1506, 1030, 1049, 1438, 14, 1436, 1598]. Solidifying [1712]. Solution [1728, 211, 1686, 1727, 1803, 1710, 1818, 936, 544, 1805]. Solutions [1518, 1558, 1730, 1788, 663, 670, 1716, 1718, 645]. Solve [1090, 1816, 1150]. Solved [1708]. Solver [1809, 303, 1706, 1597, 1826]. Solvers [1402, 1677, 1722, 1607, 1621, 1713]. Solving [620, 1116, 1817, 1822, 1367, 264, 327, 1803, 1789, 1681, 269, 1682, 1824, 271, 1684, 1227, 1821, 1659, 1704]. Some [795, 28, 333, 244, 761, 217, 784, 435, 1552]. Sonification [868]. Sort [1586]. Sound [866]. Source [1417, 108, 1410, 694]. Sources [43, 565, 821, 1718, 1099]. Space [1564, 1534, 414, 1209, 472, 327, 1015, 290, 148, 485, 518, 1844, 883, 1030]. Space-Efficient [883]. Spaces [802, 308, 916, 1208, 1313, 943, 940, 934, 873, 1262]. SPACETUB [1320]. Spanish [1218, 1217]. Spanning [506, 430, 80]. Sparse [1559, 61, 1682]. Spatial [1119, 551, 1277, 1733, 1274, 530, 724, 722, 1133, 555]. Spatio [538]. Spatio-temporal [538]. Spatiotemporal [534]. SPCONS [1069]. SPEC [1584, 1615]. Special [137, 1627]. Specific [1419, 1363]. Specification [1293, 310, 1035, 1668, 797, 955, 12, 1358, 1508, 246, 833, 304, 1314, 1405]. Specifications [145, 457, 624, 1041, 964, 1532, 18]. Specifying [165, 9]. Spectra [365]. Spectral [1827, 1826, 1842]. Spectroscopy [276]. Speculative [1605]. Speech [1093, 1216]. Speed [1104, 280]. Speedup [606]. Spelling [846]. Spi [954]. Spi-calculus [954]. SPIN [1407, 1397, 1386]. SPINning [1399]. Spite [1241]. Splitting [1767]. SPMD [1620].

Spoken [1075]. Spontaneous [1076, 1743]. Spotting [1218]. Spray [355]. Spread [1818]. Spring [102]. Springembedder [99]. SQL [1118, 521]. Square [630, 502]. Squares [1084, 1826]. Squaring [584]. SR8000 [1617]. SRN [844]. SRN/HMM [844]. SSA [995]. ST [571]. Stability [1235]. Stabilization [626]. Stabilizing [1483]. Stable [512, 1816, 1839]. Stack [588]. Stage [1141, 1755]. Stages [884]. Standard [919, 9, 1579]. Standards [1276]. Star [476]. STARK [1369]. Stars [1765]. Startup [1246]. State [1300, 1130, 1534, 1660, 165, 461, 1524, 1730, 308, 472, 264, 1760, 1451, 1458, 1030, 470]. State- [1300]. State-Level [1524]. Statecharts [1040, 1373]. States [1142, 627, 1748, 782, 193]. Static [1378, 991, 251, 1375, 742, 1625, 1603, 1714, 997, 1174]. Statics [1706]. Statistic [240]. Statistical [1325, 855, 1210, 1213, 351, 1341, 336, 538]. Statistics [1341]. Status [1483, 1626]. Steam [1246]. Stefan [1708]. Stencil [1820]. Step [1377, 450, 272, 1018, 1720]. STG [323, 1195]. Stochastic [1756, 258, 1541, 1755, 1840, 1147, 1094, 1701, 1753]. Stock [1473]. Stokes [1727, 1827]. Stony [1743]. Stopwatch [298]. Storage [1141, 1651, 744, 1272]. Storm [1602]. Straight [77]. Straight-Line [77]. Strategies [614, 1653, 622, 1339, 1445, 279, 248, 270]. Strategy [232, 551, 1698, 1139, 1269]. Stratification [1761]. Straubing [776]. Straubing-Thérien [776]. Streaming [1279, 1278, 528, 992]. StreamIt [992]. Streams [119, 557]. Street [288]. Stride [1000]. String [484, 774, 368, 71, 789]. Strings [1131, 474, 759]. Striping [1675]. Strong [497, 906, 417, 443, 454]. Structural [127, 1676, 1706, 214, 1819, 125, 1252, 130, 124, 129, 120, 128]. Structure [432, 949, 1153, 180, 230, 1371, 328, 863, 38, 347]. Structured [1529, 151, 705, 704].

Structures [459, 1685, 1577, 955, 463, 1680, 753, 967, 218, 1, 197, 1703, 1596, 982]. Structuring [1285]. Student [1079, 1354]. Student-Drawn [1354]. Students [1503]. Studies [132]. Study [1471, 696, 1235, 1358, 1815, 857, 218, 1777, 562, 1172, 1808, 407, 1225, 1315, 1366, 1253, 1359, 157, 1745, 694, 1802, 1428, 1843]. Studying [1604, 847]. Style [214, 304]. Subblock [1589]. Subgraphs [57]. Subiteration [900]. Subject [1061, 1327, 1757, 694]. Subject-Object [1757]. Subjective [1544, 1315]. Subpath [986]. Subsets [730]. Subspace [1842]. Substitutability [804]. Substitution [578]. Substitutions [497]. Subsurface [1095]. Subtree [243, 237]. Subtyping [1199]. Subunits [843]. Successes [1188]. Suffices [485]. SugiBib [100]. Suitable [1089]. Suite [537, 1627]. Sum [558, 922]. Summarization [1180, 1280]. Supercomputer [1611]. Supercomputing [1635]. Supernova [1599]. Superscalar [997, 883]. Support [114, 127, 1247, 1585, 135, 666, 1631, 529, 1092, 1537, 1540, 1093, 1014, 1740, 1168, 1749, 1163]. Supported [570]. Supporting [117, 1423, 142, 679, 660, 544]. Supports [1671, 1354]. Suppression [1329, 1339, 1328, 1326, 1330]. Supramolecules [1688]. Surface [915, 279, 231]. Surfaces [921, 895, 896, 26, 925]. Surgery [1248]. Surrogate [1135]. Surveillance [1738]. Survey [848, 712, 1381]. Sustainable [503]. SVL [322]. Swapping [1336]. Switched [625, 630, 643]. Switches [1032]. Switching [625, 1366]. Switzerland [1772]. SWN [1530]. Syllable [1138]. Symbolic [1074, 740, 1111, 323, 1534, 1401, 1542, 311, 1535, 294, 1230, 336, 233]. Symbolic-Statistical [336]. Symbols [203, 455, 782]. Symmetrically [67]. Symmetries [66]. Symmetry [632, 1388, 1637]. Symplectic [1839]. Synchronization [1634]. Synchronized [769]. Synchronizing [768]. Synchronous [1113, 254, 738]. Synonymies [380]. Synopses [556]. Syntactic [807]. Syntactical [347]. Syntax [1373]. Synthesis [866, 622, 457, 629, 850, 637, 849]. Synthesizing [313]. Synthetic [851, 811]. System [614, 416, 1048, 572, 1357, 1006, 96, 843, 1781, 132, 1655, 140, 141, 1464, 798, 1271, 845, 1240, 1459, 1075, 1465, 88, 1028, 1401, 1466, 888, 628, 1257, 1671, 1785, 142, 874, 1467, 181, 855, 808, 633, 1281, 634, 1543, 1644, 1673, 1221, 1740, 1639, 1612, 1579, 1203, 1791, 1603, 133, 814, 815, 1501, 1217, 564, 1531, 235, 1243, 1253, 1702, 1573, 1275, 1405, 860, 1615, 1554, 1354, 385, 387, 1251, 1080, 560, 1290, 1835, 134, 1030, 1049, 1428, 606, 1163, 1356, 149]. System-Level [149]. Systematic [539, 1663]. Systems [1177, 616, 1242, 321, 1503, 1247, 612, 618, 657, 1181, 619, 1829, 1623, 10, 7, 1441, 620, 621, 1471, 1297, 498, 1504, 292, 1498, 551, 1490, 1515, 1571, 1239, 1151, 623, 1235, 798, 296, 1838, 877, 818, 1301, 787, 1218, 12, 1557, 1113, 1660, 625, 626, 307, 1453, 1024, 627, 958, 801, 1500, 1250, 1540, 475, 802, 1508, 17, 1488, 1307, 803, 319, 1519, 1050, 147, 500, 1470, 542, 630, 1310, 1790, 1686, 631, 632, 753, 1572, 1516, 1520, 807, 998, 146, 633, 1469, 1454, 1662, 1492, 1311, 1390, 968, 1663, 309, 1399, 1186, 684, 1003]. Systems [887, 635, 812, 1804, 1651, 1839, 885, 1138, 637, 638, 639, 1714, 1681, 1715, 1224, 148, 1509, 1315, 1547, 1699, 1682, 640, 1801, 1458, 611, 738, 789, 975, 1703, 324, 658, 1684, 643, 1360, 644, 171, 1799, 538, 751, 1568, 1474, 19, 876, 299, 470, 1805].

Tableau [975]. Tables [205, 1327, 1339, 1330]. Tabu [1653]. Tabular [1328, 1359]. Tactile [31]. Tailoring [138]. Tails [1523]. Taking [1462]. Talk [1626]. Tamper [1456]. Target [1593]. Task [1362, 260, 1641, 1815, 1646, 1560, 1625, 1674, 1841, 1225, 1568]. Task-Oriented [1841]. Tasks [1643, 1229, 278, 34, 1609]. Taxonomies [668]. Taxonomy [182]. Taylor [1792]. TDB [560]. TDRBF [869]. TDRGBN [869]. Teachers [1379]. Teaching [1379, 172, 184]. Teams [1319]. Technique [983, 1632, 588, 1333, 1090, 355, 26, 1311, 1283, 975, 445]. Techniques [1787, 1177, 1728, 1045, 620, 1781, 1575, 274, 1447, 1103, 1456, 1796, 1445, 1505, 1739, 1148, 200, 1507, 412, 1554, 847]. Technologies [611]. Technology [1415, 1549, 291, 198, 199, 1497]. Telecom [1182]. Telecommunications [1163]. Telegraphic [1076]. Telephone [1164]. Teller [1692]. Temporal [310, 1247, 1119, 1125, 1793, 1093, 319, 312, 806, 835, 1171, 1064, 119, 974, 447, 1071, 1291, 557, 538]. Term [699, 703, 694]. Termination [458, 740, 832, 741]. Terms [1726, 694]. Terrain [1761]. Terrains [1768]. Tertiary [1651]. Test [323, 274, 1104, 312, 1014, 1572, 1144, 1406, 1443]. Testing [1669, 1050, 102, 712, 183, 1403, 681, 470]. Testing/Correcting [712]. Tetrahedral [1768]. Text [1421, 331, 1282, 1180, 701, 1220, 703, 1077]. Texts [370, 364]. Their [1372, 1663, 377, 299]. Theorem [436, 499, 213, 1193, 1348]. Theorems [895]. Theoretic [739]. Theoretical [795, 916, 854, 715, 1578]. Theories [451, 401]. Theory [207, 1728, 757, 1385, 413, 758, 211, 1237, 1013, 1183, 312, 216, 713, 450, 217, 218, 1432, 220, 151, 1326, 330, 341]. Thérien [776]. Thermal [1761, 1712]. Thermo [1809]. Thermodynamics [353]. Thinning [900, 923]. Thread [1021]. Thread-Modular [1021]. Threads [1022]. Three [1629, 1656, 77, 67, 483, 1692, 1755, 1716, 1831, 76]. Three-Dimensional [1629, 1692, 1716, 1831, 76]. Three-Stage [1755]. Threshold [600, 1214]. Tier [1304]. Tight [416, 1820, 513]. Tiling [903, 1015]. Time [621, 1533, 1498, 924, 296, 845, 274, 840, 1793, 1040, 160, 1817, 1648, 627, 801, 1250, 1642, 1643, 353, 1392, 807, 358, 1816, 503, 1085, 563, 1609, 483, 1732, 1216, 1750, 351, 1804, 1791, 293, 1083, 1547, 485, 642, 324, 27, 1647, 1291, 553]. Time-Dependent [1642, 1816, 503]. Time-Domain [1216]. Time-Sensitive [1750]. Timed [952, 292, 623, 457, 739, 295, 629, 738]. Times [296, 277, 321]. Timing [296, 1606, 849, 1022]. Tissue [1248]. Today [1713]. Together [1127]. Tokens [1361]. Tolerance [1800, 1554]. Tolerated [441]. Tool [321, 170, 323, 1544, 256, 569, 1472, 1530, 801, 472, 1041, 1527, 1014, 1539, 1037, 1538, 1547, 1536, 1405, 365, 1049]. Toolbox [1038, 1833]. Toolkit [1611, 1537]. Tools [1045, 1576, 182, 108, 1339, 375, 1382, 290, 994, 120, 149]. Toolset [1525]. Top [1566, 348, 665]. Top-Down [348]. TOPAS [1825]. Topographic [354, 907]. Topological [905, 911, 896, 893, 923, 724, 1073]. Topologies [893]. Totals [1327]. Touch [25]. Tour [209]. Tourist [138]. Town [1761]. Trace [2, 449]. Traceability [685, 679]. Tracers [1450]. Traces [320]. Tracing [169]. Tracking [26, 22]. Tractability [726]. Tractable [730]. Trade [1359]. Trade-Off [1359]. Tradeoff [1337]. Tradeoffs [147, 882]. Trading [1236, 1555, 1473]. Traffic [1812, 1539, 1528, 684, 1772, 1773, 1227]. Trailer [615]. Trained [1834]. Training [795, 1092, 1245]. Trajectories [1153]. Transaction [1543]. Transactions [525]. Transcoder [1277]. Transducers [774]. Transfer [709, 1688]. Transform [928, 1096, 362]. Transform-Based [1096]. Transformation [1441, 11, 956, 246, 1663, 994].

Transformations [949, 1420, 982]. Transforming [1076]. Transforms [913, 718]. Transiently [1313]. Transition [958, 753, 672, 1400, 309, 975]. Transitions [465, 1692]. Transitivity [460]. Translating [150]. Translation [1419, 757, 707, 1391, 1099]. Translation-Based [707]. Transparent [1313]. Transportation [489, 1734]. Traps [988]. Traveling [1211, 271]. Treatment [673, 1223]. Tree [516, 1564, 549, 459, 1558, 774, 462, 1641, 244, 1410, 415, 456, 968, 426, 345, 270, 550, 901, 348, 749]. Tree-Decomposable [462]. Tree-Interpretable [459]. Tree-Pattern [550]. Treecodes [1752]. Trees [514, 78, 505, 229, 414, 1013, 463, 1657, 963, 1658, 228, 430, 80, 337, 1751]. Trepan [1230]. Trial [1090]. Triality [499]. Triangle [428]. Triangles [912]. Triangulated [896]. Trigger [1785]. Trihedral [910]. TriMedia [157]. TriMedia/CPU64 [157]. trivial [1815]. Tropos [11]. Truck [615]. Truncated [579]. Trust [596]. Truth [1124]. Truth-Values [1124]. Tulip [85]. tuning [551, 373]. Tuple [1313, 1321]. Tuple-Based [1321]. Turing [782, 448]. Tutorial [1387, 1121, 1386, 1440]. Tutoring [1240]. TV(R) [665]. Twilight [423]. Two [1829, 1471, 1141, 1301, 52, 1304, 884, 626, 77, 903, 475, 776, 1469, 1691, 775, 1701, 1720]. Two-Dimensional [1829]. Two-Input [626]. Two-Layer [52]. Two-Letter [776]. Two-Level [884]. Two-Phase [1691]. Two-Step [1720]. Two-Tier [1304]. Two-Way [775]. Type [207, 1006, 498, 210, 211, 748, 1527, 962, 216, 217, 218, 1003, 362, 220, 1703, 388, 380, 1010]. Types [1033, 1295, 1005, 1361, 1191, 405, 1190, 1011]. Typical [275, 1403]. Typing [1190]. Typology [852].

U.S [1340]. Ubiquitous [886, 1238, 871, 1316, 876]. UBSTRING [431]. UDP [1022]. Uintah [1598]. UK [1736, 1583]. Ultimate [1792]. UMiner [572]. UML [1442, 10, 7, 1040, 1358, 1414, 97, 1373, 9, 176, 813, 1, 1532]. UML-Based [176]. Uncertain [917]. Uncertainty [1177, 572, 756, 1707, 1086, 1170, 1172, 703, 1370, 1166, 1316, 533, 1071]. Uncertainty-Based [703]. Unconditionally [1816]. Undecidability [781]. Underlying [334]. Understanding [1378, 183, 1375, 1062]. Undirected [363]. Unexpected [1241]. Unfolding [316, 315]. Unification [1432]. Unified [10, 893]. Uniform [237, 1104]. Unions [523]. Uniqueness [442, 1342]. Unit [157]. Units [1608, 1663, 997]. Universal [782, 476, 546, 822]. Unix [1411]. Unknown [1719]. Unprovable [217]. Unsafe [620]. Unsignalized [1773]. Unstructured [1711, 1621, 1694]. Unsupervised [1273]. Untangling [64]. Until [447]. Update [1006]. Updates [1451]. Updating [258]. Upwind [641]. Urban [275, 1773, 1227]. Usability [1078]. Usage [674]. Use [1781, 1423, 1812, 1044, 13, 248, 1531, 1253, 1265, 693, 570]. Used [1741, 1650]. User [138, 861, 141, 1075, 808, 1382, 1164, 68, 1565, 704, 126, 1428]. User-Level [1565]. User-Sensitive [1075]. Users [142, 1185]. Uses [1465]. Using [298, 1564, 114, 127, 1628, 1398, 1100, 843, 620, 695, 696, 232, 292, 1039, 1490, 247, 924, 1629, 8, 1828, 1379, 583, 1282, 996, 711, 1333, 1104, 258, 1090, 285, 227, 1751, 276, 169, 981, 1234, 213, 308, 252, 1788, 850, 1470, 896, 277, 865, 753, 1448, 1727, 1352, 1610, 304, 1672, 1280, 1528, 332, 701, 1096, 1690, 322, 99, 1311, 1832, 915, 1221, 1740, 426, 279, 1397, 1078, 223, 875, 811, 1560, 1220, 1621, 1625, 1043, 1208, 177, 1652, 814, 1370, 1755, 1314, 697, 1831, 1366, 910, 249, 1206, 1202, 1749, 1592, 974, 1077, 913].

Using [1150, 1317, 271, 578, 1264, 925, 1405, 1267, 1745, 1617, 326, 347, 1354, 1179, 690, 366, 288, 392, 1474, 1231, 523, 1835, 568, 1348, 592, 1843, 557, 1161]. Usually [863]. Utilising [1810]. Utility [825]. Utility-Based [825].

Vacillating [1470]. Valence [779, 780]. Validating [296, 1403]. Validation [677, 682, 1530, 1508, 684, 681]. Valley [1761]. Valuated [779]. Value [1728, 1004, 1592, 1000]. Value-Profile [1000]. Values [531, 1178, 1124]. Variabilities [651]. Variability [673, 650, 675, 674, 1416]. Variable [296, 996, 243]. Variable-Length [243]. Variables [361]. Variant [460]. Variation [226, 1416]. Varieties [959]. Variety [1680]. Vector [1092, 1093, 1749]. Vectors [1572, 1231]. Vehicle [615, 257]. Velocity [352]. Vendors [1417]. Verboseness [448]. Verification [950, 1398, 1027, 754, 305, 292, 622, 1040, 302, 1660, 1021, 629, 748, 1249, 500, 1396, 304, 322, 1390, 1395, 1405, 747]. Verified [1723]. Verify [1397]. Verifying [745, 407, 974]. Versatile [1752]. Version [239, 221]. Versioning [137, 999]. Versions [438, 485]. versus [1620, 1375, 1144, 14]. Vertex [266, 1637]. Vertex-Biconnectivity [266]. Vertex-Symmetry [1637]. Vertices [62]. Very [1092, 308, 530, 1057]. VHDL [1663]. VHDL-Descriptions [1663]. via [616, 699, 1304, 239, 80]. Viability [620]. Vibrations [1706]. ViCo [1368]. Video [905, 1284, 1276, 1285, 1282, 1279, 1278, 1277, 1738, 1269, 564, 1283]. View [1779, 779, 714, 328, 854, 233]. Viewer [1366]. Views [541, 1469, 523]. Virtual [735, 1636, 253, 1775, 1588, 1550, 137, 1479, 842, 885, 565, 1806, 839, 377, 1319]. Viscosity [645]. Visibility [920, 84]. Vision [24, 1098, 278, 23, 842, 1203, 1263, 715, 1095, 27, 25]. Vision-Estimated [842]. Visone [98]. ViSta [107]. Visual [1038, 1372, 1282, 1363, 28, 1268, 1306, 1093, 859, 1541, 374, 1097, 1778, 1274, 863, 1366, 38, 938, 124, 27, 1356]. Visualisation [88, 194]. Visualising [168]. Visualization [191, 89, 170, 105, 95, 182, 1358, 179, 186, 189, 79, 375, 372, 192, 174, 176, 185, 1287, 177, 1637, 376, 55, 93]. Visualizations [1378, 165, 1368, 181, 1365]. Visualize [1370]. Visualized [183]. Visualizer [890]. Visualizing [175, 193, 1286, 178]. Visually [865]. Viterbi [581]. VLIW [1591]. Vmgen} [980]. Vocabulary [845, 1218]. Voice [1544]. Voice-over-IP [1544]. Volcano [817]. Volume [1817, 1096]. Volumes [1711]. Voting [501]. Vowels [817]. Vowels-Oriented [817]. Voxel [913]. VPP [1628]. vs [768, 1336, 1358, 859, 1315].

W [1720]. W-Methods [1720]. Walking [247]. Walksat [1133]. WAN [1567]. Want [1185]. Warehouse [539, 386]. Warehouses [540]. Watermarking [1283]. Watermarks [592]. Wave [1557, 1601, 357, 1095, 87]. Wavelet [362, 558]. Wavelet-Based [558]. Wavepacket [1791]. Waves [356]. Way [423, 775, 1593]. Ways [1604]. Weak [749]. Weakly [1729, 477]. Wearable [880]. Weather [1611]. Web [1293, 1045, 1100, 141, 1363, 160, 571, 1546, 1549, 1414, 375, 372, 519, 573, 1823, 1539, 1803, 1406, 1256, 706, 1791, 1545, 466, 1786, 184, 690, 693, 128, 1161]. Web-Based [1823, 1803, 1406]. Web3D [850]. Weight [670, 602]. Weighted [913]. Weil [479]. Well [1076, 923, 345]. Well-Composed [923]. Well-Formed [1076]. WfMS [542]. Where [1713]. Which [960]. Wide [886, 1414, 128]. Wide-Area [886]. Width [60, 456]. Will [320]. Williams [221]. WilmaScope [88]. Wind [1831]. Winning [669]. Wireless [1182, 1567, 159, 881, 552]. within [899, 2, 806]. without [213, 82]. WML [1403]. Word [1218]. Words [768, 762, 764]. Work [1127, 1364, 1256, 1178, 1059]. Workbench [570]. Workflows [1044]. Workload [1800, 1545]. Workshop [1503, 1433, 1510, 1515, 1434, 1519, 1435, 1514, 1438, 1436]. Workstations [1562]. World [571, 14, 1414, 128]. Worlds [1733]. Would [1159]. wp [1026]. Writing [124]. WWW [375, 853].

XAHM [136]. XCS [1471, 1472, 1468]. XCSI [1475]. XML [1100, 136, 529, 569, 1741, 571, 537, 528, 527, 548, 554]. XML-Based [136]. XPMaps [893]. XQuery [566]. XXL [548].

Y-Chart [146]. Yao [438]. yFiles [93].

Young [1433].

Zero [297]. Zero-Clairvoyant [297]. ZEUS [1496].

References

Parunak:2002:RSS

[1] H. Van Dyke Parunak and James J. Odell. Representing social structures in UML. Lecture Notes in Computer Science, 2222:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2222/22220001.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2222/22220001.
                                              pdf.

Jonker:2002:DDW

[2] Catholijn Jonker, Ioan Alfred Letia, and Jan Treur. Diagnosis of the dynamics within an organization by trace checking of behavioural requirements. Lecture Notes in Computer Science, 2222:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220017.
pdf.

Dignum:2002:AST

[3] Virginia Dignum, Hans Weigand, and Lai Xu. Agent societies: Towards frameworks-based design. Lecture Notes in Computer Science, 2222:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220033.
pdf.

Flores:2002:BCA

[4] Roberto A. Flores and Robert C. Kremer. Bringing coherence to agent conversations. Lecture Notes in Computer Science, 2222:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220050.
pdf.

Koning:2002:EML

[5] Jean-Luc Koning, Marc-Philippe Huget, Jun Wei, and Xu Wang. Extended modeling languages for interaction protocol design. Lecture Notes in Computer Science, 2222:68– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220068.
                                        pdf.

Dulay:2002:PLM

[6] Naranker Dulay, Nicodemos Damianou, Emil Lupu, and Morris Sloman. A policy language for the management of distributed agents. Lecture Notes in Computer Science, 2222:84– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220084.
                                        pdf.

Bauer:2002:UCD

[7] Bernhard Bauer. UML class diagrams revisited in the context of agent-based systems. Lecture Notes in Computer Science, 2222:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220101.
                                        pdf.

Caire:2002:AOA

[8] Giovanni Caire, Wim Coulier, Francisco Garijo, Jorge Gomez, Juan Pavon, Francisco Leal, Paulo Chainho, Paul Kearney, Jamie Stark, Richard Evans, and Philippe Massonet. Agent oriented analysis using Message/UML. Lecture Notes in Computer Science, 2222:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2222/22220119.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2222/22220119.
pdf.

Lind:2002:SAI

[9] Jürgen Lind. Specifying agent interaction protocols with standard UML. Lecture Notes in Computer Science, 2222:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2222/22220136.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2222/22220136.
pdf.

Bauer:2002:AUU

[10] Bernhard Bauer, Federico Bergenti, Philippe Massonet, and James J. Odell. Agents and the UML: a unified notation for agents and multiagent systems? Lecture Notes in Computer Science, 2222:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220148.
pdf.

Bresciani:2002:MER

[11] Paolo Bresciani, Anna Perini, Paolo Giorgini, Fausto Giunchiglia, and John Mylopoulos. Modeling early requirements in Tropos: a transformation based approach. Lecture Notes in Computer Science, 2222:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220151.
                                        pdf.

Dastani:2002:RSL

[12] Mehdi Dastani, Catholijn Jonker, and Jan Treur. A requirement specification language for configuration dynamics of multi-agent systems. Lecture Notes in Computer Science, 2222:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220169.
                                        pdf.

OMalley:2002:DWU

[13] Scott A. O’Malley and Scott A. DeLoach. Determining when to use an agent-oriented software engineering paradigm. Lecture Notes in Computer Science, 2222:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220188.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2222/22220188.
pdf.

Yu:2002:AOM

[14] Eric Yu. Agent-oriented modelling: Software versus the world. Lecture Notes in Computer Science, 2222:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220206.
pdf.

Brauer:2002:EOA

[15] Wilfried Brauer, Matthias Nickles, Michael Rovatsos, Gerhard Weiß, and Kai F. Lorentzen. Expectationoriented analysis and design. Lecture Notes in Computer Science, 2222:226– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220226.
pdf.

Zambonelli:2002:AID

[16] Franco Zambonelli. Abstractions and infrastructures for the design and development of mobile agent organizations. Lecture Notes in Computer Science, 2222:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2222/22220245.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2222/22220245.
                                        pdf.

Gervais:2002:TAD

[17] Marie-Pierre Gervais and Florin Muscutariu. Towards an ADL for designing agent-based systems. Lecture Notes in Computer Science, 2222:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220263.
                                        pdf.

Sparkman:2002:ADC

[18] Clint H. Sparkman, Scott A. DeLoach, and Athie L. Self. Automated derivation of complex agent architectures from analysis specifications. Lecture Notes in Computer Science, 2222:278– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2222/22220278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2222/22220278.
                                        pdf.

Vasconcelos:2002:LML

[19] Wamberto Vasconcelos, David Robertson, Jaume Agustí, Carles Sierra, Michael Wooldridge, Simon Parsons, Christopher Walton, and Jordi Sabater. A lifecycle for models of large multi-agent systems. Lecture Notes in Computer Science, 2222:297– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2222/22220297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2222/22220297.
pdf.

Anonymous:2002:AIa

[20] Anonymous. Author index. Lecture Notes in Computer Science, 2222:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2222/2222auth.pdf.

Keselman:2002:GMA

[21] Yakov Keselman and Sven Dickinson. Generic model abstraction from examples. Lecture Notes in Computer Science, 2238:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380001.
pdf.

Kluge:2002:TMM

[22] Boris Kluge. Tracking multiple moving objects in populated, public environments. Lecture Notes in Computer Science, 2238:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380025.
pdf.

Krose:2002:OVA

[23] B. J. A. Kröse, N. Vlassis, and R. Bunschoten. Omnidirectional vision for appearance-based robot localization. Lecture Notes in Computer Science, 2238:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2238/22380039.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380039.
                                        pdf.

Christensen:2002:VI

[24] H. I. Christensen, D. Kragic, and F. Sandberg. Vision for interaction. Lecture Notes in Computer Science, 2238:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2238/22380051.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380051.
                                        pdf.

Wurtz:2002:VTG

[25] Rolf P. Würtz. Vision and touch for grasping. Lecture Notes in Computer Science, 2238:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2238/22380074.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2238/22380074.pdf.

Jiang:2002:NTE

[26] X. Jiang, S. Hofer, T. Stahs, I. Ahrns, and H. Bunke. A new technique for the extraction and tracking of surfaces in range image sequences. Lecture Notes in Computer Science, 2238: 87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380087.
pdf.

Vincze:2002:DAV

[27] Markus Vincze, Minu Ayromlou, Stefan Chroust, Michael Zillich, Wolfgang Ponweiser, and Dietmar Legenstein. Dynamic aspects of visual servoing and a framework for realtime 3D vision for robotics. Lecture Notes in Computer Science, 2238:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380101.
pdf.

Corke:2002:PIB

[28] Peter Corke, Seth Hutchinson, and Nicholas R. Gans. Partitioned image-based visual servo control: Some new results. Lecture Notes in Computer Science, 2238:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380122.
pdf.

Schiele:2002:TRP

[29] Bernt Schiele, Martin Spengler, and Hannes Kruppa. Towards robust perception and model integration. Lecture Notes in Computer Science, 2238:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380141.
                                        pdf.

Rencken:2002:LCG

[30] Wolfgang Rencken, Wendelin Feiten, and Martin Soika. Large consistent geometric landmark maps. Lecture Notes in Computer Science, 2238:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380159.
                                        pdf.

Kristensen:2002:TMR

[31] Steen Kristensen, Mathias Neumann, Sven Horstmann, Frieder Lohnert, and Andreas Stopp. Tactile man-robot interaction for an industrial service robot. Lecture Notes in Computer Science, 2238:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380177.
                                        pdf.

Hert:2002:MRM

[32] Susan Hert and Brad Richards. Multiple-robot motion planning = parallel processing + geometry. Lecture Notes in Computer Science, 2238:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380195.
pdf.

Elfes:2002:MCP

[33] Alberto Elfes, Samuel S. Bueno, Josué J. G. Ramos, Ely C. de Paiva, Marcel Bergerman, José R. H. Carvalho, Silvio M. Maeta, Luiz G. B. Mirisola, Bruno G. Faria, and José R. Azinheira. Modelling, control and perception for an autonomous robotic airship. Lecture Notes in Computer Science, 2238: 216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380216.
pdf.

Icking:2002:CCN

[34] Christian Icking, Thomas Kamphans, Rolf Klein, and Elmar Langetepe. On the competitive complexity of navigation tasks. Lecture Notes in Computer Science, 2238:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380245.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380245.
                                        pdf.

Stappen:2002:GPF

[35] A. Frank van der Stappen, Robert-Paul Berretty, Ken Goldberg, and Mark H. Overmars. Geometry and part feeding. Lecture Notes in Computer Science, 2238:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380259.
                                        pdf.

Cabrera-Gamez:2002:CCO

[36] Jorge Cabrera-Gámez, Antonio Carlos Domínguez-Brito, and Daniel Hernández-Sosa. CoolBOT: a componentoriented programming framework for robotics. Lecture Notes in Computer Science, 2238:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2238/22380282.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2238/22380282.pdf.

Eveland:2002:PFE

[37] Christopher K. Eveland. Particle filtering with evidential reasoning. Lecture Notes in Computer Science, 2238: 305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380305.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2238/22380305.
pdf.

Sagerer:2002:SPL

[38] G. Sagerer, C. Bauckhage, E. Braun, J. Fritsch, F. Kummert, F. Lömker, and S. Wachsmuth. Structure and process: Learning of visual models and construction plans for complex objects. Lecture Notes in Computer Science, 2238:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380317.
pdf.

Maass:2002:AFL

[39] Wolfgang Maass, Gerald Steinbauer, and Roland Koholka. Autonomous fast learning in a mobile robot. Lecture Notes in Computer Science, 2238:345– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2238/22380345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2238/22380345.
pdf.

Sutton:2002:ECF

[40] Melanie A. Sutton, Louise Stark, and Ken Hughes. Exploiting context in function-based reasoning. Lecture Notes in Computer Science, 2238:357– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2238/22380357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2238/22380357.
                                        pdf.

Anonymous:2002:AIb

[41] Anonymous. Author index. Lecture Notes in Computer Science, 2238:375– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2238/2238auth.pdf.

Naumann:2002:I

[42] Felix Naumann. 1 introduction. Lecture Notes in Computer Science, 2261: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2261/22610003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2261/22610003.
                                        pdf.

Naumann:2002:IAI

[43] Felix Naumann. 2 integrating autonomous information sources. Lecture Notes in Computer Science, 2261:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2261/22610011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2261/22610011.
                                        pdf.

Naumann:2002:IQC

[44] Felix Naumann. 3 information quality criteria. Lecture Notes in Computer Science, 2261:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2261/22610029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2261/22610029.
pdf.

Naumann:2002:QRM

[45] Felix Naumann. 4 quality ranking methods. Lecture Notes in Computer Science, 2261:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2261/22610051.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2261/22610051.pdf.

Naumann:2002:QDQ

[46] Felix Naumann. 5 quality-driven query planning. Lecture Notes in Computer Science, 2261:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2261/22610069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2261/22610069.
pdf.

Naumann:2002:QPR

[47] Felix Naumann. 6 query planning revisited. Lecture Notes in Computer Science, 2261:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2261/22610089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2261/22610089.
                                        pdf.

Naumann:2002:CD

[48] Felix Naumann. 7 completeness of data. Lecture Notes in Computer Science, 2261:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2261/22610101.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2261/22610101.pdf.

Naumann:2002:CDQ

[49] Felix Naumann. 8 completenessdriven query optimization. Lecture Notes in Computer Science, 2261:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2261/22610123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2261/22610123.
                                        pdf.

Naumann:2002:C

[50] Felix Naumann. 9 conclusion. Lecture Notes in Computer Science, 2261:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2261/22610153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2261/22610153.
                                        pdf.

Naumann:2002:R

[51] Felix Naumann. References. Lecture Notes in Computer Science, 2261:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2261/22610159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2261/22610159.
pdf.

Dujmovic:2002:FPA

[52] V. Dujmović, M. Fellows, M. Hallett, M. Kitching, Giuseppe Liotta, C. McCartin, N. Nishimura, P. Ragde, F. Rosamond, M. Suderman, S. Whitesides, and David R. Wood. A fixed-parameter approach to two-layer planarization. Lecture Notes in Computer Science, 2265:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650001.
pdf.

Healy:2002:HLD

[53] Patrick Healy and Nikola S. Nikolov. How to layer a directed acyclic graph. Lecture Notes in Computer Science, 2265:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2265/22650016.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2265/22650016.
pdf.

Brandes:2002:FSH

[54] Ulrik Brandes and Boris Köpf. Fast and simple horizontal coordinate assignment. Lecture Notes in Computer Science, 2265:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650031.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650031.
                                        pdf.

Six:2002:AVP

[55] Janet M. Six and Ioannis G. Tollis. Automated visualization of process diagrams. Lecture Notes in Computer Science, 2265:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650045.
                                        pdf.

Battista:2002:PCG

[56] Giuseppe Di Battista, Walter Didimo, and A. Marcandalli. Planarization of clustered graphs. Lecture Notes in Computer Science, 2265:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650060.
                                        pdf.

Edwards:2002:AFL

[57] Keith Edwards and Graham Farr. An algorithm for finding large induced planar subgraphs. Lecture Notes in Computer Science, 2265:75– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650075.
pdf.

Fraysseix:2002:CDC

[58] Hubert de Fraysseix and Patrice Ossona de Mendez. A characterization of DFS cotree critical graphs. Lecture Notes in Computer Science, 2265:84– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650084.
pdf.

Djidjev:2002:ILB

[59] Hristo Djidjev and Imrich Vrt’o. An improved lower bound for crossing numbers. Lecture Notes in Computer Science, 2265:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2265/22650096.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2265/22650096.pdf.

Hlineny:2002:CCG

[60] Petr Hliněný. Crossing-critical graphs and path-width. Lecture Notes in Computer Science, 2265:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650102.
pdf.

Munoz:2002:OSC

[61] Xavier Muñoz, W. Unger, and Imrich Vrt’o. One sided crossing minimization is NP-hard for sparse graphs. Lecture Notes in Computer Science, 2265: 115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650115.
pdf.

Eiglsperger:2002:FCO

[62] Markus Eiglsperger and Michael Kaufmann. Fast compaction for orthogonal drawings with vertices of prescribed size. Lecture Notes in Computer Science, 2265:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650124.
pdf.

Binucci:2002:LHO

[63] Carla Binucci, Walter Didimo, Giuseppe Liotta, and Maddalena Nonato. Labeling heuristics for orthogonal drawings. Lecture Notes in Computer Science, 2265:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2265/22650139.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2265/22650139.
pdf.

Pach:2002:UP

[64] János Pach and Gábor Tardos. Untangling a polygon. Lecture Notes in Computer Science, 2265:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650154.
pdf.

Duncan:2002:DFE

[65] Christian A. Duncan, Alon Efrat, Stephen G. Kobourov, and Carola Wenk. Drawing with fat edges. Lecture Notes in Computer Science, 2265: 162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650162.
pdf.

Buchheim:2002:DSB

[66] Christoph Buchheim and Michael Jünger. Detecting symmetries by branch & cut. Lecture Notes in Computer Science, 2265:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650178.
                                        pdf.

Hong:2002:DGS

[67] Seok-Hee Hong. Drawing graphs symmetrically in three dimensions. Lecture Notes in Computer Science, 2265: 189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650189.
                                        pdf.

Nascimento:2002:UHD

[68] Hugo A. D. do Nascimento and Peter Eades. User hints for directed graph drawing. Lecture Notes in Computer Science, 2265:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650205.
                                        pdf.

Friedrich:2002:GDM

[69] Carsten Friedrich and Michael E. Houle. Graph drawing in motion II. Lecture Notes in Computer Science, 2265:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2265/22650220.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2265/22650220.
pdf.

North:2002:OHG

[70] Stephen C. North and Gordon Woodhull. Online hierarchical graph drawing. Lecture Notes in Computer Science, 2265:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2265/22650232.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2265/22650232.pdf.

Pach:2002:RSG

[71] János Pach and Géza Tóth. Recognizing string graphs is decidable. Lecture Notes in Computer Science, 2265: 247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650247.
pdf.

Cerny:2002:IGS

[72] Jakub Černý, Daniel Král’, Helena Nyklová, and Ondřej Pangrác. On intersection graphs of segments with prescribed slopes. Lecture Notes in Computer Science, 2265:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650261.
pdf.

Kruja:2002:SNH

[73] Eriola Kruja, Joe Marks, Ann Blair, and Richard Waters. A short note on the history of graph drawing. Lecture Notes in Computer Science, 2265:272– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650272.
pdf.

Adamec:2002:TAI

[74] Jan Adamec and Jaroslav Nešetřil. Towards an aesthetic invariant for graph drawing. Lecture Notes in Computer Science, 2265:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650287.
pdf.

Biedl:2002:ODF

[75] Therese Biedl, John R. Johansen, Thomas Shermer, and David R. Wood.

Orthogonal drawings with few layers. Lecture Notes in Computer Science, 2265:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2265/22650297.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2265/22650297.
pdf.

Wood:2002:BDB

[76] David R. Wood. Bounded degree book embeddings and three-dimensional orthogonal graph drawing. Lecture Notes in Computer Science, 2265:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650312.
pdf.

Felsner:2002:SLD

[77] Stefan Felsner, Giuseppe Liotta, and Stephen Wismath. Straight-line drawings on restricted integer grids in two and three dimensions. Lecture Notes in Computer Science, 2265:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650328.
pdf.

Babilon:2002:LDE

[78] Robert Babilon, Jiří Matoušek, Jana Maxová, and Pavel Valtr. Lowdistortion embeddings of trees. Lecture Notes in Computer Science, 2265: 343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650343.
                                        pdf.

Groller:2002:IDT

[79] Eduard Gröller. Insight into data through visualization. Lecture Notes in Computer Science, 2265:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650352.
                                        pdf.

Liao:2002:FPO

[80] Chien-Chih Liao, Hsueh-I. Lu, and Hsu-Chun Yen. Floor-planning via orderly spanning trees. Lecture Notes in Computer Science, 2265:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650367.
                                        pdf.

Freivalds:2002:DGL

[81] Karlis Freivalds, Ugur Dogrusoz, and Paulis Kikusts. Disconnected graph layout and the polyomino packing approach. Lecture Notes in Computer Science, 2265:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650378.
pdf.

Rahman:2002:ODP

[82] Md. Saidur Rahman, Mahmuda Naznin, and Takao Nishizeki. Orthogonal drawings of plane graphs without bends. Lecture Notes in Computer Science, 2265:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650392.
pdf.

Duncan:2002:PCD

[83] Christian A. Duncan and Stephen G. Kobourov. Polar coordinate drawing of planar graphs with good angular resolution. Lecture Notes in Computer Science, 2265:407–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650407.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650407.
pdf.

Hutchinson:2002:PVR

[84] Joan P. Hutchinson. On polar visibility representations of graphs. Lecture Notes in Computer Science, 2265:422– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650422.
                                        pdf.

David:2002:T

[85] Auber David. Tulip. Lecture Notes in Computer Science, 2265:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650435.
                                        pdf.

Sander:2002:IJG

[86] Georg Sander and Adrian Vasiliu. The ILOG JViews graph layout module. Lecture Notes in Computer Science, 2265:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2265/22650438.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650438.
                                        pdf.

Giacomo:2002:W

[87] Emilio di Giacomo and Giuseppe Liotta. WAVE. Lecture Notes in Computer Science, 2265:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2265/22650440.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650440.
pdf.

Dwyer:2002:WIG

[88] Tim Dwyer and Peter Eckersley. WilmaScope — an interactive 3D graph visualisation system. Lecture Notes in Computer Science, 2265:442– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650442.
pdf.

Barbagallo:2002:EVC

[89] Gabriele Barbagallo, Andrea Carmignani, Giuseppe Di Battista, Walter Didimo, and Maurizio Pizzonia. Exploration and visualization of computer networks: Polyphemus and hermes. Lecture Notes in Computer Science, 2265: 444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650444.
pdf.

Lewerentz:2002:C

[90] Claus Lewerentz, Frank Simon, and Frank Steinbrückner. CrocoCosmos. Lecture Notes in Computer Science, 2265:446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2265/22650446.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650446.
                                        pdf.

Bridgeman:2002:GDS

[91] Stina Bridgeman and Roberto Tamassia. The graph drawing server. Lecture Notes in Computer Science, 2265: 448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650448.
                                        pdf.

Battista:2002:DDS

[92] Giuseppe Di Battista, Walter Didimo, Maurizio Patrignani, and Maurizio Pizzonia. Drawing database schemas with DBdraw. Lecture Notes in Computer Science, 2265:451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650451.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650451.
                                        pdf.

Wiese:2002:YVA

[93] Roland Wiese, Markus Eiglsperger, and Michael Kaufmann. yFiles: Visualization and automatic layout of graphs. Lecture Notes in Computer Science, 2265:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2265/22650453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650453.
pdf.

Brandenburg:2002:B

[94] Franz J. Brandenburg, Michael Forster, Andreas Pick, Marcus Raitner, and Falk Schreiber. BioPath. Lecture Notes in Computer Science, 2265:455– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650455.
pdf.

Bertault:2002:GVA

[95] François Bertault, Wendy Feng, Uli Fößmeier, Gabe Grigorescu, and Brendan Madden. Graph visualization API library for application builders. Lecture Notes in Computer Science, 2265: 457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650457.
pdf.

Bagga:2002:JJB

[96] Jay Bagga and Adrian Heinz. JGraph — A Java based system for drawing graphs and running graph algorithms. Lecture Notes in Computer Science, 2265:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2265/22650459.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650459.
                                        pdf.

Gutwenger:2002:CAL

[97] Carsten Gutwenger, Michael Jünger, Karsten Klein, Joachim Kupke, Sebastian Leipert, and Petra Mutzel. Caesar automatic layout of UML class diagrams. Lecture Notes in Computer Science, 2265:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650461.
                                        pdf.

Baur:2002:V

[98] Michael Baur, Marc Benkert, Ulrik Brandes, Sabine Cornelsen, Marco Gaertler, Boris Köpf, Jürgen Lerner, and Dorothea Wagner. Visone. Lecture Notes in Computer Science, 2265: 463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650463.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650463.
                                        pdf.

Lauther:2002:GSC

[99] Ulrich Lauther and Andreas Stübinger. Generating schematic cable plans using springembedder methods. Lecture Notes in Computer Science, 2265:465– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650465.
pdf.

Eichelberger:2002:S

[100] Holger Eichelberger. SugiBib. Lecture Notes in Computer Science, 2265:467– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650467.
pdf.

Delahousse:2002:KIM

[101] Jean Delahousse and Pascal Auillans. Knowledge index manager. Lecture Notes in Computer Science, 2265:469– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650469.
pdf.

Hotz:2002:PTG

[102] Günter Hotz and Steffen Lohse. Planarity testing of graphs on base of a spring model. Lecture Notes in Computer Science, 2265:471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650471.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2265/22650471.
                                        pdf.

Gutwenger:2002:A

[103] Carsten Gutwenger, Michael Jünger, Gunnar W. Klau, Sebastian Leipert, Petra Mutzel, and René Weiskircher. AGD. Lecture Notes in Computer Science, 2265:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2265/22650473.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2265/22650473.pdf.

Didimo:2002:IPD

[104] Walter Didimo, Maurizio Patrignani, and Maurizio Pizzonia. Industrial plant drawer. Lecture Notes in Computer Science, 2265:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650475.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650475.
                                        pdf.

Batagelj:2002:PAV

[105] Vladimir Batagelj and Andrej Mrvar. Pajek — analysis and visualization of large networks. Lecture Notes in Computer Science, 2265:477– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650477.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2265/22650477.
pdf.

Ryall:2002:G

[106] Kathy Ryall. GLIDE. Lecture Notes in Computer Science, 2265:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650479.
pdf.

Castello:2002:V

[107] Rodolfo Castelló, Rym Mili, and Ioannis G. Tollis. ViSta. Lecture Notes in Computer Science, 2265:481–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650481.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650481.
pdf.

Ellson:2002:GOS

[108] John Ellson, Emden Gansner, Lefteris Koutsofios, Stephen C. North, and Gordon Woodhull. Graphviz — open source graph drawing tools. Lecture Notes in Computer Science, 2265:483– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2265/22650483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2265/22650483.
pdf.

Winter:2002:EGG

[109] Andreas Winter. Exchanging graphs with GXL. Lecture Notes in Computer Science, 2265:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2265/22650485.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2265/22650485.pdf.

Brandes:2002:GPR

[110] Ulrik Brandes, Markus Eiglsperger, Ivan Herman, Michael Himsolt, and M. Scott Marshall. GraphML progress report. Lecture Notes in Computer Science, 2265:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2265/22650501.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650501.
                                        pdf.

Biedl:2002:GDC

[111] Therese Biedl and Franz J. Brandenburg. Graph-drawing contest report. Lecture Notes in Computer Science, 2265:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2265/22650513.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2265/22650513.
                                        pdf.

Anonymous:2002:AIc

[112] Anonymous. Author index. Lecture Notes in Computer Science, 2265:523– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2265/2265auth.pdf.

Reich:2002:IO

[113] Sigi Reich. Introduction to OHS7. Lecture Notes in Computer Science, 2266: 4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660004.
pdf.

Anderson:2002:UOH

[114] Kenneth M. Anderson and Susanne A. Sherba. Using open hypermedia to support information integration. Lecture Notes in Computer Science, 2266:8– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660008.
pdf.

Haake:2002:ACO

[115] Jörg M. Haake. Applying collaborative open hypermedia concepts to extended enterprise engineering and operation. Lecture Notes in Computer Science, 2266:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660017.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660017.
                                        pdf.

Tanaka:2002:MMM

[116] Yuzuru Tanaka, Jun Fujima, and Tsuyoshi Sugibuchi. Meme media and meme pools for re-editing and redistributing intellectual assets. Lecture Notes in Computer Science, 2266:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660028.
                                        pdf.

Beales:2002:PES

[117] Richard Beales, Don Cruickshank, David DeRoure, Nick Gibbins, Ben Juby, Danius T. Michaelides, and Kevin R. Page. The pipeline of enrichment: Supporting link creation for continuous media. Lecture Notes in Computer Science, 2266:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660047.
                                        pdf.

Michaelides:2002:ALC

[118] Danius T. Michaelides, David E. Millard, Mark J. Weal, and David DeRoure. Auld leaky: a contextual open hypermedia link server. Lecture Notes in Computer Science, 2266:59– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660059.
pdf.

Ridgway:2002:FAO

[119] Neil Ridgway and David DeRoure. FOHM+RTSP: Applying open hypermedia and temporal linking to audio streams. Lecture Notes in Computer Science, 2266:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660071.
pdf.

Wiil:2002:DTC

[120] Uffe Kock Wiil. Development tools in component-based structural computing environments. Lecture Notes in Computer Science, 2266:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660082.
pdf.

Nurnberg:2002:PRP

[121] Peter J. Nürnberg and David L. Hicks. Peer-reviewed, publishable hypertexts: a first look. Lecture Notes in Computer Science, 2266:94– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660094.
                                        pdf.

Neumuller:2002:BSI

[122] Moritz Neumüller. Because I seek an image, not a book. Lecture Notes in Computer Science, 2266:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660104.
                                        pdf.

Tzagarakis:2002:IS

[123] Manolis M. Tzagarakis. Introduction to SC3. Lecture Notes in Computer Science, 2266:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660118.
                                        pdf.

Shapiro:2002:WHS

[124] Saul Shapiro. Writing the holes: “Structural” reflections of a visual artist. Lecture Notes in Computer Science, 2266:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2266/22660120.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2266/22660120.pdf.

Kyriakopoulou:2002:BSC

[125] Maria Kyriakopoulou, Dimitris Avramidis, Michalis Vaitis, Manolis M. Tzagarakis, and Dimitris Christodoulakis. Broadening structural computing towards hypermedia development. Lecture Notes in Computer Science, 2266: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660131.
pdf.

Wang:2002:GUI

[126] Weigang Wang and Alejandro Fernández. A graphical user interface integrating features from different hypertext domains. Lecture Notes in Computer Science, 2266:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660141.
pdf.

Anderson:2002:USC

[127] Kenneth M. Anderson and Susanne A. Sherba. Using structural computing to support information integration. Lecture Notes in Computer Science, 2266: 151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660151.
                                        pdf.

Wiil:2002:PSC

[128] Uffe Kock Wiil and David L. Hicks. Providing structural computing services on the World Wide Web. Lecture Notes in Computer Science, 2266: 160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660160.
                                        pdf.

Tata:2002:CSS

[129] Samir Tata, David L. Hicks, and Uffe Kock Wiil. Cooperation services in a structural computing environment. Lecture Notes in Computer Science, 2266:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660172.
                                        pdf.

Nurnberg:2002:SCR

[130] Peter J. Nürnberg and Monica M. C. Schraefel. Structural computing and its relationships to other fields. Lecture Notes in Computer Science, 2266: 183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2266/22660183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660183.
pdf.

Bra:2002:IA

[131] Paul M. E. De Bra. Introduction to AH3. Lecture Notes in Computer Science, 2266:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660199.
pdf.

Bontcheva:2002:IES

[132] Kalina Bontcheva. The impact of empirical studies on the design of an adaptive hypertext generation system. Lecture Notes in Computer Science, 2266: 201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660201.
pdf.

Papanikolaou:2002:IIS

[133] Kyparisia A. Papanikolaou, Maria Grigoriadou, Harry Kornilakis, and George D. Magoulas. INSPIRE: An INtelligent system for personalized instruction in a remote environment. Lecture Notes in Computer Science, 2266: 215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660215.
                                        pdf.

Weber:2002:DAI

[134] Gerhard Weber, Hans-Christian Kuhl, and Stephan Weibelzahl. Developing adaptive Internet based courses with the authoring system NetCoach. Lecture Notes in Computer Science, 2266: 226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660226.
                                        pdf.

Bailey:2002:LAC

[135] Christopher Bailey, Samhaa R. El-Beltagy, and Wendy Hall. Link augmentation: a context-based approach to support adaptive hypermedia. Lecture Notes in Computer Science, 2266: 239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660239.
                                        pdf.

Cannataro:2002:XXB

[136] Mario Cannataro and Andrea Pugliese. XAHM: An XML-based adaptive hypermedia model and its implementation. Lecture Notes in Computer Science, 2266:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2266/22660252.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2266/22660252.pdf.

Iksal:2002:RVV

[137] Sébastien Iksal and Serge Garlatti. Revisiting and versioning in virtual special reports. Lecture Notes in Computer Science, 2266:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660264.
pdf.

Ardissono:2002:TRT

[138] L. Ardissono, A. Goy, G. Petrone, M. Segnan, and P. Torasso. Tailoring the recommendation of tourist information to heterogeneous user groups. Lecture Notes in Computer Science, 2266: 280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660280.
pdf.

Graef:2002:AAN

[139] Guntram Graef and Christian Schaefer. Application of ART2 networks and self-organizing maps to collaborative filtering. Lecture Notes in Computer Science, 2266:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660296.
                                        pdf.

Bueno:2002:MOO

[140] David Bueno, Ricardo Conejo, and Amos A. David. METIOREW: An objective oriented content based and collaborative recommending system. Lecture Notes in Computer Science, 2266: 310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660310.
                                        pdf.

Buono:2002:IUD

[141] Paolo Buono, Maria Francesca Costabile, Stefano Guida, and Antonio Piccinno. Integrating user data and collaborative filtering in a Web recommendation system. Lecture Notes in Computer Science, 2266:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2266/22660315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2266/22660315.
                                        pdf.

Hijikata:2002:AHS

[142] Yoshinori Hijikata, Tetsuya Yoshida, and Shogo Nishida. Adaptive hypermedia system for supporting information providers in directing users through hyperspace. Lecture Notes in Computer Science, 2266:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660322.
pdf.

Wadge:2002:CAA

[143] William W. Wadge and Monica M. C. Schraefel. A complementary approach for adaptive and adaptable hypermedia: Intensional hypertext. Lecture Notes in Computer Science, 2266:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2266/22660327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2266/22660327.
pdf.

Anonymous:2002:AId

[144] Anonymous. Author index. Lecture Notes in Computer Science, 2266:335– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2266/2266auth.pdf.

Bhattacharya:2002:CAR

[145] Bishnupriya Bhattacharya and Shuvra S. Bhattacharyya. Consistency analysis of reconfigurable dataflow specifications. Lecture Notes in Computer Science, 2268:1–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2268/22680001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2268/22680001.pdf.

Kienhuis:2002:MDP

[146] Bart Kienhuis, Ed F. Deprettere, Pieter van der Wolf, and Kees Vissers. A methodology to design programmable embedded systems — the Y-chart approach. Lecture Notes in Computer Science, 2268:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2268/22680018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2268/22680018.
                                        pdf.

Haubelt:2002:FCT

[147] Christian Haubelt, Jürgen Teich, Kai Richter, and Rolf Ernst. Flexibility/ cost-tradeoffs of platform-based systems. Lecture Notes in Computer Science, 2268:38–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2268/22680038.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2268/22680038.pdf.

Pimentel:2002:TED

[148] A. D. Pimentel, S. Polstra, F. Terpstra, A. W. van Halderen, J. E. Coffland, and L. O. Hertzberger. Towards efficient design space exploration of heterogeneous embedded media systems. Lecture Notes in Computer Science, 2268:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680057.
pdf.

Zivkovic:2002:OMT

[149] Vladimir D. Živković and Paul Lieverse. An overview of methodologies and tools in the field of system-level design. Lecture Notes in Computer Science, 2268:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680074.
pdf.

Deprettere:2002:TIA

[150] Ed F. Deprettere, Edwin Rijpkema, and Bart Kienhuis. Translating imperative affine nested loop programs into process networks. Lecture Notes in Computer Science, 2268:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680089.
pdf.

Quinton:2002:SSR

[151] Patrice Quinton and Tanguy Risset.

Structured scheduling of recurrence equations: Theory and practice. Lecture Notes in Computer Science, 2268: 112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2268/22680112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2268/22680112.
                                        pdf.

Teich:2002:EPA

[152] Jürgen Teich and Lothar Thiele. Exact partitioning of affine dependence algorithms. Lecture Notes in Computer Science, 2268:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2268/22680135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2268/22680135.
                                        pdf.

Bednara:2002:GDL

[153] Marcus Bednara, Frank Hannig, and Jürgen Teich. Generation of distributed loop control. Lecture Notes in Computer Science, 2268:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2268/22680154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2268/22680154.
                                        pdf.

Knijnenburg:2002:IC

[154] P. M. W. Knijnenburg, T. Kisuki, and M. F. P. O’Boyle. Iterative compilation. Lecture Notes in Computer Science, 2268:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680171.
pdf.

Pirsch:2002:PAM

[155] P. Pirsch, A. Freimann, C. Klar, and J. P. Wittenburg. Processor architectures for multimedia applications. Lecture Notes in Computer Science, 2268: 188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680188.
pdf.

Wong:2002:MRE

[156] Stephan Wong, Stamatis Vassiliadis, and Sorin Cotofana. Microcoded reconfigurable embedded processors: Current developments. Lecture Notes in Computer Science, 2268:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680207.
pdf.

Sima:2002:RFU

[157] Mihai Sima, Sorin Cotofana, Stamatis Vassiliadis, Jos T. J. van Eijndhoven, and Kees Vissers. A reconfigurable functional unit for TriMedia/CPU64. A case study. Lecture Notes in Computer Science, 2268:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680224.
pdf.

Muller:2002:CCP

[158] Henk Muller, Dan Page, James Irwin, and David May. Caches with compositional performance. Lecture Notes in Computer Science, 2268:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680242.
pdf.

Taylor:2002:DAA

[159] Clark N. Taylor, Debashis Panigrahi, and Sujit Dey. Design of an adaptive architecture for energy efficient wireless image communication. Lecture Notes in Computer Science, 2268: 260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680260.
pdf.

Desmet:2002:DCL

[160] Dirk Desmet, Prabhat Avasare, Paul Coene, Stijn Decneut, Filip Hendrickx, Théodore Marescaux, Jean-Yves Mignolet, Robert Pasko, Patrick Schaumont, and Diederik Verkest. Design of Cam-E-leon, a run-time reconfigurable Web camera. Lecture Notes in Computer Science, 2268:274– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680274.
pdf.

Kuzmanov:2002:AMM

[161] Georgi Kuzmanov, Stamatis Vassiliadis, and Jos T. J. van Eijndhoven. A 2D addressing mode for multimedia applications. Lecture Notes in Computer Science, 2268:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2268/22680291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2268/22680291.
pdf.

Glossner:2002:JED

[162] C. John Glossner, Michael Schulte, and Stamatis Vassiliadis. A Java-enabled DSP. Lecture Notes in Computer Science, 2268:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2268/22680307.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2268/22680307.pdf;   http://mesa.
                                        ece.wisc.edu/publications/cp_2002-
                                        01.pdf.

Anonymous:2002:AIe

[163] Anonymous. Author index. Lecture Notes in Computer Science, 2268:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2268/2268auth.pdf.

Kerren:2002:AAI

[164] Andreas Kerren and John T. Stasko. Algorithm animation — introduction. Lecture Notes in Computer Science, 2269:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2269/22690001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2269/22690001.
                                        pdf.

Demetrescu:2002:SAV

[165] Camil Demetrescu, Irene Finocchi, and John T. Stasko. Specifying algorithm visualizations: Interesting events or state mapping? Lecture Notes in Computer Science, 2269:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2269/22690016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2269/22690016.
                                        pdf.

Ben-Ari:2002:PPA

[166] Mordechai Ben-Ari, Niko Myller, Erkki Sutinen, and Jorma Tarhio. Perspectives on program animation with jeliot. Lecture Notes in Computer Science, 2269:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690031.
pdf.

Diehl:2002:AAL

[167] Stephan Diehl, Carsten Görg, and Andreas Kerren. Animating algorithms live and post mortem. Lecture Notes in Computer Science, 2269:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690046.
pdf.

Noble:2002:VOA

[168] James Noble. Visualising objects: Abstraction, encapsulation, aliasing, and ownership. Lecture Notes in Computer Science, 2269:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690058.
pdf.

Francik:2002:AAU

[169] Jaroslaw Francik. Algorithm animation using data flow tracing. Lecture Notes in Computer Science, 2269:73– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2269/22690073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2269/22690073.
                                        pdf.

Basken:2002:GGT

[170] Matthias Bäsken and Stefan Näher. GeoWin A generic tool for interactive visualization of geometric algorithms. Lecture Notes in Computer Science, 2269:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2269/22690088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2269/22690088.
                                        pdf.

Tal:2002:AAS

[171] Ayellet Tal. Algorithm animation systems for constrained domains. Lecture Notes in Computer Science, 2269:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2269/22690101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2269/22690101.
                                        pdf.

Fleischer:2002:AAT

[172] Rudolf Fleischer and Luděk Kučera. Algorithm animation for teaching. Lecture Notes in Computer Science, 2269: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690113.
pdf.

Mili:2002:SEI

[173] Rym Mili and Renee Steiner. Software engineering — introduction. Lecture Notes in Computer Science, 2269:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690129.
pdf.

Koschke:2002:SVR

[174] Rainer Koschke. Software visualization for reverse engineering. Lecture Notes in Computer Science, 2269:138– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690138.
pdf.

Pauw:2002:VEJ

[175] Wim De Pauw, Erik Jensen, Nick Mitchell, Gary Sevitsky, John Vlissides, and Jeaha Yang. Visualizing the execution of Java programs. Lecture Notes in Computer Science, 2269:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690151.
pdf.

Mehner:2002:JUB

[176] Katharina Mehner. JaVis: a UML-based visualization and debugging environment for concurrent Java programs. Lecture Notes in Computer Science, 2269:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690163.
pdf.

Oechsle:2002:JAP

[177] Rainer Oechsle and Thomas Schmitt. JAVAVIS: Automatic program visualization with object and sequence diagrams using the Java debug interface (JDI). Lecture Notes in Computer Science, 2269:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690176.
pdf.

Zimmermann:2002:VMG

[178] Thomas Zimmermann and Andreas Zeller. Visualizing memory graphs. Lecture Notes in Computer Science, 2269:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2269/22690191.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2269/22690191.
pdf.

Domingue:2002:SVE

[179] John Domingue. Software visualization and education — introduction. Lecture Notes in Computer Science, 2269: 205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690205.
pdf.

Faltin:2002:SCI

[180] Nils Faltin. Structure and constraints in interactive exploratory algorithm learning. Lecture Notes in Computer Science, 2269:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690213.
pdf.

Hundhausen:2002:LSC

[181] Christopher Hundhausen and Sarah Douglas. A language and system for constructing and presenting low fidelity algorithm visualizations. Lecture Notes in Computer Science, 2269:227– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690227.
pdf.

Crescenzi:2002:TTN

[182] Pilu Crescenzi and Gaia Innocenti. Towards a taxonomy of network protocol visualization tools. Lecture Notes in Computer Science, 2269:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690241.
pdf.

Korhonen:2002:UAM

[183] Ari Korhonen, Erkki Sutinen, and Jorma Tarhio. Understanding algorithms by means of visualized path testing. Lecture Notes in Computer Science, 2269:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2269/22690256.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2269/22690256.pdf.

Ross:2002:HAA

[184] Rockford J. Ross and Michael T. Grinder. Hypertextbooks: Animated, active learning, comprehensive teaching and learning resources for the Web. Lecture Notes in Computer Science, 2269:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690269.
pdf.

Mutzel:2002:GSV

[185] Petra Mutzel and Peter Eades. Graphs in software visualization — introduction. Lecture Notes in Computer Science, 2269:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690285.
pdf.

Eichelberger:2002:VJP

[186] H. Eichelberger and J. Wolff von Gudenberg. On the visualization of Java programs. Lecture Notes in Computer Science, 2269:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690295.
pdf.

Gutwenger:2002:GDA

[187] Carsten Gutwenger, Michael Jünger, Gunnar W. Klau, Sebastian Leipert, and Petra Mutzel. Graph drawing algorithm engineering with AGD. Lecture Notes in Computer Science, 2269: 307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690307.
pdf.

Winter:2002:OGG

[188] Andreas Winter, Bernt Kullbach, and Volker Riediger. An overview of the GXL graph exchange language. Lecture Notes in Computer Science, 2269: 324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690324.
pdf.

Evstiougov-Babaev:2002:CGC

[189] Alexander A. Evstiougov-Babaev. Call graph and control flow graph visualization for developers of embedded applications. Lecture Notes in Computer Science, 2269:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690337.
pdf.

Diehl:2002:FPI

[190] Stephan Diehl. Future perspectives — introduction. Lecture Notes in Computer Science, 2269:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690347.
pdf.

Baloian:2002:VME

[191] Nelson Baloian and Wolfram Luther. Visualization for the mind’s eye. Lecture Notes in Computer Science, 2269: 354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690354.
pdf.

Hopkins:2002:RFP

[192] John F. Hopkins and Paul A. Fishwick. The rube framework for personalized 3-D software visualization. Lecture Notes in Computer Science, 2269: 368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690368.
pdf.

Wilhelm:2002:AEV

[193] Reinhard Wilhelm, Tomasz Müldner, and Raimund Seidel. Algorithm explanation: Visualizing abstract states and invariants. Lecture Notes in Computer Science, 2269:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690381.
pdf.

Lengen:2002:VDD

[194] Rolf Hendrik van Lengen and Jan-Thies Bähr. Visualisation and debugging of decentralised information ecosystems. Lecture Notes in Computer Science, 2269:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2269/22690395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2269/22690395.
pdf.

Anonymous:2002:AIf

[195] Anonymous. Author index. Lecture Notes in Computer Science, 2269:405– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2269/2269auth.pdf.

Pflanz:2002:I

[196] Matthias Pflanz. 1. introduction. Lecture Notes in Computer Science, 2270: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2270/22700001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2270/22700001.
pdf.

Pflanz:2002:FMF

[197] Matthias Pflanz. 2. fault models and fault-behavior of processor structures. Lecture Notes in Computer Science, 2270:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2270/22700013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2270/22700013.
pdf.

Pflanz:2002:LCTa

[198] Matthias Pflanz. 3. on-line check technology for processor components. Lecture Notes in Computer Science, 2270: 27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2270/22700027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2270/22700027.
pdf.

Pflanz:2002:LCTb

[199] Matthias Pflanz. 4. on-line check technology for processor control signals. Lecture Notes in Computer Science, 2270:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2270/22700055.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2270/22700055.
                                        pdf.

Pflanz:2002:FPR

[200] Matthias Pflanz. 5. fast processor recover techniques with micro rollback. Lecture Notes in Computer Science, 2270:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2270/22700065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2270/22700065.
                                        pdf.

Pflanz:2002:CO

[201] Matthias Pflanz. 6. conclusion and outlook. Lecture Notes in Computer Science, 2270:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2270/22700081.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2270/22700081.pdf.

Pflanz:2002:ADP

[202] Matthias Pflanz. Appendix — demonstration processors. Lecture Notes in Computer Science, 2270:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2270/22700087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2270/22700087.
                                        pdf.

Anonymous:2002:ASI

[203] Anonymous. Abbreviations, symbols and identifiers. Lecture Notes in Computer Science, 2270:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2270/22700117.pdf.

Anonymous:2002:LF

[204] Anonymous. List of figures. Lecture Notes in Computer Science, 2270:119– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2270/22700119.pdf.

Anonymous:2002:LT

[205] Anonymous. List of tables. Lecture Notes in Computer Science, 2270:121– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2270/22700121.pdf.

Anonymous:2002:Ra

[206] Anonymous. References. Lecture Notes in Computer Science, 2270:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2270/22700123.pdf.

Aczel:2002:CPD

[207] Peter Aczel and Nicola Gambino. Collection principles in dependent type theory. Lecture Notes in Computer Science, 2277:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2277/22770001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2277/22770001.pdf.

Berghofer:2002:EHO

[208] Stefan Berghofer and Tobias Nipkow. Executing higher order logic. Lecture Notes in Computer Science, 2277:24– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770024.
pdf.

Ciaffaglione:2002:TCR

[209] Alberto Ciaffaglione and Pietro Di Gianantonio. A tour with constructive real numbers. Lecture Notes in Computer Science, 2277:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770041.
pdf.

Coquand:2002:ITT

[210] Thierry Coquand and Makoto Takeyama. An implementation of type:type. Lecture Notes in Computer Science, 2277: 53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2277/22770053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770053.
pdf.

Fairtlough:2002:LCC

[211] Matt Fairtlough and Michael Mendler. On the logical content of computational type theory: A solution to curry’s problem. Lecture Notes in Computer Science, 2277:63–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770063.
pdf.

Geuvers:2002:CRC

[212] Herman Geuvers and Milad Niqui. Constructive reals in Coq: Axioms and categoricity. Lecture Notes in Computer Science, 2277:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770079.
pdf.

Geuvers:2002:CPF

[213] Herman Geuvers, Freek Wiedijk, and Jan Zwanenburg. A constructive proof of the fundamental theorem of algebra without using the rationals. Lecture Notes in Computer Science, 2277:96– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2277/22770096.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2277/22770096.
                                        pdf.

Goguen:2002:KSM

[214] Healfdene Goguen. A Kripke-style model for the admissibility of structural rules (extended abstract). Lecture Notes in Computer Science, 2277: 112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2277/22770112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2277/22770112.
                                        pdf.

Hayashi:2002:TLC

[215] Susumu Hayashi and Masahiro Nakata. Towards limit computable mathematics. Lecture Notes in Computer Science, 2277:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2277/22770125.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2277/22770125.pdf.

Johannisson:2002:FHP

[216] Kristofer Johannisson. Formalizing the halting problem in a constructive type theory. Lecture Notes in Computer Science, 2277:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2277/22770145.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2277/22770145.
pdf.

Longo:2002:PSF

[217] Giuseppe Longo. On the proofs of some formally unprovable propositions and prototype proofs in type theory. Lecture Notes in Computer Science, 2277: 160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770160.
pdf.

Magaud:2002:CDS

[218] Nicolas Magaud and Yves Bertot. Changing data structures in type theory: a study of natural numbers. Lecture Notes in Computer Science, 2277: 181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2277/22770181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2277/22770181.
pdf.

McBride:2002:EM

[219] Conor McBride. Elimination with a motive. Lecture Notes in Computer Science, 2277:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2277/22770197.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2277/22770197.pdf.

Pons:2002:GTT

[220] Olivier Pons. Generalization in type theory based proof assistants. Lecture Notes in Computer Science, 2277:217– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2277/22770217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2277/22770217.
                                        pdf.

Seisenberger:2002:IVN

[221] Monika Seisenberger. An inductive version of Nash-Williams’ minimalbad-sequence argument for Higman’s lemma. Lecture Notes in Computer Science, 2277:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2277/22770233.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2277/22770233.pdf.

Anonymous:2002:AIg

[222] Anonymous. Author index. Lecture Notes in Computer Science, 2277:243– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2277/22770243.pdf.

Martin:2002:PHI

[223] Peter Martin. A pipelined hardware implementation of genetic programming using FPGAs and handel-C. Lecture Notes in Computer Science, 2278: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780001.
pdf.

Yu:2002:FNH

[224] Tina Yu and Julian Miller. Finding needles in haystacks is not hard with neutrality. Lecture Notes in Computer Science, 2278:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780013.
pdf.

Streeter:2002:RDP

[225] Matthew J. Streeter, Martin A. Keane, and John R. Koza. Routine duplication of post-2000 patented inventions by means of genetic programming. Lecture Notes in Computer Science, 2278: 26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780026.
pdf.

Brameier:2002:ECD

[226] Markus Brameier and Wolfgang Banzhaf. Explicit control of diversity and effective variation distance in linear genetic programming. Lecture Notes in Computer Science, 2278:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780037.
                                        pdf.

Ferreira:2002:DBF

[227] Cândida Ferreira. Discovery of the Boolean functions to the best densityclassification rules using gene expression programming. Lecture Notes in Computer Science, 2278:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780050.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780050.
                                        pdf.

Langdon:2002:CDT

[228] William B. Langdon, S. J. Barrett, and B. F. Buxton. Combining decision trees and neural networks for drug discovery. Lecture Notes in Computer Science, 2278:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780060.
                                        pdf.

Eggermont:2002:EFD

[229] Jeroen Eggermont. Evolving fuzzy decision trees with genetic programming and clustering. Lecture Notes in Computer Science, 2278:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780071.
pdf.

Kantschik:2002:LGG

[230] Wolfgang Kantschik and Wolfgang Banzhaf. Linear-graph GP — A new GP structure. Lecture Notes in Computer Science, 2278:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780083.
pdf.

Weinert:2002:PSR

[231] Klaus Weinert, Tobias Surmann, and Jörn Mehnen. Parallel surface reconstruction. Lecture Notes in Computer Science, 2278:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780093.
pdf.

Brabazon:2002:ECM

[232] Anthony Brabazon, Michael O’Neill, Conor Ryan, and Robin Matthews. Evolving classifiers to model the relationship between strategy and corporate performance using grammatical evolution. Lecture Notes in Computer Science, 2278:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780103.
                                        pdf.

Weinert:2002:NVS

[233] Klaus Weinert and Marc Stautner. A new view on symbolic regression. Lecture Notes in Computer Science, 2278: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780113.
                                        pdf.

Keijzer:2002:GER

[234] Maarten Keijzer, Michael O’Neill, Conor Ryan, and Mike Cattolico. Grammatical evolution rules: The mod and the bucket rule. Lecture Notes in Computer Science, 2278:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780123.
                                        pdf.

Ryan:2002:NCN

[235] Conor Ryan, Atif Azad, Alan Sheahan, and Michael O’Neill. No coercion and no prohibition, a position independent encoding scheme for evolutionary algorithms — the chorus system. Lecture Notes in Computer Science, 2278:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780131.
pdf.

Soule:2002:ECG

[236] Terence Soule. Exons and code growth in genetic programming. Lecture Notes in Computer Science, 2278:142– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780142.
pdf.

VanBelle:2002:USM

[237] Terry Van Belle and David H. Ackley. Uniform subtree mutation. Lecture Notes in Computer Science, 2278: 152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780152.
pdf.

Ekart:2002:MDG

[238] Anikó Ekárt and Sandor Z. Németh. Maintaining the diversity of genetic programs. Lecture Notes in Computer Science, 2278:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780162.
                                        pdf.

Imamura:2002:VGP

[239] Kosuke Imamura, Robert B. Heckendorn, Terence Soule, and James A. Foster. N-version genetic programming via fault masking. Lecture Notes in Computer Science, 2278:172– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780172.
                                        pdf.

Christensen:2002:AKC

[240] Steffen Christensen and Franz Oppacher. An analysis of Koza’s computational effort statistic for genetic programming. Lecture Notes in Computer Science, 2278:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780182.
                                        pdf.

Werner:2002:GCA

[241] James Cunha Werner and Terence C. Fogarty. Genetic control applied to asset managements. Lecture Notes in Computer Science, 2278:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780192.
pdf.

Baber:2002:EAA

[242] Vinaysheel Baber, Rema Ananthanarayanan, and Krishna Kummamuru. Evolutionary algorithm approach to bilateral negotiations. Lecture Notes in Computer Science, 2278:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780202.
pdf.

Poli:2002:ADL

[243] Riccardo Poli, Jonathan E. Rowe, Christopher R. Stephens, and Alden H. Wright. Allele diffusion in linear genetic programming and variablelength genetic algorithms with subtree crossover. Lecture Notes in Computer Science, 2278:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780212.
pdf.

Hoai:2002:SER

[244] Nguyen Xuan Hoai, R. I. McKay, and D. Essam. Some experimental results with tree adjunct grammar guided genetic programming. Lecture Notes in Computer Science, 2278:228– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780228.
                                        pdf.

Burke:2002:PCG

[245] Edmund Burke, Steven Gustafson, and Graham Kendall. A puzzle to challenge genetic programming. Lecture Notes in Computer Science, 2278:238– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780238.
                                        pdf.

Ibarra:2002:TES

[246] Aitor Ibarra, J. Lanchares, J. M. Mendias, J. I. Hidalgo, and R. Hermida. Transformation of equational specification by means of genetic programming. Lecture Notes in Computer Science, 2278:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780248.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2278/22780248.
pdf.

Busch:2002:AGC

[247] Jens Busch, Jens Ziegler, Christian Aue, Andree Ross, Daniel Sawitzki, and Wolfgang Banzhaf. Automatic generation of control programs for walking robots using genetic programming. Lecture Notes in Computer Science, 2278:258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780258.
pdf.

OSullivan:2002:IUD

[248] John O’Sullivan and Conor Ryan. An investigation into the use of different search strategies with grammatical evolution. Lecture Notes in Computer Science, 2278:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780268.
pdf.

Ryan:2002:GAU

[249] Conor Ryan, Miguel Nicolau, and Michael O’Neill. Genetic algorithms using grammatical evolution. Lecture Notes in Computer Science, 2278:278– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780278.
                                        pdf.

Kuhling:2002:BFA

[250] Felix Kühling, Krister Wolff, and Peter Nordin. A brute-force approach to automatic induction of machine code on CISC architectures. Lecture Notes in Computer Science, 2278:288– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780288.
                                        pdf.

Johnson:2002:DGP

[251] Colin G. Johnson. Deriving genetic programming fitness properties by static analysis. Lecture Notes in Computer Science, 2278:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2278/22780298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2278/22780298.
                                        pdf.

Golubski:2002:NRF

[252] Wolfgang Golubski. New results on fuzzy regression by using genetic programming. Lecture Notes in Computer Science, 2278:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2278/22780308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780308.
pdf.

Ebner:2002:CPA

[253] Marc Ebner, Adrian Grigore, Alexander Heffner, and Jürgen Albert. Coevolution produces an arms race among virtual plants. Lecture Notes in Computer Science, 2278:316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780316.
pdf.

Fernandez:2002:CSA

[254] Francisco Fernández, G. Galeano, and J. A. Gómez. Comparing synchronous and asynchronous parallel and distributed genetic programming models. Lecture Notes in Computer Science, 2278:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2278/22780326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2278/22780326.
pdf.

Anonymous:2002:AIh

[255] Anonymous. Author index. Lecture Notes in Computer Science, 2278:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2278/2278auth.pdf.

Cowling:2002:HTR

[256] Peter Cowling, Graham Kendall, and Eric Soubeiga. Hyperheuristics: a tool for rapid prototyping in scheduling and optimisation. Lecture Notes in Computer Science, 2279:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790001.
                                        pdf.

Doerner:2002:SVR

[257] Karl Doerner, Manfred Gronalt, Richard F. Hartl, Marc Reimann, Christine Strauss, and Michael Stummer. SavingsAnts for the vehicle routing problem. Lecture Notes in Computer Science, 2279:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790011.
                                        pdf.

Dorigo:2002:UAP

[258] Marco Dorigo, Mark Zlochin, Nicolas Meuleau, and Mauro Birattari. Updating ACO pheromones using stochastic gradient ascent and crossentropy methods. Lecture Notes in Computer Science, 2279:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790021.
pdf.

Eremeev:2002:NPE

[259] Anton Eremeev and Colin R. Reeves. Non-parametric estimation of properties of combinatorial landscapes. Lecture Notes in Computer Science, 2279: 31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790031.
pdf.

Esquivel:2002:PEA

[260] Susana Esquivel, Claudia Gatica, and Raúl Gallard. Performance of evolutionary approaches for parallel task scheduling under different representations. Lecture Notes in Computer Science, 2279:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790041.
pdf.

Esquivel:2002:PCA

[261] Susana Esquivel, Guillermo Leguizamón, Federico Zuppa, and Raúl Gallard. A performance comparison of alternative heuristics for the flow shop scheduling problem. Lecture Notes in Computer Science, 2279:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790051.
                                        pdf.

Finger:2002:EFD

[262] Markus Finger, Thomas Stützle, and Helena Lourenço. Exploiting fitness distance correlation of set covering problems. Lecture Notes in Computer Science, 2279:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790061.
                                        pdf.

Guntsch:2002:PBA

[263] Michael Guntsch and Martin Middendorf. A population based approach for ACO. Lecture Notes in Computer Science, 2279:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790072.
                                        pdf.

Hemert:2002:CCM

[264] Jano I. van Hemert. Comparing classical methods for solving binary constraint satisfaction problems with state of the art evolutionary computation. Lecture Notes in Computer Science, 2279:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790082.
pdf.

Johnston:2002:AGA

[265] Roy L. Johnston, Thomas V. Mortimer-Jones, Christopher Roberts, Sarah Darby, and Frederick R. Manby. Application of genetic algorithms in nanoscience: Cluster geometry optimization. Lecture Notes in Computer Science, 2279:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790092.
pdf.

Kersting:2002:MAV

[266] Sandor Kersting, Günther R. Raidl, and Ivana Ljubic. A memetic algorithm for vertex-biconnectivity augmentation. Lecture Notes in Computer Science, 2279:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2279/22790102.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2279/22790102.pdf.

Marchiori:2002:GIM

[267] Elena Marchiori. Genetic, iterated and multistart local search for the maximum clique problem. Lecture Notes in Computer Science, 2279:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790112.
                                        pdf.

Paquete:2002:EII

[268] Luis Paquete and Thomas Stützle. An experimental investigation of iterated local search for coloring graphs. Lecture Notes in Computer Science, 2279: 122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790122.
                                        pdf.

Puchta:2002:SCS

[269] Markus Puchta and Jens Gottlieb. Solving car sequencing problems by local optimization. Lecture Notes in Computer Science, 2279:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790132.
                                        pdf.

Schindler:2002:ESN

[270] Barbara Schindler, Franz Rothlauf, and Hans-Josef Pesch. Evolution strategies, network random keys, and the one-max tree problem. Lecture Notes in Computer Science, 2279:143– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790143.
pdf.

Sofge:2002:ECA

[271] Donald Sofge, Alan Schultz, and Kenneth De Jong. Evolutionary computational approaches to solving the multiple traveling salesman problem using a neighborhood attractor schema. Lecture Notes in Computer Science, 2279: 153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790153.
pdf.

Solnon:2002:BAP

[272] Christine Solnon. Boosting ACO with a preprocessing step. Lecture Notes in Computer Science, 2279:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790163.
pdf.

Torres-Velazquez:2002:MAG

[273] Rodolfo Torres-Velázquez and Vladimir Estivill-Castro. A memetic algorithm guided by quicksort for the error-correcting graph isomorphism problem. Lecture Notes in Computer Science, 2279:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790173.
                                        pdf.

Corno:2002:ETM

[274] Fulvio Corno, Matteo Sonza Reorda, and Giovanni Squillero. Evolutionary techniques for minimizing test signals application time. Lecture Notes in Computer Science, 2279:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790183.
                                        pdf.

Dorado:2002:PMF

[275] Julian Dorado, Juan R. Rabuñal, Jerónimo Puertas, Antonino Santos, and Daniel Rivero. Prediction and modelling of the flow of a typical urban basin through genetic programming. Lecture Notes in Computer Science, 2279:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2279/22790190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790190.
pdf.

Fonlupt:2002:UEE

[276] Cyril Fonlupt, Sébastien Cahon, Denis Robilliard, El-Ghazali Talbi, and Ludovic Duponchel. Using EAs for error prediction in near infrared spectroscopy. Lecture Notes in Computer Science, 2279:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790202.
pdf.

Howard:2002:PJT

[277] Daniel Howard and Simon C. Roberts. The prediction of journey times on motorways using genetic programming. Lecture Notes in Computer Science, 2279:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2279/22790210.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2279/22790210.
pdf.

Howard:2002:BDC

[278] Daniel Howard, Simon C. Roberts, and Conor Ryan. The boru data crawler for object detection tasks in machine vision. Lecture Notes in Computer Science, 2279:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790222.
                                        pdf.

Macias:2002:SPR

[279] Demetrio Macías, Gustavo Olague, and Eugenio R. Méndez. Surface profile reconstruction from scattered intensity data using evolutionary strategies. Lecture Notes in Computer Science, 2279: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790233.
                                        pdf.

Roberts:2002:DIM

[280] Simon C. Roberts and Daniel Howard. Detection of incidents on motorways in low flow high speed conditions by genetic programming. Lecture Notes in Computer Science, 2279:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790245.
                                        pdf.

Sekanina:2002:IFD

[281] Lukás Sekanina. Image filter design with evolvable hardware. Lecture Notes in Computer Science, 2279:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790255.
pdf.

Torresen:2002:DFF

[282] Jim Torresen. A dynamic fitness function applied to improve the generalisation when evolving a signal processing hardware architecture. Lecture Notes in Computer Science, 2279:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790267.
pdf.

Vanyi:2002:ECF

[283] Róbert Ványi. Efficiently computable fitness functions for binary image evolution. Lecture Notes in Computer Science, 2279:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790280.
pdf.

Whitehead:2002:EBA

[284] Anthony Whitehead and Gerhard Roth. Evolutionary based autocalibration from the fundamental matrix. Lecture Notes in Computer Science, 2279: 292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790292.
                                        pdf.

Fan:2002:MIR

[285] Yong Fan, Tianzi Jiang, and David J. Evans. Medical image registration using parallel genetic algorithms. Lecture Notes in Computer Science, 2279: 304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790304.
                                        pdf.

Love:2002:DMA

[286] Michael Løve, Kim Riis Sørensen, Jesper Larsen, and Jens Clausen. Disruption management for an airline — rescheduling of aircraft. Lecture Notes in Computer Science, 2279:315– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2279/22790315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2279/22790315.
                                        pdf.

Merkle:2002:ACO

[287] Daniel Merkle and Martin Middendorf. Ant colony optimization with the relative pheromone evaluation method. Lecture Notes in Computer Science, 2279:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790325.
pdf.

Urquhart:2002:ISB

[288] Neil Urquhart, Peter Ross, Ben Paechter, and Kenneth Chisholm. Improving street based routing using building block mutations. Lecture Notes in Computer Science, 2279:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2279/22790334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2279/22790334.
pdf.

Anonymous:2002:AIi

[289] Anonymous. Author index. Lecture Notes in Computer Science, 2279:343– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2279/2279auth.pdf.

Lowry:2002:SCA

[290] Michael R. Lowry. Software construction and analysis tools for future space missions. Lecture Notes in Computer Science, 2280:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2280/22800001.
                                        pdf.

Jackson:2002:ANT

[291] Daniel Jackson. Alloy: a new technology for software modelling. Lecture Notes in Computer Science, 2280:20– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800020.
                                        pdf.

Braberman:2002:IVT

[292] Víctor Braberman, Diego Garbervetsky, and Alfredo Olivero. Improving the verification of timed systems using influence information. Lecture Notes in Computer Science, 2280:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800021.
                                        pdf.

Ouaknine:2002:DFA

[293] Joël Ouaknine. Digitisation and full abstraction for dense-time model checking. Lecture Notes in Computer Science, 2280:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2280/22800037.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2280/22800037.pdf.

Kwiatkowska:2002:PSM

[294] Marta Kwiatkowska, Gethin Norman, and David Parker. Probabilistic symbolic model checking with PRISM: A hybrid approach. Lecture Notes in Computer Science, 2280:52– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800052.
pdf.

Fersman:2002:TAA

[295] Elena Fersman, Paul Pettersson, and Wang Yi. Timed automata with asynchronous processes: Schedulability and decidability. Lecture Notes in Computer Science, 2280:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800067.
pdf.

Cha:2002:VTC

[296] Hojung Cha and Rhan Ha. Validating timing constraints of dependent jobs with variable execution times in distributed real-time systems. Lecture Notes in Computer Science, 2280:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800083.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2280/22800083.
                                        pdf.

Subramani:2002:AZC

[297] K. Subramani. An analysis of zeroclairvoyant scheduling. Lecture Notes in Computer Science, 2280:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800098.
                                        pdf.

Abdeddaim:2002:PJS

[298] Yasmina Abdeddaïm and Oded Maler. Preemptive job-shop scheduling using stopwatch automata. Lecture Notes in Computer Science, 2280:113– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800113.
                                        pdf.

Wedde:2002:EMI

[299] Horst F. Wedde and Arnim Wedig. Explicit modeling of influences, and of their absence, in distributed systems. Lecture Notes in Computer Science, 2280:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2280/22800127.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800127.
                                        pdf.

Backhouse:2002:FSA

[300] Kevin Backhouse. A functional semantics of attribute grammars. Lecture Notes in Computer Science, 2280:142– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800142.
pdf.

Ball:2002:RCA

[301] Thomas Ball, Andreas Podelski, and Sriram K. Rajamani. Relative completeness of abstraction refinement for software model checking. Lecture Notes in Computer Science, 2280:158– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800158.
pdf.

Delzanno:2002:TAV

[302] Giorgio Delzanno, Jean-François Raskin, and Laurent Van Begin. Towards the automated verification of multithreaded Java programs. Lecture Notes in Computer Science, 2280:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800173.
pdf.

Bouquet:2002:CBC

[303] Fabrice Bouquet, Bruno Legeard, and Fabien Peureux. CLPS-B — A constraint solver for B. Lecture Notes in Computer Science, 2280:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800188.
                                        pdf.

Kim:2002:FVF

[304] Taeho Kim, David Stringer-Calvert, and Sungdeok Cha. Formal verification of functional properties of an SCR-style software requirements specification using PVS. Lecture Notes in Computer Science, 2280:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800205.
                                        pdf.

Bozzano:2002:BPV

[305] Marco Bozzano and Giorgio Delzanno. Beyond parameterized verification. Lecture Notes in Computer Science, 2280:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2280/22800221.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800221.
                                        pdf.

Basu:2002:RCM

[306] Samik Basu, K. Narayan Kumar, L. Robert Pokorny, and C. R. Ramakrishnan. Resource-constrained model checking of recursive programs. Lecture Notes in Computer Science, 2280: 236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800236.
pdf.

Emerson:2002:MCL

[307] E. Allen Emerson and Vineet Kahlon. Model checking large-scale and parameterized resource allocation systems. Lecture Notes in Computer Science, 2280:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2280/22800251.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2280/22800251.
pdf.

Godefroid:2002:EVL

[308] Patrice Godefroid and Sarfraz Khurshid. Exploring very large state spaces using genetic algorithms. Lecture Notes in Computer Science, 2280:266– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800266.
pdf.

Mateescu:2002:LMC

[309] Radu Mateescu. Local model-checking of modal mu-calculus on acyclic labeled transition systems. Lecture Notes in Computer Science, 2280:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800281.
pdf.

Armoni:2002:FTL

[310] Roy Armoni, Limor Fix, Alon Flaisher, Rob Gerth, Boris Ginsburg, Tomer Kanza, Avner Landver, Sela Mador-Haim, Eli Singerman, Andreas Tiemeyer, Moshe Y. Vardi, and Yael Zbar. The For-Spec temporal logic: a new temporal property-specification language. Lecture Notes in Computer Science, 2280: 296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800296.
pdf.

Jin:2002:FGC

[311] HoonSang Jin, Andreas Kuehlmann, and Fabio Somenzi. Fine-grain conjunction scheduling for symbolic reachability analysis. Lecture Notes in Computer Science, 2280:312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2280/22800312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800312.
pdf.

Hong:2002:TLB

[312] Hyoung Seok Hong, Insup Lee, Oleg Sokolsky, and Hasan Ural. A temporal logic based theory of test coverage and generation. Lecture Notes in Computer Science, 2280:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800327.
pdf.

Havelund:2002:SMS

[313] Klaus Havelund and Grigore Rosu. Synthesizing monitors for safety properties. Lecture Notes in Computer Science, 2280:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2280/22800342.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2280/22800342.pdf.

Groce:2002:AMC

[314] Alex Groce, Doron Peled, and Mihalis Yannakakis. Adaptive model checking. Lecture Notes in Computer Science, 2280:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800357.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800357.
                                        pdf.

Heljanko:2002:PPN

[315] Keijo Heljanko, Victor Khomenko, and Maciej Koutny. Parallelisation of the Petri net unfolding algorithm. Lecture Notes in Computer Science, 2280:371– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800371.
                                        pdf.

Bornot:2002:BBU

[316] Sebastien Bornot, Remi Morin, Peter Niebert, and Sarah Zennou. Black box unfolding with local first search. Lecture Notes in Computer Science, 2280: 386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800386.
                                        pdf.

Bustan:2002:AFS

[317] Doron Bustan and Orna Grumberg. Applicability of fair simulation. Lecture Notes in Computer Science, 2280: 401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800401.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2280/22800401.
pdf.

Gentilini:2002:SCP

[318] Raffaella Gentilini, Carla Piazza, and Alberto Policriti. Simulation as coarsest partition problem. Lecture Notes in Computer Science, 2280:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800415.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800415.
pdf.

Gunter:2002:TDC

[319] Elsa Gunter and Doron Peled. Temporal debugging for concurrent systems. Lecture Notes in Computer Science, 2280:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2280/22800431.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2280/22800431.
pdf.

Jin:2002:FFW

[320] HoonSang Jin, Kavita Ravi, and Fabio Somenzi. Fate and free will in error traces. Lecture Notes in Computer Science, 2280:445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800445.
pdf.

Amnell:2002:TTM

[321] Tobias Amnell, Elena Fersman, Leonid Mokrushin, Paul Pettersson, and Wang Yi. TIMES — A tool for modelling and implementation of embedded systems. Lecture Notes in Computer Science, 2280:460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800460.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800460.
                                        pdf.

Lang:2002:CVU

[322] Frédéric Lang. Compositional verification using SVL scripts. Lecture Notes in Computer Science, 2280:465– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800465.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800465.
                                        pdf.

Clarke:2002:SST

[323] Duncan Clarke, Thierry Jéron, Vlad Rusu, and Elena Zinovieva. STG: a symbolic test generation tool. Lecture Notes in Computer Science, 2280:470– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2280/22800470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2280/22800470.
                                        pdf.

Stehno:2002:RTS

[324] Christian Stehno. Real-time systems design with PEP. Lecture Notes in Computer Science, 2280:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2280/22800476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2280/22800476.
pdf.

Anonymous:2002:AIj

[325] Anonymous. Author index. Lecture Notes in Computer Science, 2280:481– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2280/2280auth.pdf.

Takahashi:2002:SME

[326] Koichi Takahashi and Masami Hagiya. Searching for mutual exclusion algorithms using BDDs. Lecture Notes in Computer Science, 2281:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810001.
pdf.

Ida:2002:RSS

[327] Tetsuo Ida, Mircea Marin, and Taro Suzuki. Reducing search space in solving higher-order equations. Lecture Notes in Computer Science, 2281:19– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810019.
pdf.

Noe:2002:SSD

[328] Keiichi Noé. The structure of scientific discovery: From a philosophical point of view. Lecture Notes in Computer Science, 2281:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810031.
pdf.

Okada:2002:ICI

[329] Mitsuhiro Okada. Ideal concepts, intuitions, and mathematical knowledge acquisitions in husserl and hilbert. Lecture Notes in Computer Science, 2281: 40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810040.
pdf.

Sato:2002:TJD

[330] Masahiko Sato. Theory of judgments and derivations. Lecture Notes in Computer Science, 2281:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2281/22810078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810078.
pdf.

Arimura:2002:EDM

[331] Hiroki Arimura, Hiroshi Sakamoto, and Setsuo Arikawa. Efficient data mining from large text databases. Lecture Notes in Computer Science, 2281:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810123.
pdf.

Kobayashi:2002:CMC

[332] Ikuo Kobayashi, Koichi Furukawa, Tomonobu Ozaki, and Mutsumi Imai. A computational model for children’s language acquisition using inductive logic programming. Lecture Notes in Computer Science, 2281:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810140.
pdf.

Haraguchi:2002:SCS

[333] Makoto Haraguchi and Yoshimitsu Kudoh. Some criterions for selecting the best data abstractions. Lecture Notes in Computer Science, 2281:156– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810156.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810156.
                                        pdf.

Ohsawa:2002:DCU

[334] Yukio Ohsawa. Discovery of chances underlying real data. Lecture Notes in Computer Science, 2281:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810168.
                                        pdf.

Sakama:2002:TII

[335] Chiaki Sakama. Towards the integration of inductive and nonmonotonic logic programming. Lecture Notes in Computer Science, 2281:178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810178.
                                        pdf.

Sato:2002:ELS

[336] Taisuke Sato. EM learning for symbolic-statistical models in statistical abduction. Lecture Notes in Computer Science, 2281:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2281/22810189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810189.
pdf.

Sato:2002:RIL

[337] Masako Sato, Yasuhito Mukouchi, and Mikiharu Terada. Refutable/ inductive learning from neighbor examples and its application to decision trees over patterns. Lecture Notes in Computer Science, 2281:201– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810201.
pdf.

Satoh:2002:CCC

[338] Ken Satoh. Constructing a critical casebase to represent a latticebased relation. Lecture Notes in Computer Science, 2281:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810214.
pdf.

Shinohara:2002:DRM

[339] Takeshi Shinohara and Hiroki Ishizaka. On dimension reduction mappings for approximate retrieval of multidimensional data. Lecture Notes in Computer Science, 2281:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810224.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810224.
                                        pdf.

Tsukimoto:2002:RDF

[340] Hiroshi Tsukimoto, Mitsuru Kakimoto, Chie Morita, and Yoshiaki Kikuchi. Rule discovery from fMRI brain images by logical regression analysis. Lecture Notes in Computer Science, 2281:232– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810232.
                                        pdf.

Yamamoto:2002:THF

[341] Akihiro Yamamoto and Bertram Fronhöfer. A theory of hypothesis finding in clausal logic. Lecture Notes in Computer Science, 2281:246– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810246.
                                        pdf.

Mamitsuka:2002:EDM

[342] Hiroshi Mamitsuka and Naoki Abe. Efficient data mining by active learning. Lecture Notes in Computer Science, 2281:258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2281/22810258.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2281/22810258.
pdf.

Okazaki:2002:DCM

[343] Takumi Okazaki, Kunihiko Sadakane, and Hiroshi Imai. Data compression method combining properties of PPM and CTW. Lecture Notes in Computer Science, 2281:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810268.
pdf.

Tsuchiya:2002:DDP

[344] Masatoshi Tsuchiya, Sadao Kurohashi, and Satoshi Sato. Discovery of definition patterns by compressing dictionary sentences. Lecture Notes in Computer Science, 2281:284– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810284.
pdf.

Maruoka:2002:LAP

[345] Akira Maruoka and Eiji Takimoto. On-line algorithm to predict nearly as well as the best pruning of a decision tree. Lecture Notes in Computer Science, 2281:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810296.
                                        pdf.

Shinohara:2002:FBP

[346] Ayumi Shinohara, Masayuki Takeda, Setsuo Arikawa, Masahiro Hirao, Hiromasa Hoshino, and Shunsuke Inenaga. Finding best patterns practically. Lecture Notes in Computer Science, 2281: 307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810307.
                                        pdf.

Takasu:2002:COS

[347] Atsuhiro Takasu. Classification of object sequences using syntactical structure. Lecture Notes in Computer Science, 2281:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810318.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810318.
                                        pdf.

Takimoto:2002:TDT

[348] Eiji Takimoto and Akira Maruoka. Top-down decision tree boosting and its applications. Lecture Notes in Computer Science, 2281:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810327.
pdf.

Uehara:2002:EPM

[349] Kuniaki Uehara and Mitsuomi Shimada. Extraction of primitive motion and discovery of association rules from human motion data. Lecture Notes in Computer Science, 2281:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810338.
pdf.

Watanabe:2002:AAB

[350] Osamu Watanabe. Algorithmic aspects of boosting. Lecture Notes in Computer Science, 2281:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810349.
pdf.

Nagao:2002:ADG

[351] Hiromichi Nagao, Tomoyuki Higuchi, Toshihiko Iyemori, and Tohru Araki. Automatic detection of geomagnetic jerks by applying a statistical time series model to geomagnetic monthly means. Lecture Notes in Computer Science, 2281:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810360.
                                        pdf.

Ueno:2002:AMM

[352] Genta Ueno, Nagatomo Nakamura, Tomoyuki Higuchi, Takashi Tsuchiya, Shinobu Machida, and Tohru Araki. Application of multivariate Maxwellian mixture model to plasma velocity distribution. Lecture Notes in Computer Science, 2281:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810372.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810372.
                                        pdf.

Hasegawa:2002:ITT

[353] Hiroshi H. Hasegawa, Takashi Washio, and Yukari Ishimiya. Inductive thermodynamics from time series data analysis. Lecture Notes in Computer Science, 2281:384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2281/22810384.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810384.
                                        pdf.

Honda:2002:MTF

[354] Rie Honda, Yuichi Iijima, and Osamu Konishi. Mining of topographic feature from heterogeneous imagery and its application to lunar craters. Lecture Notes in Computer Science, 2281: 395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810395.
pdf.

Ikeda:2002:ANN

[355] Yuji Ikeda and Dariusz Mazurkiewicz. Application of neural network technique to combustion spray dynamics analysis. Lecture Notes in Computer Science, 2281:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810408.
pdf.

Kasahara:2002:CAP

[356] Yoshiya Kasahara, Ryotaro Niitsu, and Toru Sato. Computational analysis of plasma waves and particles in the auroral region observed by scientific satellite. Lecture Notes in Computer Science, 2281:426–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810426.
pdf.

Goto:2002:FMG

[357] Yoshitaka Goto, Yoshiya Kasahara, and Toru Sato. A flexible modeling of global plasma profile deduced from wave data. Lecture Notes in Computer Science, 2281:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810438.
                                        pdf.

Kitagawa:2002:ESH

[358] Genshiro Kitagawa, Tetsuo Takanami, Asako Kuwano, Yoshio Murai, and Hideki Shimamura. Extraction of signal from high dimensional time series: Analysis of ocean bottom seismograph data. Lecture Notes in Computer Science, 2281:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810449.
                                        pdf.

Tamada:2002:FDC

[359] Yoshinori Tamada, Hideo Bannai, Osamu Maruyama, and Satoru Miyano. Foundations of designing computational knowledge discovery processes. Lecture Notes in Computer Science, 2281:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2281/22810459.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2281/22810459.
pdf.

Morishita:2002:COH

[360] Shinichi Morishita. Computing optimal hypotheses efficiently for boosting. Lecture Notes in Computer Science, 2281: 471–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810471.
pdf.

Nakano:2002:DPF

[361] Ryohei Nakano and Kazumi Saito. Discovering polynomials to fit multivariate data having numeric and nominal variables. Lecture Notes in Computer Science, 2281:482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810482.
pdf.

Niijima:2002:FSI

[362] Koichi Niijima and Shigeru Takano. Finding of signal and image by integertype Haar lifting wavelet transform. Lecture Notes in Computer Science, 2281:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2281/22810494.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2281/22810494.
                                        pdf.

Suzuki:2002:PIP

[363] Einoshin Suzuki. In pursuit of interesting patterns with undirected discovery of exception rules. Lecture Notes in Computer Science, 2281:504– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810504.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810504.
                                        pdf.

Takeda:2002:MLT

[364] Masayuki Takeda, Tomoko Fukuda, and Ichiro Nanri. Mining from literary texts: Pattern discovery and similarity computation. Lecture Notes in Computer Science, 2281:518–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810518.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810518.
                                        pdf.

Tanaka:2002:SDM

[365] Takehiko Tanaka. Second difference method reinforced by grouping: a new tool for assistance in assignment of complex molecular spectra. Lecture Notes in Computer Science, 2281:532– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810532.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2281/22810532.
pdf.

Tsumoto:2002:DPN

[366] Shusaku Tsumoto. Discovery of positive and negative knowledge in medical databases using rough sets. Lecture Notes in Computer Science, 2281: 543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810543.
pdf.

Washio:2002:TDF

[367] Takashi Washio and Hiroshi Motoda. Toward the discovery of first principle based scientific law equations. Lecture Notes in Computer Science, 2281:553– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810553.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810553.
pdf.

Hamuro:2002:MLA

[368] Yukinobu Hamuro, Hideki Kawata, Naoki Katoh, and Katsutoshi Yada. A machine learning algorithm for analyzing string patterns helps to discover simple and interpretable business rules from purchase history. Lecture Notes in Computer Science, 2281:565– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810565.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810565.
                                        pdf.

Abe:2002:CIA

[369] Hidenao Abe and Takahira Yamaguchi. Constructing inductive applications by meta-learning with method repositories. Lecture Notes in Computer Science, 2281:576–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810576.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810576.
                                        pdf.

Sakamoto:2002:KDS

[370] Hiroshi Sakamoto, Hiroki Arimura, and Setsuo Arikawa. Knowledge discovery from semistructured texts. Lecture Notes in Computer Science, 2281:586– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810586.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810586.
                                        pdf.

Fukushima:2002:PAC

[371] Masaki Fukushima and Shigeki Goto. Packet analysis in congested networks. Lecture Notes in Computer Science, 2281:600–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2281/22810600.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2281/22810600.
pdf.

Hirokawa:2002:VAW

[372] Sachio Hirokawa and Daisuke Ikeda. Visualization and analysis of Web graphs. Lecture Notes in Computer Science, 2281:616–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2281/22810616.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2281/22810616.pdf.

Kuroda:2002:KDA

[373] Hisayasu Kuroda, Takahiro Katagiri, and Yasumasa Kanada. Knowledge discovery in auto-tuning parallel numerical library. Lecture Notes in Computer Science, 2281:628–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810628.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810628.
pdf.

Kawano:2002:EAA

[374] Hiroyuki Kawano and Minoru Kawahara. Extended association algorithm based on ROC analysis for visual information navigator. Lecture Notes in Computer Science, 2281:640– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2281/22810640.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810640.
                                        pdf.

Hiraishi:2002:WVT

[375] Hironori Hiraishi and Fumio Mizoguchi. WWW visualization tools for discovering interesting Web pages. Lecture Notes in Computer Science, 2281: 650–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810650.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810650.
                                        pdf.

Shibayama:2002:SCV

[376] Etsuya Shibayama, Masashi Toyoda, Jun Yabe, and Shin Takahashi. Scalable and comprehensible visualization for discovery of knowledge from the Internet. Lecture Notes in Computer Science, 2281:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2281/22810661.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2281/22810661.
                                        pdf.

Tanaka:2002:MMR

[377] Yuzuru Tanaka. Meme media for re-editing and redistributing intellectual assets and their application to interactive virtual information materialization. Lecture Notes in Computer Science, 2281:672–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2281/22810672.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2281/22810672.
pdf.

Anonymous:2002:AIk

[378] Anonymous. Author index. Lecture Notes in Computer Science, 2281:683– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2281/2281auth.pdf.

Ursino:2002:I

[379] Domenico Ursino. 1 introduction. Lecture Notes in Computer Science, 2282: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820001.
pdf.

Ursino:2002:ESH

[380] Domenico Ursino. 2 extraction of synonymies, homonymies, and type conflicts. Lecture Notes in Computer Science, 2282:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2282/22820027.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2282/22820027.pdf.

Ursino:2002:EOC

[381] Domenico Ursino. 3 extraction of object cluster similarities. Lecture Notes in Computer Science, 2282:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820063.
                                        pdf.

Ursino:2002:EHO

[382] Domenico Ursino. 4 extraction of hyponymies and overlappings. Lecture Notes in Computer Science, 2282:71– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820071.
                                        pdf.

Ursino:2002:EAB

[383] Domenico Ursino. 5 extraction of assertions between knowledge patterns. Lecture Notes in Computer Science, 2282: 87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820087.
                                        pdf.

Ursino:2002:CDR

[384] Domenico Ursino. 6 construction of a data repository. Lecture Notes in Computer Science, 2282:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820113.
pdf.

Ursino:2002:CCI

[385] Domenico Ursino. 7 construction of a cooperative information system. Lecture Notes in Computer Science, 2282: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820143.
pdf.

Ursino:2002:CDW

[386] Domenico Ursino. 8 construction of a data warehouse. Lecture Notes in Computer Science, 2282:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820161.
pdf.

Ursino:2002:SDK

[387] Domenico Ursino. 9 the system D.I.K.E. Lecture Notes in Computer Science, 2282:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2282/22820175.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2282/22820175.pdf.

Ursino:2002:EDS

[388] Domenico Ursino. 10 experiments on the derivation of similarities and type conflicts. Lecture Notes in Computer Science, 2282:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820189.
                                        pdf.

Ursino:2002:EEH

[389] Domenico Ursino. 11 experiments on the extraction of hyponymies. Lecture Notes in Computer Science, 2282:215– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820215.
                                        pdf.

Ursino:2002:EEA

[390] Domenico Ursino. 12 experiments on the extraction of assertions between knowledge patterns. Lecture Notes in Computer Science, 2282:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820221.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2282/22820221.
pdf.

Ursino:2002:ECD

[391] Domenico Ursino. 13 experiments on the construction of a data repository. Lecture Notes in Computer Science, 2282:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820225.
pdf.

Ursino:2002:UCR

[392] Domenico Ursino. 14 using the CIS relative to ICGO databases. Lecture Notes in Computer Science, 2282:243– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820243.
pdf.

Ursino:2002:LF

[393] Domenico Ursino. 15 A look at the future. Lecture Notes in Computer Science, 2282:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2282/22820255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2282/22820255.
pdf.

Ursino:2002:C

[394] Domenico Ursino. 16 conclusions. Lecture Notes in Computer Science, 2282: 267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820267.
                                        pdf.

Anonymous:2002:Rb

[395] Anonymous. References. Lecture Notes in Computer Science, 2282:269– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2282/22820269.pdf.

Ursino:2002:A

[396] Domenico Ursino. A appendix. Lecture Notes in Computer Science, 2282: 279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2282/22820279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2282/22820279.
                                        pdf.

Anonymous:2002:Ia

[397] Anonymous. Index. Lecture Notes in Computer Science, 2282:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2282/2282ind.pdf.

Nipkow:2002:Ba

[398] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 1. the basics. Lecture Notes in Computer Science, 2283: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830003.
pdf.

Nipkow:2002:FPH

[399] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 2. functional programming in HOL. Lecture Notes in Computer Science, 2283:9– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830009.
pdf.

Nipkow:2002:MFP

[400] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 3. more functional programming. Lecture Notes in Computer Science, 2283:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830027.
pdf.

Nipkow:2002:PT

[401] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 4. presenting theories. Lecture Notes in Computer Science, 2283:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830053.
pdf.

Nipkow:2002:RG

[402] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 5. the rules of the game. Lecture Notes in Computer Science, 2283:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830067.
pdf.

Nipkow:2002:SFR

[403] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 6. sets, functions, and relations. Lecture Notes in Computer Science, 2283:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830105.
pdf.

Nipkow:2002:IDS

[404] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 7. inductively defined sets. Lecture Notes in Computer Science, 2283:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830127.
pdf.

Nipkow:2002:MAT

[405] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 8. more about types. Lecture Notes in Computer Science, 2283:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830149.
pdf.

Nipkow:2002:ASR

[406] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 9. advanced simplification, recursion, and induction. Lecture Notes in Computer Science, 2283:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2283/22830175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2283/22830175.
pdf.

Nipkow:2002:CSV

[407] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. 10. case study: Verifying a security protocol. Lecture Notes in Computer Science, 2283:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2283/22830195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2283/22830195.
                                        pdf.

Nipkow:2002:A

[408] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. A. appendix. Lecture Notes in Computer Science, 2283: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2283/22830207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2283/22830207.
                                        pdf.

Nipkow:2002:Bb

[409] Tobias Nipkow, Lawrence C. Paulson, and Markus Wenzel. Bibliography. Lecture Notes in Computer Science, 2283: 209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2283/22830209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2283/22830209.
                                        pdf.

Anonymous:2002:Ib

[410] Anonymous. Index. Lecture Notes in Computer Science, 2283:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2283/2283indx.pdf.

Ding:2002:HEE

[411] Yan Zong Ding and Michael O. Rabin. Hyper-encryption and everlasting security. Lecture Notes in Computer Science, 2285:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850001.
pdf.

Scheideler:2002:MTC

[412] Christian Scheideler. Models and techniques for communication in dynamic networks. Lecture Notes in Computer Science, 2285:27–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850027.
pdf.

Dowek:2002:WT

[413] Gilles Dowek. What is a theory? Lecture Notes in Computer Science, 2285: 50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850050.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2285/22850050.
                                        pdf.

Fraigniaud:2002:SLB

[414] Pierre Fraigniaud and Cyril Gavoille. A space lower bound for routing in trees. Lecture Notes in Computer Science, 2285:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850065.
                                        pdf.

Korman:2002:LSD

[415] Amos Korman, David Peleg, and Yoav Rodeh. Labeling schemes for dynamic tree networks. Lecture Notes in Computer Science, 2285:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850076.
                                        pdf.

Adler:2002:TBP

[416] Micah Adler and Adi Rosén. Tight bounds for the performance of longestin-system on DAGs. Lecture Notes in Computer Science, 2285:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850088.
                                        pdf.

Jansen:2002:ASS

[417] Klaus Jansen. Approximate strong separation with application in fractional graph coloring and preemptive scheduling. Lecture Notes in Computer Science, 2285:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850100.
pdf.

Doerr:2002:BCE

[418] Benjamin Doerr. Balanced coloring: Equally easy for all numbers of colors? Lecture Notes in Computer Science, 2285:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850112.
pdf.

Kobler:2002:CGI

[419] Johannes Köbler and Jacobo Torán. The complexity of graph isomorphism for colored graphs with color classes of size 2 and 3. Lecture Notes in Computer Science, 2285:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850121.
pdf.

Boros:2002:CGM

[420] E. Boros, V. Gurvich, L. Khachiyan, and K. Makino. On the complexity of generating maximal frequent and minimal infrequent sets. Lecture Notes in Computer Science, 2285:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850133.
                                        pdf.

Elbassioni:2002:DPF

[421] Khaled M. Elbassioni. On dualization in products of forests. Lecture Notes in Computer Science, 2285:142– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850142.
                                        pdf.

Lepere:2002:AAA

[422] R. Lepere and C. Rapine. An asymptotic O(ln ρ/ ln ln ρ)-approximation algorithm for the scheduling problem with duplication on large communication delay graphs. Lecture Notes in Computer Science, 2285:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850154.
                                        pdf.

Bast:2002:STE

[423] Hannah Bast. Scheduling at twilight the easy way. Lecture Notes in Computer Science, 2285:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850166.
pdf.

Darte:2002:CMD

[424] Alain Darte and Guillaume Huard. Complexity of multi-dimensional loop alignment. Lecture Notes in Computer Science, 2285:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2285/22850179.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2285/22850179.pdf.

Hofmeister:2002:PSA

[425] Thomas Hofmeister, Uwe Schöning, Rainer Schuler, and Osamu Watanabe. A probabilistic 3-SAT algorithm further improved. Lecture Notes in Computer Science, 2285:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850192.
pdf.

Lorenz:2002:SSG

[426] U. Lorenz and B. Monien. The secret of selective game tree search, when using random-error evaluations. Lecture Notes in Computer Science, 2285:203– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850203.
                                        pdf.

Pan:2002:RAF

[427] Victor Y. Pan. Randomized acceleration of fundamental matrix computations. Lecture Notes in Computer Science, 2285:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2285/22850215.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2285/22850215.pdf.

Chandran:2002:AAP

[428] L. Sunil Chandran and L. Shankar Ram. Approximations for ATSP with parametrized triangle inequality. Lecture Notes in Computer Science, 2285: 227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850227.
                                        pdf.

Giesen:2002:NDD

[429] Joachim Giesen and Matthias John. A new diagram from disks in the plane. Lecture Notes in Computer Science, 2285:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2285/22850238.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2285/22850238.
pdf.

Langerman:2002:CMD

[430] Stefan Langerman, Pat Morin, and Michael Soss. Computing the maximum detour and spanning ratio of planar paths, trees, and cycles. Lecture Notes in Computer Science, 2285:250– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850250.
pdf.

Fellows:2002:PIC

[431] Michael R. Fellows, Jens Gramm, and Rolf Niedermeier. On the parameterized intractability of C LOS-EST S UBSTRING size and related problems. Lecture Notes in Computer Science, 2285:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850262.
pdf.

Backofen:2002:CPS

[432] Rolf Backofen, N. S. Narayanaswamy, and Firas Swidan. On the complexity of protein similarity search under mRNA structure constraints. Lecture Notes in Computer Science, 2285:274– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850274.
                                        pdf.

Bodirsky:2002:PDC

[433] Manuel Bodirsky and Martin Kutz. Pure dominance constraints. Lecture Notes in Computer Science, 2285:287– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850287.
                                        pdf.

Hoyer:2002:IQC

[434] Peter Høyer and Ronald de Wolf. Improved quantum communication complexity bounds for disjointness and equality. Lecture Notes in Computer Science, 2285:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2285/22850299.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2285/22850299.pdf.

Nishimura:2002:QCS

[435] Harumichi Nishimura. On quantum computation with some restricted amplitudes. Lecture Notes in Computer Science, 2285:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850311.
pdf.

Adcock:2002:QGL

[436] Mark Adcock and Richard Cleve. A quantum Goldreich–Levin theorem with cryptographic applications. Lecture Notes in Computer Science, 2285: 323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850323.
pdf.

Klauck:2002:QAP

[437] Hartmut Klauck. On quantum and approximate privacy. Lecture Notes in Computer Science, 2285:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850335.
pdf.

Graaf:2002:QVY

[438] Mart de Graaf and Ronald de Wolf. On quantum versions of the yao principle. Lecture Notes in Computer Science, 2285:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850347.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850347.
                                        pdf.

Flum:2002:DPC

[439] Jörg Flum and Martin Grohe. Describing parameterized complexity classes. Lecture Notes in Computer Science, 2285:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850359.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850359.
                                        pdf.

Krause:2002:CPB

[440] Matthias Krause. On the computational power of Boolean decision lists. Lecture Notes in Computer Science, 2285:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850372.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850372.
                                        pdf.

Simon:2002:HMM

[441] Hans Ulrich Simon. How many missing answers can be tolerated by query learners? Lecture Notes in Computer Science, 2285:384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850384.
pdf.

Aida:2002:GUP

[442] Shin Aida, Marcel Crasmaru, Kenneth Regan, and Osamu Watanabe. Games with a uniqueness property. Lecture Notes in Computer Science, 2285:396– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850396.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850396.
pdf.

Pavan:2002:BIS

[443] A. Pavan and Alan L. Selman. Bi-immunity separates strong NPcompleteness notions. Lecture Notes in Computer Science, 2285:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850408.
pdf.

Grigoriev:2002:CSA

[444] Dima Grigoriev, Edward A. Hirsch, and Dmitrii V. Pasechnik. Complexity of semi-algebraic proofs. Lecture Notes in Computer Science, 2285:419– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850419.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850419.
                                        pdf.

Woelfel:2002:LBT

[445] Philipp Woelfel. A lower bound technique for restricted branching programs and applications. Lecture Notes in Computer Science, 2285:431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850431.
                                        pdf.

Krokhin:2002:CCI

[446] Andrei Krokhin, Peter Jeavons, and Peter Jonsson. The complexity of constraints on intervals and lengths. Lecture Notes in Computer Science, 2285: 443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850443.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850443.
                                        pdf.

Therien:2002:NUS

[447] Denis Thérien and Thomas Wilke. Nesting until and since in linear temporal logic. Lecture Notes in Computer Science, 2285:455–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850455.
pdf.

Tantau:2002:CVF

[448] Till Tantau. Comparing verboseness for finite automata and Turing machines. Lecture Notes in Computer Science, 2285:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850465.
pdf.

Krob:2002:APT

[449] Daniel Krob, Jean Mairesse, and Ioannis Michos. On the average parallelism in trace monoids. Lecture Notes in Computer Science, 2285:477– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850477.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850477.
pdf.

Kuske:2002:FST

[450] Dietrich Kuske. A further step towards a theory of regular MSC languages. Lecture Notes in Computer Science, 2285:489–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850489.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850489.
                                        pdf.

Diekert:2002:EPT

[451] Volker Diekert and Markus Lohrey. Existential and positive theories of equations in graph products. Lecture Notes in Computer Science, 2285:501– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850501.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850501.
                                        pdf.

Petersen:2002:MPR

[452] Holger Petersen. The membership problem for regular expressions with intersection is complete in LOGCFL. Lecture Notes in Computer Science, 2285:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850513.
                                        pdf.

Morin:2002:RSM

[453] Rémi Morin. Recognizable sets of message sequence charts. Lecture Notes in Computer Science, 2285:523– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2285/22850523.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850523.
pdf.

Srba:2002:SBR

[454] Jiří Srba. Strong bisimilarity and regularity of basic parallel processes is PSPACE-hard. Lecture Notes in Computer Science, 2285:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850535.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850535.
pdf.

Beal:2002:ESR

[455] Marie-Pierre Béal and Dominique Perrin. On the enumerative sequences of regular languages on k symbols. Lecture Notes in Computer Science, 2285: 547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850547.
pdf.

Loding:2002:GTR

[456] Christof Löding. Ground tree rewriting graphs of bounded tree width. Lecture Notes in Computer Science, 2285: 559–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850559.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850559.
                                        pdf.

DSouza:2002:TCS

[457] Deepak D’Souza and P. Madhusudan. Timed control synthesis for external specifications. Lecture Notes in Computer Science, 2285:571–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850571.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850571.
                                        pdf.

Baeten:2002:AGT

[458] J. C. M. Baeten and E. P. de Vink. Axiomatizing GSOS with termination. Lecture Notes in Computer Science, 2285:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2285/22850583.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850583.
                                        pdf.

Blumensath:2002:ATI

[459] Achim Blumensath. Axiomatising tree-interpretable structures. Lecture Notes in Computer Science, 2285:596– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850596.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850596.
                                        pdf.

Kieronski:2002:ECV

[460] Emanuel Kieronski. EXPSPACEcomplete variant of guarded fragment with transitivity. Lecture Notes in Computer Science, 2285:608–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850608.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850608.
pdf.

Demri:2002:PAS

[461] S. Demri, F. Laroussinie, and P. Schnoebelen. A parametric analysis of the state explosion problem in model checking. Lecture Notes in Computer Science, 2285:620–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2285/22850620.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2285/22850620.
pdf.

Frick:2002:GMC

[462] Markus Frick. Generalized modelchecking over locally tree-decomposable classes. Lecture Notes in Computer Science, 2285:632–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2285/22850632.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2285/22850632.pdf.

Grohe:2002:LDT

[463] Martin Grohe and Gyorgy Turán. Learnability and definability in trees and similar structures. Lecture Notes in Computer Science, 2285:645–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2285/22850645.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2285/22850645.
                                        pdf.

Anonymous:2002:AIl

[464] Anonymous. Author index. Lecture Notes in Computer Science, 2285:659– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2285/2285auth.pdf.

Chayes:2002:PTC

[465] Jennifer Chayes. Phase transitions in computer science. Lecture Notes in Computer Science, 2286:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860001.
                                        pdf.

Papadimitriou:2002:IWA

[466] Christos H. Papadimitriou. The Internet, the Web, and algorithms. Lecture Notes in Computer Science, 2286: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860002.
pdf.

Spencer:2002:EM

[467] Joel Spencer. Erdos magic. Lecture Notes in Computer Science, 2286:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860003.
pdf.

Urrutia:2002:OPC

[468] Jorge Urrutia. Open problems in computational geometry. Lecture Notes in Computer Science, 2286:4– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860004.
pdf.

Vazirani:2002:QA

[469] Umesh Vazirani. Quantum algorithms. Lecture Notes in Computer Science, 2286:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2286/22860012.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2286/22860012.
pdf.

Yannakakis:2002:TCF

[470] Mihalis Yannakakis. Testing and checking of finite state systems. Lecture Notes in Computer Science, 2286:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860014.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860014.
                                        pdf.

Luccio:2002:AC

[471] Fabrizio Luccio and Linda Pagli. From algorithms to cryptography. Lecture Notes in Computer Science, 2286:15– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860015.
                                        pdf.

Goubault:2002:DTS

[472] Éric Goubault and Martin Raussen. Dihomotopy as a tool in state space analysis. Lecture Notes in Computer Science, 2286:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860016.
                                        pdf.

Arslan:2002:ALA

[473] Abdullah N. Arslan and Ömer Egecioglu. Algorithms for local alignment with length constraints. Lecture Notes in Computer Science, 2286:38– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860038.
pdf.

Braga:2002:ABS

[474] Marília D. V. Braga and João Meidanis. An algorithm that builds a set of strings given its overlap graph. Lecture Notes in Computer Science, 2286: 52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860052.
pdf.

Frougny:2002:CBT

[475] Christiane Frougny. Conversion between two multiplicatively dependent linear numeration systems. Lecture Notes in Computer Science, 2286:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860064.
pdf.

Lombardy:2002:SHR

[476] Sylvain Lombardy and Jacques Sakarovitch. Star height of reversible languages and universal automata. Lecture Notes in Computer Science, 2286:76– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860076.
                                        pdf.

Straubing:2002:WIB

[477] Howard Straubing and Denis Thérien. Weakly iterated block products of finite monoids. Lecture Notes in Computer Science, 2286:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860091.
                                        pdf.

Vasco:2002:HNP

[478] María Isabel González Vasco, Mats Näslund, and Igor E. Shparlinski. The hidden number problem in extension fields and its applications. Lecture Notes in Computer Science, 2286:105– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860105.
                                        pdf.

Garefalakis:2002:GWP

[479] Theodoulos Garefalakis. The generalized Weil pairing and the discrete logarithm problem on elliptic curves. Lecture Notes in Computer Science, 2286: 118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860118.
pdf.

Canfield:2002:RPN

[480] Rod Canfield, Sylvie Corteel, and Pawel Hitczenko. Random partitions with non negative r^{th} differences. Lecture Notes in Computer Science, 2286: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860131.
pdf.

Bassino:2002:BEC

[481] Frédérique Bassino. Beta-expansions for cubic Pisot numbers. Lecture Notes in Computer Science, 2286:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860141.
pdf.

Bose:2002:FLC

[482] Prosenjit Bose and Qingda Wang. Facility location constrained to a polygonal domain. Lecture Notes in Computer Science, 2286:153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860153.
                                        pdf.

Lefmann:2002:DPT

[483] Hanno Lefmann and Niels Schmitt. A deterministic polynomial time algorithm for heilbronn’s problem in dimension three. Lecture Notes in Computer Science, 2286:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860165.
                                        pdf.

Chavez:2002:MIA

[484] Edgar Chávez and Gonzalo Navarro. A metric index for approximate string matching. Lecture Notes in Computer Science, 2286:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860181.
                                        pdf.

Rytter:2002:MSC

[485] Wojciech Rytter. On maximal suffices and constant-space linear-time versions of KMP algorithm. Lecture Notes in Computer Science, 2286:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860196.
pdf.

Corneil:2002:PBD

[486] Derek G. Corneil, Feodor F. Dragan, and Ekkehard Köhler. On the power of BFS to determine a graphs diameter. Lecture Notes in Computer Science, 2286:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860209.
pdf.

Matamala:2002:PLG

[487] Martín Matamala, Erich Prisner, and Ivan Rapaport. k-pseudosnakes in large grids. Lecture Notes in Computer Science, 2286:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860224.
pdf.

Calamoneri:2002:CMG

[488] Tiziana Calamoneri and Rossella Petreschi. L (2,1)-coloring matrogenic graphs. Lecture Notes in Computer Science, 2286:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860236.
                                        pdf.

Milidiu:2002:PTP

[489] Ruy Luiz Milidiú, Artur Alves Pessoa, and Eduardo Sany Laber. Pipeline transportation of petroleum products with no due dates. Lecture Notes in Computer Science, 2286:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860248.
                                        pdf.

Pontelli:2002:APP

[490] Enrico Pontelli and Desh Ranjan. Ancestor problems on pure pointer machines. Lecture Notes in Computer Science, 2286:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860263.
                                        pdf.

Carmo:2002:SRP

[491] Renato Carmo, Jair Donadelli, Yoshiharu Kohayakawa, and Eduardo Laber. Searching in random partially ordered sets. Lecture Notes in Computer Science, 2286:278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860278.
pdf.

Stevens:2002:PA

[492] Brett Stevens and Eric Mendelsohn. Packing arrays. Lecture Notes in Computer Science, 2286:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860293.
pdf.

Drmota:2002:GSC

[493] Michael Drmota and Wojciech Szpankowski. Generalized Shannon code minimizes the maximal redundancy. Lecture Notes in Computer Science, 2286:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2286/22860306.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2286/22860306.
pdf.

Muthukrishnan:2002:IAS

[494] S. Muthukrishnan and S. Cenk Sahinalp. An improved algorithm for sequence comparison with block reversals. Lecture Notes in Computer Science, 2286:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860319.
                                        pdf.

Genest:2002:PMM

[495] Blaise Genest and Anca Muscholl. Pattern matching and membership for hierarchical message sequence charts. Lecture Notes in Computer Science, 2286:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2286/22860326.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860326.
                                        pdf.

Chen:2002:IEA

[496] Jianer Chen and Iyad A. Kanj. Improved exact algorithms for MAX-SAT. Lecture Notes in Computer Science, 2286:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2286/22860341.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860341.
                                        pdf.

Bakel:2002:CSN

[497] Steffen van Bakel and Mariangiola Dezani-Ciancaglini. Characterising strong normalisation for explicit substitutions. Lecture Notes in Computer Science, 2286:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860356.
pdf.

Bloo:2002:PPT

[498] Roel Bloo, Fairouz Kamareddine, Twan Laan, and Rob Nederpelt. Parameters in pure type systems. Lecture Notes in Computer Science, 2286: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860371.
pdf.

Freivalds:2002:CMI

[499] Rusins Freivalds and Carl H. Smith. Category, measure, inductive inference: a triality theorem and its applications. Lecture Notes in Computer Science, 2286:386–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860386.
pdf.

Herbreteau:2002:VER

[500] Frédéric Herbreteau, Franck Cassez, Alain Finkel, Olivier Roux, and Grégoire Sutre. Verification of embedded reactive fiffo systems. Lecture Notes in Computer Science, 2286:400– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860400.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860400.
                                        pdf.

Hevia:2002:EJV

[501] Alejandro Hevia and Marcos Kiwi. Electronic jury voting protocols. Lecture Notes in Computer Science, 2286: 415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860415.
                                        pdf.

Tornaria:2002:SRM

[502] Gonzalo Tornaría. Square roots modulo p. Lecture Notes in Computer Science, 2286:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2286/22860430.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2286/22860430.pdf.

Konjevod:2002:FMS

[503] Goran Konjevod, Soohyun Oh, and Andréa W. Richa. Finding most sustainable paths in networks with timedependent edge reliabilities. Lecture Notes in Computer Science, 2286:435– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2286/22860435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860435.
pdf.

Dubacq:2002:SCA

[504] Jean-Christophe Dubacq and Véronique Terrier. Signals for cellular automata in dimension 2 or higher. Lecture Notes in Computer Science, 2286:451– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860451.
pdf.

Boldi:2002:HT

[505] Paolo Boldi and Sebastiano Vigna. Holographic trees. Lecture Notes in Computer Science, 2286:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860465.
pdf.

Bose:2002:SRG

[506] Prosenjit Bose, Luc Devroye, William Evans, and David Kirkpatrick. On the spanning ratio of Gabriel graphs and β-skeletons. Lecture Notes in Computer Science, 2286:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2286/22860479.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860479.
                                        pdf.

Bronnimann:2002:PPC

[507] Hervé Brönnimann, John Iacono, Jyrki Katajainen, Pat Morin, Jason Morrison, and Godfried Toussaint. In-place planar convex hull algorithms. Lecture Notes in Computer Science, 2286:494– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860494.
                                        pdf.

Bender:2002:LAP

[508] Michael A. Bender and Martín Farach-Colton. The level ancestor problem simplified. Lecture Notes in Computer Science, 2286:508–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860508.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860508.
                                        pdf.

Bornstein:2002:FM

[509] Claudson F. Bornstein and Santosh Vempala. Flow metrics. Lecture Notes in Computer Science, 2286:516– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860516.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2286/22860516.
pdf.

Straubing:2002:LDR

[510] Howard Straubing. On logical descriptions of regular languages. Lecture Notes in Computer Science, 2286:528– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860528.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860528.
pdf.

Szegedy:2002:CBF

[511] Mario Szegedy and Xiaomin Chen. Computing Boolean functions from multiple faulty copies of input bits. Lecture Notes in Computer Science, 2286: 539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860539.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2286/22860539.
pdf.

Halldorsson:2002:IRS

[512] Magnús Halldórsson, Kazuo Iwama, Shuichi Miyazaki, and Yasufumi Morita. Inapproximability results on stable marriage problems. Lecture Notes in Computer Science, 2286:554– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2286/22860554.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2286/22860554.
                                        pdf.

Shachnai:2002:TBO

[513] Hadas Shachnai and Tami Tamir. Tight bounds for online class-constrained packing. Lecture Notes in Computer Science, 2286:569–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2286/22860569.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2286/22860569.pdf.

Anand:2002:LAE

[514] R. Sai Anand and Thomas Erlebach. On-line algorithms for edge-disjoint paths in trees of rings. Lecture Notes in Computer Science, 2286:584– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860584.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860584.
                                        pdf.

Abello:2002:MQC

[515] James Abello, Mauricio G. C. Resende, and Sandra Sudarsky. Massive quasi-clique detection. Lecture Notes in Computer Science, 2286:598– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2286/22860598.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2286/22860598.
                                        pdf.

Alber:2002:ITD

[516] Jochen Alber and Rolf Niedermeier. Improved tree decomposition based algorithms for domination-like problems. Lecture Notes in Computer Science, 2286:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2286/22860613.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2286/22860613.
pdf.

Anonymous:2002:AIm

[517] Anonymous. Author index. Lecture Notes in Computer Science, 2286:629– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2286/2286indx.pdf.

Schek:2002:HII

[518] Hans-Jörg Schek. Hyperdatabases: Infrastructure for the information space. Lecture Notes in Computer Science, 2287:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2287/22870001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2287/22870001.
pdf.

Horrocks:2002:DRA

[519] Ian Horrocks. DAML+OIL: a reasonable Web ontology language. Lecture Notes in Computer Science, 2287:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870002.
pdf.

Ahola:2002:AIP

[520] Jari Ahola. Ambient intelligence: Plenty of challenges by 2010. Lecture Notes in Computer Science, 2287:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870014.
pdf.

Ortega-Binderberger:2002:AIQ

[521] Michael Ortega-Binderberger, Kaushik Chakrabarti, and Sharad Mehrotra. An approach to integrating query refinement in SQL. Lecture Notes in Computer Science, 2287:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870015.
pdf.

Chomicki:2002:QIP

[522] Jan Chomicki. Querying with intrinsic preferences. Lecture Notes in Computer Science, 2287:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2287/22870034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870034.
pdf.

Wang:2002:RUG

[523] Junhu Wang, Michael Maher, and Rodney Topor. Rewriting unions of general conjunctive queries using views. Lecture Notes in Computer Science, 2287: 52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870052.
pdf.

Wang:2002:PMP

[524] Ke Wang, Senqiang Zhou, and Jiawei Han. Profit mining: From patterns to actions. Lecture Notes in Computer Science, 2287:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870070.
pdf.

Lou:2002:CPT

[525] Wenwu Lou, Guimei Liu, Hongjun Lu, and Qiang Yang. Cut-and-pick transactions for proxy log mining. Lecture Notes in Computer Science, 2287:88– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870088.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870088.
                                        pdf.

Diop:2002:CMC

[526] Cheikh Talibouya Diop, Arnaud Giacometti, Dominique Laurent, and Nicolas Spyratos. Composition of mining contexts for efficient extraction of association rules. Lecture Notes in Computer Science, 2287:106– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870106.
                                        pdf.

Lee:2002:DFD

[527] Mong Li Lee, Tok Wang Ling, and Wai Lup Low. Designing functional dependencies for XML. Lecture Notes in Computer Science, 2287:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870124.
                                        pdf.

Lakshmanan:2002:EMS

[528] Laks V. S. Lakshmanan and Sailaja Parthasarathy. On efficient matching of streaming XML documents and queries. Lecture Notes in Computer Science, 2287:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/

series/0558/bibs/2287/22870142.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2287/22870142.pdf.

Chien:2002:ECQ

[529] Shu-Yao Chien, Vassilis J. Tsotras, Carlo Zaniolo, and Donghui Zhang. Efficient complex query support for multiversion XML documents. Lecture Notes in Computer Science, 2287:161– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870161.
pdf.

Papadias:2002:APM

[530] Dimitris Papadias and Dinos Arkoumanis. Approximate processing of multiway spatial joins in very large databases. Lecture Notes in Computer Science, 2287:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870179.
pdf.

Kang:2002:IVC

[531] Myoung-Ah Kang, Christos Faloutsos, Robert Laurini, and Sylvie Servigne. Indexing values in continuous field databases. Lecture Notes in Computer Science, 2287:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870197.
                                        pdf.

Tok:2002:EAP

[532] Wee Hyong Tok and Stéphane Bressan. Efficient and adaptive processing of multiple continuous queries. Lecture Notes in Computer Science, 2287:215– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870215.
                                        pdf.

Trajcevski:2002:GUM

[533] Goce Trajcevski, Ouri Wolfson, Fengli Zhang, and Sam Chamberlain. The geometry of uncertainty in moving objects databases. Lecture Notes in Computer Science, 2287:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870233.
                                        pdf.

Hadjieleftheriou:2002:EIS

[534] Marios Hadjieleftheriou, George Kollios, Vassilis J. Tsotras, and Dimitrios Gunopulos. Efficient indexing of spatiotemporal objects. Lecture Notes in Computer Science, 2287:251– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870251.
pdf.

Lazaridis:2002:DQM

[535] Iosif Lazaridis, Kriengkrai Porkaew, and Sharad Mehrotra. Dynamic queries over mobile objects. Lecture Notes in Computer Science, 2287:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870269.
pdf.

Casati:2002:SAB

[536] Fabio Casati and Ming-Chien Shan. Semantic analysis of business process executions. Lecture Notes in Computer Science, 2287:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870287.
pdf.

Gardarin:2002:IXD

[537] Georges Gardarin, Antoine Mensch, and Anthony Tomasic. An introduction to the e-XML data integration suite. Lecture Notes in Computer Science, 2287:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870297.
                                        pdf.

Tininini:2002:STI

[538] Leonardo Tininini, Mario Paolucci, Giuseppe Sindoni, and Stefano De Francisci. Spatio-temporal information systems in a statistical context. Lecture Notes in Computer Science, 2287: 307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870307.
                                        pdf.

Engstrom:2002:SAS

[539] Henrik Engström, Sharma Chakravarthy, and Brian Lings. A systematic approach to selecting maintenance policies in a data warehouse environment. Lecture Notes in Computer Science, 2287:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870317.
                                        pdf.

Akinde:2002:EOQ

[540] Michael O. Akinde, Michael H. Böhlen, Theodore Johnson, Laks V. S. Lakshmanan, and Divesh Srivastava. Efficient OLAP query processing in distributed data warehouses. Lecture Notes in Computer Science, 2287:336– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870336.
pdf.

Koeller:2002:IMS

[541] Andreas Koeller and Elke A. Rundensteiner. Incremental maintenance of schema-restructuring views. Lecture Notes in Computer Science, 2287:354– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870354.
pdf.

Hergula:2002:CFW

[542] Klaudia Hergula and Theo Härder. Coupling of FDBS and WfMS for integrating database and application systems: Architecture, complexity, performance. Lecture Notes in Computer Science, 2287:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870372.
pdf.

Stolte:2002:OSD

[543] Etzard Stolte and Gustavo Alonso. Optimizing scientific databases for client side data processing. Lecture Notes in Computer Science, 2287:390– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870390.
                                        pdf.

Wang:2002:SEP

[544] Min Wang, Yuan-Chi Chang, and Sriram Padmanabhan. Supporting efficient parametric search of E-commerce data: a loosely-coupled solution. Lecture Notes in Computer Science, 2287: 409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870409.
                                        pdf.

Melnik:2002:DCA

[545] Sergey Melnik and Hector Garcia-Molina. Divide-and-conquer algorithm for computing set containment joins. Lecture Notes in Computer Science, 2287:427–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870427.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870427.
                                        pdf.

Rantzau:2002:UQR

[546] Ralf Rantzau, Leonard Shapiro, Bernhard Mitschang, and Quan Wang. Universal quantification in relational databases: a classification of data and algorithms. Lecture Notes in Computer Science, 2287:445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870445.
pdf.

Marchi:2002:EAM

[547] Fabien De Marchi, Stéphane Lopes, and Jean-Marc Petit. Efficient algorithms for mining inclusion dependencies. Lecture Notes in Computer Science, 2287:464–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2287/22870464.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2287/22870464.pdf.

Theobald:2002:IBX

[548] Anja Theobald and Gerhard Weikum. The index-based XXL search engine for querying XML data with relevance ranking. Lecture Notes in Computer Science, 2287:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870477.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2287/22870477.
                                        pdf.

Amer-Yahia:2002:TPR

[549] Sihem Amer-Yahia, SungRan Cho, and Divesh Srivastava. Tree pattern relaxation. Lecture Notes in Computer Science, 2287:496–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870496.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870496.
                                        pdf.

Schlieder:2002:SDE

[550] Torsten Schlieder. Schema-driven evaluation of approximate tree-pattern queries. Lecture Notes in Computer Science, 2287:514–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2287/22870514.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2287/22870514.pdf.

Brinkhoff:2002:RST

[551] Thomas Brinkhoff. A robust and self-tuning page-replacement strategy for spatial database systems. Lecture Notes in Computer Science, 2287:533– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870533.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870533.
                                        pdf.

Yang:2002:BBD

[552] Xu Yang and Athman Bouguettaya. Broadcast-based data access in wireless environments. Lecture Notes in Computer Science, 2287:553–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870553.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870553.
pdf.

Yee:2002:BGB

[553] Wai Gen Yee, Shamkant B. Navathe, Edward Omiecinski, and Christopher Jermaine. Bridging the gap between response time and energy-efficiency in broadcast schedule design. Lecture Notes in Computer Science, 2287:572– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870572.
pdf.

Wu:2002:EAS

[554] Yuqing Wu, Jignesh M. Patel, and H. V. Jagadish. Estimating answer sizes for XML queries. Lecture Notes in Computer Science, 2287:590– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870590.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870590.
pdf.

Sun:2002:SES

[555] Chengyu Sun, Divyakant Agrawal, and Amr El Abbadi. Selectivity estimation for spatial joins with geometric selections. Lecture Notes in Computer Science, 2287:609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870609.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870609.
                                        pdf.

Konig:2002:FPD

[556] Arnd Christian König and Gerhard Weikum. A framework for the physical design problem for data synopses. Lecture Notes in Computer Science, 2287: 627–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870627.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870627.
                                        pdf.

Zhang:2002:TAD

[557] Donghui Zhang, Dimitrios Gunopulos, Vassilis J. Tsotras, and Bernhard Seeger. Temporal aggregation over data streams using multiple granularities. Lecture Notes in Computer Science, 2287:646–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870646.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2287/22870646.
pdf.

Schmidt:2002:PFW

[558] Rolfe R. Schmidt and Cyrus Shahabi. ProPolyne: a fast wavelet-based algorithm for progressive evaluation of polynomial range-sum queries. Lecture Notes in Computer Science, 2287:664– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870664.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870664.
pdf.

Tao:2002:APP

[559] Yufei Tao, Dimitris Papadias, and Jun Zhang. Aggregate processing of planar points. Lecture Notes in Computer Science, 2287:682–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870682.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870682.
pdf.

Vingralek:2002:TDS

[560] Radek Vingralek, Umesh Maheshwari, and William Shapiro. TDB: a database system for digital rights management. Lecture Notes in Computer Science, 2287:701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2287/22870701.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2287/22870701.
                                        pdf.

Wienberg:2002:CSE

[561] Axel Wienberg, Matthias Ernst, Andreas Gawecki, Olaf Kummer, Frank Wienberg, and Joachim W. Schmidt. Content schema evolution in the CoreMedia(R) content application platform CAP. Lecture Notes in Computer Science, 2287:712–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870712.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870712.
                                        pdf.

Markowitz:2002:GED

[562] Victor M. Markowitz, I-Min A. Chen, and Anthony Kosky. Gene expression data management: a case study. Lecture Notes in Computer Science, 2287: 722–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870722.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870722.
                                        pdf.

Kraiss:2002:HTR

[563] Achim Kraiss, Frank Schoen, Gerhard Weikum, and Uwe Deppisch. With HEART towards response time guarantees for message-based e-services. Lecture Notes in Computer Science, 2287: 732–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2287/22870732.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870732.
pdf.

Petkovic:2002:CCB

[564] Milan Petkovic and Willem Jonker. Cobra: a content-based video retrieval system. Lecture Notes in Computer Science, 2287:736–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870736.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870736.
pdf.

Qian:2002:NVI

[565] Xufei Qian, Bertram Ludäscher, Maryann E. Martone, and Amarnath Gupta. Navigating virtual information sources with know-ME. Lecture Notes in Computer Science, 2287:739– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870739.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870739.
pdf.

Fankhauser:2002:XBI

[566] Peter Fankhauser, Tobias Groh, and Sven Overhage. XQuery by the book: The IPSI XQuery demonstrator. Lecture Notes in Computer Science, 2287: 742–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870742.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870742.
                                        pdf.

Mahnke:2002:OBS

[567] Wolfgang Mahnke and Norbert Ritter. The ORDB-based SFB-501-reuse-repository. Lecture Notes in Computer Science, 2287:745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870745.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870745.
                                        pdf.

Wiesner:2002:BDM

[568] Christian Wiesner, Peter Winklhofer, and Alfons Kemper. Building dynamic market places using HyperQueries. Lecture Notes in Computer Science, 2287:749–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870749.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870749.
                                        pdf.

Damiani:2002:PXT

[569] Ernesto Damiani, Nico Lavarini, Stefania Marrara, Barbara Oliboni, Daniele Pasini, Letizia Tanca, and Giuseppe Viviani. The A PPRO XML tool demonstration. Lecture Notes in Computer Science, 2287:753– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870753.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870753.
pdf.

Dunemann:2002:DSW

[570] Oliver Dunemann, Ingolf Geist, Roland Jesse, Kai-Uwe Sattler, and Andreas Stephanik. A databasesupported workbench for information fusion: I N F USE. Lecture Notes in Computer Science, 2287:756–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870756.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870756.
pdf.

Fundulaki:2002:ECX

[571] Irini Fundulaki, Bernd Amann, Catriel Beeri, Michel Scholl, and Anne-Marie Vercoustre. ST_{YX}: Connecting the XML Web to the world of semantics. Lecture Notes in Computer Science, 2287:759–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870759.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870759.
pdf.

Amanatidis:2002:UDM

[572] Christos Amanatidis, Maria Halkidi, and Michalis Vazirgiannis. UMiner: a data mining system handling uncertainty and quality. Lecture Notes in Computer Science, 2287:762–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870762.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870762.
                                        pdf.

Jin:2002:MWS

[573] Yuhui Jin, Sichun Xu, Stefan Decker, and Gio Wiederhold. Managing Web sites with OntoWebber. Lecture Notes in Computer Science, 2287:766–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870766.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870766.
                                        pdf.

Wolfson:2002:MDL

[574] Ouri Wolfson, Hu Cao, Hai Lin, Goce Trajcevski, Fengli Zhang, and N. Rishe. Management of dynamic location information in DOMINO. Lecture Notes in Computer Science, 2287: 769–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2287/22870769.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2287/22870769.
                                        pdf.

Haase:2002:SAM

[575] Peter Haase. Situation aware mobile access to digital libraries. Lecture Notes in Computer Science, 2287:772– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2287/22870772.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2287/22870772.
pdf.

Anonymous:2002:AIn

[576] Anonymous. Author index. Lecture Notes in Computer Science, 2287:776– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2287/2287auth.pdf.

Pointcheval:2002:PSP

[577] David Pointcheval. Practical security in public-key cryptography. Lecture Notes in Computer Science, 2288:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880001.
pdf.

Song:2002:NCM

[578] Beomsik Song, Huaxiong Wang, and Jennifer Seberry. A new cryptanalytic method using the distribution characteristics of substitution distances. Lecture Notes in Computer Science, 2288: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2288/22880018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880018.
                                        pdf.

Lee:2002:TDC

[579] Seonhee Lee, Seokhie Hong, Sangjin Lee, Jongin Lim, and Seonhee Yoon. Truncated differential cryptanalysis of camellia. Lecture Notes in Computer Science, 2288:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880032.
                                        pdf.

Cheon:2002:IID

[580] Jung Hee Cheon, MunJu Kim, Kwangjo Kim, Jung-Yeun Lee, and SungWoo Kang. Improved impossible differential cryptanalysis of Rijndael and Crypton. Lecture Notes in Computer Science, 2288:39–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880039.
                                        pdf.

Leveiller:2002:CNF

[581] Sabine Leveiller, Joseph Boutros, Philippe Guillot, and Gilles Zémor. Cryptanalysis of nonlinear filter generators with {0, 1}-metric Viterbi decoding. Lecture Notes in Computer Science, 2288:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2288/22880050.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2288/22880050.pdf.

Buchmann:2002:ICP

[582] Johannes Buchmann, Kouichi Sakurai, and Tsuyoshi Takagi. An IND-CCA2 public-key cryptosystem with fast decryption. Lecture Notes in Computer Science, 2288:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880051.
pdf.

Choi:2002:IPP

[583] Dug-Hwan Choi, Seungbok Choi, and Dongho Won. Improvement of probabilistic public key cryptosystems using discrete logarithm. Lecture Notes in Computer Science, 2288:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880072.
pdf.

Nishioka:2002:DAF

[584] Mototsugu Nishioka, Hisayoshi Satoh, and Kouichi Sakurai. Design and analysis of fast provably secure public-key cryptosystems based on a modular squaring. Lecture Notes in Computer Science, 2288:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880081.
                                        pdf.

Sung:2002:CSA

[585] Jaechul Sung, Sangjin Lee, Jongin Lim, Wonil Lee, and Okyeon Yi. Concrete security analysis of CTR-OFB and CTR-CFB modes of operation. Lecture Notes in Computer Science, 2288: 103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880103.
                                        pdf.

Krugel:2002:DEC

[586] Christopher Krügel, Thomas Toth, and Clemens Kerer. Decentralized event correlation for intrusion detection. Lecture Notes in Computer Science, 2288: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880114.
                                        pdf.

Khu-smith:2002:ESC

[587] Vorapranee Khu-smith and Chris Mitchell. Enhancing the security of cookies. Lecture Notes in Computer Science, 2288:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880132.
pdf.

Choi:2002:NSB

[588] Yang-Seo Choi, Dong il Seo, and Sung-Won Sohn. A new stack buffer overflow hacking defense technique with memory address confirmation. Lecture Notes in Computer Science, 2288:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880146.
pdf.

Kurnio:2002:ERS

[589] Hartono Kurnio, Rei Safavi-Naini, and Huaxiong Wang. Efficient revocation schemes for secure multicast. Lecture Notes in Computer Science, 2288:160– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880160.
pdf.

Cohen:2002:BCC

[590] Gérard Cohen, Simon Litsyn, and Gilles Zémor. Binary codes for collusion-secure fingerprinting. Lecture Notes in Computer Science, 2288:178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880178.
                                        pdf.

Pastuszak:2002:CPO

[591] Jarek Pastuszak, Darek Michałek, and Josef Pieprzyk. Copyright protection of object-oriented software. Lecture Notes in Computer Science, 2288:186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880186.
                                        pdf.

Yoo:2002:LAU

[592] Hyejoung Yoo, Kwangsoo Lee, Sangjin Lee, and Jongin Lim. Off-line authentication using watermarks. Lecture Notes in Computer Science, 2288:200– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880200.
                                        pdf.

Furuya:2002:SAK

[593] Soichi Furuya. Slide attacks with a known-plaintext cryptanalysis. Lecture Notes in Computer Science, 2288:214– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880214.
pdf.

Pieprzyk:2002:CCI

[594] Josef Pieprzyk and Xian-Mo Zhang. Constructions of cheating immune secret sharing. Lecture Notes in Computer Science, 2288:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880226.
pdf.

Galdi:2002:PCS

[595] Clemente Galdi and Pino Persiano. Private computation with shared randomness over broadcast channel. Lecture Notes in Computer Science, 2288:244– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880244.
pdf.

Gonzalez-Deleito:2002:OMP

[596] Nicolás González-Deleito and Olivier Markowitch. An optimistic multiparty fair exchange protocol with reduced trust requirements. Lecture Notes in Computer Science, 2288:258– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880258.
                                        pdf.

Kungpisdan:2002:PRA

[597] Supakorn Kungpisdan and Yongyuth Permpoontanalarp. Practical reasoning about accountability in electronic commerce protocols. Lecture Notes in Computer Science, 2288:268– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880268.
                                        pdf.

Steinfeld:2002:CES

[598] Ron Steinfeld, Laurence Bull, and Yuliang Zheng. Content extraction signatures. Lecture Notes in Computer Science, 2288:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880285.
                                        pdf.

Yum:2002:NSS

[599] Dae Hyun Yum and Pil Joong Lee. New signcryption schemes based on KCDSA. Lecture Notes in Computer Science, 2288:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880305.
pdf.

Kim:2002:EPS

[600] Jinho Kim, Kwangjo Kim, and Chulsoo Lee. An efficient and provably secure threshold blind signature. Lecture Notes in Computer Science, 2288:318– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880318.
pdf.

Kawauchi:2002:MSS

[601] Kei Kawauchi, Hiroshi Minato, Atsuko Miyaji, and Mitsuru Tada. A multisignature scheme with signers’ intentions secure against active attacks. Lecture Notes in Computer Science, 2288: 328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880328.
pdf.

Weimerskirch:2002:DLW

[602] André Weimerskirch and Gilles Thonet. A distributed light-weight authentication model for ad-hoc networks. Lecture Notes in Computer Science, 2288:

341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880341.
                                        pdf.

Cimato:2002:DAP

[603] Stelvio Cimato. Design of an authentication protocol for Gsm Javacards. Lecture Notes in Computer Science, 2288: 355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880355.
                                        pdf.

Au:2002:SAA

[604] Richard Au, Mark Looi, Paul Ashley, and Loo Tang Seet. Secure authorisation agent for cross-domain access control in a mobile computing environment. Lecture Notes in Computer Science, 2288:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880369.
                                        pdf.

Mir:2002:PGF

[605] Joan Mir and Joan Borrell. Protecting general flexible itineraries of mobile agents. Lecture Notes in Computer Science, 2288:382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880382.
pdf.

Yen:2002:RSR

[606] Sung-Ming Yen, Seungjoo Kim, Seongan Lim, and Sangjae Moon. RSA speedup with residue number system immune against hardware fault cryptanalysis. Lecture Notes in Computer Science, 2288:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2288/22880397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2288/22880397.
pdf.

Yen:2002:CAO

[607] Sung-Ming Yen, Seungjoo Kim, Seongan Lim, and Sangjae Moon. A countermeasure against one physical cryptanalysis may benefit another attack. Lecture Notes in Computer Science, 2288:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2288/22880414.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2288/22880414.
pdf.

Okeya:2002:FSM

[608] Katsuyuki Okeya, Kunihiko Miyazaki, and Kouichi Sakurai. A fast scalar multiplication method with randomized projective coordinates on a Montgomery-form elliptic curve secure against side channel attacks. Lecture Notes in Computer Science, 2288:428– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880428.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880428.
                                        pdf.

Itoh:2002:DCB

[609] Kouichi Itoh, Masahiko Takenaka, and Naoya Torii. DPA countermeasure based on the “Masking Method”. Lecture Notes in Computer Science, 2288: 440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2288/22880440.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2288/22880440.
                                        pdf.

Anonymous:2002:AIo

[610] Anonymous. Author index. Lecture Notes in Computer Science, 2288:457– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2288/2288auth.pdf.

Sharp:2002:HES

[611] David Sharp. Hybrid and embedded software technologies for production large-scale systems. Lecture Notes in Computer Science, 2289:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890001.
pdf.

Ascher:2002:NMD

[612] Uri M. Ascher. Numerical methods for differential systems with algebraic equality and inequality constraints. Lecture Notes in Computer Science, 2289:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2289/22890003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2289/22890003.
pdf.

Henzinger:2002:MCM

[613] Thomas A. Henzinger. From models to code: The missing link in embedded software. Lecture Notes in Computer Science, 2289:5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890005.
pdf.

Aaron:2002:HSM

[614] Eric Aaron, Franjo Ivančić, and Dimitris Metaxas. Hybrid system models of navigation strategies for games and animations. Lecture Notes in Computer Science, 2289:7– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890007.
                                        pdf.

Altafini:2002:HCT

[615] Claudio Altafini, Alberto Speranzon, and Karl Henrik Johansson. Hybrid control of a truck and trailer vehicle. Lecture Notes in Computer Science, 2289:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890021.
                                        pdf.

Alur:2002:RAH

[616] Rajeev Alur, Thao Dang, and Franjo Ivančić. Reachability analysis of hybrid systems via predicate abstraction. Lecture Notes in Computer Science, 2289: 35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890035.
                                        pdf.

Asarin:2002:TCP

[617] Eugene Asarin, Gerardo Schneider, and Sergio Yovine. Towards computing phase portraits of polygonal differential inclusions. Lecture Notes in Computer Science, 2289:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890049.
pdf.

Aubin:2002:DQA

[618] Jean-Pierre Aubin and Olivier Dordan. Dynamical qualitative analysis of evolutionary systems. Lecture Notes in Computer Science, 2289:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890062.
pdf.

Balluchi:2002:DOH

[619] Andrea Balluchi, Luca Benvenuti, Maria D. Di Benedetto, and Alberto L. Sangiovanni-Vincentelli. Design of observers for hybrid systems. Lecture Notes in Computer Science, 2289:76– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890076.
pdf.

Bayen:2002:GOU

[620] Alexandre M. Bayen, Eva Crück, and Claire J. Tomlin. Guaranteed overapproximations of unsafe sets for continuous and hybrid systems: Solving the Hamilton–Jacobi equation using viability techniques. Lecture Notes in Computer Science, 2289:90– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890090.
pdf.

Bemporad:2002:OCL

[621] Alberto Bemporad, Francesco Borrelli, and Manfred Morari. On the optimal control law for linear discrete time hybrid systems. Lecture Notes in Computer Science, 2289:105– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890105.
pdf.

Branicky:2002:CFV

[622] Michael S. Branicky and Siddharth R. Chhatpar. A computational framework for the verification and synthesis of force-guided robotic assembly strategies. Lecture Notes in Computer Science, 2289:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890120.
pdf.

Cassez:2002:CCP

[623] Franck Cassez, Thomas A. Henzinger, and Jean-François Raskin. A comparison of control problems for timed and hybrid systems. Lecture Notes in Computer Science, 2289:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890134.
pdf.

Davoren:2002:HCL

[624] J. M. Davoren, T. Moor, and A. Nerode. Hybrid control loops, A/D maps, and dynamic specifications. Lecture Notes in Computer Science, 2289: 149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890149.
pdf.

El-Farra:2002:SFL

[625] Nael H. El-Farra and Panagiotis D. Christofides. Switching and feedback laws for control of constrained switched nonlinear systems. Lecture Notes in Computer Science, 2289:164– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890164.
pdf.

Elia:2002:QST

[626] Nicola Elia and Emilio Frazzoli. Quantized stabilization of two-input linear systems: A lower bound on the minimal quantization density. Lecture Notes in Computer Science, 2289:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890179.
                                        pdf.

Ferrari-Trecate:2002:ADT

[627] Giancarlo Ferrari-Trecate, Francesco A. Cuzzola, and Manfred Morari. Analysis of discrete-time PWA systems with logic states. Lecture Notes in Computer Science, 2289:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890194.
                                        pdf.

Ferrari-Trecate:2002:MCC

[628] Giancarlo Ferrari-Trecate, Eduardo Gallestey, Paolo Letizia, Matteo Spedicato, Manfred Morari, and Marc Antoine. Modeling and control of co-generation power plants: A hybrid system approach. Lecture Notes in Computer Science, 2289:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890209.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2289/22890209.
pdf.

Goldman:2002:EIR

[629] Robert P. Goldman, David J. Musliner, and Michael J. S. Pelican. Exploiting implicit representations in timed automaton verification for controller synthesis. Lecture Notes in Computer Science, 2289:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890225.
pdf.

Hespanha:2002:CRM

[630] João P. Hespanha. Computation of root-mean-square gains of switched linear systems. Lecture Notes in Computer Science, 2289:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890239.
pdf.

Hofbaur:2002:MEP

[631] Michael W. Hofbaur and Brian C. Williams. Mode estimation of probabilistic hybrid systems. Lecture Notes in Computer Science, 2289:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2289/22890253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890253.
                                        pdf.

Hu:2002:SRC

[632] Jianghai Hu and Shankar Sastry. Symmetry reduction of a class of hybrid systems. Lecture Notes in Computer Science, 2289:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890267.
                                        pdf.

Koo:2002:BBH

[633] T. John Koo and Shankar Sastry. Bisimulation based hierarchical system architecture for singleagent multi-modal systems. Lecture Notes in Computer Science, 2289:281– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890281.
                                        pdf.

Kuipers:2002:QMH

[634] Benjamin Kuipers and Subramanian Ramamoorthy. Qualitative modeling and heterogeneous control of global system behavior. Lecture Notes in Computer Science, 2289:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2289/22890294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890294.
pdf.

Narasimhan:2002:AMB

[635] Sriram Narasimhan and Gautam Biswas. An approach to model-based diagnosis of hybrid systems. Lecture Notes in Computer Science, 2289:308– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890308.
pdf.

Neller:2002:IBA

[636] Todd W. Neller. Informationbased alpha-beta search and the homicidal chauffeur. Lecture Notes in Computer Science, 2289:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890323.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890323.
pdf.

Palopoli:2002:SRC

[637] Luigi Palopoli, Claudio Pinello, Alberto Sangiovanni Vincentelli, Laurent Elghaoui, and Antonio Bicchi. Synthesis of robust control systems under resource constraints. Lecture Notes in Computer Science, 2289:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890337.
                                        pdf.

Pancanti:2002:OCQ

[638] Stefania Pancanti, Laura Leonardi, Lucia Pallottino, and Antonio Bicchi. Optimal control of quantized input systems. Lecture Notes in Computer Science, 2289:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890351.
                                        pdf.

Pasternak:2002:RHC

[639] Tal Pasternak. Reconfiguration in hierarchical control of piecewiseaffine systems. Lecture Notes in Computer Science, 2289:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890364.
                                        pdf.

Saint-Pierre:2002:HKC

[640] Patrick Saint-Pierre. Hybrid kernels and capture basins for impulse constrained systems. Lecture Notes in Computer Science, 2289:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890378.
pdf.

Sethian:2002:OUM

[641] James A. Sethian and Alexander Vladimirsky. Ordered upwind methods for hybrid control. Lecture Notes in Computer Science, 2289:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890393.
pdf.

Stauner:2002:DTR

[642] Thomas Stauner. Discrete-time refinement of hybrid automata. Lecture Notes in Computer Science, 2289:407– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2289/22890407.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2289/22890407.
pdf.

Stursberg:2002:CSH

[643] Olaf Stursberg and Sebastian Panek. Control of switched hybrid systems based on disjunctive formulations. Lecture Notes in Computer Science, 2289: 421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2289/22890421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890421.
                                        pdf.

Tabuada:2002:CAH

[644] Paulo Tabuada, George J. Pappas, and Pedro Lima. Composing abstractions of hybrid systems. Lecture Notes in Computer Science, 2289:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890436.
                                        pdf.

Tan:2002:OCH

[645] Xiaobo Tan and John S. Baras. Optimal control of hysteresis in smart actuators: a viscosity solutions approach. Lecture Notes in Computer Science, 2289:451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2289/22890451.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2289/22890451.
                                        pdf.

Tiwari:2002:SAH

[646] Ashish Tiwari and Gaurav Khanna. Series of abstractions for hybrid automata. Lecture Notes in Computer Science, 2289:465–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2289/22890465.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2289/22890465.pdf.

Anonymous:2002:AIp

[647] Anonymous. Author index. Lecture Notes in Computer Science, 2289:479– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2289/2289indx.pdf.

Linden:2002:I

[648] Frank van der Linden. Introduction. Lecture Notes in Computer Science, 2290:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2290/22900001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2290/22900001.
pdf.

Knauber:2002:SRP

[649] Peter Knauber and Steffen Thiel. Session report on product issues in product family engineering. Lecture Notes in Computer Science, 2290:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900003.
pdf.

Bosch:2002:VIS

[650] Jan Bosch, Gert Florijn, Danny Greefhorst, Juha Kuusela, J. Henk Obbink, and Klaus Pohl. Variability issues in software product lines. Lecture Notes in Computer Science, 2290:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900013.
pdf.

Pohl:2002:CVD

[651] Klaus Pohl and Andreas Reuys. Considering variabilities during component selection in product family development. Lecture Notes in Computer Science, 2290:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900022.
pdf.

Schmid:2002:IMP

[652] Klaus Schmid. An initial model of product line economics. Lecture Notes in Computer Science, 2290:38– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900038.
pdf.

Ommering:2002:RPP

[653] Rob van Ommering. Roadmapping a product population architecture. Lecture Notes in Computer Science, 2290:

51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900051.
pdf.

Maccari:2002:AEL

[654] Alessandro Maccari and Claudio Riva. Architectural evolution of legacy product families. Lecture Notes in Computer Science, 2290:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900064.
pdf.

Clements:2002:IPL

[655] Paul C. Clements. On the importance of product line scope. Lecture Notes in Computer Science, 2290:70– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900070.
pdf.

Bayer:2002:SRS

[656] Joachim Bayer and Günter Böckle. Session report for session 2: Process issues. Lecture Notes in Computer Science, 2290:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2290/22900079.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900079.
                                        pdf.

Auerswald:2002:ROP

[657] Marko Auerswald, Martin Herrmann, Stefan Kowalewski, and Vincent Schulte-Coerne. Reliability-oriented product line engineering of embedded systems. Lecture Notes in Computer Science, 2290:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900083.
                                        pdf.

Stoermer:2002:IPL

[658] Christoph Stoermer and Markus Roeddiger. Introducing product lines in small embedded systems. Lecture Notes in Computer Science, 2290:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900101.
                                        pdf.

John:2002:ILD

[659] Isabel John. Integrating legacy documentation assets into a product line. Lecture Notes in Computer Science, 2290:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2290/22900113.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2290/22900113.
pdf.

Thiel:2002:DFA

[660] Steffen Thiel. On the definition of a framework for an architecting process supporting product family development. Lecture Notes in Computer Science, 2290:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900125.
pdf.

Svahnberg:2002:CRP

[661] Mikael Svahnberg and Michael Mattsson. Conditions and restrictions for product line generation migration. Lecture Notes in Computer Science, 2290: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900143.
pdf.

Knauber:2002:QPL

[662] Peter Knauber, Jesus Bermejo, Günter Böckle, Julio Cesar Sampaio do Prado Leite, Frank van der Linden, Linda Northrop, Michael Stark, and David M. Weiss. Quantifying product line benefits. Lecture Notes in Computer Science, 2290: 155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900155.
                                        pdf.

Hallsteinsen:2002:PQS

[663] Svein Hallsteinsen and Len Bass. Platform & quality solutions. Lecture Notes in Computer Science, 2290:164– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900164.
                                        pdf.

Bass:2002:QAD

[664] Len Bass, Mark Klein, and Felix Bachmann. Quality attribute design primitives and the attribute driven design method. Lecture Notes in Computer Science, 2290:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900169.
                                        pdf.

deLange:2002:POT

[665] Fons de Lange. The Philips-Open TV(R) product family architecture for interactive set-top boxes. Lecture Notes in Computer Science, 2290:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900187.
pdf.

Blin:2002:PCB

[666] Marie-José Blin, Françoise Fabret, Olga Kapitskaia, and François Llirbat. ProjectLeader: a constraint-based process support for the distributed design of component-based products. Lecture Notes in Computer Science, 2290:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900207.
pdf.

Linden:2002:PEM

[667] Frank van der Linden and Jan Gerben Wijnstra. Platform engineering for the medical domain. Lecture Notes in Computer Science, 2290:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900224.
pdf.

Purhonen:2002:QAT

[668] Anu Purhonen. Quality attribute taxonomies for DSP software architecture design. Lecture Notes in Computer Science, 2290:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900238.
                                        pdf.

Bauman:2002:PAN

[669] Joe Bauman. The perfect architecture is non-optimal — winning with chaos. Lecture Notes in Computer Science, 2290:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900248.
                                        pdf.

Northrop:2002:RDS

[670] Linda Northrop. Report on discussion sessions “Diversity Solutions” and “Light-Weight Processes”. Lecture Notes in Computer Science, 2290:258– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900258.
                                        pdf.

Hallsteinsen:2002:HDN

[671] Svein Hallsteinsen and Eric Swane. Handling the diversity of networked devices by means of a product family approach. Lecture Notes in Computer Science, 2290:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900264.
pdf.

Krueger:2002:ETS

[672] Charles W. Krueger. Easing the transition to software mass customization. Lecture Notes in Computer Science, 2290:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2290/22900282.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2290/22900282.
pdf.

Becker:2002:CVM

[673] Martin Becker, Lars Geyer, Andreas Gilbert, and Karsten Becker. Comprehensive variability modelling to facilitate efficient variability treatment. Lecture Notes in Computer Science, 2290: 294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900294.
pdf.

Salicki:2002:EUV

[674] Serge Salicki and Nicolas Farcet. Expression and usage of the variability in the software product lines. Lecture Notes in Computer Science, 2290:304– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900304.
                                        pdf.

Capilla:2002:MVF

[675] Rafael Capilla and Juan C. Dueñas. Modelling variability with features in distributed architectures. Lecture Notes in Computer Science, 2290:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900319.
                                        pdf.

Dashofy:2002:RPF

[676] Eric M. Dashofy and André van der Hoek. Representing product family architectures in an extensible architecture description language. Lecture Notes in Computer Science, 2290:330– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900330.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900330.
                                        pdf.

Bockle:2002:PVD

[677] Günter Böckle and Klaus Schmid. Product validation: Discussion report. Lecture Notes in Computer Science, 2290:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2290/22900342.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2290/22900342.
pdf.

Dincel:2002:MPL

[678] Ebru Dincel, Nenad Medvidovic, and André van der Hoek. Measuring product line architectures. Lecture Notes in Computer Science, 2290:346– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900346.
pdf.

Ramesh:2002:SIP

[679] Balasubramaniam Ramesh, Amrit Tiwana, and Kannan Mohan. Supporting information product and service families with traceability. Lecture Notes in Computer Science, 2290:353– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900353.
pdf.

Ferber:2002:RPL

[680] Stefan Ferber, Peter Heidl, and Peter Lutz. Reviewing product line architectures: Experience report of ATAM in an automotive context. Lecture Notes in Computer Science, 2290:364– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900364.
                                        pdf.

Weingartner:2002:PFE

[681] Josef Weingärtner. Product family engineering and testing in the medical domain — validation aspects. Lecture Notes in Computer Science, 2290:383– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900383.
                                        pdf.

Clements:2002:PVS

[682] Paul C. Clements. Process validation, session report. Lecture Notes in Computer Science, 2290:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2290/22900388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2290/22900388.
                                        pdf.

Postema:2002:PBP

[683] Herman Postema and J. Henk Obbink. Platform based product development. Lecture Notes in Computer Science, 2290:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2290/22900390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900390.
pdf.

Mellado:2002:AVE

[684] Julio Mellado and Juan C. Dueñas. Automated validation environment for a product line of railway traffic control systems. Lecture Notes in Computer Science, 2290:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2290/22900401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2290/22900401.
pdf.

Bayer:2002:ITP

[685] Joachim Bayer and Tanya Widen. Introducing traceability to product lines. Lecture Notes in Computer Science, 2290:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2290/22900409.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2290/22900409.
pdf.

Anonymous:2002:AIq

[686] Anonymous. Author index. Lecture Notes in Computer Science, 2290:417– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2290/2290auth.pdf.

Sodring:2002:EME

[687] Thomas Sødring and Alan F. Smeaton. Evaluating a melody extraction engine. Lecture Notes in Computer Science, 2291:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2291/22910001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2291/22910001.
pdf.

Duffing:2002:OSP

[688] Gérald Duffing and Malika Smaïl. Organising and searching partially indexed image databases. Lecture Notes in Computer Science, 2291:22– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910022.
pdf.

Heesch:2002:CFC

[689] Daniel Heesch and Stefan Rüger. Combining features for content-based sketch retrieval — A comparative evaluation of retrieval performance. Lecture Notes in Computer Science, 2291:41– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910041.
pdf.

Tsikrika:2002:CWD

[690] Theodora Tsikrika and Mounia Lalmas. Combining Web document representations in a Bayesian inference network model using link and content-based evidence. Lecture Notes in Computer Science, 2291:53–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910053.
pdf.

Kim:2002:ICP

[691] Sung Jin Kim and Sang Ho Lee. An improved computation of the PageRank algorithm. Lecture Notes in Computer Science, 2291:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910073.
pdf.

Lepouras:2002:SEH

[692] George Lepouras, Costas Vassilakis, and George R. S. Weir. Serving enhanced hypermedia information. Lecture Notes in Computer Science, 2291: 86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910086.
pdf.

White:2002:UIE

[693] Ryen W. White, Ian Ruthven, and Joemon M. Jose. The use of implicit evidence for relevance feedback in Web retrieval. Lecture Notes in Computer Science, 2291:93– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910093.
                                        pdf.

Vakkari:2002:SKS

[694] Pertti Vakkari. Subject knowledge, source of terms, and term selection in query expansion: An analytical study. Lecture Notes in Computer Science, 2291:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910110.
                                        pdf.

Benammar:2002:APR

[695] Anis Benammar, Gilles Hubert, and Josiane Mothe. Automatic profile reformulation using a local document analysis. Lecture Notes in Computer Science, 2291:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910124.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2291/22910124.
pdf.

Boughanem:2002:SUG

[696] Mohand Boughanem and Lynda Tamine. A study on using genetic niching for query optimisation in document retrieval. Lecture Notes in Computer Science, 2291:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910135.
pdf.

Rajapakse:2002:CBA

[697] R. K. Rajapakse and M. Denham. Concept based adaptive IR model using FCA-BAM combination for concept representation and encoding. Lecture Notes in Computer Science, 2291:150– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910150.
pdf.

Campos:2002:LBN

[698] Luis M. de Campos, Juan M. Fernández-Luna, and Juan F. Huete. A layered Bayesian network model for document retrieval. Lecture Notes in Computer Science, 2291:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2291/22910169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910169.
                                        pdf.

Amati:2002:TFN

[699] Gianni Amati and Cornelis Joost van Rijsbergen. Term frequency normalization via Pareto distributions. Lecture Notes in Computer Science, 2291: 183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910183.
                                        pdf.

Lavrenko:2002:OMM

[700] Victor Lavrenko. Optimal mixture models in IR. Lecture Notes in Computer Science, 2291:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910193.
                                        pdf.

Kongovi:2002:TCE

[701] Madhusudhan Kongovi, Juan Carlos Guzman, and Venu Dasigi. Text categorization: An experiment using phrases. Lecture Notes in Computer Science, 2291:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910213.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2291/22910213.
pdf.

Gaussier:2002:HMC

[702] E. Gaussier, C. Goutte, K. Popat, and F. Chen. A hierarchical model for clustering and categorising documents. Lecture Notes in Computer Science, 2291:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910229.
pdf.

Peters:2002:UBN

[703] C. Peters and C. H. A. Koster. Uncertainty-based noise reduction and term selection in text categorization. Lecture Notes in Computer Science, 2291:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2291/22910248.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2291/22910248.
pdf.

Vegas:2002:GUI

[704] Jesús Vegas, Pablo de la Fuente, and Fabio Crestani. A graphical user interface for structured document retrieval. Lecture Notes in Computer Science, 2291:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2291/22910268.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910268.
                                        pdf.

Roelleke:2002:ADS

[705] Thomas Roelleke, Mounia Lalmas, Gabriella Kazai, Ian Ruthven, and Stefan Quicker. The accessibility dimension for structured document retrieval. Lecture Notes in Computer Science, 2291:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910284.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910284.
                                        pdf.

McEwan:2002:BBD

[706] Craig J. A. McEwan, Iadh Ounis, and Ian Ruthven. Building bilingual dictionaries from parallel Web documents. Lecture Notes in Computer Science, 2291:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2291/22910303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2291/22910303.
                                        pdf.

Oard:2002:TBI

[707] Douglas W. Oard and Funda Ertunc. Translation-based indexing for cross-language retrieval. Lecture Notes in Computer Science, 2291:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910324.
pdf.

Sormunen:2002:REM

[708] Eero Sormunen. A retrospective evaluation method for exact-match and bestmatch queries applying an interactive query performance analyser. Lecture Notes in Computer Science, 2291:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910334.
pdf.

Finn:2002:GCD

[709] Aidan Finn, Nicholas Kushmerick, and Barry Smyth. Genre classification and domain transfer for information filtering. Lecture Notes in Computer Science, 2291:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2291/22910353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2291/22910353.
pdf.

Anonymous:2002:AIr

[710] Anonymous. Author index. Lecture Notes in Computer Science, 2291:363– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2291/2291auth.pdf.

Colbourn:2002:MAC

[711] Charles J. Colbourn. Multiple access communications using combinatorial designs. Lecture Notes in Computer Science, 2292:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2292/22920001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2292/22920001.
                                        pdf.

Kiwi:2002:EAT

[712] Marcos Kiwi, Frédéric Magniez, and Miklos Santha. Exact and approximate testing/correcting of algebraic functions: a survey. Lecture Notes in Computer Science, 2292:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2292/22920030.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2292/22920030.
                                        pdf.

Komlos:2002:RLA

[713] János Komlós, Ali Shokoufandeh, Miklós Simonovits, and Endre Szemerédi. The regularity lemma and its applications in graph theory. Lecture Notes in Computer Science, 2292:84– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2292/22920084.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2292/22920084.
pdf.

Lellahi:2002:MDO

[714] Kazem Lellahi. Modeling data and objects: An algebraic view point. Lecture Notes in Computer Science, 2292: 113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2292/22920113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2292/22920113.
pdf.

Shokoufandeh:2002:GTM

[715] Ali Shokoufandeh and Sven Dickinson. Graph-theoretical methods in computer vision. Lecture Notes in Computer Science, 2292:148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2292/22920148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2292/22920148.
pdf.

Shokrollahi:2002:ILD

[716] Amin Shokrollahi. An introduction to low-density parity-check codes. Lecture Notes in Computer Science, 2292:175– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2292/22920175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2292/22920175.
pdf.

Vazirani:2002:PDS

[717] Vijay V. Vazirani. Primal-dual schema based approximation algorithms. Lecture Notes in Computer Science, 2292: 198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2292/22920198.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2292/22920198.
                                        pdf.

Vazirani:2002:FTQ

[718] Umesh Vazirani. Fourier transforms and quantum computation. Lecture Notes in Computer Science, 2292:208– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2292/22920208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2292/22920208.
                                        pdf.

Anonymous:2002:AIs

[719] Anonymous. Author index. Lecture Notes in Computer Science, 2292:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2292/2292auth.pdf.

Renz:2002:I

[720] J. Renz. 1. introduction. Lecture Notes in Computer Science, 2293:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2293/22930001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930001.
pdf.

Renz:2002:B

[721] J. Renz. 2. background. Lecture Notes in Computer Science, 2293:13– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930013.
pdf.

Renz:2002:QSR

[722] J. Renz. 3. qualitative spatial representation and reasoning. Lecture Notes in Computer Science, 2293:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930031.
pdf.

Renz:2002:RCC

[723] J. Renz. 4. the region connection calculus. Lecture Notes in Computer Science, 2293:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930041.
pdf.

Renz:2002:CPT

[724] J. Renz. 5. cognitive properties of topological spatial relations. Lecture Notes in Computer Science, 2293:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2293/22930051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2293/22930051.
                                        pdf.

Renz:2002:CPR

[725] J. Renz. 6. computational properties of RCC-8. Lecture Notes in Computer Science, 2293:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2293/22930065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2293/22930065.
                                        pdf.

Renz:2002:CAT

[726] J. Renz. 7. A complete analysis of tractability in RCC-8. Lecture Notes in Computer Science, 2293:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2293/22930117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2293/22930117.
                                        pdf.

Renz:2002:EER

[727] J. Renz. 8. empirical evaluation of reasoning with RCC-8. Lecture Notes in Computer Science, 2293:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930131.
pdf.

Renz:2002:RPR

[728] J. Renz. 9. representational properties of RCC-8. Lecture Notes in Computer Science, 2293:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930155.
pdf.

Renz:2002:C

[729] J. Renz. 10. conclusions. Lecture Notes in Computer Science, 2293:173– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2293/22930173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2293/22930173.
pdf.

Renz:2002:ERM

[730] J. Renz. A. enumeration of the relations of the maximal tractable subsets of RCC-8. Lecture Notes in Computer Science, 2293:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2293/22930179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2293/22930179.
                                        pdf.

Anonymous:2002:Rc

[731] Anonymous. References. Lecture Notes in Computer Science, 2293:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2293/2293ref.pdf.

Anonymous:2002:Ic

[732] Anonymous. Index. Lecture Notes in Computer Science, 2293:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2293/2293ind.pdf.

Bernardeschi:2002:CAI

[733] Cinzia Bernardeschi and Nicoletta De Francesco. Combining abstract interpretation and model checking for analysing security properties of Java bytecode. Lecture Notes in Computer Science, 2294:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940001.
                                        pdf.

Focardi:2002:PMB

[734] Riccardo Focardi, Carla Piazza, and Sabina Rossi. Proofs methods for bisimulation based information flow security. Lecture Notes in Computer Science, 2294:16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940016.
pdf.

Barthe:2002:FCB

[735] Gilles Barthe, Guillaume Dufay, Line Jakubiec, and Simão Melo de Sousa. A formal correspondence between offensive and defensive JavaCard virtual machines. Lecture Notes in Computer Science, 2294:32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940032.
pdf.

Shyamasundar:2002:ACP

[736] R. K. Shyamasundar. Analyzing cryptographic protocols in a reactive framework. Lecture Notes in Computer Science, 2294:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940046.
pdf.

Tan:2002:ASE

[737] Li Tan. An abstract schema for equivalence-checking games. Lecture Notes in Computer Science, 2294:65– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940065.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940065.
                                        pdf.

Sidorova:2002:SCT

[738] Natalia Sidorova and Martin Steffen. Synchronous closing of timed SDL systems for model checking. Lecture Notes in Computer Science, 2294:79– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940079.
                                        pdf.

Faella:2002:ATD

[739] Marco Faella, Salvatore La Torre, and Aniello Murano. Automata-theoretic decision of timed games. Lecture Notes in Computer Science, 2294:94– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940094.
                                        pdf.

Charatonik:2002:CTA

[740] Witold Charatonik, Supratik Mukhopadhyay, and Andreas Podelski. Compositional termination analysis of symbolic forward analysis. Lecture Notes in Computer Science, 2294:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940109.
pdf.

Genaim:2002:CNP

[741] Samir Genaim, Michael Codish, John Gallagher, and Vitaly Lagoon. Combining norms to prove termination. Lecture Notes in Computer Science, 2294:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2294/22940126.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2294/22940126.
pdf.

Murawski:2002:SMA

[742] Andrzej S. Murawski and Kwangkeun Yi. Static monotonicity analysis for λdefinable functions over lattices. Lecture Notes in Computer Science, 2294: 139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940139.
pdf.

Hill:2002:REP

[743] Patricia M. Hill and Fausto Spoto. A refinement of the escape property. Lecture Notes in Computer Science, 2294: 154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940154.
                                        pdf.

Troncon:2002:SSR

[744] Remko Tronçon, Maurice Bruynooghe, Gerda Janssens, and Francky Catthoor. Storage size reduction by in-place mapping of arrays. Lecture Notes in Computer Science, 2294:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940167.
                                        pdf.

Krstic:2002:VBA

[745] Sava Krstic and John Matthews. Verifying BDD algorithms through monadic interpretation. Lecture Notes in Computer Science, 2294:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940182.
                                        pdf.

Cimatti:2002:IEL

[746] Alessandro Cimatti, Marco Pistore, Marco Roveri, and Roberto Sebastiani. Improving the encoding of LTL model checking into SAT. Lecture Notes in Computer Science, 2294:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940196.
pdf.

Zuck:2002:AVP

[747] Lenore Zuck, Amir Pnueli, and Yonit Kesten. Automatic verification of probabilistic free choice. Lecture Notes in Computer Science, 2294:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940208.
pdf.

Gori:2002:ETI

[748] Roberta Gori and Giorgio Levi. An experiment in type inference and verification by abstract interpretation. Lecture Notes in Computer Science, 2294: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940225.
pdf.

Torre:2002:WMA

[749] Salvatore La Torre, Aniello Murano, and Margherita Napoli. Weak Muller acceptance conditions for tree automata. Lecture Notes in Computer Science, 2294:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940240.
                                        pdf.

Coppo:2002:FAM

[750] Mario Coppo and Mariangiola Dezani-Ciancaglini. A fully abstract model for higher-order mobile ambients. Lecture Notes in Computer Science, 2294:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940255.
                                        pdf.

Tiplea:2002:SPA

[751] Ferucio Laurentiu Tiplea and Aurora Tiplea. A simulation preorder for abstraction of reactive systems. Lecture Notes in Computer Science, 2294:272– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2294/22940272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2294/22940272.
                                        pdf.

Harding:2002:AAA

[752] Aidan Harding, Mark Ryan, and Pierre-Yves Schobbens. Approximating ATL^{∗} in ATL. Lecture Notes in Computer Science, 2294:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940289.
pdf.

Huth:2002:MCM

[753] Michael Huth. Model checking modal transition systems using Kripke structures. Lecture Notes in Computer Science, 2294:302–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2294/22940302.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2294/22940302.pdf.

Baukus:2002:PVC

[754] Kai Baukus, Yassine Lakhnech, and Karsten Stahl. Parameterized verification of a cache coherence protocol: Safety and liveness. Lecture Notes in Computer Science, 2294:317– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2294/22940317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2294/22940317.
pdf.

Anonymous:2002:AIt

[755] Anonymous. Author index. Lecture Notes in Computer Science, 2294:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2294/2294indx.pdf.

Calude:2002:AUQ

[756] Cristian S. Calude and Elena Calude. Automata: From uncertainty to quantum. Lecture Notes in Computer Science, 2295:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2295/22950001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2295/22950001.pdf.

Choffrut:2002:ETO

[757] Christian Choffrut. Elementary theory of ordinals with addition and left translation by ω. Lecture Notes in Computer Science, 2295:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950015.
                                        pdf.

Esik:2002:ETF

[758] Z. Ésik. The equational theory of fixed points with applications to generalized language theory. Lecture Notes in Computer Science, 2295:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950021.
pdf.

Eiter:2002:SOL

[759] Thomas Eiter, Georg Gottlob, and Thomas Schwentick. Second-order logic over strings: Regular and nonregular fragments. Lecture Notes in Computer Science, 2295:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950037.
pdf.

Harju:2002:DQI

[760] Tero Harju. Decision questions on integer matrices. Lecture Notes in Computer Science, 2295:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950057.
pdf.

Ito:2002:SPN

[761] Masami Ito and Yoshiyuki Kunimoch. Some Petri net languages and codes. Lecture Notes in Computer Science, 2295:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950069.
                                        pdf.

Prodinger:2002:WPR

[762] Helmut Prodinger. Words, permutations, and representations of numbers. Lecture Notes in Computer Science, 2295:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2295/22950081.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950081.
                                        pdf.

Razborov:2002:PCP

[763] Alexander A. Razborov. Proof complexity of pigeonhole principles. Lecture Notes in Computer Science, 2295: 100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950100.
                                        pdf.

Restivo:2002:WP

[764] Antonio Restivo and Sergio Salemi. Words and patterns. Lecture Notes in Computer Science, 2295:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950117.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2295/22950117.
pdf.

Thomas:2002:SII

[765] Wolfgang Thomas. A short introduction to infinite automata. Lecture Notes in Computer Science, 2295:130– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950130.
pdf.

Cachat:2002:POL

[766] Thierry Cachat. The power of oneletter rational languages. Lecture Notes in Computer Science, 2295:145– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950145.
pdf.

Staiger:2002:ELL

[767] Ludwig Staiger. The entropy of Lukasiewicz-languages. Lecture Notes in Computer Science, 2295:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950155.
pdf.

Ananichev:2002:CWV

[768] D. S. Ananichev and M. V. Volkov. Collapsing words vs. synchronizing words. Lecture Notes in Computer Science, 2295:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2295/22950166.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2295/22950166.
pdf.

Culik:2002:NSA

[769] Karel Culik, Juhani Karhumäki, and Jarkko Kari. A note on synchronized automata and road coloring problem. Lecture Notes in Computer Science, 2295:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950175.
pdf.

Campeanu:2002:SQD

[770] C. Câmpeanu, K. Salomaa, and S. Vágvölgyi. Shuffle quotient and decompositions. Lecture Notes in Computer Science, 2295:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950186.
pdf.

Niemann:2002:GCS

[771] Gundula Niemann and Jens R. Woinowski. The growing contextsensitive languages are the acyclic context-sensitive languages. Lecture Notes in Computer Science, 2295:197– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950197.
pdf.

Kuske:2002:RSF

[772] Dietrich Kuske. Recognizable sets of N-free pomsets are monadically axiomatizable. Lecture Notes in Computer Science, 2295:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950206.
pdf.

Esik:2002:ASP

[773] Z. Ésik and Z. L. Németh. Automata on series-parallel biposets. Lecture Notes in Computer Science, 2295:217– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950217.
pdf.

Engelfriet:2002:HSL

[774] Joost Engelfriet and Sebastian Maneth. Hierarchies of string languages generated by deterministic tree transducers. Lecture Notes in Computer Science, 2295:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2295/22950228.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2295/22950228.
pdf.

Schwentick:2002:POT

[775] Thomas Schwentick, Denis Thérien, and Heribert Vollmer. Partiallyordered two-way automata: a new characterization of DA. Lecture Notes in Computer Science, 2295:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950239.
pdf.

Glasser:2002:LST

[776] Christian Glaßer and Heinz Schmitz. Level 5/2 of the Straubing-Thérien hierarchy for two-letter alphabets. Lecture Notes in Computer Science, 2295: 251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950251.
pdf.

Hromkovic:2002:PRP

[777] Juraj Hromkovič and Georg Schnitger. On the power of randomized pushdown automata. Lecture Notes in Computer Science, 2295:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950262.
pdf.

Lischke:2002:RLC

[778] Gerhard Lischke. The root of a language and its complexity. Lecture Notes in Computer Science, 2295:272– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950272.
pdf.

Fernau:2002:VVG

[779] Henning Fernau and Ralf Stiebe. Valuated and valence grammars: An algebraic view. Lecture Notes in Computer Science, 2295:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950281.
pdf.

Hoogeboom:2002:CFV

[780] Hendrik Jan Hoogeboom. Context-free valence grammars — revisited. Lecture Notes in Computer Science, 2295: 293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950293.
                                        pdf.

Halava:2002:URC

[781] Vesa Halava and Tero Harju. An undecidability result concerning periodic morphisms. Lecture Notes in Computer Science, 2295:304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950304.
                                        pdf.

Kudlek:2002:UTM

[782] Manfred Kudlek and Yurii Rogozhin. A universal Turing machine with 3 states and 9 symbols. Lecture Notes in Computer Science, 2295:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950311.
                                        pdf.

Domaratzki:2002:MCF

[783] Michael Domaratzki, Jeffrey Shallit, and Sheng Yu. Minimal covers of formal languages. Lecture Notes in Computer Science, 2295:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950319.
pdf.

Niemann:2002:SRL

[784] Gundula Niemann and Johannes Waldmann. Some regular languages that are Church–Rosser congruential. Lecture Notes in Computer Science, 2295: 330–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950330.
pdf.

Beaudry:2002:RBM

[785] M. Beaudry, M. Holzer, Gundula Niemann, and F. Otto. On the relationship between the Mc-Naughton families of languages and the Chomsky hierarchy. Lecture Notes in Computer Science, 2295:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950340.
pdf.

Mignosi:2002:FFF

[786] Filippo Mignosi, Antonio Restivo, and Marinella Sciortino. Forbidden factors and fragment assembly. Lecture Notes in Computer Science, 2295:349– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950349.
                                        pdf.

Csuhaj-Varju:2002:PCG

[787] Erzsébet Csuhaj-Varjú and György Vaszil. Parallel communicating grammar systems with incomplete information communication. Lecture Notes in Computer Science, 2295:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950359.
                                        pdf.

Rovan:2002:ECP

[788] Branislav Rovan and Marián Slašt’an. Eliminating communication by parallel rewriting. Lecture Notes in Computer Science, 2295:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2295/22950369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2295/22950369.
                                        pdf.

Sosik:2002:SRS

[789] Petr Sosík and Rudolf Freund. String rewriting sequential P-systems and regulated rewriting. Lecture Notes in Computer Science, 2295:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2295/22950379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2295/22950379.
pdf.

Anonymous:2002:AIu

[790] Anonymous. Author index. Lecture Notes in Computer Science, 2295:389– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2295/2295auth.pdf.

Skowron:2002:ARA

[791] Andrzej Skowron. Approximate reasoning by agents. Lecture Notes in Computer Science, 2296:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960003.
pdf.

Zielinski:2002:NSC

[792] Krzysztof Zieliński. Network services in context of pervasive mobile Internet. Lecture Notes in Computer Science, 2296:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2296/22960015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2296/22960015.
pdf.

Agray:2002:BLI

[793] Nesria Agray, Wiebe van der Hoek, and Erik de Vink. On BAN logics for industrial security protocols. Lecture Notes in Computer Science, 2296:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960029.
pdf.

Ambroszkiewicz:2002:CAL

[794] Stanislaw Ambroszkiewicz, Tomasz Nowak, Dariusz Mikulowski, and Leszek Rozwadowski. A concept of agent language in agentspace. Lecture Notes in Computer Science, 2296:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960037.
pdf.

Bielecki:2002:AAS

[795] Andrzej Bielecki. Agent’s adaptivity — some aspects of theoretical foundations of a neural agent training process. Lecture Notes in Computer Science, 2296: 47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960047.
pdf.

Burgett:2002:MPA

[796] J. Burgett, S. Hussain, C. M. Jonker, A. Razaq, and K. Silz. Mobie: Personalized automated recharge of prepaid mobile phones. Lecture Notes in Computer Science, 2296:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960054.
pdf.

Cetnarowicz:2002:FSM

[797] Krzysztof Cetnarowicz, Pablo Gruer, Vincent Hilaire, and Abder Koukam. A formal specification of M-agent architecture. Lecture Notes in Computer Science, 2296:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2296/22960062.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2296/22960062.pdf.

Cetnarowicz:2002:MAS

[798] Krzysztof Cetnarowicz and Jarosław Koźlak. Multi-agent system for flexible manufacturing systems management. Lecture Notes in Computer Science, 2296:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2296/22960073.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960073.
                                        pdf.

Dignum:2002:MAA

[799] Frank Dignum, David Kinny, and Liz Sonenberg. Motivational attitudes of agents: On desires, obligations, and norms. Lecture Notes in Computer Science, 2296:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960083.
                                        pdf.

Fasli:2002:CRO

[800] Maria Fasli. On commitments, roles, and obligations. Lecture Notes in Computer Science, 2296:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960093.
                                        pdf.

Flasinski:2002:ABM

[801] Mariusz Flasinski. Automata-based multi-agent model as a tool for constructing real-time intelligent control systems. Lecture Notes in Computer Science, 2296:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2296/22960103.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2296/22960103.pdf.

Gajos:2002:RRM

[802] Krzysztof Gajos. Rascal — A resource manager for multi agent systems in smart spaces. Lecture Notes in Computer Science, 2296:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960111.
pdf.

Gorodetski:2002:SDK

[803] Vladimir Gorodetski, Oleg Karsayev, Igor Kotenko, and Alexey Khabalov. Software development kit for multi-agent systems design and implementation. Lecture Notes in Computer Science, 2296:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960121.
pdf.

Hameurlain:2002:FSB

[804] Nabil Hameurlain. Formal semantics for behavioural substitutability of agent components: Application to interaction protocols. Lecture Notes in Computer Science, 2296:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960131.
                                        pdf.

Jonker:2002:MEM

[805] Catholijn M. Jonker, Jan Treur, and Wouter C. A. Wijngaards. A modelling environment for mind and matter aspects of intentional behaviour. Lecture Notes in Computer Science, 2296: 141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960141.
                                        pdf.

Jonker:2002:TLS

[806] Catholijn M. Jonker, Jan Treur, and Wouter C. A. Wijngaards. Temporal languages for simulation and analysis of the dynamics within an organisation. Lecture Notes in Computer Science, 2296:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960151.
                                        pdf.

Jurek:2002:SPR

[807] Janusz Jurek. Syntactic pattern recognition-based agents for realtime expert systems. Lecture Notes in Computer Science, 2296:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960161.
pdf.

Kiss:2002:LUP

[808] Adorjan Kiss and Joël Quinqueton. Learning user preferences in multi-agent system. Lecture Notes in Computer Science, 2296:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960169.
pdf.

Lioncourt:2002:TRA

[809] Sorabain Wolfheart de Lioncourt and Michael Luck. Towards requirements analysis for autonomous agent behaviour. Lecture Notes in Computer Science, 2296:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960179.
pdf.

Mazurkiewicz:2002:MCG

[810] Antoni Mazurkiewicz and Dobiesław Wróblewski. Messages, clocks, and gravitation. Lecture Notes in Computer Science, 2296:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960187.
                                        pdf.

Monekosso:2002:IQL

[811] Ndedi Monekosso, Paolo Remagnino, and Adam Szarowicz. An improved Q-learning algorithm using synthetic pheromones. Lecture Notes in Computer Science, 2296:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960197.
                                        pdf.

Nawarecki:2002:OPC

[812] Edward Nawarecki, Marek Kisiel-Dorohinicki, and Grzegorz Dobrowolski. Organisations in the particular class of multi-agent systems. Lecture Notes in Computer Science, 2296:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960207.
                                        pdf.

Oechslein:2002:UBO

[813] Christoph Oechslein, Franziska Klügl, Rainer Herrler, and Frank Puppe.

UML for behavior-oriented multiagent simulations. Lecture Notes in Computer Science, 2296:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960217.
pdf.

Pavliska:2002:RDP

[814] Ales Pavliska and Vilem Srovnal. Robot disassembly process using multi-agent system. Lecture Notes in Computer Science, 2296:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960227.
pdf.

Pechoucek:2002:CAM

[815] Michal Pechoucek, Vladimir Marik, and Jaroslav Barta. CPlanT: An acquaintance model-based coalition formation multi-agent system. Lecture Notes in Computer Science, 2296:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960234.
pdf.

Penczek:2002:EMC

[816] Wojciech Penczek. Efficient model checking of causal-knowledge protocols. Lecture Notes in Computer Science, 2296:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2296/22960242.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960242.
                                        pdf.

Ricordel:2002:VVO

[817] Pierre-Michel Ricordel and Yves Demazeau. Volcano, a vowels-oriented multi-agent platform. Lecture Notes in Computer Science, 2296:253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960253.
                                        pdf.

Cortes:2002:MCM

[818] José C. Romero Cortés and Leonid B. Sheremetov. Model of cooperation in multi-agent systems with fuzzy coalitions. Lecture Notes in Computer Science, 2296:263–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960263.
                                        pdf.

Sabouret:2002:LCB

[819] Nicolas Sabouret and Jean-Paul Sansonnet. Learning collective behaviour from local interactions. Lecture Notes in Computer Science, 2296:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960273.
pdf.

Sauvage:2002:MOP

[820] Sylvain Sauvage. MAS oriented patterns. Lecture Notes in Computer Science, 2296:283–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2296/22960283.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2296/22960283.pdf.

Smirnov:2002:MAA

[821] Alexander Smirnov, Mikhail Pashkin, Nikolai Chilov, and Tatiana Levashova. Multi-agent architecture for knowledge fusion from distributed sources. Lecture Notes in Computer Science, 2296: 293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960293.
pdf.

Szuba:2002:UFM

[822] Tadeusz Szuba. Universal formal model of collective intelligence and its IQ measure. Lecture Notes in Computer Science, 2296:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960303.
                                        pdf.

Vries:2002:PLC

[823] Wieke de Vries, Frank S. de Boer, Koen V. Hindriks, Wiebe van der Hoek, and John-Jules C. Meyer. A programming language for coordinating group actions. Lecture Notes in Computer Science, 2296:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960313.
                                        pdf.

Zaikin:2002:ORA

[824] Oleg Zaikin, Alexandre Dolgui, and Przemyslaw Korytkowski. Optimization of resource allocation in distributed production networks. Lecture Notes in Computer Science, 2296:322– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2296/22960322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2296/22960322.
                                        pdf.

Zhang:2002:UBR

[825] Xin Zhang and Henry Hexmoor. Utility-based role exchange. Lecture Notes in Computer Science, 2296:332– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2296/22960332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2296/22960332.
pdf.

Anonymous:2002:AIv

[826] Anonymous. Author index. Lecture Notes in Computer Science, 2296:341– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2296/2296auth.pdf.

Crole:2002:CI

[827] Roy Crole. Chapter 1 introduction. Lecture Notes in Computer Science, 2297:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2297/22970001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2297/22970001.
pdf.

Priestley:2002:COS

[828] Hilary A. Priestley. Chapter 2 ordered sets and complete lattices. Lecture Notes in Computer Science, 2297:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2297/22970021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2297/22970021.
pdf.

Aczel:2002:CAC

[829] Peter Aczel. Chapter 3 algebras and coalgebras. Lecture Notes in Computer Science, 2297:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2297/22970079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2297/22970079.
                                        pdf.

Backhouse:2002:CGC

[830] Roland Backhouse. Chapter 4 Galois connections and fixed point calculus. Lecture Notes in Computer Science, 2297:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2297/22970089.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2297/22970089.
                                        pdf.

Gibbons:2002:CCF

[831] Jeremy Gibbons. Chapter 5 calculating functional programs. Lecture Notes in Computer Science, 2297:149– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2297/22970149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2297/22970149.
                                        pdf.

Doornbos:2002:CAP

[832] Henk Doornbos and Roland Backhouse. Chapter 6 algebra of program termination. Lecture Notes in Computer Science, 2297:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2297/22970203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2297/22970203.
pdf.

Jacobs:2002:CEC

[833] Bart Jacobs. Chapter 7 exercises in coalgebraic specification. Lecture Notes in Computer Science, 2297:237– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2297/22970237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2297/22970237.
pdf.

Bird:2002:CAM

[834] Richard Bird, Jeremy Gibbons, and Shin-Cheng Mu. Chapter 8 algebraic methods for optimization problems. Lecture Notes in Computer Science, 2297:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2297/22970281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2297/22970281.
pdf.

Karger:2002:CTA

[835] Burghard von Karger. Chapter 9 temporal algebra. Lecture Notes in Computer Science, 2297:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2297/22970309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2297/22970309.
                                        pdf.

Anonymous:2002:AIw

[836] Anonymous. Author index. Lecture Notes in Computer Science, 2297:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2297/2297ind.pdf.

Braffort:2002:RCS

[837] Annelies Braffort. Research on computer science and sign language: Ethical aspects. Lecture Notes in Computer Science, 2298:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980001.
                                        pdf.

Benbasat:2002:IMF

[838] Ari Y. Benbasat and Joseph A. Paradiso. An inertial measurement framework for gesture recognition and applications. Lecture Notes in Computer Science, 2298:9– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980009.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2298/22980009.
pdf.

Sowa:2002:ISR

[839] Timo Sowa and Ipke Wachsmuth. Interpretation of shape-related iconic gestures in virtual environments. Lecture Notes in Computer Science, 2298:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980021.
pdf.

Corradini:2002:RTG

[840] Andrea Corradini. Real-time gesture recognition by means of hybrid recognizers. Lecture Notes in Computer Science, 2298:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980034.
pdf.

Munk:2002:DGP

[841] Karin Husballe Munk. Development of a gesture plug-in for natural dialogue interfaces. Lecture Notes in Computer Science, 2298:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980047.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2298/22980047.
pdf.

Moeslund:2002:NIV

[842] Thomas B. Moeslund, Moritz Störring, and Erik Granum. A natural interface to a virtual environment through computer vision-estimated pointing gestures. Lecture Notes in Computer Science, 2298:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980059.
pdf.

Bauer:2002:TAS

[843] Britta Bauer and Karl-Friedrich Kraiss. Towards an automatic sign language recognition system using subunits. Lecture Notes in Computer Science, 2298: 64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980064.
pdf.

Fang:2002:SIC

[844] Gaolin Fang, Wen Gao, Xilin Chen, Chunli Wang, and Jiyong Ma. Signerindependent continuous sign language recognition based on SRN/HMM. Lecture Notes in Computer Science, 2298: 76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2298/22980076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980076.
pdf.

Chunli:2002:RTL

[845] Wang Chunli, Gao Wen, and Ma Jiyong. A real-time large vocabulary recognition system for Chinese Sign Language. Lecture Notes in Computer Science, 2298:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980086.
pdf.

Jiangqin:2002:RFS

[846] Wu Jiangqin and Gao Wen. The recognition of finger-spelling for Chinese Sign Language. Lecture Notes in Computer Science, 2298:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980096.
pdf.

Tyrone:2002:OCT

[847] Martha E. Tyrone. Overview of capture techniques for studying sign language phonetics. Lecture Notes in Computer Science, 2298:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2298/22980101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980101.
                                        pdf.

Gibet:2002:MBR

[848] Sylvie Gibet, Pierre-François Marteau, and Frédéric Julliard. Models with biological relevance to control anthropomorphic limbs: a survey. Lecture Notes in Computer Science, 2298:105– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980105.
                                        pdf.

Wachsmuth:2002:LGS

[849] Ipke Wachsmuth and Stefan Kopp. Lifelike gesture synthesis and timing for conversational agents. Lecture Notes in Computer Science, 2298:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980120.
                                        pdf.

Grieve-Smith:2002:SSL

[850] Angus B. Grieve-Smith. SignSynth: a sign language synthesis application using Web3D and Perl. Lecture Notes in Computer Science, 2298:134– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2298/22980134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980134.
pdf.

Kennaway:2002:SAD

[851] Richard Kennaway. Synthetic animation of deaf signing gestures. Lecture Notes in Computer Science, 2298:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980146.
pdf.

Poggi:2002:TGP

[852] Isabella Poggi. From a typology of gestures to a procedure for gesture production. Lecture Notes in Computer Science, 2298:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980158.
pdf.

Verlinden:2002:SAW

[853] Margriet Verlinden, Corrie Tijsseling, and Han Frowein. A signing avatar on the WWW. Lecture Notes in Computer Science, 2298:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980169.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2298/22980169.
                                        pdf.

Sallandre:2002:ISL

[854] Marie-Anne Sallandre and Christian Cuxac. Iconicity in sign language: a theoretical and methodological point of view. Lecture Notes in Computer Science, 2298:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980173.
                                        pdf.

Kanda:2002:NSS

[855] Kazuyuki Kanda, Akira Ichikawa, Yuji Nagashima, Yushi Kato, Mina Terauchi, Daisuke Hara, and Masanobu Sato. Notation system and statistical analysis of NMS in JSL. Lecture Notes in Computer Science, 2298:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980181.
                                        pdf.

Antzakas:2002:HMN

[856] Klimis Antzakas and Bencie Woll. Head movements and negation in Greek Sign Language. Lecture Notes in Computer Science, 2298:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2298/22980193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980193.
pdf.

Lejeune:2002:SSR

[857] Fanch Lejeune, Annelies Braffort, and Jean-Pierre Desclés. Study on semantic representations of French Sign Language sentences. Lecture Notes in Computer Science, 2298:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980197.
pdf.

Costa:2002:SBS

[858] Antônio Carlos da Rocha Costa and Graçaliz Pereira Dimuro. SignWritingbased sign language processing. Lecture Notes in Computer Science, 2298: 202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980202.
pdf.

Gullberg:2002:VAT

[859] Marianne Gullberg and Kenneth Holmqvist. Visual attention towards gestures in face-to-face interaction vs. on screen. Lecture Notes in Computer Science, 2298:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980206.
                                        pdf.

Steininger:2002:LGS

[860] Silke Steininger, Bernd Lindemann, and Thorsten Paetzold. Labeling of gestures in SmartKom — the coding system. Lecture Notes in Computer Science, 2298:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980215.
                                        pdf.

Beringer:2002:EGS

[861] Nicole Beringer. Evoking gestures in SmartKom — design of the graphical user interface. Lecture Notes in Computer Science, 2298:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980228.
                                        pdf.

Wanderley:2002:QAN

[862] Marcelo M. Wanderley. Quantitative analysis of non-obvious performer gestures. Lecture Notes in Computer Science, 2298:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2298/22980241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980241.
pdf.

Ogden:2002:ISA

[863] Bernard Ogden, Kerstin Dautenhahn, and Penny Stribling. Interactional structure applied to the identification and generation of visual interactive behavior: Robots that (usually) follow the rules. Lecture Notes in Computer Science, 2298:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980254.
pdf.

Dumas:2002:PGS

[864] Isabelle Dumas. Are praxical gestures semiotised in service encounters? Lecture Notes in Computer Science, 2298: 268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980268.
pdf.

Howell:2002:VMI

[865] A. Jonathan Howell and Hilary Buxton. Visually mediated interaction using learnt gestures and camera control. Lecture Notes in Computer Science, 2298:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980272.
                                        pdf.

Arfib:2002:GCS

[866] Daniel Arfib and Loïc Kessous. Gestural control of sound synthesis and processing algorithms. Lecture Notes in Computer Science, 2298:285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980285.
                                        pdf.

Willier:2002:JGA

[867] Aymeric Willier and Catherine Marque. Juggling gestures analysis for music control. Lecture Notes in Computer Science, 2298:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2298/22980296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2298/22980296.
                                        pdf.

Hermann:2002:HPS

[868] Thomas Hermann, Claudia Nölker, and Helge Ritter. Hand postures for sonification control. Lecture Notes in Computer Science, 2298:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980307.
pdf.

Vassilakis:2002:CFT

[869] Helen Vassilakis, A. Jonathan Howell, and Hilary Buxton. Comparison of feedforward (TDRBF) and generative (TDRGBN) network for gesture based control. Lecture Notes in Computer Science, 2298:317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2298/22980317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2298/22980317.
pdf.

Anonymous:2002:AIx

[870] Anonymous. Author index. Lecture Notes in Computer Science, 2298:323– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2298/2298ind.pdf.

Herrtwich:2002:KCC

[871] Ralf Guido Herrtwich. Keynote: Communicating cars: a case for ubiquitous computing in the automotive domain. Lecture Notes in Computer Science, 2299:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990003.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990003.
                                        pdf.

Fleischmann:2002:KMA

[872] Marc Fleischmann. Keynote: Microprocessor architectures for the mobile Internet era. Lecture Notes in Computer Science, 2299:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990004.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990004.
                                        pdf.

Thayer:2002:AIP

[873] Scott M. Thayer and Peter Steenkiste. An architecture for the integration of physical and informational spaces. Lecture Notes in Computer Science, 2299: 7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990007.
                                        pdf.

Hohl:2002:CSM

[874] Fritz Hohl, Lars Mehrmann, and Amen Hamdan. A context system for a mobile service platform. Lecture Notes in Computer Science, 2299:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990021.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2299/22990021.
pdf.

Michahelles:2002:DCD

[875] Florian Michahelles, Michael Samulowitz, and Bernt Schiele. Detecting context in distributed sensor networks by using smart contextaware packets. Lecture Notes in Computer Science, 2299:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990034.
pdf.

Weatherall:2002:UCA

[876] James “Wez” Weatherall and Duncan Grisby. A ubiquitous control architecture for low power systems. Lecture Notes in Computer Science, 2299:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990051.
pdf.

Cheng:2002:SAB

[877] Shang-Wen Cheng, David Garlan, Bradley Schmerl, João Pedro Sousa, Bridget Spitznagel, Peter Steenkiste, and Ningning Hu. Software architecturebased adaptation for pervasive systems. Lecture Notes in Computer Science, 2299:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2299/22990067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990067.
                                        pdf.

Poellabauer:2002:KEC

[878] Christian Poellabauer, Karsten Schwan, Greg Eisenhauer, and Jiantao Kong. KECho — event communication for distributed kernel services. Lecture Notes in Computer Science, 2299:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990083.
                                        pdf.

Coschurba:2002:FGA

[879] Peter Coschurba, Kurt Rothermel, and Frank Dürr. A fine-grained addressing concept for GeoCast. Lecture Notes in Computer Science, 2299:101– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990101.
                                        pdf.

Decker:2002:DPW

[880] Christian Decker and Michael Beigl. Data paths in wearable communication networks. Lecture Notes in Computer Science, 2299:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2299/22990114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990114.
pdf.

Tonev:2002:LNQ

[881] Georgi Tonev, Vaidy Sunderam, Roger Loader, and James Pascoe. Location and network quality issues in local area wireless networks. Lecture Notes in Computer Science, 2299:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990131.
pdf.

Wolf:2002:DTE

[882] Tilman Wolf and Mark A. Franklin. Design tradeoffs for embedded network processors. Lecture Notes in Computer Science, 2299:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990149.
pdf.

Wennekers:2002:RRN

[883] Sascha Wennekers and Christian Siemers. Reconfigurable RISC — A new approach for space-efficient superscalar microprocessor architecture. Lecture Notes in Computer Science, 2299:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2299/22990165.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990165.
                                        pdf.

Egan:2002:CTL

[884] Colin Egan, Gordon Steven, and Lucian Vintan. Cached two-level adaptive branch predictors with multiple stages. Lecture Notes in Computer Science, 2299:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990179.
                                        pdf.

Oodes:2002:CAV

[885] Tim Oodes, Holger Krisp, and Christian Müller-Schloer. On the combination of assertions and virtual prototyping for the design of safetycritical systems. Lecture Notes in Computer Science, 2299:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990195.
                                        pdf.

Burchert:2002:UAW

[886] Frank Burchert, Stephan Gatzka, Christian Hochberger, Chang-Kun Lee, Ulrike Lucke, and Djamshid Tavangarian. Ubiquitous access to wide-area high-performance computing. Lecture Notes in Computer Science, 2299:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990209.
pdf.

Muhl:2002:FSC

[887] Gero Mühl, Ludger Fiege, and Alejandro Buchmann. Filter similarities in content-based publish/ subscribe systems. Lecture Notes in Computer Science, 2299:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990224.
pdf.

Feldbusch:2002:BRC

[888] Fridtjof Feldbusch, Alexander Paar, Manuel Odendahl, and Ivan Ivanov. A Bluetooth remote control system. Lecture Notes in Computer Science, 2299: 241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2299/22990241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2299/22990241.
pdf.

Siegemund:2002:RLP

[889] Frank Siegemund and Michael Rohs. Rendezvous layer protocols for bluetoothenabled smart devices. Lecture Notes in Computer Science, 2299:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990256.
                                        pdf.

Gu:2002:RHC

[890] Xiaoyuan Gu, Hannes Hartenstein, and Stefan Fischer. A robust header compression simulator & visualizer. Lecture Notes in Computer Science, 2299: 274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2299/22990274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2299/22990274.
                                        pdf.

Anonymous:2002:AIy

[891] Anonymous. Author index. Lecture Notes in Computer Science, 2299:287– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2299/2299ind.pdf.

Kropatsch:2002:APD

[892] Walter G. Kropatsch. Abstraction pyramids on discrete representations. Lecture Notes in Computer Science, 2301:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2301/23010001.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2301/23010001.
pdf.

Kothe:2002:XTS

[893] Ullrich Köthe. XPMaps and topological segmentation — A unified approach to finite topologies in the plane. Lecture Notes in Computer Science, 2301: 22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010022.
pdf.

Thurmer:2002:C

[894] Grit Thürmer. Curves in ZZ^{n}. Lecture Notes in Computer Science, 2301: 34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010034.
pdf.

Ciria:2002:STS

[895] J. C. Ciria, E. Domínguez, and A. R. Francés. Separation theorems for simplicity 26-surfaces. Lecture Notes in Computer Science, 2301:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010045.
pdf.

Hetroy:2002:TQC

[896] Franck Hétroy and Dominique Attali. Topological quadrangulations of closed triangulated surfaces using the Reeb graph. Lecture Notes in Computer Science, 2301:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010057.
                                        pdf.

Floriani:2002:NMD

[897] Leila De Floriani, Mostefa Mohammed Mesmoudi, Franco Morando, and Enrico Puppo. Non-manifold decomposition in arbitrary dimensions. Lecture Notes in Computer Science, 2301:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010069.
                                        pdf.

Gau:2002:MNS

[898] C. J. Gau and T. Yung Kong. 4D minimal non-simple sets. Lecture Notes in Computer Science, 2301:81– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010081.
                                        pdf.

Brun:2002:RFW

[899] Luc Brun and Walter G. Kropatsch. Receptive fields within the combinatorial pyramid framework. Lecture Notes in Computer Science, 2301:92– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010092.
pdf.

Lohou:2002:NST

[900] Christophe Lohou and Gilles Bertrand. A new 3D 6-subiteration thinning algorithm based on P -simple points. Lecture Notes in Computer Science, 2301: 102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010102.
pdf.

Song:2002:MT

[901] Yuqing Song and Aidong Zhang. Monotonic tree. Lecture Notes in Computer Science, 2301:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010114.
pdf.

Chastel:2002:DIN

[902] S. Chastel, P. Colantoni, and A. Bretto. Displaying image neighborhood hypergraphs line-graphs. Lecture Notes in Computer Science, 2301:124–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010124.
                                        pdf.

Frosini:2002:RBD

[903] A. Frosini and G. Simi. The reconstruction of a bicolored domino tiling from two projections. Lecture Notes in Computer Science, 2301:136– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010136.
                                        pdf.

Bruckstein:2002:DGI

[904] Alfred M. Bruckstein. Digital geometry for image-based metrology. Lecture Notes in Computer Science, 2301: 145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010145.
                                        pdf.

Agnus:2002:TRO

[905] Vincent Agnus and Christian Ronse. Topological reconstruction of occluded objects in video sequences. Lecture Notes in Computer Science, 2301:155– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010155.
pdf.

Crespo:2002:SPC

[906] Jose Crespo, Victor Maojo, José A. Sanandrés, Holger Billhardt, and Alberto Muñoz. On the strong property of connected open-close and close-open filters. Lecture Notes in Computer Science, 2301:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010165.
pdf.

Soille:2002:AAT

[907] Pierre Soille. Advances in the analysis of topographic features on discrete images. Lecture Notes in Computer Science, 2301:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010175.
pdf.

Jonker:2002:MOR

[908] Pieter P. Jonker. Morphological operations in recursive neighbourhoods. Lecture Notes in Computer Science, 2301: 187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010187.
                                        pdf.

Malandain:2002:CDP

[909] Grégoire Malandain and Jean-Daniel Boissonnat. Computing the diameter of a point set. Lecture Notes in Computer Science, 2301:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010197.
                                        pdf.

Ros:2002:SRU

[910] Lluís Ros, Kokichi Sugihara, and Federico Thomas. Shape representation using trihedral mesh projections. Lecture Notes in Computer Science, 2301: 209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010209.
                                        pdf.

Damiand:2002:TMB

[911] Guillaume Damiand and Patrick Resch. Topological map based algorithms for 3D image segmentation. Lecture Notes in Computer Science, 2301: 220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010220.
pdf.

Zunic:2002:CDT

[912] Joviša Žunić. On characterization of discrete triangles by discrete moments. Lecture Notes in Computer Science, 2301:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010232.
pdf.

Sintorn:2002:WDT

[913] Ida-Maria Sintorn and Gunilla Borgefors. Weighted distance transforms for images using elongated voxel grids. Lecture Notes in Computer Science, 2301:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2301/23010244.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2301/23010244.
pdf.

Cortadellas:2002:RNS

[914] Javier Cortadellas, Josep Amat, and Manel Frigola. Robust normalization of shapes. Lecture Notes in Computer Science, 2301:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010255.
                                        pdf.

Lindblad:2002:SAE

[915] Joakim Lindblad and Ingela Nyström. Surface area estimation of digitized 3D objects using local computations. Lecture Notes in Computer Science, 2301: 267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010267.
                                        pdf.

Herman:2002:ATF

[916] Gabor T. Herman. An abstract theoretical foundation of the geometry of digital spaces. Lecture Notes in Computer Science, 2301:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010279.
                                        pdf.

Veelaert:2002:CLS

[917] Peter Veelaert. Concurrency of line segments in uncertain geometry. Lecture Notes in Computer Science, 2301:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010289.
pdf.

Couprie:2002:DO

[918] Michel Couprie, Gilles Bertrand, and Yukiko Kenmochi. Discretization in 2D and 3D orders. Lecture Notes in Computer Science, 2301:301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010301.
pdf.

Andres:2002:DDO

[919] Eric Andres. Defining discrete objects for polygonalization: The standard model. Lecture Notes in Computer Science, 2301:313–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010313.
pdf.

Coeurjolly:2002:VDG

[920] David Coeurjolly. Visibility in discrete geometry: An application to discrete geodesic paths. Lecture Notes in Computer Science, 2301:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010326.
                                        pdf.

Burguet:2002:MSD

[921] Jasmine Burguet and Rémy Malgouyres. Multi-scale discrete surfaces. Lecture Notes in Computer Science, 2301:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2301/23010338.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010338.
                                        pdf.

Sugihara:2002:IMS

[922] Kokichi Sugihara. Invertible Minkowski sum of polygons. Lecture Notes in Computer Science, 2301:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010350.
                                        pdf.

Marchadier:2002:TGW

[923] Jocelyn Marchadier, Didier Arquès, and Sylvain Michelin. Thinning grayscale well-composed images: a new approach for topological coherent image segmentation. Lecture Notes in Computer Science, 2301:360– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010360.
pdf.

Buzer:2002:ILT

[924] Lilian Buzer. An incremental linear time algorithm for digital line and plane recognition using a linear incremental feasibility problem. Lecture Notes in Computer Science, 2301:372– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010372.
pdf.

Starck:2002:RAM

[925] J. Starck, A. Hilton, and J. Illingworth. Reconstruction of animated models from images using constrained deformable surfaces. Lecture Notes in Computer Science, 2301:382–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010382.
pdf.

Balogh:2002:RBM

[926] E. Balogh, A. Kuba, A. Del Lungo, and M. Nivat. Reconstruction of binary matrices from absorbed projections. Lecture Notes in Computer Science, 2301: 392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010392.
                                        pdf.

Mesmoudi:2002:SRA

[927] Mohammed Mostefa Mesmoudi. A simplified recognition algorithm of digital planes pieces. Lecture Notes in Computer Science, 2301:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2301/23010404.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2301/23010404.
                                        pdf.

Carre:2002:RTB

[928] Philippe Carré and Eric Andres. Ridgelet transform based on reveillès discrete lines. Lecture Notes in Computer Science, 2301:417–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/2301/23010417.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2301/23010417.pdf.

Malgouyres:2002:DRM

[929] Rémy Malgouyres. A discrete radiosity method. Lecture Notes in Computer Science, 2301:428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2301/23010428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2301/23010428.
pdf.

Anonymous:2002:AIz

[930] Anonymous. Author index. Lecture Notes in Computer Science, 2301:439– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2301/2301auth.pdf.

Schulte:2002:I

[931] Christian Schulte. 1. introduction. Lecture Notes in Computer Science, 2302: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020001.
pdf.

Schulte:2002:CP

[932] Christian Schulte. 2. constraint programming. Lecture Notes in Computer Science, 2302:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2302/23020009.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2302/23020009.pdf.

Schulte:2002:IOLa

[933] Christian Schulte. 3. introducing Oz Light. Lecture Notes in Computer Science, 2302:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2302/23020015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2302/23020015.
                                        pdf.

Schulte:2002:SS

[934] Christian Schulte. 4. spaces for search. Lecture Notes in Computer Science, 2302:29–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2302/23020029.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2302/23020029.
                                        pdf.

Schulte:2002:SE

[935] Christian Schulte. 5. search engines. Lecture Notes in Computer Science, 2302:45–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2302/23020045.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2302/23020045.
                                        pdf.

Schulte:2002:BSS

[936] Christian Schulte. 6. best-solution search. Lecture Notes in Computer Science, 2302:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2302/23020055.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2302/23020055.pdf.

Schulte:2002:R

[937] Christian Schulte. 7. recomputation. Lecture Notes in Computer Science, 2302:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2302/23020059.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2302/23020059.
pdf.

Schulte:2002:OEV

[938] Christian Schulte. 8. Oz Explorer: Visual search. Lecture Notes in Computer Science, 2302:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020069.
pdf.

Schulte:2002:DS

[939] Christian Schulte. 9. distributed search. Lecture Notes in Computer Science, 2302:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2302/23020079.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2302/23020079.pdf.

Schulte:2002:SC

[940] Christian Schulte. 10. spaces for combinators. Lecture Notes in Computer Science, 2302:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020093.
pdf.

Schulte:2002:CC

[941] Christian Schulte. 11. constraint combinators. Lecture Notes in Computer Science, 2302:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2302/23020105.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2302/23020105.pdf.

Schulte:2002:IOLb

[942] Christian Schulte. 12. implementing Oz Light. Lecture Notes in Computer Science, 2302:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020117.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2302/23020117.
pdf.

Schulte:2002:IS

[943] Christian Schulte. 13. implementing spaces. Lecture Notes in Computer Science, 2302:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020121.
pdf.

Schulte:2002:OAS

[944] Christian Schulte. 14. other approaches to search. Lecture Notes in Computer Science, 2302:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020143.
pdf.

Schulte:2002:C

[945] Christian Schulte. 15. conclusion. Lecture Notes in Computer Science, 2302: 153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2302/23020153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2302/23020153.
pdf.

Schulte:2002:BPP

[946] Christian Schulte. A. benchmark problems and platforms. Lecture Notes in Computer Science, 2302:157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2302/23020157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2302/23020157.
                                        pdf.

Anonymous:2002:Rd

[947] Anonymous. References. Lecture Notes in Computer Science, 2302:161– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2302/2302refs.pdf.

Anonymous:2002:Id

[948] Anonymous. Index. Lecture Notes in Computer Science, 2302:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2302/2302ind.pdf.

Courcelle:2002:SEM

[949] Bruno Courcelle. Semantical evaluations as monadic second-order compatible structure transformations. Lecture Notes in Computer Science, 2303:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2303/23030001.
pdf.

Abraham-Mumm:2002:VJR

[950] Erika Ábrahám-Mumm, Frank S. de Boer, Willem-Paul de Roever, and Martin Steffen. Verification for Java’s reentrant multithreading concept. Lecture Notes in Computer Science, 2303: 5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030005.
pdf.

Bidoit:2002:IOR

[951] Michel Bidoit and Rolf Hennicker. On the integration of observability and reachability concepts. Lecture Notes in Computer Science, 2303:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030021.
pdf.

Boer:2002:PCT

[952] Frank S. de Boer, Maurizio Gabbrielli, and Maria Chiara Meo. Proving correctness of timed concurrent constraint programs. Lecture Notes in Computer Science, 2303:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030037.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030037.
                                        pdf.

Bollig:2002:GRM

[953] Benedikt Bollig, Martin Leucker, and Thomas Noll. Generalised regular MSC languages. Lecture Notes in Computer Science, 2303:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030052.
                                        pdf.

Boreale:2002:CRS

[954] Michele Boreale and Daniele Gorla. On compositional reasoning in the spicalculus. Lecture Notes in Computer Science, 2303:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2303/23030067.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2303/23030067.pdf.

Cirstea:2002:SLA

[955] Corina Cîrstea. On specification logics for algebra-coalgebra structures: Reconciling reachability and observability. Lecture Notes in Computer Science, 2303:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030082.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2303/23030082.
pdf.

Danvy:2002:FOO

[956] Olivier Danvy and Lasse R. Nielsen. A first-order one-pass CPS transformation. Lecture Notes in Computer Science, 2303:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2303/23030098.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2303/23030098.pdf.

Doberkat:2002:DPP

[957] Ernst-Erich Doberkat. The demonic product of probabilistic relations. Lecture Notes in Computer Science, 2303: 114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030114.
pdf.

Ferrari:2002:MTS

[958] Gianluigi Ferrari, Ugo Montanari, and Marco Pistore. Minimizing transition systems for name passing calculi: a co-algebraic formulation. Lecture Notes in Computer Science, 2303:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030129.
pdf.

Fuhrmann:2002:VE

[959] Carsten Führmann. Varieties of effects. Lecture Notes in Computer Science, 2303:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2303/23030144.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030144.
                                        pdf.

Godard:2002:CFG

[960] Emmanuel Godard and Yves Métivier. A characterization of families of graphs in which election is possible. Lecture Notes in Computer Science, 2303:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030159.
                                        pdf.

Jancar:2002:ECO

[961] Petr Jančar, Antonín Kučera, Faron Moller, and Zdenék Sawa. Equivalencechecking with one-counter automata: a generic method for proving lower bounds. Lecture Notes in Computer Science, 2303:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2303/23030172.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2303/23030172.pdf.

Jha:2002:ETM

[962] Somesh Jha, Jens Palsberg, and Tian Zhao. Efficient type matching. Lecture Notes in Computer Science, 2303: 187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030187.
pdf.

Knapik:2002:HOP

[963] Teodor Knapik, Damian Niwinski, and Paweł Urzyczyn. Higher-order pushdown trees are easy. Lecture Notes in Computer Science, 2303:205– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030205.
pdf.

Koch:2002:CDR

[964] Manuel Koch, Luigi V. Mancini, and Francesco Parisi-Presicce. Conflict detection and resolution in access control policy specifications. Lecture Notes in Computer Science, 2303:223– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030223.
pdf.

Kurz:2002:LAF

[965] Alexander Kurz. Logics admitting final semantics. Lecture Notes in Computer Science, 2303:238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030238.
                                        pdf.

Lange:2002:MCF

[966] Martin Lange and Colin Stirling. Model checking fixed point logic with chop. Lecture Notes in Computer Science, 2303:250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2303/23030250.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2303/23030250.pdf.

Laroussinie:2002:MCD

[967] François Laroussinie, Nicolas Markey, and Philippe Schnoebelen. On model checking durational Kripke structures. Lecture Notes in Computer Science, 2303:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2303/23030264.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030264.
                                        pdf.

Loding:2002:MCI

[968] Christof Löding. Model-checking infinite systems generated by ground tree rewriting. Lecture Notes in Computer Science, 2303:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030280.
pdf.

Lohrey:2002:BMC

[969] Markus Lohrey and Anca Muscholl. Bounded MSC communication. Lecture Notes in Computer Science, 2303: 295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030295.
pdf.

Martin:2002:IDC

[970] Keye Martin. The informatic derivative at a compact element. Lecture Notes in Computer Science, 2303:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030310.
pdf.

Mossakowski:2002:HDG

[971] Till Mossakowski. Heterogeneous development graphs and heterogeneous borrowing. Lecture Notes in Computer Science, 2303:326–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030326.
                                        pdf.

Plotkin:2002:NCD

[972] Gordon Plotkin and John Power. Notions of computation determine monads. Lecture Notes in Computer Science, 2303:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2303/23030342.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2303/23030342.pdf.

Santocanale:2002:CCP

[973] Luigi Santocanale. A calculus of circular proofs and its categorical semantics. Lecture Notes in Computer Science, 2303:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2303/23030357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030357.
                                        pdf.

Schopp:2002:VTP

[974] Ulrich Schöpp and Alex Simpson. Verifying temporal properties using explicit approximants: Completeness for context-free processes. Lecture Notes in Computer Science, 2303:372– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030372.
pdf.

Srba:2002:NTT

[975] Jiří Srba. Note on the tableau technique for commutative transition systems. Lecture Notes in Computer Science, 2303:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030387.
pdf.

Yang:2002:SBL

[976] Hongseok Yang and Peter O’Hearn. A semantic basis for local reasoning. Lecture Notes in Computer Science, 2303: 402–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2303/23030402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2303/23030402.
pdf.

Yoshida:2002:LB

[977] Nobuko Yoshida, Kohei Honda, and Martin Berger. Linearity and bisimulation. Lecture Notes in Computer Science, 2303:417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2303/23030417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2303/23030417.
                                        pdf.

Anonymous:2002:AIaa

[978] Anonymous. Author index. Lecture Notes in Computer Science, 2303:435– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2303/2303auth.pdf.

Mernik:2002:LIE

[979] Marjan Mernik, Mitja Lenic, Enis Avdicauševic, and Viljem Zumer. LISA: An interactive environment for programming language development. Lecture Notes in Computer Science, 2304:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2304/23040001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040001.
                                        pdf.

Ertl:2002:BIV

[980] M. Anton Ertl and David Gregg. Building an interpreter with Vmgen. Lecture Notes in Computer Science, 2304: 5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040005.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040005.
                                        pdf.

Garavel:2002:CCU

[981] Hubert Garavel, Frédéric Lang, and Radu Mateescu. Compiler construction using LOTOS NT. Lecture Notes in Computer Science, 2304:9– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040009.
pdf.

Zhang:2002:DCT

[982] Youtao Zhang and Rajiv Gupta. Data compression transformations for dynamically allocated data structures. Lecture Notes in Computer Science, 2304:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2304/23040014.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2304/23040014.
pdf.

Agrawal:2002:EDD

[983] Gagan Agrawal, Jinqian Li, and Qi Su. Evaluating a demand driven technique for call graph construction. Lecture Notes in Computer Science, 2304:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040029.
pdf.

Mohnen:2002:GFA

[984] Markus Mohnen. A graph-free approach to data-flow analysis. Lecture Notes in Computer Science, 2304:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040046.
                                        pdf.

Gupta:2002:RBS

[985] Rajiv Gupta, Eduard Mehofer, and Youtao Zhang. A representation for bit section based analysis and optimization. Lecture Notes in Computer Science, 2304:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040062.
                                        pdf.

Oren:2002:OSP

[986] David Oren, Yossi Matias, and Mooly Sagiv. Online subpath profiling. Lecture Notes in Computer Science, 2304: 78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040078.
                                        pdf.

Gschwind:2002:PES

[987] Michael Gschwind and Erik Altman. Precise exception semantics in dynamic compilation. Lecture Notes in Computer Science, 2304:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040095.
pdf.

Miecznikowski:2002:DJB

[988] Jerome Miecznikowski and Laurie Hendren. Decompiling Java bytecode: Problems, traps and pitfalls. Lecture Notes in Computer Science, 2304:111– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040111.
pdf.

VanWyk:2002:FAG

[989] Eric Van Wyk, Oege de Moor, Kevin Backhouse, and Paul Kwiatkowski. Forwarding in attribute grammars for modular language design. Lecture Notes in Computer Science, 2304:128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040128.
pdf.

Brand:2002:DFS

[990] Mark G. J. van den Brand, Jeroen Scheerder, Jurgen J. Vinju, and Eelco Visser. Disambiguation filters for scannerless generalized LR parsers. Lecture Notes in Computer Science, 2304:143– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040143.
                                        pdf.

Cousot:2002:MSP

[991] Patrick Cousot and Radhia Cousot. Modular static program analysis. Lecture Notes in Computer Science, 2304: 159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040159.
                                        pdf.

Thies:2002:SLS

[992] William Thies, Michal Karczmarek, and Saman Amarasinghe. StreamIt: a language for streaming applications. Lecture Notes in Computer Science, 2304:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2304/23040179.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040179.
                                        pdf.

Henderson:2002:CMH

[993] Fergus Henderson and Zoltan Somogyi. Compiling Mercury to high-level C code. Lecture Notes in Computer Science, 2304:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040197.
pdf.

Necula:2002:CIL

[994] George C. Necula, Scott McPeak, Shree P. Rahul, and Westley Weimer. CIL: Intermediate language and tools for analysis and transformation of C programs. Lecture Notes in Computer Science, 2304:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040213.
pdf.

Mossenbock:2002:LSR

[995] Hanspeter Mössenböck and Michael Pfeiffer. Linear scan register allocation in the context of SSA form and register constraints. Lecture Notes in Computer Science, 2304:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040229.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2304/23040229.
                                        pdf.

Cilio:2002:GVP

[996] Andrea G. M. Cilio and Henk Corporaal. Global variable promotion: Using registers to reduce cache power dissipation. Lecture Notes in Computer Science, 2304:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040247.
                                        pdf.

Rele:2002:OSP

[997] Siddharth Rele, Santosh Pande, Soner Onder, and Rajiv Gupta. Optimizing static power dissipation by functional units in superscalar processors. Lecture Notes in Computer Science, 2304: 261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040261.
                                        pdf.

Kandemir:2002:ILO

[998] Mahmut Kandemir, Ibrahim Kolcu, and Ismail Kadayif. Influence of loop optimizations on energy consumption of multi-bank memory systems. Lecture Notes in Computer Science, 2304: 276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2304/23040276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040276.
pdf.

Mikheev:2002:EEL

[999] Vitaly V. Mikheev, Stanislav A. Fedoseev, Vladimir V. Sukharev, and Nikita V. Lipsky. Effective enhancement of loop versioning in Java. Lecture Notes in Computer Science, 2304: 293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040293.
pdf.

Wu:2002:VPG

[1000] Youfeng Wu, Mauricio Serrano, Rakesh Krishnaiyer, Wei Li, and Jesse Fang. Value-profile guided stride prefetching for irregular code. Lecture Notes in Computer Science, 2304:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2304/23040307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2304/23040307.
pdf.

Qian:2002:CAA

[1001] Feng Qian, Laurie Hendren, and Clark Verbrugge. A comprehensive approach to array bounds check elimination for Java. Lecture Notes in Computer Science, 2304:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2304/23040325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2304/23040325.
                                        pdf.

Anonymous:2002:AIab

[1002] Anonymous. Author index. Lecture Notes in Computer Science, 2304:343– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2304/2304ind.pdf.

Morrisett:2002:TCS

[1003] Greg Morrisett. Type checking systems code. Lecture Notes in Computer Science, 2305:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2305/23050001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2305/23050001.
                                        pdf.

Hirschowitz:2002:MMC

[1004] Tom Hirschowitz and Xavier Leroy. Mixin modules in a call-by-value setting. Lecture Notes in Computer Science, 2305:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2305/23050006.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2305/23050006.pdf.

Grossman:2002:ETI

[1005] Dan Grossman. Existential types for imperative languages. Lecture Notes in Computer Science, 2305:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050021.
pdf.

Aspinall:2002:ATS

[1006] David Aspinall and Martin Hofmann. Another type system for in-place update. Lecture Notes in Computer Science, 2305:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050036.
pdf.

Bistarelli:2002:SCC

[1007] Stefano Bistarelli, Ugo Montanari, and Francesca Rossi. Soft concurrent constraint programming. Lecture Notes in Computer Science, 2305:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050053.
pdf.

McGuire:2002:PLC

[1008] Morgan McGuire, Shriram Krishnamurthi, and John F. Hughes. Programming languages for compressing graphics. Lecture Notes in Computer Science, 2305:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050068.
pdf.

Hu:2002:APS

[1009] Zhenjiang Hu, Hideya Iwasaki, and Masato Takeichi. An accumulative parallel skeleton for all. Lecture Notes in Computer Science, 2305:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050083.
pdf.

Weirich:2002:HOI

[1010] Stephanie Weirich. Higher-order intensional type analysis. Lecture Notes in Computer Science, 2305:98– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050098.
pdf.

Wells:2002:BT

[1011] Joe B. Wells and Christian Haack.

Branching types. Lecture Notes in Computer Science, 2305:115–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050115.
pdf.

Laird:2002:ECM

[1012] James Laird. Exceptions, continuations and macro-expressiveness. Lecture Notes in Computer Science, 2305: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050133.
pdf.

Glew:2002:TSO

[1013] Neal Glew. A theory of second-order trees. Lecture Notes in Computer Science, 2305:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2305/23050147.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2305/23050147.pdf.

Horwitz:2002:TSI

[1014] Susan Horwitz. Tool support for improving test coverage. Lecture Notes in Computer Science, 2305:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050162.
pdf.

Kandemir:2002:DSO

[1015] Mahmut Kandemir. Data space oriented tiling. Lecture Notes in Computer Science, 2305:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050178.
pdf.

Martel:2002:PRE

[1016] Matthieu Martel. Propagation of roundoff errors in finite precision computations: a semantics approach. Lecture Notes in Computer Science, 2305: 194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050194.
pdf.

Goubault:2002:APF

[1017] Eric Goubault, Matthieu Martel, and Sylvie Putot. Asserting the precision of floating-point computations: a simple abstract interpreter. Lecture Notes in Computer Science, 2305:209– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050209.
pdf.

Wand:2002:MEP

[1018] Mitchell Wand and Galen B. Williamson.

A modular, extensible proof method for small-step flow analyses. Lecture Notes in Computer Science, 2305:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050213.
pdf.

Thiemann:2002:PDC

[1019] Peter Thiemann. A prototype dependency calculus. Lecture Notes in Computer Science, 2305:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050228.
pdf.

Nielson:2002:ACA

[1020] Flemming Nielson, Hanne Riis Nielson, and Helmut Seidl. Automatic complexity analysis. Lecture Notes in Computer Science, 2305:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2305/23050243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2305/23050243.
                                        pdf.

Flanagan:2002:TMV

[1021] Cormac Flanagan, Stephen N. Freund, and Shaz Qadeer. Threadmodular verification for sharedmemory programs. Lecture Notes in Computer Science, 2305:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2305/23050262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2305/23050262.
                                        pdf.

Wansbrough:2002:TUM

[1022] Keith Wansbrough, Michael Norrish, Peter Sewell, and Andrei Serjantov. Timing UDP: Mechanized semantics for sockets, threads, and failures. Lecture Notes in Computer Science, 2305: 278–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2305/23050278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2305/23050278.
                                        pdf.

Charatonik:2002:FCM

[1023] Witold Charatonik, Andrew D. Gordon, and Jean-Marc Talbot. Finitecontrol mobile ambients. Lecture Notes in Computer Science, 2305:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050295.
pdf.

Feret:2002:DAM

[1024] Jérôme Feret. Dependency analysis of mobile systems. Lecture Notes in Computer Science, 2305:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2305/23050314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2305/23050314.
pdf.

Anonymous:2002:AIac

[1025] Anonymous. Author index. Lecture Notes in Computer Science, 2305:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2305/2305auth.pdf.

Charpentier:2002:ACM

[1026] Michel Charpentier. An approach to composition motivated by wp. Lecture Notes in Computer Science, 2306: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060001.
pdf.

Barthe:2002:CVS

[1027] Gilles Barthe, Dilian Gurov, and Marieke Huisman. Compositional verification of secure applet interactions. Lecture Notes in Computer Science, 2306:15–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2306/23060015.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2306/23060015.
pdf.

Ehrig:2002:GCF

[1028] Hartmut Ehrig, Fernando Orejas, Benjamin Braatz, Markus Klein, and Martti Piirainen. A generic component framework for system modeling. Lecture Notes in Computer Science, 2306: 33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060033.
pdf.

Cerone:2002:ICE

[1029] Antonio Cerone. Implementing condition/event nets in the circal process algebra. Lecture Notes in Computer Science, 2306:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060049.
pdf.

Xie:2002:ISS

[1030] Fei Xie and James C. Browne. Integrated state space reduction for model checking executable object-oriented software system designs. Lecture Notes in Computer Science, 2306:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060064.
pdf.

Hagerer:2002:MGM

[1031] Andreas Hagerer, Hardi Hungar, Oliver Niese, and Bernhard Steffen. Model generation by moderated regular extrapolation. Lecture Notes in Computer Science, 2306:80– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060080.
pdf.

Kumar:2002:BSP

[1032] Sumit Kumar and Susan Horwitz.

Better slicing of programs with jumps and switches. Lecture Notes in Computer Science, 2306:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060096.
pdf.

Bernardo:2002:ATR

[1033] Marco Bernardo and Francesco Franzè. Architectural types revisited: Extensible and/or connections. Lecture Notes in Computer Science, 2306:113– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2306/23060113.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2306/23060113.
                                        pdf.

Paula:2002:MAC

[1034] Virgínia C. C. de Paula and Thais V. Batista. Mapping an ADL to a component-based application development environment. Lecture Notes in Computer Science, 2306:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2306/23060128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2306/23060128.
                                        pdf.

Belaunde:2002:ECC

[1035] Mariano Belaunde and Mikael Peltier. From EDOC components to CCM components: a precise mapping specification. Lecture Notes in Computer Science, 2306:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2306/23060143.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2306/23060143.pdf.

Clark:2002:EML

[1036] Tony Clark, Andy Evans, and Stuart Kent. Engineering modelling languages: a precise metamodelling approach. Lecture Notes in Computer Science, 2306:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060159.
pdf.

Lara:2002:ATM

[1037] Juan de Lara and Hans Vangheluwe.

AToM^{3}: a tool for multi-formalism and meta-modelling. Lecture Notes in Computer Science, 2306:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060174.
pdf.

Baresi:2002:TAV

[1038] Luciano Baresi and Mauro Pezzè. A toolbox for automating visual software engineering. Lecture Notes in Computer Science, 2306:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060189.
pdf.

Bradfield:2002:EOU

[1039] Julian Bradfield, Juliana Küster Filipe, and Perdita Stevens. Enriching OCL using observational mu-calculus. Lecture Notes in Computer Science, 2306: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060203.
pdf.

David:2002:FVU

[1040] Alexandre David, M. Oliver Möller, and Wang Yi. Formal verification of UML statecharts with realtime extensions. Lecture Notes in Computer Science, 2306:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060218.
pdf.

Hahnle:2002:ATI

[1041] Reiner Hähnle, Kristofer Johannisson, and Aarne Ranta. An authoring tool for informal and formal requirements specifications. Lecture Notes in Computer Science, 2306:233– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060233.
pdf.

Bubl:2002:ICB

[1042] Felix Bübl. Introducing contextbased constraints. Lecture Notes in Computer Science, 2306:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060249.
pdf.

Nonnengart:2002:FRE

[1043] Andreas Nonnengart, Georg Rock, and Werner Stephan. Formal requirements engineering using observer models. Lecture Notes in Computer Science, 2306: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060264.
pdf.

Lopez:2002:AGU

[1044] Oscar López, Miguel A. Laguna, and Francisco J. García. Automatic generation of use cases from workflows: a petri net based approach. Lecture Notes in Computer Science, 2306:279– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060279.
pdf.

Baresi:2002:MMT

[1045] Luciano Baresi, Franca Garzotto, Luca Mainetti, and Paolo Paolini. Metamodeling techniques meet Web application design tools. Lecture Notes in Computer Science, 2306:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060294.
pdf.

Navarro:2002:FDC

[1046] Antonio Navarro, Baltasar Fernandez-Manjon, Alfredo Fernandez-Valmayor, and Jose Luis Sierra. Formal-driven conceptualization and prototyping of hypermedia applications. Lecture Notes in Computer Science, 2306:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060308.
pdf.

Gouveia:2002:CDE

[1047] João Gouveia, Georgios Koutsoukos, Michel Wermelinger, Luís Andrade, and José Luiz Fiadeiro. The coordination development environment. Lecture Notes in Computer Science, 2306: 323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060323.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2306/23060323.
pdf.

Ahrendt:2002:KSI

[1048] Wolfgang Ahrendt, Thomas Baar, Bernhard Beckert, Martin Giese, Elmar Habermalz, Reiner Hähnle, Wolfram Menzel, Wojciech Mostowski, and Peter H. Schmitt. The KEY system: Integrating object-oriented design and formal methods. Lecture Notes in Computer Science, 2306:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060327.
pdf.

Xie:2002:OMC

[1049] Fei Xie, Vladimir Levin, and James C.

Browne. ObjectCheck: a model checking tool for executable object-oriented software system designs. Lecture Notes in Computer Science, 2306:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2306/23060331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2306/23060331.
pdf.

Hagerer:2002:DOP

[1050] Andreas Hagerer, Hardi Hungar, Tiziana Margaria, Oliver Niese, Bernhard Steffen, and Hans-Dieter Ide. Demonstration of an operational procedure for the model-based testing of CTI systems. Lecture Notes in Computer Science, 2306:336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2306/23060336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2306/23060336.
                                        pdf.

Anonymous:2002:AIad

[1051] Anonymous. Author index. Lecture Notes in Computer Science, 2306:341– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2306/23060341.pdf.

Zhang:2002:I

[1052] Chengqi Zhang and Shichao Zhang. 1. introduction. Lecture Notes in Computer Science, 2307:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2307/23070001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2307/23070001.
                                        pdf.

Zhang:2002:AR

[1053] Chengqi Zhang and Shichao Zhang. 2. association rule. Lecture Notes in Computer Science, 2307:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2307/23070025.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2307/23070025.
pdf.

Zhang:2002:NAR

[1054] Chengqi Zhang and Shichao Zhang.

3. negative association rule. Lecture Notes in Computer Science, 2307:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2307/23070047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2307/23070047.
pdf.

Zhang:2002:CD

[1055] Chengqi Zhang and Shichao Zhang.

4. causality in databases. Lecture Notes in Computer Science, 2307:85– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2307/23070085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2307/23070085.
pdf.

Zhang:2002:CRA

[1056] Chengqi Zhang and Shichao Zhang. 5.

causal rule analysis. Lecture Notes in Computer Science, 2307:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2307/23070121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2307/23070121.
pdf.

Zhang:2002:ARV

[1057] Chengqi Zhang and Shichao Zhang. 6. association rules in very large databases. Lecture Notes in Computer Science, 2307:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2307/23070161.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2307/23070161.pdf.

Zhang:2002:ARS

[1058] Chengqi Zhang and Shichao Zhang. 7. association rules in small databases. Lecture Notes in Computer Science, 2307:199–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2307/23070199.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2307/23070199.
                                        pdf.

Zhang:2002:CFW

[1059] Chengqi Zhang and Shichao Zhang. 8. conclusion and future work. Lecture Notes in Computer Science, 2307:225– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2307/23070225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2307/23070225.
                                        pdf.

Zhang:2002:R

[1060] Chengqi Zhang and Shichao Zhang. References. Lecture Notes in Computer Science, 2307:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2307/23070229.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2307/23070229.pdf.

Anonymous:2002:SI

[1061] Anonymous. Subject index. Lecture Notes in Computer Science, 2307:237– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2307/2307indx.pdf.

Papadimitriou:2002:UI

[1062] Christos H. Papadimitriou. Understanding the Internet. Lecture Notes in Computer Science, 2308:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080001.
pdf.

Mylopoulos:2002:AOS

[1063] John Mylopoulos, Manuel Kolp, and Paolo Giorgini. Agent-oriented software development. Lecture Notes in Computer Science, 2308:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080003.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2308/23080003.
pdf.

Papadakis:2002:RQP

[1064] Nick Papadakis and Dimitris Plexousakis. The ramification and qualification problems in temporal databases. Lecture Notes in Computer Science, 2308:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2308/23080018.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2308/23080018.
pdf.

Hatzilygeroudis:2002:MIM

[1065] Ioannis Hatzilygeroudis and Jim Prentzas. Multi-inference with multineurules. Lecture Notes in Computer Science, 2308:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2308/23080030.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2308/23080030.pdf.

Stamelos:2002:DMB

[1066] Ioannis Stamelos and Ioannis Refanidis. Decision making based on past problem cases. Lecture Notes in Computer Science, 2308:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080042.
pdf.

Antoniou:2002:RDL

[1067] George Antoniou. Relating defeasible logic to extended logic programs. Lecture Notes in Computer Science, 2308: 54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080054.
pdf.

Stergiou:2002:ADC

[1068] Kostas Stergiou. On algorithms for decomposable constraints. Lecture Notes in Computer Science, 2308:65– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080065.
pdf.

Vlahavas:2002:CSC

[1069] Ioannis P. Vlahavas, Ilias Sakellariou, Ivan Futo, Zoltan Pasztor, and Janos Szeredi. C SPCONS: a Communicating Sequential Prolog with constraints. Lecture Notes in Computer Science, 2308:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080072.
pdf.

Panayiotopoulos:2002:GES

[1070] Themistoklis Panayiotopoulos, Harry Kalogirou, Anthony Petropoulos, and Dionisis Dimopoulos. Genetic evolution of software microorganisms. Lecture Notes in Computer Science, 2308: 85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080085.
pdf.

Tsamardinos:2002:PAR

[1071] Ioannis Tsamardinos. A probabilistic approach to robust execution of temporal plans with uncertainty. Lecture Notes in Computer Science, 2308:97– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080097.
pdf.

Kornilakis:2002:CPO

[1072] Harry Kornilakis and Panagiotis Stamatopoulos. Crew pairing optimization with genetic algorithms. Lecture Notes in Computer Science, 2308:109– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080109.
pdf.

Zavlangas:2002:ITM

[1073] Panagiotis G. Zavlangas and Spyros G. Tzafestas. Integration of topological and metric maps for indoor mobile robot path planning and navigation. Lecture Notes in Computer Science, 2308:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080121.
pdf.

Androutsopoulos:2002:SAM

[1074] Ion Androutsopoulos, Dimitris Spiliotopoulos, Konstantinos Stamatakis, Aggeliki Dimitromanolaki, Vangelis Karkaletsis, and Constantine D. Spyropoulos. Symbolic authoring for multilingual natural language generation. Lecture Notes in Computer Science, 2308:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080131.
pdf.

Dabiri:2002:USS

[1075] Gloria Dabiri, Michael Brown, Maria Aretoulaki, and Matthias Nitzsche. A user-sensitive spoken dialogue system incorporating emotional responsiveness. Lecture Notes in Computer Science, 2308:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2308/23080143.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2308/23080143.pdf.

Karberis:2002:TST

[1076] Georgios Karberis and Georgios Kouroupetroglou. Transforming spontaneous telegraphic language to wellformed Greek sentences for alternative and augmentative communication. Lecture Notes in Computer Science, 2308: 155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080155.
                                        pdf.

Sigletos:2002:RIF

[1077] Georgios Sigletos, Georgios Paliouras, and Vangelis Karkaletsis. Role identification from free text using hidden Markov models. Lecture Notes in Computer Science, 2308:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080167.
                                        pdf.

Maragoudakis:2002:ISU

[1078] Manolis Maragoudakis, Nikolaos K. Tselios, Nikolaos Fakotakis, and Nikolaos M. Avouris. Improving SMS usability using Bayesian networks. Lecture Notes in Computer Science, 2308: 179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080179.
pdf.

Grigoriadou:2002:FIS

[1079] Maria Grigoriadou, Harry Kornilakis, Kyparisia A. Papanikolaou, and George D. Magoulas. Fuzzy inference for student diagnosis in adaptive educational hypermedia. Lecture Notes in Computer Science, 2308:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080191.
pdf.

Vassilas:2002:MGS

[1080] Nikolaos Vassilas, George Miaoulis, Dionysios Chronopoulos, Elias Konstantinidis, Ioanna Ravani, Dimitrios Makris, and Dimitri Plemenos. MultiCAD-GA: a system for the design of 3D forms based on genetic algorithms and human evaluation. Lecture Notes in Computer Science, 2308: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080203.
pdf.

Avrithis:2002:ISA

[1081] Yannis Avrithis, Giorgos Stamou, Anastasios Delopoulos, and Stefanos Kollias. Intelligent semantic access to audiovisual content. Lecture Notes in Computer Science, 2308:215–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080215.
                                        pdf.

Frossyniotis:2002:MCF

[1082] Dimitrios Frossyniotis, Minas Pertselakis, and Andreas Stafylopatis. A multi-clustering fusion algorithm. Lecture Notes in Computer Science, 2308: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080225.
                                        pdf.

Potamias:2002:DFB

[1083] George Potamias. Distance and feature-based clustering of time series: An application on neurophysiology. Lecture Notes in Computer Science, 2308:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2308/23080237.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080237.
                                        pdf.

Lagoudakis:2002:LSM

[1084] Michail G. Lagoudakis, Ronald Parr, and Michael L. Littman. Leastsquares methods in reinforcement learning for control. Lecture Notes in Computer Science, 2308:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080249.
pdf.

Koundourakis:2002:ARE

[1085] George Koundourakis and Babis Theodoulidis. Association rules & evolution in time. Lecture Notes in Computer Science, 2308:261–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080261.
pdf.

Halkidi:2002:MUQ

[1086] Maria Halkidi and Michalis Vazirgiannis. Managing uncertainty and quality in the classification process. Lecture Notes in Computer Science, 2308: 273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080273.
pdf.

Kopanas:2002:RDK

[1087] Ioannis Kopanas, Nikolaos M. Avouris, and Sophia Daskalaki. The role of domain knowledge in a large scale data mining project. Lecture Notes in Computer Science, 2308:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080288.
                                        pdf.

Neocleous:2002:ANN

[1088] Costas Neocleous and Christos Schizas. Artificial neural network learning: a comparative review. Lecture Notes in Computer Science, 2308:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080300.
                                        pdf.

Tsoulos:2002:PNN

[1089] Ioannis G. Tsoulos, Isaac E. Lagaris, and Aristidis C. Likas. Piecewise neural networks for function approximation, cast in a form suitable for parallel computation. Lecture Notes in Computer Science, 2308:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080314.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2308/23080314.
pdf.

Douligeris:2002:UHN

[1090] Christos Douligeris and Gang Feng.

Using Hopfield networks to solve assignment problem and N-queen problem: An application of guided trial and error technique. Lecture Notes in Computer Science, 2308:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080325.
pdf.

Constantinopoulos:2002:BRM

[1091] Constantinos Constantinopoulos, Michalis K.

Titsias, and Aristidis Likas. A Bayesian regularization method for the probabilistic RBF network. Lecture Notes in Computer Science, 2308:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080337.
pdf.

Evgeniou:2002:SVM

[1092] Theodoros Evgeniou and Massimiliano Pontil. Support vector machines with clustering for training with very large datasets. Lecture Notes in Computer Science, 2308:346– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080346.
                                        pdf.

Gordan:2002:TNS

[1093] Mihaela Gordan, Constantine Kotropoulos, and Ioannis Pitas. A temporal network of support vector machine classifiers for the recognition of visual speech. Lecture Notes in Computer Science, 2308:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080355.
                                        pdf.

Rigatos:2002:FSA

[1094] Gerasimos G. Rigatos. Fuzzy stochastic automata for reactive learning and hybrid control. Lecture Notes in Computer Science, 2308:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080366.
                                        pdf.

Tsihrintzis:2002:OWP

[1095] George A. Tsihrintzis and Konstantinos G. Girtis. Overview of wave probebased high-resolution subsurface sensing, imaging, and vision. Lecture Notes in Computer Science, 2308:378– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080378.
pdf.

Krinidis:2002:VRS

[1096] Stelios Krinidis, Christophoros Nikou, and Ioannis Pitas. 3D volume reconstruction by serially acquired 2D slices using a distance transform-based global cost function. Lecture Notes in Computer Science, 2308:390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080390.
pdf.

Kosmopoulos:2002:DEV

[1097] Dimitrios I. Kosmopoulos and Konstantinos V. Chandrinos. Definition and extraction of visual landmarks for indoor robot navigation. Lecture Notes in Computer Science, 2308:401– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080401.
pdf.

Gasteratos:2002:FAA

[1098] Antonios Gasteratos and Giulio Sandini. Factors affecting the accuracy of an active vision head. Lecture Notes in Computer Science, 2308:413– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080413.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080413.
                                        pdf.

Tzitzikas:2002:QTM

[1099] Yannis Tzitzikas, Nicolas Spyratos, and Panos Constantopoulos. Query translation for mediators over ontologybased information sources. Lecture Notes in Computer Science, 2308:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080423.
                                        pdf.

Bassiliades:2002:IQW

[1100] Nick Bassiliades and Ioannis Vlahavas. Intelligent querying of Web documents using a deductive XML repository. Lecture Notes in Computer Science, 2308: 437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080437.
                                        pdf.

Partsakoulakis:2002:RCA

[1101] Ioannis Partsakoulakis and George Vouros. Roles in collaborative activity. Lecture Notes in Computer Science, 2308:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2308/23080449.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2308/23080449.
pdf.

Kefalas:2002:FMR

[1102] Petros Kefalas. Formal modelling of reactive agents as an aggregation of simple behaviours. Lecture Notes in Computer Science, 2308:461– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080461.
pdf.

Georgilakis:2002:AAI

[1103] Pavlos Georgilakis and Nikos Hatziargyriou. On the application of artificial intelligence techniques to the quality improvement of industrial processes. Lecture Notes in Computer Science, 2308:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2308/23080473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2308/23080473.
pdf.

Dimopoulos:2002:UNU

[1104] Michael Dimopoulos and Panagiotis Linardis. Using non-uniform crossover in genetic algorithm methods to speed up the generation of test patterns for sequential circuits. Lecture Notes in Computer Science, 2308:485–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080485.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080485.
                                        pdf.

Tsakonas:2002:HCI

[1105] Athanasios Tsakonas and George Dounias. Hybrid computational intelligence schemes in complex domains: An extended review. Lecture Notes in Computer Science, 2308:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2308/23080494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2308/23080494.
                                        pdf.

Anonymous:2002:AIae

[1106] Anonymous. Author index. Lecture Notes in Computer Science, 2308:513– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2308/2308auth.pdf.

Nelson:2002:FCB

[1107] Greg Nelson. Foundations of a constraint-based illustrator. Lecture Notes in Computer Science, 2309:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2309/23090001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090001.
pdf.

Autexier:2002:IHO

[1108] Serge Autexier and Till Mossakowski.

Integrating Hol-Casl into the development graph manager Maya. Lecture Notes in Computer Science, 2309: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090002.
pdf.

Luth:2002:MM

[1109] Christoph Lüth and Neil Ghani. Monads and modularity. Lecture Notes in Computer Science, 2309:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090018.
pdf.

Marte:2002:MAP

[1110] Michael Marte. A modular approach to proving confluence. Lecture Notes in Computer Science, 2309:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090033.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2309/23090033.
                                        pdf.

Cimatti:2002:IBB

[1111] Alessandro Cimatti, Enrico Giunchiglia, Marco Pistore, Marco Roveri, Roberto Sebastiani, and Armando Tacchella. Integrating BDD-based and SAT-based symbolic model checking. Lecture Notes in Computer Science, 2309:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2309/23090049.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090049.
                                        pdf.

Yavuz-Kahveci:2002:HEM

[1112] Tuba Yavuz-Kahveci and Tevfik Bultan. Heuristics for efficient manipulation of composite constraints. Lecture Notes in Computer Science, 2309:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2309/23090057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090057.
                                        pdf.

Delzanno:2002:CBM

[1113] Giorgio Delzanno. Constraint-based model checking for parameterized synchronous systems. Lecture Notes in Computer Science, 2309:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2309/23090072.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2309/23090072.
pdf.

Kapur:2002:RRB

[1114] Deepak Kapur. A rewrite rule based framework for combining decision procedures. Lecture Notes in Computer Science, 2309:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090087.
pdf.

Zarba:2002:CSI

[1115] Calogero G. Zarba. Combining sets with integers. Lecture Notes in Computer Science, 2309:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090103.
pdf.

Ceberio:2002:SNE

[1116] Martine Ceberio and Laurent Granvilliers. Solving nonlinear equations by abstraction, Gaussian elimination, and interval methods. Lecture Notes in Computer Science, 2309:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090117.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2309/23090117.
                                        pdf.

Barrett:2002:GSM

[1117] Clark W. Barrett, David L. Dill, and Aaron Stump. A generalization of Shostak’s method for combining decision procedures. Lecture Notes in Computer Science, 2309:132– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2309/23090132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090132.
                                        pdf.

Cadoli:2002:CRA

[1118] Marco Cadoli and Toni Mancini. Combining relational algebra, SQL, and constraint programming. Lecture Notes in Computer Science, 2309:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2309/23090147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090147.
                                        pdf.

Balbiani:2002:CCP

[1119] Philippe Balbiani and Jean-François Condotta. Computational complexity of propositional linear temporal logics based on qualitative spatial or temporal reasoning. Lecture Notes in Computer Science, 2309:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2309/23090162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090162.
pdf.

Gavanelli:2002:ECD

[1120] Marco Gavanelli, Evelina Lamma, Paola Mello, and Michela Milano. Exploiting constraints for domain managing in CLP(FD). Lecture Notes in Computer Science, 2309:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090177.
pdf.

Fruehwirth:2002:TRA

[1121] Thom Fruehwirth and Slim Abdennadher. Tutorial: Reasoning with, about and for constraint handling rules. Lecture Notes in Computer Science, 2309: 192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2309/23090192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2309/23090192.
pdf.

Melham:2002:PRI

[1122] T. F. Melham. P ROSPER an investigation into software architecture for embedded proof engines. Lecture Notes in Computer Science, 2309:193– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2309/23090193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090193.
                                        pdf.

Holzl:2002:CLC

[1123] Matthias Hölzl and John N. Crossley. Constraint-lambda calculi. Lecture Notes in Computer Science, 2309: 207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2309/23090207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090207.
                                        pdf.

Rasga:2002:LDA

[1124] João Rasga, Amílcar Sernadas, Cristina Sernadas, and Luca Viganò. Labelled deduction over algebras of truth-values. Lecture Notes in Computer Science, 2309:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2309/23090222.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2309/23090222.
                                        pdf.

Burrieza:2002:TMA

[1125] Alfredo Burrieza and Inma P. de Guzmán. A temporal × modal approach to the definability of properties of functions. Lecture Notes in Computer Science, 2309:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2309/23090239.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2309/23090239.
pdf.

Anonymous:2002:AIaf

[1126] Anonymous. Author index. Lecture Notes in Computer Science, 2309:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2309/2309auth.pdf.

Bentley:2002:WBC

[1127] Peter J. Bentley. Why biologists and computer scientists should work together. Lecture Notes in Computer Science, 2310:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100003.
pdf.

Bienvenue:2002:NMC

[1128] Alexis Bienvenüe, Marc Joannides, Jean Bérard, ÉricFontenas, and Olivier François. Niching in Monte Carlo filtering algorithms. Lecture Notes in Computer Science, 2310:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100019.
pdf.

Morrison:2002:MPD

[1129] Ronald W. Morrison and Kenneth A. De Jong. Measurement of population diversity. Lecture Notes in Computer Science, 2310:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100031.
pdf.

Cerruti:2002:PBS

[1130] Umberto Cerruti, Mario Giacobini, and Pierre Liardet. Prediction of binary sequences by evolving finite state machines. Lecture Notes in Computer Science, 2310:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100042.
pdf.

Berny:2002:ESL

[1131] Arnaud Berny. Extending selection learning toward fixed-length d-ary strings. Lecture Notes in Computer Science, 2310:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2310/23100054.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2310/23100054.pdf.

Brown:2002:MRF

[1132] D. F. Brown, A. B. Garmendia-Doval, and J. A. W. McCall. Markov random field modelling of royal road genetic algorithms. Lecture Notes in Computer Science, 2310:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100065.
pdf.

Sidaner:2002:MSD

[1133] Alain Sidaner, Olivier Bailleux, and Jean-Jacques Chabrier. Measuring the spatial dispersion of evolutionary search processes: Application to walksat. Lecture Notes in Computer Science, 2310:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100077.
pdf.

Johnson:2002:ISM

[1134] Andrew Johnson and Jonathan Shapiro.

The importance of selection mechanisms in distribution estimation algorithms. Lecture Notes in Computer Science, 2310:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2310/23100091.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2310/23100091.pdf.

Abboud:2002:SDM

[1135] K. Abboud and Marc Schoenauer. Surrogate deterministic mutation: Preliminary results. Lecture Notes in Computer Science, 2310:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100104.
                                        pdf.

Tendresse:2002:EPR

[1136] Ingo la Tendresse, Jens Gottlieb, and Odej Kao. The effects of partial restarts in evolutionary search. Lecture Notes in Computer Science, 2310: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100117.
                                        pdf.

Leblanc:2002:HIE

[1137] Benoit Leblanc, Evelyne Lutton, Bertrand Braunschweig, and Hervé Toulhoat. History and immortality in evolutionary computation. Lecture Notes in Computer Science, 2310:128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100128.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2310/23100128.
pdf.

Oudeyer:2002:OLS

[1138] Pierre yves Oudeyer. Origins and learnability of syllable systems: a cultural evolutionary model. Lecture Notes in Computer Science, 2310:143– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100143.
pdf.

Korczak:2002:ESP

[1139] Jerzy J. Korczak, Piotr Lipiński, and Patrick Roger. Evolution strategy in portfolio optimization. Lecture Notes in Computer Science, 2310:156– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100156.
pdf.

Hamiez:2002:SSG

[1140] Jean-Philippe Hamiez and Jin-Kao Hao. Scatter search for graph coloring. Lecture Notes in Computer Science, 2310:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2310/23100168.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2310/23100168.
pdf.

Bousonville:2002:TSC

[1141] Thomas Bousonville. The two stage continuous parallel flow shop problem with limited storage: Modeling and algorithms. Lecture Notes in Computer Science, 2310:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100180.
pdf.

Belaidouni:2002:SLS

[1142] Mériéma Bélaidouni and Jin-Kao Hao. SAT, local search dynamics and density of states. Lecture Notes in Computer Science, 2310:192– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100192.
pdf.

Rudenko:2002:MEA

[1143] Olga Rudenko, Marc Schoenauer, Tiziana Bosio, and Roberto Fontana. A multiobjective evolutionary algorithm for car front end design. Lecture Notes in Computer Science, 2310: 205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100205.
pdf.

Lutton:2002:ECT

[1144] Evelyne Lutton, Pierre Collet, and Jean Louchet. EASEA comparisons on test functions: GALib versus EO. Lecture Notes in Computer Science, 2310: 219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100219.
pdf.

Keijzer:2002:EOG

[1145] M. Keijzer, J. J. Merelo, G. Romero, and Marc Schoenauer. Evolving objects: a general purpose evolutionary computation library. Lecture Notes in Computer Science, 2310:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100231.
pdf.

Robilliard:2002:BOC

[1146] Denis Robilliard and Cyril Fonlupt.

Backwarding : An overfitting control for genetic programming in a remote sensing application. Lecture Notes in Computer Science, 2310:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100245.
pdf.

Ratle:2002:ABS

[1147] Alain Ratle and Michèle Sebag. Avoiding the bloat with stochastic grammarbased genetic programming. Lecture Notes in Computer Science, 2310:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100255.
                                        pdf.

Paris:2002:ABT

[1148] Gregory Paris, Denis Robilliard, and Cyril Fonlupt. Applying boosting techniques to genetic programming. Lecture Notes in Computer Science, 2310: 267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100267.
                                        pdf.

Riche:2002:DEO

[1149] Rodolphe Le Riche and Frédéric Guyon. Dual evolutionary optimization. Lecture Notes in Computer Science, 2310:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2310/23100281.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2310/23100281.pdf.

Smith:2002:UEA

[1150] Stephen Smith. Using evolutionary algorithms incorporating the augmented Lagrangian penalty function to solve discrete and continuous constrained non-linear optimal control problems. Lecture Notes in Computer Science, 2310:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100295.
pdf.

Casillas:2002:CCL

[1151] Jorge Casillas, O. Cordón, F. Herrera, and J. J. Merelo. Cooperative coevolution for learning fuzzy rule-based systems. Lecture Notes in Computer Science, 2310:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2310/23100311.
pdf.

Srivastava:2002:ECE

[1152] Ravi Srivastava and Amit Kaldate.

Evolving cooperative ecosystems: a multi-agent simulation of deforestation activities. Lecture Notes in Computer Science, 2310:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2310/23100323.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100323.
                                        pdf.

Edmonds:2002:IES

[1153] Ian R. Edmonds. The impact of environmental structure on the evolutionary trajectories of a foraging agent. Lecture Notes in Computer Science, 2310:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2310/23100338.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100338.
                                        pdf.

Delepoulle:2002:LCS

[1154] Samuel Delepoulle, Philippe Preux, and Jean-Claude Darcheville. Learning as a consequence of selection. Lecture Notes in Computer Science, 2310: 350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2310/23100350.
                                        pdf.

Seredynski:2002:CEP

[1155] Franciszek Seredynski and Albert Y. Zomaya. Coevolution and evolving parallel cellular automata — based scheduling algorithms. Lecture Notes in Computer Science, 2310:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2310/23100362.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2310/23100362.
pdf.

Anonymous:2002:AIag

[1156] Anonymous. Author index. Lecture Notes in Computer Science, 2310:375– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2310/2310auth.pdf.

Rossides:2002:OFR

[1157] L. Rossides, C. Chrysostomou, A. Pitsillides, and A. Sekercioglu. Overview of fuzzy-RED in diff-serv networks. Lecture Notes in Computer Science, 2311:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2311/23110001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2311/23110001.
pdf.

Chieng:2002:AAE

[1158] David Chieng, Ivan Ho, Alan Marshall, and Gerard Parr. An architecture for agent-enhanced network service provisioning through SLA negotiation. Lecture Notes in Computer Science, 2311: 14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110014.
pdf.

Sterritt:2002:FFM

[1159] Roy Sterritt. Facing fault management as it is, aiming for what you would like it to be. Lecture Notes in Computer Science, 2311:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110031.
pdf.

Bontempi:2002:EMQ

[1160] Gianluca Bontempi and Gauthier Lafruit. Enabling multimedia QoS control with black-box modelling. Lecture Notes in Computer Science, 2311:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110046.
pdf.

Zhu:2002:UMC

[1161] Jianhan Zhu, Jun Hong, and John G. Hughes. Using Markov chains for link prediction in adaptive Web sites. Lecture Notes in Computer Science, 2311: 60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110060.
pdf.

Black:2002:CCC

[1162] Michaela Black and Ray Hickey. Classification of customer call data in the presence of concept drift and noise. Lecture Notes in Computer Science, 2311:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2311/23110074.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2311/23110074.
pdf.

Zelezny:2002:LSD

[1163] Filip Železný, Jiří Zídgek, and Olga Štěpánková. A learning system for decision support in telecommunications. Lecture Notes in Computer Science, 2311:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110088.
pdf.

Martin:2002:AUM

[1164] Trevor P. Martin and Benham Azvine.

Adaptive user modelling in an intelligent telephone assistant. Lecture Notes in Computer Science, 2311:102– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110102.
pdf.

Haenni:2002:QDA

[1165] Rolf Haenni. A query-driven anytime algorithm for argumentative and abductive reasoning. Lecture Notes in Computer Science, 2311:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110114.
                                        pdf.

Richard:2002:PLU

[1166] Gilles Richard and Fatima Zohra Kettaf. Proof length as an uncertainty factor in ILP. Lecture Notes in Computer Science, 2311:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110128.
                                        pdf.

Bagai:2002:POO

[1167] Rajiv Bagai and Shellene J. Kelley. Paraconsistency in objectoriented databases. Lecture Notes in Computer Science, 2311:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110141.
                                        pdf.

Lukacs:2002:DSI

[1168] Gergely Lukács. Decision support with imprecise data for consumers. Lecture Notes in Computer Science, 2311:151– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110151.
pdf.

Golubski:2002:GPP

[1169] Wolfgang Golubski. Genetic programming: a parallel approach. Lecture Notes in Computer Science, 2311:166– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110166.
pdf.

Lehman:2002:SU

[1170] Manny M. Lehman and J. F. Ramil.

Software uncertainty. Lecture Notes in Computer Science, 2311:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110174.
pdf.

McClean:2002:TPC

[1171] Sally McClean, Bryan Scotney, and Fiona Palmer. Temporal probabilistic concepts from heterogeneous data sequences. Lecture Notes in Computer Science, 2311:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110191.
pdf.

Marshall:2002:HUM

[1172] Adele Marshall, David Bell, and Roy Sterritt. Handling uncertainty in a medical study of dietary intake during pregnancy. Lecture Notes in Computer Science, 2311:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110206.
pdf.

McSherry:2002:SDI

[1173] David McSherry. Sequential diagnosis in the independence Bayesian framework. Lecture Notes in Computer Science, 2311:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2311/23110217.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2311/23110217.pdf.

Ruta:2002:SFA

[1174] Dymitr Ruta and Bogdan Gabrys. Static field approach for pattern classification. Lecture Notes in Computer Science, 2311:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2311/23110232.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2311/23110232.pdf.

Kryszkiewicz:2002:IKF

[1175] Marzena Kryszkiewicz. Inferring knowledge from frequent patterns. Lecture Notes in Computer Science, 2311: 247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110247.
pdf.

Amor:2002:APP

[1176] Nahla Ben Amor, Salem Benferhat, and Khaled Mellouli. Anytime possibilistic propagation algorithm. Lecture Notes in Computer Science, 2311:263– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110263.
pdf.

Adams:2002:MAT

[1177] Carl Adams and David E. Avison.

Macro analysis of techniques to deal with uncertainty in information systems development: Mapping representational framing influences. Lecture Notes in Computer Science, 2311:280– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110280.
                                        pdf.

Ramos:2002:REV

[1178] Isabel Ramos, Daniel M. Berry, and João Á. Carvalho. The role of emotion, values, and beliefs in the construction of innovative work realities. Lecture Notes in Computer Science, 2311: 300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110300.
                                        pdf.

Tomayko:2002:MER

[1179] Jim Tomayko. Managing evolving requirements using eXtreme programming. Lecture Notes in Computer Science, 2311:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2311/23110315.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2311/23110315.pdf.

Crangle:2002:TSD

[1180] Colleen E. Crangle. Text summarization in data mining. Lecture Notes in Computer Science, 2311:332– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110332.
pdf.

Azvine:2002:IAI

[1181] Benham Azvine. Industrial applications of intelligent systems at BTexact. Lecture Notes in Computer Science, 2311:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2311/23110348.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2311/23110348.
pdf.

Bigham:2002:ICW

[1182] John Bigham. Intelligent control of wireless and fixed telecom networks. Lecture Notes in Computer Science, 2311:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110349.
pdf.

Hoare:2002:APS

[1183] Professor Sir Tony Hoare. Assertions in programming: From scientific theory to engineering practice. Lecture Notes in Computer Science, 2311:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2311/23110350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110350.
                                        pdf.

Bonissone:2002:HSC

[1184] Piero Bonissone. Hybrid soft computing for classification and prediction applications. Lecture Notes in Computer Science, 2311:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110352.
                                        pdf.

Paul:2002:WUC

[1185] Ray Paul. Why users cannot ’get what they want’. Lecture Notes in Computer Science, 2311:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110354.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2311/23110354.
                                        pdf.

McAuley:2002:SDR

[1186] Derek McAuley. Systems design with the reverend bayes. Lecture Notes in Computer Science, 2311:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2311/23110355.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2311/23110355.
pdf.

Jackson:2002:FIS

[1187] Michael Jackson. Formalism and informality in software development. Lecture Notes in Computer Science, 2311: 356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110356.
pdf.

Bell:2002:IPS

[1188] Gordon Bell, Bob Barbour, Paul Mc-Menamin, Dave Allen, and Maurice Mulvenna. An industrial perspective on soft issues: Successes, opportunities, and challenges. Lecture Notes in Computer Science, 2311:357– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2311/23110357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2311/23110357.
pdf.

Anonymous:2002:AIah

[1189] Anonymous. Author index. Lecture Notes in Computer Science, 2311:359– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2311/2311auth.pdf.

Pena:2002:STT

[1190] Ricardo Peña and Clara Segura. Sized types for typing Eden skeletons. Lecture Notes in Computer Science, 2312: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120001.
pdf.

Kreye:2002:CSH

[1191] Dietmar Kreye. A compilation scheme for a hierarchy of array types. Lecture Notes in Computer Science, 2312:18– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120018.
pdf.

Grelck:2002:OAS

[1192] Clemens Grelck. Optimizations on array skeletons in a shared memory environment. Lecture Notes in Computer Science, 2312:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120036.
pdf.

Mol:2002:TPF

[1193] Maarten de Mol, Marko van Eekelen, and Rinus Plasmeijer. Theorem proving for functional programmers. Lecture Notes in Computer Science, 2312: 55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120055.
pdf.

Butterfield:2002:PCP

[1194] Andrew Butterfield and Glenn Strong.

Proving correctness of programs with IO — A paradigm comparison. Lecture Notes in Computer Science, 2312: 72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120072.
pdf.

delaEncina:2002:PCS

[1195] Alberto de la Encina and Ricardo Peña. Proving the correctness of the STG machine. Lecture Notes in Computer Science, 2312:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120088.
pdf.

Faxen:2002:DCE

[1196] Karl-Filip Faxén. Dynamic cheap eagerness. Lecture Notes in Computer Science, 2312:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2312/23120105.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2312/23120105.pdf.

Pena:2002:PCN

[1197] Ricardo Peña and Clara Segura. A polynomial-cost non-determinism analysis. Lecture Notes in Computer Science, 2312:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2312/23120121.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2312/23120121.pdf.

Scaife:2002:CCP

[1198] Norman Scaife, Greg Michaelson, and Susumu Horiguchi. Comparative crossplatform performance results from a parallelizing SML compiler. Lecture Notes in Computer Science, 2312:138– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2312/23120138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2312/23120138.
                                        pdf.

Minamide:2002:RBC

[1199] Yasuhiko Minamide. Runtime behavior of conversion interpretation of subtyping. Lecture Notes in Computer Science, 2312:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2312/23120155.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2312/23120155.pdf.

Alimarine:2002:GPE

[1200] Artem Alimarine and Rinus Plasmeijer. A generic programming extension for clean. Lecture Notes in Computer Science, 2312:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2312/23120168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2312/23120168.
pdf.

Anonymous:2002:AIai

[1201] Anonymous. Author index. Lecture Notes in Computer Science, 2312:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2312/2312auth.pdf.

Sanchez:2002:MPC

[1202] Abraham Sánchez L., René Zapata, and J. Abraham Arenas B. Motion planning for car-like robots using lazy probabilistic roadmap method. Lecture Notes in Computer Science, 2313: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130001.
pdf.

Murrieta-Cid:2002:VSE

[1203] Rafael Murrieta-Cid, Carlos Parra, Michel Devy, Benjamín Tovar, and Claudia Esteves. A vision system for environment representation: From landscapes to landmarks. Lecture Notes in Computer Science, 2313:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130011.
                                        pdf.

Cueva:2002:AMG

[1204] Victor de la Cueva and Fernando Ramos. Adapting the messy genetic algorithm for path planning in redundant and non-redundant manipulators. Lecture Notes in Computer Science, 2313: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130021.
                                        pdf.

Gomez:2002:NAH

[1205] G. Gomez, L. Enrique Sucar, and Duncan F. Gillies. Navigation advice from pq-histograms. Lecture Notes in Computer Science, 2313:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130031.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2313/23130031.
pdf.

Sanchez-Ante:2002:PPU

[1206] Gildardo Sánchez-Ante. Path planning using a single-query bi-directional lazy collision checking planner. Lecture Notes in Computer Science, 2313: 41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130041.
pdf.

Romero:2002:EAI

[1207] Leonardo Romero, Eduardo F. Morales, and L. Enrique Sucar. An exploration approach for indoor mobile robots reducing odometric errors. Lecture Notes in Computer Science, 2313:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130051.
pdf.

Nuno:2002:FMU

[1208] Jesús A. Martínez Nuño and Juan Humberto Sossa Azuela. Feature matching using accumulation spaces. Lecture Notes in Computer Science, 2313:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2313/23130061.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2313/23130061.
                                        pdf.

Gomez:2002:SAC

[1209] G. Gomez, M. Sanchez, and L. Enrique Sucar. On selecting an appropriate colour space for skin detection. Lecture Notes in Computer Science, 2313: 69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130069.
                                        pdf.

Kuri-Morales:2002:MSC

[1210] Angel Fernando Kuri-Morales. A methodology for the statistical characterization of genetic algorithms. Lecture Notes in Computer Science, 2313: 79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130079.
                                        pdf.

Sanvicente-Sanchez:2002:MMP

[1211] Héctor Sanvicente-Sánchez and Juan Frausto-Solís. MPSA: a methodology to parallelize simulated annealing and its application to the traveling salesman problem. Lecture Notes in Computer Science, 2313:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130089.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2313/23130089.
pdf.

Coello:2002:CAC

[1212] Carlos A. Coello Coello and Ricardo Landa Becerra. A cultural algorithm for constrained optimization. Lecture Notes in Computer Science, 2313:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2313/23130098.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2313/23130098.
pdf.

Kuri-Morales:2002:PFM

[1213] Angel Fernando Kuri-Morales and Jesús Gutiérrez-García. Penalty function methods for constrained optimization with genetic algorithms: a statistical analysis. Lecture Notes in Computer Science, 2313:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130108.
pdf.

Perez:2002:AGC

[1214] Joaquín Pérez, Rodolfo Pazos, Laura Velez, and Guillermo Rodríguez. Automatic generation of control parameters for the threshold accepting algorithm. Lecture Notes in Computer Science, 2313:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130118.
                                        pdf.

Perez:2002:GAC

[1215] E. Islas Pérez, Carlos A. Coello Coello, Arturo Hernández-Aguirre, and Alejandro Villavicencio Ramírez. Genetic algorithms and case-based reasoning as a discovery and learning machine in the optimization of combinational logic circuits. Lecture Notes in Computer Science, 2313:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130128.
                                        pdf.

Mayora-Ibarra:2002:TDS

[1216] O. Mayora-Ibarra and F. Curatelli. Time-domain segmentation and labelling of speech with fuzzy-logic post-correction rules. Lecture Notes in Computer Science, 2313:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130138.
                                        pdf.

Peral:2002:MSE

[1217] Jesús Peral and Antonio Ferrández. IL MT system. evaluation for Spanish– English pronominal anaphora generation. Lecture Notes in Computer Science, 2313:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2313/23130146.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2313/23130146.pdf.

Cuayahuitl:2002:VWM

[1218] Heriberto Cuayáhuitl and Ben Serridge. Out-of-vocabulary word modeling and rejection for Spanish keyword spotting systems. Lecture Notes in Computer Science, 2313:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130156.
pdf.

Pineda:2002:DP

[1219] L. A. Pineda, A. Massé, I. Meza, M. Salas, E. Schwarz, E. Uraga, and L. Villaseñor. The DIME project. Lecture Notes in Computer Science, 2313: 166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130166.
pdf.

Montes-y-Gomez:2002:DDT

[1220] M. Montes y Gómez, A. Gelbukh, and A. López-López. Detecting deviations in text collections: An approach using conceptual graphs. Lecture Notes in Computer Science, 2313: 176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130176.
                                        pdf.

Llopis:2002:ULQ

[1221] Fernando Llopis, Antonio Ferrández, and José Luis Vicedo. Using long queries in a passage retrieval system. Lecture Notes in Computer Science, 2313:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2313/23130185.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130185.
                                        pdf.

Ringwelski:2002:OOC

[1222] Georg Ringwelski. Object-oriented constraint programming with J.CP. Lecture Notes in Computer Science, 2313:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2313/23130194.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130194.
                                        pdf.

Heinemann:2002:HTE

[1223] Bernhard Heinemann. A hybrid treatment of evolutionary sets. Lecture Notes in Computer Science, 2313:204– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130204.
pdf.

Pietarinen:2002:GLK

[1224] Ahti Pietarinen. Games and logics of knowledge for multi-agent systems. Lecture Notes in Computer Science, 2313:214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2313/23130214.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2313/23130214.
pdf.

Quintana:2002:MLC

[1225] Ma. Guadalupe Quintana and Rafael Morales. Modelling learners of a control task with inductive logic programming: a case study. Lecture Notes in Computer Science, 2313:224– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130224.
pdf.

Morales-Luna:2002:SEL

[1226] Guillermo Morales-Luna. Simple epistemic logic for relational database. Lecture Notes in Computer Science, 2313: 234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130234.
                                        pdf.

Tomas:2002:SOL

[1227] Ana Paula Tomás. Solving optimal location of traffic counting points at urban intersections in CLP(FD). Lecture Notes in Computer Science, 2313:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130242.
                                        pdf.

Palacios:2002:FAP

[1228] José Juan Palacios. Flexible agent programming in linear logic. Lecture Notes in Computer Science, 2313:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130252.
                                        pdf.

Hernandez-Aguirre:2002:SCF

[1229] Arturo Hernández-Aguirre, Cris Koutsougeras, and Bill P. Buckles. Sample complexity for function learning tasks through linear neural networks. Lecture Notes in Computer Science, 2313: 262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2313/23130262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130262.
pdf.

Milare:2002:EKA

[1230] Claudia Regina Milaré, André Carlos Ponce Leon Ferreira de Carvalho, and Maria Carolina Monard. Extracting knowledge from artificial neural networks: An empirical comparison of trepan and symbolic learning algorithms. Lecture Notes in Computer Science, 2313:272– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130272.
pdf.

Villela:2002:IPR

[1231] Patricia Rayón Villela and Juan Humberto Sossa Azuela. Improving pattern recognition using several feature vectors. Lecture Notes in Computer Science, 2313:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130282.
pdf.

Cruz:2002:LOM

[1232] Isis Bonet Cruz, Adolfo Díaz Sardiñas, Rafael Bello Pérez, and Yanetsy Sardiñas Oliva. Learning optimization in a MLP neural network applied to OCR. Lecture Notes in Computer Science, 2313: 292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130292.
                                        pdf.

Barros:2002:ACL

[1233] B. Barros, M. F. Verdejo, T. Read, and R. Mizoguchi. Applications of a collaborative learning ontology. Lecture Notes in Computer Science, 2313:301– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130301.
                                        pdf.

Geng:2002:ACG

[1234] Liqiang Geng and Christine W. Chan. Automated case generation from databases using similarity-based rough approximation. Lecture Notes in Computer Science, 2313:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130311.
                                        pdf.

Castillo:2002:SED

[1235] Lourdes Peña Castillo and Stefan Wrobel. On the stability of exampledriven learning systems: A case study in multirelational learning. Lecture Notes in Computer Science, 2313:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130321.
pdf.

Liu:2002:SRO

[1236] Yang Liu, Xiaohui Yu, and Jiqing Han. Sharpe ratio-oriented active trading: a learning approach. Lecture Notes in Computer Science, 2313:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130331.
pdf.

Fallah-Seghrouchni:2002:FSA

[1237] Amal El Fallah-Seghrouchni and Christian Lemaître. A framework for social agents’ interaction based on communicative action theory and dynamic deontic logic. Lecture Notes in Computer Science, 2313:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130340.
pdf.

Favela:2002:AAU

[1238] Jesús Favela, Manuel Alba, and Marcela Rodríguez. Autonomous agents for ubiquitous collaborative environments. Lecture Notes in Computer Science, 2313:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130351.
                                        pdf.

Caballero:2002:FAB

[1239] Alberto Caballero and Ronny Alfonso. A framework for agent based network management systems. Lecture Notes in Computer Science, 2313:361– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130361.
                                        pdf.

Costa:2002:MAC

[1240] Evandro de Barros Costa, Hyggo Oliveira de Almeida, Emerson Ferreira de Araújo Lima, Ricardo Rubens Gomes Nunes Filho, Klebson dos Santos Silva, and Fernando Maia Assunção. A multiagent cooperative intelligent tutoring system: The case of musical harmony domain. Lecture Notes in Computer Science, 2313:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2313/23130367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130367.
pdf.

Guzman:2002:RUE

[1241] Adolfo Guzmán, Carmen Domínguez, and Jesús Olivares. Reacting to unexpected events and communicating in spite of mixed ontologies. Lecture Notes in Computer Science, 2313:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130377.
pdf.

Alvarado:2002:LIM

[1242] Matías Alvarado, Leonid Sheremetov, Ernesto Germán, and Erick Alva. Logic of interaction for multiagent systems. Lecture Notes in Computer Science, 2313:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130387.
pdf.

Saab:2002:TMI

[1243] Rosa Saab, Raúl Monroy, and Fernando Godínez. Towards a model for an immune system. Lecture Notes in Computer Science, 2313:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130401.
                                        pdf.

Pappas:2002:NMA

[1244] Alexandros Pappas and Duncan F. Gillies. A new measure for the accuracy of a Bayesian network. Lecture Notes in Computer Science, 2313:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130411.
                                        pdf.

Puch:2002:FTP

[1245] Roberto O. Puch and Jim Q. Smith. FINDS: a training package to assess forensic fibre evidence. Lecture Notes in Computer Science, 2313:420– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130420.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130420.
                                        pdf.

Hernandez:2002:PCB

[1246] José Antonio Ruz Hernández, Dionisio A. Suárez Cerda, Evgen Shelomov, and Alejandro Villavicencio Ramírez. Predictive control based on an autoregressive neuro-fuzzy model applied to the steam generator startup process at a fossil power plant. Lecture Notes in Computer Science, 2313:430– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130430.
pdf.

Arroyo-Figueroa:2002:MDC

[1247] Gustavo Arroyo-Figueroa and Raúl Herrera-Avelar. Modeling dynamical causal interactions with fuzzy temporal networks for process operation support systems. Lecture Notes in Computer Science, 2313:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130440.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130440.
pdf.

Castaneda:2002:RSL

[1248] Miguel A. Padilla Castañeda and Fernando Arámbula Cosío. Resection simulation with local tissue deformations for computer assisted surgery of the prostate. Lecture Notes in Computer Science, 2313:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130450.
pdf.

Gutierrez:2002:BDM

[1249] Francisco J. Gutiérrez, Margarita M.

Lerma-Rascón, Luis R. Salgado-Garza, and Francisco J. Cantú. Biometrics and data mining: Comparison of data mining-based keystroke dynamics methods for identity verification. Lecture Notes in Computer Science, 2313: 460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130460.
pdf.

Flores:2002:QSI

[1250] Juan J. Flores and Nelio Pastor. Qualitative systems identification for linear time invariant dynamic systems. Lecture Notes in Computer Science, 2313: 470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130470.
pdf.

Valdez:2002:MEI

[1251] José Guillermo Valdéz and Pablo H. Ibargüengoytia. Monitoring the execution of an intelligent planning system. Lecture Notes in Computer Science, 2313:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2313/23130477.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2313/23130477.
pdf.

Li:2002:SME

[1252] Xiaoou Li, Joselito Medina Marín, and Sergio V. Chapa. A structural model of ECA rules in active database. Lecture Notes in Computer Science, 2313: 486–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130486.
pdf.

Sanchez-Parra:2002:URB

[1253] Marino Sánchez-Parra and René Vite-Hernández. Use of a rule-based system for process control: Flow instabilities in axial compressors case study. Lecture Notes in Computer Science, 2313: 494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130494.
pdf.

Decouchant:2002:DES

[1254] Dominique Decouchant, Ana María Martínez-Enríquez, Jesus Favela, Alberto L. Morán, Sonia Mendoza, and Samir Jafar. A distributed event service for adaptive group awareness. Lecture Notes in Computer Science, 2313: 506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2313/23130506.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2313/23130506.
                                        pdf.

Morales:2002:MRA

[1255] Eduardo F. Morales, Dulce Ma. Heredia, and Andrés F. Rodríguez. Mining road accidents. Lecture Notes in Computer Science, 2313:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130516.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130516.
                                        pdf.

Martinez-Enriquez:2002:IEW

[1256] Ana María Martínez-Enríquez, Aslam Muhammad, Dominique Decouchant, and Jesús Favela. An inference engine for Web adaptive cooperative work. Lecture Notes in Computer Science, 2313:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2313/23130526.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2313/23130526.
                                        pdf.

Garza:2002:FDI

[1257] Luis E. Garza, Francisco J. Cantú, and Salvador Acevedo. Faults diagnosis in industrial processes with a hybrid diagnostic system. Lecture Notes in Computer Science, 2313:536– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2313/23130536.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2313/23130536.
pdf.

Anonymous:2002:AIaj

[1258] Anonymous. Author index. Lecture Notes in Computer Science, 2313:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2313/2313auth.pdf.

Chang:2002:MSI

[1259] Shi-Kuo Chang, Gennaro Costagliola, and Erland Jungert. Multi-sensor information fusion by query refinement. Lecture Notes in Computer Science, 2314:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2314/23140001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2314/23140001.
pdf.

Stehling:2002:MMD

[1260] Renato O. Stehling, Mario A. Nascimento, and Alexandre X. Falcão. Mi-CRoM: a metric distance to compare segmented images. Lecture Notes in Computer Science, 2314:12– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140012.
pdf.

Fauqueur:2002:IRR

[1261] Julien Fauqueur and Nozha Boujemaa. Image retrieval by regions: Coarse segmentation and fine color description. Lecture Notes in Computer Science, 2314:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140024.
pdf.

Volmer:2002:FAN

[1262] Stephan Volmer. Fast approximate nearest-neighbor queries in metric feature spaces by buoy indexing. Lecture Notes in Computer Science, 2314:36– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140036.
pdf.

Qiu:2002:BCV

[1263] Guoping Qiu and S. Sudirman. A binary color vision framework for content-based image indexing. Lecture Notes in Computer Science, 2314:50– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140050.
pdf.

Sridhar:2002:RBI

[1264] Veena Sridhar, Mario A. Nascimento, and Xiaobo Li. Region-based image retrieval using multiple-features. Lecture Notes in Computer Science, 2314: 61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140061.
pdf.

Tao:2002:BMC

[1265] Ju-Lan Tao and Yi-Ping Hung.

A Bayesian method for contentbased image retrieval by use of relevance feedback. Lecture Notes in Computer Science, 2314:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140076.
pdf.

Shih:2002:CIR

[1266] Jau-Ling Shih and Ling-Hwei Chen.

Color image retrieval based on primitives of color moments. Lecture Notes in Computer Science, 2314:88– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140088.
pdf.

Sun:2002:IFE

[1267] Shu-Kuo Sun, Zen Chen, and Tsorng-Lin Chia. Invariant feature extraction and object shape matching using Gabor filtering. Lecture Notes in Computer Science, 2314:95– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140095.
pdf.

Eidenberger:2002:FVI

[1268] Horst Eidenberger, Christian Breiteneder, and Martin Hitz. A framework for visual information retrieval. Lecture Notes in Computer Science, 2314: 105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140105.
pdf.

Oostveen:2002:FED

[1269] Job Oostveen, Ton Kalker, and Jaap Haitsma. Feature extraction and a database strategy for video fingerprinting. Lecture Notes in Computer Science, 2314:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140117.
pdf.

Nakazato:2002:ISA

[1270] Munehiro Nakazato, Lubomir Manola, and Thomas S. Huang. Image-Grouper: Search, annotate and organize images by groups. Lecture Notes in Computer Science, 2314:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140129.
pdf.

Chiu:2002:TPC

[1271] Chih-Yi Chiu, Hsin-Chih Lin, and Shi-Nine Yang. Toward a personalized CBIR system. Lecture Notes in Computer Science, 2314:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140143.
pdf.

Tse:2002:ESO

[1272] Philip K. C. Tse and Clement H. C.

Leung. An efficient storage organization for multimedia databases. Lecture Notes in Computer Science, 2314:152– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140152.
pdf.

Saux:2002:UCI

[1273] Bertrand Le Saux and Nozha Boujemaa. Unsupervised categorization for image database overview. Lecture Notes in Computer Science, 2314:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140163.
                                        pdf.

Morris:2002:DFA

[1274] Andrew J. Morris, Alia I. Abdelmoty, Baher A. El-Geresy, and Christopher B. Jones. A data-flow approach to visual querying in large spatial databases. Lecture Notes in Computer Science, 2314:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140175.
                                        pdf.

Stanchev:2002:MMD

[1275] Peter L. Stanchev and Farshad Fotouhi. MEDIMAGE — A multimedia database management system for alzheimer’s disease patients. Lecture Notes in Computer Science, 2314:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140187.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2314/23140187.
pdf.

Chen:2002:LAV

[1276] Trista Pei chun Chen, Tsuhan Chen, and Yuh-Feng Hsu. Life after video coding standards: Rate shaping and error concealment. Lecture Notes in Computer Science, 2314:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140194.
pdf.

Lee:2002:DDV

[1277] Yuh-Reuy Lee, Chia-Wen Lin, and Cheng-Chien Kao. A DCT-domain video transcoder for spatial resolution downconversion. Lecture Notes in Computer Science, 2314:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140207.
pdf.

Kuo:2002:RDC

[1278] Chin-Ying Kuo, Chen-Lung Chan, Vincent Hsu, and Jia-Shung Wang. A receiver-driven channel adjustment scheme for periodic broadcast of streaming video. Lecture Notes in Computer Science, 2314:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140219.
                                        pdf.

Gatica-Perez:2002:VOH

[1279] Daniel Gatica-Perez, Zhi Zhou, Ming-Ting Sun, and Vincent Hsu. Video object hyper-links for streaming applications. Lecture Notes in Computer Science, 2314:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2314/23140229.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2314/23140229.pdf.

Kim:2002:SHS

[1280] Jung-Rim Kim, Seong Soo Chun, Seok jin Oh, and Sanghoon Sull. Scalable hierarchical summarization of news using fidelity in MPEG-7 description scheme. Lecture Notes in Computer Science, 2314:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140239.
                                        pdf.

Koskela:2002:MDC

[1281] Markus Koskela, Jorma Laaksonen, and Erkki Oja. MPEG-7 descriptors in content-based image retrieval with PicSOM system. Lecture Notes in Computer Science, 2314:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140247.
pdf.

Chun:2002:FTC

[1282] Seong Soo Chun, Hyeokman Kim, Jung-Rim Kim, Sangwook Oh, and Sanghoon Sull. Fast text caption localization on video using visual rhythm. Lecture Notes in Computer Science, 2314:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140259.
pdf.

Shen:2002:NDW

[1283] Kuan-Ting Shen and Ling-Hwei Chen.

A new digital watermarking technique for video. Lecture Notes in Computer Science, 2314:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140269.
pdf.

Chen:2002:ACC

[1284] Duan-Yu Chen, Ming-Ho Hsiao, and Suh-Yin Lee. Automatic closed caption detection and font size differentiation in MPEG video. Lecture Notes in Computer Science, 2314:276– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140276.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140276.
                                        pdf.

Chen:2002:MAB

[1285] Duan-Yu Chen, Shu-Jiuan Lin, and Suh-Yin Lee. Motion activity based shot identification and closed caption detection for video structuring. Lecture Notes in Computer Science, 2314: 288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140288.
                                        pdf.

Wu:2002:VCG

[1286] Peter Y. Wu, Kai A. Olsen, and Per Saetre. Visualizing the construction of generic bills of material. Lecture Notes in Computer Science, 2314:302– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2314/23140302.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2314/23140302.
                                        pdf.

Nguyen:2002:DKV

[1287] TrongDung Nguyen, TuBao Ho, and DucDung Nguyen. Data and knowledge visualization in knowledge discovery process. Lecture Notes in Computer Science, 2314:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2314/23140311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2314/23140311.
pdf.

Anonymous:2002:AIak

[1288] Anonymous. Author index. Lecture Notes in Computer Science, 2314:323– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2314/2314auth.pdf.

Stevens:2002:PGS

[1289] Perdita Stevens. Playing games with software design. Lecture Notes in Computer Science, 2315:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150001.
pdf.

Waldo:2002:CSD

[1290] Jim Waldo. Coordination and system design in a network-centric age. Lecture Notes in Computer Science, 2315: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150002.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2315/23150002.
                                        pdf.

Wooldridge:2002:TKC

[1291] Michael Wooldridge and Wiebe van der Hoek. Time, knowledge, and cooperation: Alternating-time temporal epistemic logic and its applications. Lecture Notes in Computer Science, 2315: 4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150004.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150004.
                                        pdf.

Andrade:2002:CO

[1292] Luis F. Andrade, Jose L. Fiadeiro, Joao Gouveia, Georgios Koutsoukos, and Michael Wermelinger. Coordination for orchestration. Lecture Notes in Computer Science, 2315:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150005.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150005.
                                        pdf.

Ankolekar:2002:CSW

[1293] Anupriya Ankolekar, Frank Huch, and Katia Sycara. Concurrent semantics for the Web services specification language DAML-S. Lecture Notes in Computer Science, 2315:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2315/23150014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150014.
pdf.

Arbab:2002:CTC

[1294] Farhad Arbab and Farhad Mavaddat. Coordination through channel composition. Lecture Notes in Computer Science, 2315:22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150022.
pdf.

Bernardo:2002:EEE

[1295] Marco Bernardo and Francesco Franzè.

Exogenous and endogenous extensions of architectural types. Lecture Notes in Computer Science, 2315:40– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150040.
pdf.

Bettini:2002:CMO

[1296] Lorenzo Bettini, Viviana Bono, and Betti Venneri. Coordinating mobile object-oriented code. Lecture Notes in Computer Science, 2315:56– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150056.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150056.
                                        pdf.

Bettini:2002:FPM

[1297] Lorenzo Bettini, Rocco De Nicola, and Michele Loreti. Formalizing properties of mobile agent systems. Lecture Notes in Computer Science, 2315:72– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150072.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150072.
                                        pdf.

Bracciali:2002:DAB

[1298] Andrea Bracciali, Antonio Brogi, and Carlos Canal. Dynamically adapting the behaviour of software components. Lecture Notes in Computer Science, 2315:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150088.
                                        pdf.

Browne:2002:ABB

[1299] James C. Browne, Kevin Kane, and Hongxia Tian. An associative broadcast based coordination model for distributed processes. Lecture Notes in Computer Science, 2315:96–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2315/23150096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150096.
pdf.

Busi:2002:SEB

[1300] Nadia Busi, Antony Rowstron, and Gianluigi Zavattaro. State- and event-based reactive programming in shared dataspaces. Lecture Notes in Computer Science, 2315:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150111.
pdf.

Costa:2002:ITO

[1301] Carlos J. Costa, Pedro Antunes, and João Ferreira Dias. Integrating two organizational systems through communication genres. Lecture Notes in Computer Science, 2315:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150125.
pdf.

Cruz:2002:OCF

[1302] Juan Carlos Cruz. OpenCoLaS a coordination framework for CoLaS dialects. Lecture Notes in Computer Science, 2315:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2315/23150133.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150133.
                                        pdf.

Cuesta:2002:CRA

[1303] Carlos E. Cuesta, Pablo de la Fuente, Manuel Barrio-Solórzano, and Encarnación Beato. Coordination in a reflective architecture description language. Lecture Notes in Computer Science, 2315:141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150141.
                                        pdf.

Eden:2002:CSE

[1304] Amnon H. Eden and Jens Jahnke. Coordinating software evolution via two-tier programming. Lecture Notes in Computer Science, 2315:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150149.
                                        pdf.

Frozza:2002:CAC

[1305] Rejane Frozza and Luis Otávio Alvares. Criteria for the analysis of coordination in multi-agent applications. Lecture Notes in Computer Science, 2315: 158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2315/23150158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150158.
pdf.

Gibbons:2002:TCB

[1306] Jeremy Gibbons. Towards a colimitbased semantics for visual programming. Lecture Notes in Computer Science, 2315:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2315/23150166.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2315/23150166.pdf.

Goossens:2002:CCP

[1307] Kees G. W. Goossens and Om Prakash Gangwal. The cost of communication protocols and coordination languages in embedded systems. Lecture Notes in Computer Science, 2315:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150174.
pdf.

Groenewegen:2002:OSC

[1308] Luuk Groenewegen and Erik de Vink.

Operational semantics for coordination in paradigm. Lecture Notes in Computer Science, 2315:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2315/23150191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150191.
                                        pdf.

Handorean:2002:SPA

[1309] Radu Handorean and Gruia-Catalin Roman. Service provision in ad hoc networks. Lecture Notes in Computer Science, 2315:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2315/23150207.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2315/23150207.pdf.

Hiraishi:2002:PEM

[1310] Kunihiko Hiraishi. PN^{2}: An elementary model for design and analysis of multi-agent systems. Lecture Notes in Computer Science, 2315:220– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150220.
                                        pdf.

Lee:2002:RTU

[1311] Hwa-Min Lee, Kwang-Sik Chung, Sang-Chul Shin, Dae-Won Lee, Won-Gyu Lee, and Heon-Chang Yu. A recovery technique using multi-agent in distributed computing systems. Lecture Notes in Computer Science, 2315: 236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2315/23150236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150236.
pdf.

Perez:2002:OBD

[1312] José Antonio Pérez, Rafael Corchuelo, David Ruiz, and Miguel Toro. An order-based, distributed algorithm for implementing multiparty interactions. Lecture Notes in Computer Science, 2315:250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2315/23150250.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2315/23150250.
pdf.

Picco:2002:ETS

[1313] Gian Pietro Picco and Marco L. Buschini. Exploiting transiently shared tuple spaces for location transparent code mobility. Lecture Notes in Computer Science, 2315:258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150258.
pdf.

Pol:2002:FSJ

[1314] Jaco van de Pol and Miguel Valero Espada. Formal specification of JavaSpaces^{TM} Architecture using μCRL. Lecture Notes in Computer Science, 2315:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2315/23150274.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2315/23150274.pdf.

Ricci:2002:OVS

[1315] Alessandro Ricci, Andrea Omicini, and Enrico Denti. Objective vs. subjective coordination in agent-based systems: a case study. Lecture Notes in Computer Science, 2315:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150291.
                                        pdf.

Sample:2002:SUU

[1316] Neal Sample, Pedram Keyani, and Gio Wiederhold. Scheduling under uncertainty: Planning for the ubiquitous grid. Lecture Notes in Computer Science, 2315:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150300.
                                        pdf.

Snyder:2002:ULO

[1317] Jim Snyder and Ronaldo Menezes. Using logical operators as an extended coordination mechanism in linda. Lecture Notes in Computer Science, 2315:317– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150317.
pdf.

Stam:2002:FCP

[1318] Andries Stam. A framework for coordinating parallel branch and bound algorithms. Lecture Notes in Computer Science, 2315:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150332.
pdf.

Tata:2002:PCV

[1319] Samir Tata. Policies for cooperative virtual teams. Lecture Notes in Computer Science, 2315:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2315/23150340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150340.
pdf.

Tolksdorf:2002:SMF

[1320] Robert Tolksdorf and Gregor Rojec-Goldmann. The SPACETUB models and framework. Lecture Notes in Computer Science, 2315:348–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2315/23150348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150348.
                                        pdf.

Viroli:2002:TBM

[1321] Mirko Viroli and Andrea Omicini. Tuple-based models in the observation framework. Lecture Notes in Computer Science, 2315:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150364.
                                        pdf.

Wells:2002:EMF

[1322] George Wells, Alan Chalmers, and Peter Clayton. Extending the matching facilities of linda. Lecture Notes in Computer Science, 2315:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2315/23150380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2315/23150380.
                                        pdf.

Wojciechowski:2002:SPM

[1323] Paweł T. Wojciechowski, Sergio Mena, and André Schiper. Semantics of protocol modules composition and interaction. Lecture Notes in Computer Science, 2315:389–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2315/23150389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2315/23150389.
pdf.

Anonymous:2002:AIal

[1324] Anonymous. Author index. Lecture Notes in Computer Science, 2315:405– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2315/2315auth.pdf.

Domingo-Ferrer:2002:AIC

[1325] Josep Domingo-Ferrer. Advances in inference control in statistical databases: An overview. Lecture Notes in Computer Science, 2316:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160001.
pdf.

Robertson:2002:CSE

[1326] Dale A. Robertson and Richard Ethier. Cell suppression: Experience and theory. Lecture Notes in Computer Science, 2316:8–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160008.
pdf.

Cox:2002:BED

[1327] Lawrence H. Cox. Bounds on entries in 3-dimensional contingency tables subject to given marginal totals. Lecture Notes in Computer Science, 2316:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160021.
pdf.

Gonzalez:2002:ECS

[1328] Juan José Salazar González. Extending cell suppression to protect tabular data against several attackers. Lecture Notes in Computer Science, 2316:34– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160034.
pdf.

Castro:2002:NFH

[1329] Jordi Castro. Network flows heuristics for complementary cell suppression: An empirical evaluation and extensions. Lecture Notes in Computer Science, 2316:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160059.
pdf.

Wolf:2002:HHA

[1330] Peter-Paul de Wolf. HiTaS: a heuristic approach to cell suppression in hierarchical tables. Lecture Notes in Computer Science, 2316:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160074.
pdf.

Polettini:2002:MBD

[1331] Silvia Polettini, Luisa Franconi, and Julian Stander. Model based disclosure protection. Lecture Notes in Computer Science, 2316:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160083.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160083.
pdf.

Brand:2002:MPT

[1332] Ruth Brand. Microdata protection through noise addition. Lecture Notes in Computer Science, 2316:97– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160097.
pdf.

Dandekar:2002:SMD

[1333] Ramesh A. Dandekar, Michael Cohen, and Nancy Kirkendall. Sensitive micro data protection using Latin hypercube sampling technique. Lecture Notes in Computer Science, 2316:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160117.
pdf.

Elliot:2002:IFR

[1334] Mark Elliot. Integrating file and record level disclosure risk assessment. Lecture Notes in Computer Science, 2316: 126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160126.
pdf.

Yancey:2002:DRA

[1335] William E. Yancey, William E. Winkler, and Robert H. Creecy. Disclosure risk assessment in perturbative microdata protection. Lecture Notes in Computer Science, 2316:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160135.
pdf.

Dandekar:2002:LBH

[1336] Ramesh A. Dandekar, Josep Domingo-Ferrer, and Francesc Sebé. LHSbased hybrid microdata vs rank swapping and microaggregation for numeric microdata protection. Lecture Notes in Computer Science, 2316:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160153.
pdf.

Sebe:2002:PMO

[1337] Francesc Sebé, Josep Domingo-Ferrer, Josep Maria Mateo-Sanz, and Vicenç Torra. Post-masking optimization of the tradeoff between information loss and disclosure risk in masked microdata sets. Lecture Notes in Computer Science, 2316:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160163.
pdf.

Hundepool:2002:CP

[1338] Anco Hundepool. The CASC project.

Lecture Notes in Computer Science, 2316:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2316/23160172.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2316/23160172.
                                        pdf.

Giessing:2002:TSP

[1339] Sarah Giessing and Dietz Repsilber. Tools and strategies to protect multiples tables with the GHQUAR cell suppression engine. Lecture Notes in Computer Science, 2316:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2316/23160181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2316/23160181.
                                        pdf.

Zayatz:2002:SUD

[1340] Laura Zayatz. SDC in the 2000 U.S. decennial census. Lecture Notes in Computer Science, 2316:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2316/23160193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2316/23160193.
                                        pdf.

Nordholt:2002:ASD

[1341] Eric Schulte Nordholt. Applications of statistical disclosure control at statistics netherlands. Lecture Notes in Computer Science, 2316:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2316/23160203.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2316/23160203.
pdf.

Urrutia:2002:EEP

[1342] Julià Urrutia and Enric Ripoll. Empirical evidences on protecting population uniqueness at idescat. Lecture Notes in Computer Science, 2316:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2316/23160213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2316/23160213.
pdf.

Anonymous:2002:AIam

[1343] Anonymous. Author index. Lecture Notes in Computer Science, 2316:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2316/23160231.pdf.

Chandrasekaran:2002:WDI

[1344] B. Chandrasekaran. What does it mean for a computer to do diagrammatic reasoning? A functional characterization of diagrammatic reasoning and its implications. Lecture Notes in Computer Science, 2317:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170001.
pdf.

Umata:2002:MCG

[1345] Ichiro Umata, Yasuhiro Katagiri, and Atsushi Shimojima. Movement conceptualizations in graphical communication. Lecture Notes in Computer Science, 2317:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170003.
pdf.

Freedman:2002:TMK

[1346] Eric G. Freedman and Priti Shah. Toward a model of knowledge-based graph comprehension. Lecture Notes in Computer Science, 2317:18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170018.
pdf.

Epstein:2002:LPD

[1347] Susan L. Epstein and J.-Holger Keibel. Learning on paper: Diagrams and discovery in game playing. Lecture Notes in Computer Science, 2317:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170031.
pdf.

Winterstein:2002:UAD

[1348] Daniel Winterstein, Alan Bundy, Corin Gurr, and Mateja Jamnik. Using animation in diagrammatic theorem proving. Lecture Notes in Computer Science, 2317:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170046.
pdf.

Flower:2002:GED

[1349] Jean Flower and John Howse. Generating Euler diagrams. Lecture Notes in Computer Science, 2317:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170061.
pdf.

Howse:2002:CRE

[1350] John Howse, Gemma Stapleton, Jean Flower, and John Taylor. Corresponding regions in Euler diagrams. Lecture Notes in Computer Science, 2317:76– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170076.
pdf.

Miller:2002:CCD

[1351] Nathaniel Miller. CDEG: Computerized diagrammatic Euclidean geometry. Lecture Notes in Computer Science, 2317:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2317/23170091.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170091.
                                        pdf.

Jayaraman:2002:CSD

[1352] Bharat Jayaraman and Pallavi Tambay. Compositional semantics for diagrams using constrained objects. Lecture Notes in Computer Science, 2317: 94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170094.
                                        pdf.

Yaner:2002:RDL

[1353] Patrick W. Yaner and Ashok K. Goel. Retrieving 2-D line drawings by example. Lecture Notes in Computer Science, 2317:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170097.
                                        pdf.

Tanimoto:2002:SSU

[1354] Steven Tanimoto, William Winn, and David Akers. A system that supports using student-drawn diagrams to assess comprehension of mathematical formulas. Lecture Notes in Computer Science, 2317:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170100.
pdf.

Healey:2002:ECA

[1355] Patrick G. T. Healey, Nik Swoboda, and James King. An environment for conducting and analysing graphical communication experiments. Lecture Notes in Computer Science, 2317: 103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170103.
pdf.

Zhang:2002:GBL

[1356] Ke-Bing Zhang, Kang Zhang, and Mehmet A. Orgun. Grammar-based layout for a visual programming language generation system. Lecture Notes in Computer Science, 2317:106– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170106.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2317/23170106.
                                        pdf.

Anderson:2002:HDQ

[1357] Michael Anderson and Brian Andersen. Heterogeneous data querying in a diagrammatic information system. Lecture Notes in Computer Science, 2317: 109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170109.
                                        pdf.

Diskin:2002:VVS

[1358] Zinovy Diskin. Visualization vs. specification in diagrammatic notations: a case study with the UML. Lecture Notes in Computer Science, 2317:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170112.
                                        pdf.

Shimojima:2002:IET

[1359] Atsushi Shimojima. The inferentialexpressive trade-off: a case study of tabular representations. Lecture Notes in Computer Science, 2317:116– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170116.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2317/23170116.
pdf.

Swoboda:2002:MHS

[1360] Nik Swoboda and Gerard Allwein.

Modeling heterogeneous systems. Lecture Notes in Computer Science, 2317: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170131.
pdf.

Howse:2002:DTT

[1361] John Howse, Fernando Molina, Sun-Joo Shin, and John Taylor. On diagram tokens and types. Lecture Notes in Computer Science, 2317:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170146.
pdf.

Bennett:2002:ENP

[1362] David J. Bennett. Effects of navigation and position on task when presenting diagrams. Lecture Notes in Computer Science, 2317:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170161.
pdf.

Collberg:2002:FVQ

[1363] Christian S. Collberg. A fuzzy visual query language for a domainspecific Web search engine. Lecture Notes in Computer Science, 2317:176– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170176.
pdf.

Blackwell:2002:DIA

[1364] Alan F. Blackwell and Hanna Wallach. Diagrammatic integration of abstract operations into software work contexts. Lecture Notes in Computer Science, 2317:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170191.
pdf.

Trafton:2002:EEI

[1365] J. Gregory Trafton, Sandra Marshall, Farilee Mintz, and Susan B. Trickett. Extracting explicit and implict information from complex visualizations. Lecture Notes in Computer Science, 2317:206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2317/23170206.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2317/23170206.
pdf.

Romero:2002:VAR

[1366] Pablo Romero, Richard Cox, Benedict du Boulay, and Rudi Lutz. Visual attention and representation switching during Java program debugging: a study using the restricted focus viewer. Lecture Notes in Computer Science, 2317:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170221.
pdf.

Grant:2002:GAP

[1367] Elizabeth R. Grant and Michael J.

Spivey. Guiding attention produces inferences in diagram-based problem solving. Lecture Notes in Computer Science, 2317:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170236.
pdf.

Gartner:2002:VMC

[1368] Johannes Gärtner, Silvia Miksch, and Stefan Carl-McGrath. ViCo: a metric for the complexity of information visualizations. Lecture Notes in Computer Science, 2317:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170249.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170249.
                                        pdf.

Cheng:2002:OIB

[1369] Peter C-H. Cheng, Rossano Barone, Peter I. Cowling, and Samad Ahmadi. Opening the information bottleneck in complex scheduling problems with a novel representation: STARK diagrams. Lecture Notes in Computer Science, 2317:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170264.
                                        pdf.

Pfeiffer:2002:UBS

[1370] Joseph J. Pfeiffer, Jr. Using brightness and saturation to visualize belief and uncertainty. Lecture Notes in Computer Science, 2317:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170279.
                                        pdf.

Koth:2002:SAD

[1371] Oliver Köth and Mark Minas. Structure, abstraction, and direct manipulation in diagram editors. Lecture Notes in Computer Science, 2317:290– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2317/23170290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170290.
pdf.

Bottoni:2002:DVL

[1372] Paolo Bottoni and Gennaro Costagliola. On the definition of visual languages and their editors. Lecture Notes in Computer Science, 2317:305– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170305.
pdf.

Jin:2002:DSS

[1373] Yan Jin, Robert Esser, and Jörn W.

Janneck. Describing the syntax and semantics of UML statecharts in a heterogeneous modelling environment. Lecture Notes in Computer Science, 2317:320–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170320.
pdf.

Irani:2002:LDS

[1374] Pourang Irani. The learnability of diagram semantics. Lecture Notes in Computer Science, 2317:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170335.
                                        pdf.

Kriz:2002:USC

[1375] Sarah Kriz. Understanding simultaneity and causality in static diagrams versus animation. Lecture Notes in Computer Science, 2317:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170338.
                                        pdf.

Suwa:2002:ERC

[1376] Masaki Suwa and Barbara Tversky. External representations contribute to the dynamic construction of ideas. Lecture Notes in Computer Science, 2317: 341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170341.
                                        pdf.

Hoffman:2002:OSS

[1377] Robert R. Hoffman, John W. Coffey, Patrick J. Hayes, Alberto J. Cañas, Kenneth M. Ford, and Mary J. Carnot. One small step for a diagram, one giant leap for meaning. Lecture Notes in Computer Science, 2317:344– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170344.
pdf.

Bogacz:2002:USD

[1378] Sally Bogacz and J. Gregory Trafton.

Understanding static and dynamic visualizations. Lecture Notes in Computer Science, 2317:347–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2317/23170347.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2317/23170347.pdf.

Cheng:2002:TST

[1379] Peter C-H. Cheng and Nigel G. Pitt.

Teaching science teachers electricity using AVOW diagrams. Lecture Notes in Computer Science, 2317:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2317/23170350.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2317/23170350.
pdf.

Dogan:2002:CDR

[1380] Fehmi Dogan and Nancy J. Nersessian. Conceptual diagrams: Representing ideas in design. Lecture Notes in Computer Science, 2317:353– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2317/23170353.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170353.
                                        pdf.

Peters:2002:SDC

[1381] Charlotte R. Peters and Patrick G. T. Healey. A survey of drawing in crosslinguistic communication. Lecture Notes in Computer Science, 2317:356– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170356.
                                        pdf.

Landay:2002:ITD

[1382] James A. Landay. Informal tools for designing anywhere, anytime, anydevice user interfaces. Lecture Notes in Computer Science, 2317:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2317/23170359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2317/23170359.
                                        pdf.

Anonymous:2002:AIan

[1383] Anonymous. Author index. Lecture Notes in Computer Science, 2317:361– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2317/2317auth.pdf.

Clarke:2002:SBC

[1384] Edmund M. Clarke. SAT-based counterexample guided abstraction refinement. Lecture Notes in Computer Science, 2318:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2318/23180001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2318/23180001.pdf.

Cousot:2002:AIT

[1385] Patrick Cousot. Abstract interpretation: Theory and practice. Lecture Notes in Computer Science, 2318:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180002.
pdf.

Ruys:2002:STH

[1386] Theo C. Ruys. SPIN tutorial:

How to become a SPIN doctor (extended abstract). Lecture Notes in Computer Science, 2318:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180006.
pdf.

Dams:2002:ASM

[1387] Dennis Dams. Abstraction in software model checking: Principles and practice (tutorial overview and bibliography). Lecture Notes in Computer Science, 2318:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180014.
pdf.

Iosif:2002:SRC

[1388] Radu Iosif. Symmetry reduction criteria for software model checking. Lecture Notes in Computer Science, 2318: 22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180022.
pdf.

Basin:2002:BMC

[1389] David Basin, Stefan Friedrich, Marek Gawkowski, and Joachim Posegga. Bytecode model checking: An experimental analysis. Lecture Notes in Computer Science, 2318:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180042.
pdf.

Li:2002:ISM

[1390] Harry C. Li, Kathi Fisler, and Shriram Krishnamurthi. The influence of software module systems on modular verification. Lecture Notes in Computer Science, 2318:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180060.
pdf.

Prigent:2002:ETS

[1391] Armelle Prigent, Franck Cassez, Philippe Dhaussy, and Olivier Roux. Extending the translation from SDL to Promela. Lecture Notes in Computer Science, 2318:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2318/23180079.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2318/23180079.pdf.

Hoek:2002:MCK

[1392] Wiebe van der Hoek and Michael Wooldridge. Model checking knowledge and time. Lecture Notes in Computer Science, 2318:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180095.
pdf.

Lluch-Lafuente:2002:POR

[1393] Alberto Lluch-Lafuente, Stefan Edelkamp, and Stefan Leue. Partial order reduction in directed model checking. Lecture Notes in Computer Science, 2318: 112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180112.
                                        pdf.

Bollig:2002:LPM

[1394] Benedikt Bollig, Martin Leucker, and Michael Weber. Local parallel model checking for the alternationfree μ-calculus. Lecture Notes in Computer Science, 2318:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180128.
                                        pdf.

Pascoe:2002:APP

[1395] J. S. Pascoe, R. J. Loader, and V. S. Sunderam. The agreement problem protocol verification environment. Lecture Notes in Computer Science, 2318: 148–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180148.
                                        pdf.

Jin:2002:BAG

[1396] Hyun-Wook Jin, Ki-Seok Bang, Chuck Yoo, Jin-Young Choi, and Ho jung Cha. Bottleneck analysis of a gigabit network interface card: Formal verification approach. Lecture Notes in Computer Science, 2318:170– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180170.
pdf.

Maggi:2002:USV

[1397] Paolo Maggi and Riccardo Sisto. Using SPIN to verify security properties of cryptographic protocols. Lecture Notes in Computer Science, 2318:187– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180187.
pdf.

Aygun:2002:MVI

[1398] Ramazan Savaş Aygün and Aidong Zhang. Modeling and verification of interactive flexible multimedia presentations using PROMELA/SPIN. Lecture Notes in Computer Science, 2318: 205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180205.
pdf.

Matlin:2002:SPS

[1399] Olga Shumsky Matlin, Ewing Lusk, and William McCune. SPINning parallel systems software. Lecture Notes in Computer Science, 2318:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180213.
                                        pdf.

Lugiez:2002:DBT

[1400] D. Lugiez, P. Niebert, and S. Zennou. Dynamic bounds and transition merging for local first search. Lecture Notes in Computer Science, 2318:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180221.
                                        pdf.

Eisner:2002:CSE

[1401] Cindy Eisner and Doron Peled. Comparing symbolic and explicit model checking of a software system. Lecture Notes in Computer Science, 2318: 230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2318/23180230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2318/23180230.
                                        pdf.

Bjesse:2002:IMC

[1402] Per Bjesse. Industrial model checking based on satisfiability solvers. Lecture Notes in Computer Science, 2318:240– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180240.
pdf.

Quemener:2002:TTP

[1403] Yves-Marie Quemener. A typical testing problem: Validating WML cellphones. Lecture Notes in Computer Science, 2318:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180241.
pdf.

Groce:2002:HMC

[1404] Alex Groce and Willem Visser. Heuristic model checking for Java programs. Lecture Notes in Computer Science, 2318:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2318/23180242.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2318/23180242.
pdf.

Stehno:2002:SSV

[1405] Christian Stehno. System specification and verification using high level concepts — A tool demonstration. Lecture Notes in Computer Science, 2318: 246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180246.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180246.
pdf.

Margaria:2002:DAI

[1406] Tiziana Margaria, Oliver Niese, and Bernhard Steffen. Demonstration of an automated integrated test environment for Web-based applications. Lecture Notes in Computer Science, 2318: 250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180250.
pdf.

Gallardo:2002:ESA

[1407] María del Mar Gallardo, Jesús Martínez, Pedro Merino, and Ernesto Pimentel. αSPIN: Extending SPIN with abstraction. Lecture Notes in Computer Science, 2318:254–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2318/23180254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2318/23180254.
pdf.

Anonymous:2002:AIao

[1408] Anonymous. Author index. Lecture Notes in Computer Science, 2318:259– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2318/2318auth.pdf.

Grechanik:2002:IRG

[1409] Mark Grechanik, Don Batory, and Dewayne E. Perry. Integrating and reusing GUI-driven applications. Lecture Notes in Computer Science, 2319: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190001.
pdf.

Jonge:2002:STC

[1410] Merijn de Jonge. Source tree composition. Lecture Notes in Computer Science, 2319:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2319/23190017.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2319/23190017.pdf.

Smaragdakis:2002:LDU

[1411] Yannis Smaragdakis. Layered development with (unix) dynamic libraries. Lecture Notes in Computer Science, 2319:33–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2319/23190033.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2319/23190033.
                                        pdf.

Pike:2002:ERC

[1412] Scott M. Pike and Nigamanth Sridhar. Early-reply components: Concurrent execution with sequential reasoning. Lecture Notes in Computer Science, 2319:46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2319/23190046.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190046.
                                        pdf.

Lee:2002:CGF

[1413] Kwanwoo Lee, Kyo C. Kang, and Jaejoon Lee. Concepts and guidelines of feature modeling for product line software engineering. Lecture Notes in Computer Science, 2319:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190062.
                                        pdf.

Gomaa:2002:DMW

[1414] Hassan Gomaa and Mark Gianturco. Domain modeling for World Wide Web based software product lines with UML. Lecture Notes in Computer Science, 2319:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2319/23190078.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2319/23190078.pdf.

Atkinson:2002:ECR

[1415] Colin Atkinson and Dirk Muthig. Enhancing component reusability through product line technology. Lecture Notes in Computer Science, 2319:93– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190093.
pdf.

Webber:2002:MVV

[1416] Diana L. Webber and Hassan Gomaa. Modeling variability with the variation point model. Lecture Notes in Computer Science, 2319:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190109.
pdf.

Brown:2002:ROS

[1417] Alan W. Brown and Grady Booch.

Reusing open-source software and practices: The impact of open-source on commercial vendors. Lecture Notes in Computer Science, 2319:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190123.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2319/23190123.
                                        pdf.

Schmid:2002:IRA

[1418] Klaus Schmid. Integrating reference architecture definition and reuse investment planning. Lecture Notes in Computer Science, 2319:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190137.
                                        pdf.

Biggerstaff:2002:CLD

[1419] Ted J. Biggerstaff. Control localization in domain specific translation. Lecture Notes in Computer Science, 2319:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190153.
                                        pdf.

Levendovszky:2002:MRM

[1420] Tihamer Levendovszky, Gabor Karsai, Miklos Maroti, Akos Ledeczi, and Hassan Charaf. Model reuse with metamodel-based transformations. Lecture Notes in Computer Science, 2319:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2319/23190166.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2319/23190166.pdf.

Alonso:2002:GTS

[1421] Omar Alonso. Generation of text search applications for databases. an exercise on domain engineering. Lecture Notes in Computer Science, 2319: 179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190179.
pdf.

Bergmann:2002:DNS

[1422] Ulf Bergmann and Julio Cesar Sampaio do Prado Leite. Domain networks in the software development process. Lecture Notes in Computer Science, 2319: 194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190194.
pdf.

Biddle:2002:SRU

[1423] Robert Biddle, James Noble, and Ewan Tempero. Supporting reusable use cases. Lecture Notes in Computer Science, 2319:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2319/23190210.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2319/23190210.pdf.

Barros:2002:PMK

[1424] Márcio de O. Barros, Cláudia M. L. Werner, and Guilherme H. Travassos. Project management knowledge reuse through scenario models. Lecture Notes in Computer Science, 2319: 227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190227.
                                        pdf.

Gorgonio:2002:ACP

[1425] Kyller Costa Gorgônio and Angelo Perkusich. Adaptation of coloured Petri nets models of software artifacts for reuse. Lecture Notes in Computer Science, 2319:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190240.
                                        pdf.

Smith:2002:IHC

[1426] Shamus P. Smith and Michael D. Harrison. Improving hazard classification through the reuse of descriptive arguments. Lecture Notes in Computer Science, 2319:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190255.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2319/23190255.
pdf.

Sillitti:2002:SOP

[1427] Alberto Sillitti, Tullio Vernazza, and Giancarlo Succi. Service oriented programming: a new paradigm of software reuse. Lecture Notes in Computer Science, 2319:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190269.
pdf.

Ye:2002:EUS

[1428] Yunwen Ye. An empirical user study of an active reuse repository system. Lecture Notes in Computer Science, 2319: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190281.
pdf.

Caballero:2002:TFR

[1429] Rodrigo E. Caballero, Steven A. Demurjian, and Sr. Towards the formalization of a reusability framework for refactoring. Lecture Notes in Computer Science, 2319:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190293.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2319/23190293.
                                        pdf.

Sridhar:2002:SFE

[1430] Nigamanth Sridhar, Bruce W. Weide, and Paolo Bucci. Service facilities: Extending abstract factories to decouple advanced dependencies. Lecture Notes in Computer Science, 2319:309– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190309.
                                        pdf.

Sessions:2002:SF

[1431] Roger Sessions. Software fortresses. Lecture Notes in Computer Science, 2319:327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2319/23190327.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190327.
                                        pdf.

Misra:2002:CAG

[1432] Jayadev Misra. The case against a grand unification theory. Lecture Notes in Computer Science, 2319:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190328.
                                        pdf.

Boellert:2002:IYR

[1433] Kai Boellert, Detlef Streitferdt, and Dirk Heuzeroth. ICSR7 young researchers workshop. Lecture Notes in Computer Science, 2319:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190329.
pdf.

Favaro:2002:IWR

[1434] John Favaro, Hakan Erdogmus, and Klaus Schmid. International workshop on reuse economics. Lecture Notes in Computer Science, 2319:330– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190330.
pdf.

Jonge:2002:WGP

[1435] Merijn de Jonge and Joost Visser.

Workshop on generative programming 2002 (GP2002). Lecture Notes in Computer Science, 2319:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190331.
pdf.

Zeidler:2002:IWC

[1436] Christian Zeidler. ICSR7 workshop on component-based software development processes. Lecture Notes in Computer Science, 2319:332–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190332.
                                        pdf.

Cohen:2002:IEP

[1437] Sholom Cohen. Industrial experience with product line approaches. Lecture Notes in Computer Science, 2319:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190334.
                                        pdf.

Yoder:2002:WSR

[1438] Joseph W. Yoder. Workshop on software reuse and agile approaches. Lecture Notes in Computer Science, 2319: 336–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190336.
                                        pdf.

Dobrica:2002:SAQ

[1439] Liliana Dobrica and Eila Niemelä. Software architecture quality analysis methods. Lecture Notes in Computer Science, 2319:337–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2319/23190337.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2319/23190337.pdf.

Schmid:2002:TPP

[1440] Klaus Schmid and Isabel John. Tutorial on practical product line scoping and modeling. Lecture Notes in Computer Science, 2319:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190339.
pdf.

Baxter:2002:TSG

[1441] Ira D. Baxter. Transformation systems:

Generative reuse for software generation, maintenance and reengineering. Lecture Notes in Computer Science, 2319:341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2319/23190341.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2319/23190341.
pdf.

Atkinson:2002:CBP

[1442] Colin Atkinson and Dirk Muthig.

Component-based product-line engineering with the UML. Lecture Notes in Computer Science, 2319:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190343.
                                        pdf.

McGregor:2002:BRT

[1443] John D. McGregor. Building reusable test assets for a product line. Lecture Notes in Computer Science, 2319:345– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190345.
                                        pdf.

Bosch:2002:ACS

[1444] Jan Bosch. Architecture-centric software engineering. Lecture Notes in Computer Science, 2319:347–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2319/23190347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2319/23190347.
                                        pdf.

Krueger:2002:PST

[1445] Charles W. Krueger. Practical strategies and techniques for adopting software product lines. Lecture Notes in Computer Science, 2319:349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2319/23190349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190349.
pdf.

Anonymous:2002:AIap

[1446] Anonymous. Author index. Lecture Notes in Computer Science, 2319:351– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2319/2319auth.pdf.

Czarnecki:2002:GPM

[1447] Krzysztof Czarnecki. Generative programming: Methods, techniques, and applications. Lecture Notes in Computer Science, 2319:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2319/23190351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2319/23190351.
pdf.

Jakobsson:2002:DSP

[1448] Markus Jakobsson and Michael K. Reiter. Discouraging software piracy using software aging. Lecture Notes in Computer Science, 2320:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200001.
pdf.

Mihcak:2002:NIG

[1449] M. Kivanç Mihçak and Ramarathnam Venkatesan. New iterative geometric methods for robust perceptual image hashing. Lecture Notes in Computer Science, 2320:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200013.
                                        pdf.

Kiayias:2002:CPF

[1450] Aggelos Kiayias and Moti Yung. On crafty pirates and foxy tracers. Lecture Notes in Computer Science, 2320: 22–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200022.
                                        pdf.

Pinkas:2002:ESU

[1451] Benny Pinkas. Efficient state updates for key management. Lecture Notes in Computer Science, 2320:40– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200040.
                                        pdf.

Safavi-Naini:2002:CSA

[1452] Reihaneh Safavi-Naini and Yejing Wang. Collusion secure q-ary fingerprinting for perceptual content. Lecture Notes in Computer Science, 2320: 57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200057.
pdf.

Feigenbaum:2002:PED

[1453] Joan Feigenbaum, Michael J. Freedman, Tomas Sander, and Adam Shostack. Privacy engineering for digital rights management systems. Lecture Notes in Computer Science, 2320: 76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200076.
pdf.

Kravitz:2002:SOS

[1454] David Kravitz, Kim-Ee Yeoh, and Nicol So. Secure open systems for protecting privacy and digital services. Lecture Notes in Computer Science, 2320:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2320/23200106.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2320/23200106.
pdf.

King:2002:MIE

[1455] James King and Panos Kudumakis. MPEG-4 IPMP extensions. Lecture Notes in Computer Science, 2320:126– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200126.
                                        pdf.

Horne:2002:DSC

[1456] Bill Horne, Lesley Matheson, Casey Sheehan, and Robert E. Tarjan. Dynamic self-checking techniques for improved tamper resistance. Lecture Notes in Computer Science, 2320:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200141.
                                        pdf.

Chang:2002:PSC

[1457] Hoi Chang and Mikhail J. Atallah. Protecting software code by guards. Lecture Notes in Computer Science, 2320: 160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200160.
                                        pdf.

Shapiro:2002:HMP

[1458] William Shapiro and Radek Vingralek. How to manage persistent state in DRM systems. Lecture Notes in Computer Science, 2320:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200176.
pdf.

Crosby:2002:CHB

[1459] Scott Crosby, Ian Goldberg, Robert Johnson, Dawn Song, and David Wagner. A cryptanalysis of the high-bandwidth digital content protection system. Lecture Notes in Computer Science, 2320:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200192.
pdf.

Buhse:2002:IDR

[1460] Willms Buhse. Implications of digital rights management for online music — A business perspective. Lecture Notes in Computer Science, 2320:201– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2320/23200201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2320/23200201.
pdf.

Bechtold:2002:CIL

[1461] Stefan Bechtold. From copyright to information law — implications of digital rights management. Lecture Notes in Computer Science, 2320:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200213.
                                        pdf.

Miller:2002:TCC

[1462] Ernest Miller and Joan Feigenbaum. Taking the copy out of copyright. Lecture Notes in Computer Science, 2320: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2320/23200233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2320/23200233.
                                        pdf.

Anonymous:2002:AIaq

[1463] Anonymous. Author index. Lecture Notes in Computer Science, 2320:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2320/2320ind.pdf.

Butz:2002:BEA

[1464] Martin V. Butz. Biasing exploration in an anticipatory learning classifier system. Lecture Notes in Computer Science, 2321:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210003.
pdf.

Davis:2002:IMP

[1465] Lawrence Davis, Chunsheng Fu, and Stewart W. Wilson. An incremental multiplexer problem and its uses in classifier system research. Lecture Notes in Computer Science, 2321:23– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210023.
pdf.

Enee:2002:MMC

[1466] Gilles Énée and Cathy Escazut. A minimal model of communication for a multi-agent classifier system. Lecture Notes in Computer Science, 2321: 32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210032.
pdf.

Holmes:2002:RAB

[1467] John H. Holmes. A representation for accuracy-based assessment of classifier system prediction performance. Lecture Notes in Computer Science, 2321: 43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2321/23210043.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2321/23210043.
                                        pdf.

Hurst:2002:SAX

[1468] Jacob Hurst and Larry Bull. A self-adaptive XCS. Lecture Notes in Computer Science, 2321:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2321/23210057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2321/23210057.
                                        pdf.

Kovacs:2002:TVC

[1469] Tim Kovacs. Two views of classifier systems. Lecture Notes in Computer Science, 2321:74–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2321/23210074.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2321/23210074.pdf.

Hercog:2002:SSU

[1470] Luis Miramontes Hercog and Terence C. Fogarty. Social simulation using a multi-agent model based on classifier systems: The emergence of vacillating behaviour in the “El Farol” bar problem. Lecture Notes in Computer Science, 2321:88– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210088.
pdf.

Bernado:2002:XGC

[1471] Ester Bernadó, Xavier Llorà, and Josep M. Garrell. XCS and GALE: a comparative study of two learning classifier systems on data mining. Lecture Notes in Computer Science, 2321:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210115.
pdf.

Dixon:2002:PIM

[1472] Phillip William Dixon, David W.

Corne, and Martin John Oates. A preliminary investigation of modified XCS as a generic data mining tool. Lecture Notes in Computer Science, 2321:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210133.
pdf.

Schulenburg:2002:ELM

[1473] Sonia Schulenburg and Peter Ross.

Explorations in LCS models of stock trading. Lecture Notes in Computer Science, 2321:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2321/23210151.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2321/23210151.pdf.

Vargas:2002:LAL

[1474] Patrícia Amâncio Vargas, Christiano LyraFilho, and Fernando J. Von Zuben. On-line approach for loss reduction in electric power distribution networks using learning classifier systems. Lecture Notes in Computer Science, 2321:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2321/23210181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2321/23210181.
                                        pdf.

Wilson:2002:CRX

[1475] Stewart W. Wilson. Compact rulesets from XCSI. Lecture Notes in Computer Science, 2321:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2321/23210197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2321/23210197.
                                        pdf.

Butz:2002:ADA

[1476] Martin V. Butz and Wolfgang Stolzmann. An algorithmic description of ACS2. Lecture Notes in Computer Science, 2321:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2321/23210211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2321/23210211.
pdf.

Anonymous:2002:AIar

[1477] Anonymous. Author index. Lecture Notes in Computer Science, 2321:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2321/23210231.pdf.

Wooldridge:2002:IAK

[1478] Michael Wooldridge. Intelligent agents: The key concepts. Lecture Notes in Computer Science, 2322:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220003.
pdf.

Lioncourt:2002:MAB

[1479] Sorabain Wolfheart de Lioncourt and Michael Luck. Motivated agent behaviour and requirements applied to virtual emergencies. Lecture Notes in Computer Science, 2322:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220044.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2322/23220044.
                                        pdf.

Muller:2002:SAE

[1480] Jörg P. Müller, Bernhard Bauer, and Michael Berger. Software agents for electronic business: Opportunities and challenges. Lecture Notes in Computer Science, 2322:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220061.
                                        pdf.

Andre:2002:SDI

[1481] Elisabeth André, Thomas Rist, and Stephan Baldes. From simulated dialogues to interactive performances. Lecture Notes in Computer Science, 2322:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2322/23220107.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220107.
                                        pdf.

Deen:2002:CAH

[1482] S. Misbah Deen. Cooperating agents for holonic manufacturing. Lecture Notes in Computer Science, 2322:119– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220119.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2322/23220119.
pdf.

Giacomin:2002:SSD

[1483] Massimiliano Giacomin. Self-stabilizing distributed algorithms for defeat status computation in argumentation. Lecture Notes in Computer Science, 2322: 137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220137.
pdf.

Krogt:2002:RRB

[1484] Roman van der Krogt, André Bos, and Cees Witteveen. Replanning in a resource-based framework. Lecture Notes in Computer Science, 2322:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220148.
pdf.

Barta:2002:DBB

[1485] Jaroslav Bárta, Olga Štépánková, and Michal Pěchouček. Distributed branch and bound algorithm in coalition planning. Lecture Notes in Computer Science, 2322:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220159.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2322/23220159.
                                        pdf.

Flores:2002:FCC

[1486] Roberto A. Flores and Robert C. Kremer. Formal conversations for the contract net protocol. Lecture Notes in Computer Science, 2322:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220169.
                                        pdf.

Carabelea:2002:AAA

[1487] Cosmin Carabelea. Adaptive agents in argumentation-based negotiation. Lecture Notes in Computer Science, 2322: 180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220180.
                                        pdf.

Gomoluch:2002:FLB

[1488] Jacek Gomoluch and Michael Schroeder. Flexible load balancing in distributed information agent systems. Lecture Notes in Computer Science, 2322:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220188.
                                        pdf.

Gozzi:2002:AAA

[1489] Andrea Gozzi, Massimo Paolucci, and Antonio Boccalatte. Autonomous agents applied to manufacturing scheduling problems: a negotiationbased heuristic approach. Lecture Notes in Computer Science, 2322:198– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220198.
pdf.

Briot:2002:AEU

[1490] Jean-Pierre Briot, Angélica Muñoz-Meléndez, and Alain Cardon. Adaptability and embodiment using multiagent systems. Lecture Notes in Computer Science, 2322:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220211.
pdf.

Brazier:2002:DSA

[1491] Frances Brazier, Maarten van Steen, and Niek Wijngaards. Distributed shared agent representations. Lecture Notes in Computer Science, 2322:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220213.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2322/23220213.
                                        pdf.

Landau:2002:MAM

[1492] Samuel Landau and Sébastien Picault. Modeling adaptive multi-agent systems inspired by developmental biology. Lecture Notes in Computer Science, 2322: 221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220221.
                                        pdf.

Marik:2002:HAR

[1493] Vladimír Mařík, Martyn Fletcher, and Michal Pěchouček. Holons & agents: Recent developments and mutual impacts. Lecture Notes in Computer Science, 2322:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220233.
                                        pdf.

Vrba:2002:MHP

[1494] Pavel Vrba and Václav Hrdonka. Material handling problem: FIPA compliant agent implementation. Lecture Notes in Computer Science, 2322:268– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220268.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2322/23220268.
pdf.

Tichy:2002:IMP

[1495] Pavel Tichý, Petr Šlechta, Francisco Maturana, and Sivaram Balasubramanian. Industrial MAS for planning and control. Lecture Notes in Computer Science, 2322:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220280.
pdf.

Glanzer:2002:IIM

[1496] Klaus Glanzer, Alexander Hämmerle, and Ralf Geurts. An integral implementation of a machine-holon applying the ZEUS agent framework. Lecture Notes in Computer Science, 2322:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220296.
pdf.

Riha:2002:EEA

[1497] Aleš Ríha, Michal Pěchouček, Jirí Vokřínek, and Vladimír Mařík. Ex-PlanTech: Exploitation of agent-based technology in production planning. Lecture Notes in Computer Science, 2322:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2322/23220308.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220308.
                                        pdf.

Brennan:2002:HAR

[1498] Robert W. Brennan, Martyn Fletcher, and Douglas H. Norrie. A holonic approach to reconfiguring real-time distributed control systems. Lecture Notes in Computer Science, 2322:323– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220323.
                                        pdf.

Chokshi:2002:RHA

[1499] N. N. Chokshi and D. C. McFarlane. Rationales for holonic applications in chemical process industries. Lecture Notes in Computer Science, 2322:336– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2322/23220336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2322/23220336.
                                        pdf.

Fletcher:2002:DDC

[1500] Martyn Fletcher, Robert W. Brennan, and Douglas H. Norrie. Distributed deadline control in holonic manufacturing systems. Lecture Notes in Computer Science, 2322:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2322/23220351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220351.
pdf.

Pechoucek:2002:CCP

[1501] Michal Pěchouček, Jaroslav Bárta, and Vladimír Mařík. CPlanT: Coalition planning multi-agent system for humanitarian relief operations. Lecture Notes in Computer Science, 2322:363– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2322/23220363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2322/23220363.
pdf.

Anonymous:2002:AIas

[1502] Anonymous. Author index. Lecture Notes in Computer Science, 2322:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2322/2322auth.pdf.

Ardourel:2002:WPS

[1503] Gilles Ardourel, Michael Haupt, José Luis Herrero Agustín, Rainer Ruggaber, and Charles Suscheck. 11th workshop for PhD students in objectoriented systems. Lecture Notes in Computer Science, 2323:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2323/23230001.
                                        pdf.

Boszormenyi:2002:QSD

[1504] László Böszörményi, Christian Stary, Harald Kosch, and Christian Becker. Quality of service in distributed object systems and distributed multimedia object/component systems. Lecture Notes in Computer Science, 2323: 7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230007.
                                        pdf.

Leavens:2002:FTJ

[1505] Gary T. Leavens, Sophia Drossopoulou, Susan Eisenbach, Arnd Poetzsch-Heffter, and Erik Poll. Formal techniques for Java programs. Lecture Notes in Computer Science, 2323:30– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230030.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230030.
                                        pdf.

Tekinerdogan:2002:AOO

[1506] Bedir Tekinerdoğan, Motoshi Saeki, Gerson Sunyé, Pim van den Broek, and Pavel Hruby. Automating objectoriented software development methods. Lecture Notes in Computer Science, 2323:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2323/23230041.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2323/23230041.pdf.

Revault:2002:AOM

[1507] Nicolas Revault and Joseph W. Yoder. Adaptive object-models and metamodeling techniques. Lecture Notes in Computer Science, 2323:57– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230057.
pdf.

Gerard:2002:SIV

[1508] Sébastien Gerard, Alexandre David, and François Terrier. Specification, implementation, and validation of objectoriented embedded systems. Lecture Notes in Computer Science, 2323:72– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230072.
pdf.

Pulvermueller:2002:FIC

[1509] Elke Pulvermueller, Andreas Speck, James O. Coplien, Maja D’Hondt, and Wolfgang De Meuter. Feature interaction in composed systems. Lecture Notes in Computer Science, 2323:86– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230086.
                                        pdf.

Bosch:2002:WCO

[1510] Jan Bosch, Clemens Szyperski, and Wolfgang Weck. 6th workshop on component-oriented programming. Lecture Notes in Computer Science, 2323:98–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2323/23230098.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230098.
                                        pdf.

Brichau:2002:ASC

[1511] Johan Brichau, Maurice Glandrup, Siobhan Clarke, and Lodewijk Bergmans. Advanced separation of concerns. Lecture Notes in Computer Science, 2323: 107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230107.
                                        pdf.

Davis:2002:MPO

[1512] Kei Davis, Yannis Smaragdakis, and Jörg Striegnitz. Multiparadigm programming with OO languages. Lecture Notes in Computer Science, 2323:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2323/23230131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230131.
pdf.

Barth:2002:GP

[1513] Barbara Barth, Greg Butler, Krzysztof Czarnecki, and Ulrich Eisenecker. Generative programming. Lecture Notes in Computer Science, 2323:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230135.
pdf.

Mens:2002:WOO

[1514] Tom Mens and Galal Hassan Galal.

4th workshop on object-oriented architectural evolution. Lecture Notes in Computer Science, 2323:150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230150.
pdf.

Bryce:2002:WMO

[1515] Ciarán Bryce and Alexander Romanovsky. 7th workshop on mobile object systems. Lecture Notes in Computer Science, 2323:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2323/23230165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230165.
                                        pdf.

Jul:2002:NDO

[1516] Eric Jul, Andrew Black, Anne-Marie Kermarrec, Doug Lea, and Salah Sadou. The next 700 distributed object systems. Lecture Notes in Computer Science, 2323:169– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230169.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230169.
                                        pdf.

BritoeAbreu:2002:QAO

[1517] Fernando Brito e Abreu, Brian Henderson-Sellers, Mario Piattini, Geert Poels, and Houari A. Sahraoui. Quantitative approaches in objectoriented software engineering. Lecture Notes in Computer Science, 2323:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230174.
                                        pdf.

Corchuelo:2002:OOB

[1518] Rafael Corchuelo, Antonio Ruiz, Jörg R. Mühlbacher, and Jesús D. García-Consuegra. Object-oriented business solutions. Lecture Notes in Computer Science, 2323:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230184.
pdf.

Gutierrez:2002:WOO

[1519] Darío Álvarez Gutiérrez, Francisco Ballesteros, and Paniti Netinant. 4th workshop on OO and operating systems. Lecture Notes in Computer Science, 2323:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230201.
pdf.

Jul:2002:PND

[1520] Eric Jul, Andrew Black, Doug Lea, Robert Filman, and Ian Welch. Panel: The next 700 distributed object systems. Lecture Notes in Computer Science, 2323:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2323/23230208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2323/23230208.
pdf.

Zsok:2002:PS

[1521] Viktória Zsók and István Juhász.

Poster session. Lecture Notes in Computer Science, 2323:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2323/23230213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2323/23230213.
                                        pdf.

Anonymous:2002:AIat

[1522] Anonymous. Author index. Lecture Notes in Computer Science, 2323:225– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2323/2323auth.pdf.

Borst:2002:HTE

[1523] S. C. Borst, O. J. Boxma, and R. Núñez-Queija. Heavy tails: The effect of the service discipline. Lecture Notes in Computer Science, 2324: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240001.
                                        pdf.

Derisavi:2002:MSL

[1524] Salem Derisavi, Peter Kemper, William H. Sanders, and Tod Courtney. The Möbius state-level abstract functional interface. Lecture Notes in Computer Science, 2324:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240031.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2324/23240031.
pdf.

Bause:2002:PBT

[1525] F. Bause, H. Beilner, M. Fischer, P. Kemper, and M. Völker. The ProC/B toolset for the modelling and analysis of process chains. Lecture Notes in Computer Science, 2324:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240051.
pdf.

Racz:2002:MDE

[1526] Sándor Rácz, Árpád Tari, and Miklós Telek. MRMSolve: Distribution estimation of large Markov reward models. Lecture Notes in Computer Science, 2324:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240071.
pdf.

Horvath:2002:PGP

[1527] András Horváth and Miklós Telek. Ph-Fit: a general phase-type fitting tool. Lecture Notes in Computer Science, 2324:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2324/23240082.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2324/23240082.
                                        pdf.

Klemm:2002:TMI

[1528] Alexander Klemm, Christoph Lindemann, and Marco Lohmann. Traffic modeling of IP networks using the batch Markovian arrival process. Lecture Notes in Computer Science, 2324: 92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240092.
                                        pdf.

Gilmore:2002:PNS

[1529] Stephen Gilmore, Jane Hillston, and Marina Ribaudo. PEPA nets: a structured performance modelling formalism. Lecture Notes in Computer Science, 2324:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240111.
                                        pdf.

Donatelli:2002:VGS

[1530] Susanna Donatelli and Liliana Ferro. Validation of GSPN and SWN models through the PROD tool. Lecture Notes in Computer Science, 2324:131– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240131.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2324/23240131.
pdf.

Petriu:2002:SPM

[1531] Dorin Petriu and Murray Woodside.

Software performance models from system scenarios in use case maps. Lecture Notes in Computer Science, 2324:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240141.
pdf.

Petriu:2002:AUP

[1532] Dorina C. Petriu and Hui Shen. Applying the UML performance profile: Graph grammar-based derivation of LQN models from UML specifications. Lecture Notes in Computer Science, 2324:159–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2324/23240159.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2324/23240159.
pdf.

Bradley:2002:PTP

[1533] J. T. Bradley. A passage-time preserving equivalence for semi-Markov processes. Lecture Notes in Computer Science, 2324:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240178.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240178.
                                        pdf.

Davies:2002:SMS

[1534] Ian Davies, William J. Knottenbelt, and Pieter S. Kritzinger. Symbolic methods for the state space exploration of GSPN models. Lecture Notes in Computer Science, 2324:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240188.
                                        pdf.

Kwiatkowska:2002:PPS

[1535] Marta Kwiatkowska, Gethin Norman, and David Parker. PRISM: Probabilistic symbolic model checker. Lecture Notes in Computer Science, 2324:200– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240200.
                                        pdf.

Riska:2002:MMA

[1536] Alma Riska and Evgenia Smirni. MAMSolver: a matrix analytic methods tool. Lecture Notes in Computer Science, 2324:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2324/23240205.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2324/23240205.pdf.

Festag:2002:MSE

[1537] Andreas Festag, Lars Westerhoff, and Adam Wolisz. The MOMBASA software environment — A toolkit for performance evaluation of multicastbased mobility support. Lecture Notes in Computer Science, 2324:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240212.
pdf.

Pearce:2002:GDI

[1538] David J. Pearce, Paul H. J. Kelly, Tony Field, and Uli Harder. GILK: a dynamic instrumentation tool for the Linux kernel. Lecture Notes in Computer Science, 2324:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240220.
pdf.

Kant:2002:GWT

[1539] Krishna Kant, Vijay Tewari, and Ravishankar Iyer. Geist: a Web traffic generation tool. Lecture Notes in Computer Science, 2324:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2324/23240227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240227.
                                        pdf.

Franceschinis:2002:DMO

[1540] Giuliana Franceschinis, M. Gribaudo, M. Iacono, N. Mazzocca, and V. Vittorini. DrawNET++: Model objects to support performance analysis and simulation of systems. Lecture Notes in Computer Science, 2324:233– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240233.
                                        pdf.

Hopkins:2002:VFC

[1541] Richard Hopkins and Peter King. A visual formalism for the composition of stochastic petri nets. Lecture Notes in Computer Science, 2324:239– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240239.
                                        pdf.

Gautama:2002:SPP

[1542] Hasyim Gautama and Arjan J. C. van Gemund. Symbolic performance prediction of data-dependent parallel programs. Lecture Notes in Computer Science, 2324:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240259.
pdf.

Kumar:2002:ATS

[1543] L. Kumar, M. Misra, and I. Mitrani. Analysis of a transaction system with checkpointing, failures, and rollback. Lecture Notes in Computer Science, 2324:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240279.
pdf.

Conway:2002:AVI

[1544] Adrian E. Conway and Yali Zhu.

Analyzing voice-over-IP subjective quality as a function of network QoS: a simulation-based methodology and tool. Lecture Notes in Computer Science, 2324:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2324/23240289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2324/23240289.
pdf.

Oke:2002:HWC

[1545] Adeniyi Oke and Rick Bunt. Hierarchical workload characterization for a busy Web server. Lecture Notes in Computer Science, 2324:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240309.
                                        pdf.

Garg:2002:ENC

[1546] Pankaj K. Garg, Kave Eshghi, Thomas Gschwind, Boudewijn Haverkort, and Katinka Wolter. Enabling network caching of dynamic Web objects. Lecture Notes in Computer Science, 2324: 329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2324/23240329.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240329.
                                        pdf.

Richard:2002:TCR

[1547] Pascal Richard. A tool for controlling response time in real-time systems. Lecture Notes in Computer Science, 2324:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2324/23240339.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2324/23240339.
                                        pdf.

Anonymous:2002:AIau

[1548] Anonymous. Author index. Lecture Notes in Computer Science, 2324:349– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2324/2324auth.pdf.

Gentzsch:2002:GCN

[1549] Wolfgang Gentzsch. Grid computing: a new technology for the advanced Web. Lecture Notes in Computer Science, 2326:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260001.
pdf.

Folliot:2002:BFR

[1550] B. Folliot, I. Piumarta, L. Seinturier, C. Baillarguet, C. Khoury, A. Leger, and F. Ogel. Beyond flexibility and reflection: The virtual virtual machine approach. Lecture Notes in Computer Science, 2326:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260016.
pdf.

Grigoras:2002:PMC

[1551] Dan Grigoras. Programming models for cluster computing. Lecture Notes in Computer Science, 2326:26– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260026.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2326/23260026.
                                        pdf.

Toursel:2002:ADE

[1552] Bernard Toursel. About design and efficiency of distributed programming: Some algorithmic aspects. Lecture Notes in Computer Science, 2326:36– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260036.
                                        pdf.

Cristea:2002:CEH

[1553] Valentin Cristea. A collaborative environment for high performance computing. Lecture Notes in Computer Science, 2326:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260047.
                                        pdf.

Tan:2002:LFT

[1554] Guang Tan, Hai Jin, and Liping Pang. Layer 4 fault tolerance: Reliability techniques for cluster system in Internet services. Lecture Notes in Computer Science, 2326:60– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260060.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2326/23260060.
pdf.

Ozturan:2002:CIC

[1555] Can Özturan. Computational issues in E-commerce trading exchanges. Lecture Notes in Computer Science, 2326: 69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260069.
pdf.

Tudruj:2002:ECC

[1556] Marek Tudruj. Embedded cluster computing through dynamic reconfigurability of inter-processor connections. Lecture Notes in Computer Science, 2326: 77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260077.
pdf.

Dedu:2002:PWP

[1557] Eugen Dedu, Stéphane Vialle, and Claude Timsit. Parallelisation of wave propagation algorithms for odour propagation in multi-agent systems. Lecture Notes in Computer Science, 2326:92– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260092.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2326/23260092.
pdf.

Craus:2002:PDS

[1558] Mitică Craus. Parallel and distributed solutions for the optimal binary search tree problem. Lecture Notes in Computer Science, 2326:103– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260103.
pdf.

Grigori:2002:FES

[1559] Laura Grigori. A framework for efficient sparse LU factorization in a cluster based platform. Lecture Notes in Computer Science, 2326:118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260118.
pdf.

Monteil:2002:TAU

[1560] Thierry Monteil and Patricia Pascal. Task allocation using processor load prediction on multiprocessors cluster. Lecture Notes in Computer Science, 2326:126–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260126.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2326/23260126.
pdf.

Rotaru:2002:HDLa

[1561] Tiberiu Rotaru and Hans-Heinrich Nägeli. Heterogeneous dynamic load balancing. Lecture Notes in Computer Science, 2326:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260136.
pdf.

Kebbal:2002:MAS

[1562] D. Kebbal, E.-G. Talbi, and J.-M.

Geib. Multi-application scheduling in networks of workstations and clusters of processors. Lecture Notes in Computer Science, 2326:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260145.
pdf.

Jordan:2002:PAC

[1563] Andrzej Jordan and Robert Piotr Bycul. The parallel algorithm of conjugate gradient method. Lecture Notes in Computer Science, 2326:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260156.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2326/23260156.
                                        pdf.

Alpkocak:2002:PSS

[1564] Adil Alpkocak, Taner Danisman, and Tuba Ulker. A parallel similarity search in high dimensional metric space using M-tree. Lecture Notes in Computer Science, 2326:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260166.
                                        pdf.

Petcu:2002:ULI

[1565] Dana Petcu, Horia Popa, and Dorin Ţepeneu. A user-level interface for clustering mathematical software kernels. Lecture Notes in Computer Science, 2326:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260172.
                                        pdf.

Antoniu:2002:IMP

[1566] Gabriel Antoniu and Luc Bougé. Implementing multithreaded protocols for release consistency on top of the generic DSM-PM platform. Lecture Notes in Computer Science, 2326:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2326/23260179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260179.
pdf.

Ivan:2002:CBW

[1567] Cosmina Ivan. Communication balancing for wireless nodes in a LAN-WAN cluster. Lecture Notes in Computer Science, 2326:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260189.
pdf.

Tudruj:2002:ATS

[1568] Marek Tudruj and Lukasz Masko.

An architecture and task scheduling algorithm for systems based on dynamically reconfigurable shared memory clusters. Lecture Notes in Computer Science, 2326:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260197.
pdf.

Tapus:2002:DRE

[1569] Nicolae Tapus, Emil Slusanschi, and Tudor Popescu. Distributed rendering engine. Lecture Notes in Computer Science, 2326:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260207.
                                        pdf.

Tapus:2002:DAM

[1570] Nicolae Tapus, Mihai Burcea, and Vlad Staicu. A distributed application for Monte Carlo simulations. Lecture Notes in Computer Science, 2326:216– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260216.
                                        pdf.

Buraga:2002:MAR

[1571] Sabin C. Buraga. A model for accessing resources of the distributed file systems. Lecture Notes in Computer Science, 2326:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2326/23260224.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2326/23260224.pdf.

Ionescu:2002:CGP

[1572] Felicia Ionescu, Mihail Ionescu, Cristina Coconu, and Valentin Stoica. Coarse-grain parallelization of test vectors generation on multiprocessor systems. Lecture Notes in Computer Science, 2326:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2326/23260231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260231.
pdf.

Smyk:2002:OMP

[1573] Adam Smyk and Marek Tudruj.

OpenMP/MPI programming in a multi-cluster system based on shared memory/message passing communication. Lecture Notes in Computer Science, 2326:241–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2326/23260241.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2326/23260241.pdf.

Felea:2002:EPJ

[1574] Violeta Felea, Nathalie Devesa, Bernard Toursel, and Pierre Lecouffe. Expressing parallelism in Java applications distributed on clusters. Lecture Notes in Computer Science, 2326:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260249.
pdf.

Boian:2002:ACT

[1575] Florian Mircea Boian and Corina Ferdean. Advanced collaboration techniques between Java objects distributed on clusters. Lecture Notes in Computer Science, 2326:259–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260259.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260259.
                                        pdf.

Bouchi:2002:JTM

[1576] A. Bouchi, R. Olejnik, and Bernard Toursel. Java tools for measurement of the machine loads. Lecture Notes in Computer Science, 2326:271– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260271.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260271.
                                        pdf.

Butrylo:2002:PFA

[1577] Boguslaw Butrylo. The parallel FDTD algorithm and its application to the analysis of complex structures. Lecture Notes in Computer Science, 2326: 279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2326/23260279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2326/23260279.
                                        pdf.

Tabirca:2002:TAF

[1578] Tatiana Tabirca, Len Freeman, and Sabin Tabirca. A theoretical application of feedback guided dynamic loop scheduling. Lecture Notes in Computer Science, 2326:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260287.
pdf.

Mitrea:2002:DDI

[1579] Paulina Mitrea. DIM — A distributed image processing system, based on ISO IEC 12087 image processing standard. Lecture Notes in Computer Science, 2326:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260293.
pdf.

Butincu:2002:DDA

[1580] Cristian Buţincu and Dan Grigoras.

DisMedJava — A distributed application for medical image processing. Lecture Notes in Computer Science, 2326: 308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2326/23260308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2326/23260308.
pdf.

Anonymous:2002:AIav

[1581] Anonymous. Author index. Lecture Notes in Computer Science, 2326:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2326/2326auth.pdf.

Sterling:2002:GMP

[1582] Thomas Sterling. The gilgamesh MIND processor-in-memory architecture for petaflops-scale computing. Lecture Notes in Computer Science, 2327:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270001.
                                        pdf.

Hey:2002:USP

[1583] Tony Hey. The UK e-science program and the Grid. Lecture Notes in Computer Science, 2327:6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270006.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270006.
                                        pdf.

Eigenmann:2002:SHN

[1584] Rudolf Eigenmann, Greg Gaertner, Wesley Jones, Hideki Saito, and Brian Whitney. SPEC HPC2002: The next high-performance computer benchmark. Lecture Notes in Computer Science, 2327:7–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270007.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2327/23270007.
pdf.

Award:2002:LCS

[1585] (Distinguished Paper Award). Language and compiler support for hybridparallel programming on SMP clusters. Lecture Notes in Computer Science, 2327:11–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2327/23270011.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2327/23270011.
pdf.

Award:2002:PMS

[1586] (Best Student Paper Award). Parallelizing merge sort onto distributed memory parallel computers. Lecture Notes in Computer Science, 2327:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270025.
pdf.

Baydal:2002:ANC

[1587] E. Baydal, P. López, and J. Duato. Avoiding network congestion with local information. Lecture Notes in Computer Science, 2327:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270035.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2327/23270035.
pdf.

Flich:2002:IIR

[1588] J. Flich, P. López, J. C. Sancho, A. Robles, and J. Duato. Improving InfiniBand routing through multiple virtual networks. Lecture Notes in Computer Science, 2327:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270049.
pdf.

Rothman:2002:MAS

[1589] Jeffrey B. Rothman and Alan Jay Smith. Minerva: An adaptive subblock coherence protocol for improved SMP performance. Lecture Notes in Computer Science, 2327:64–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270064.
pdf.

Heinrich:2002:AMC

[1590] Mark Heinrich, Evan Speight, and Mainak Chaudhuri. Active memory clusters: Efficient multiprocessing on commodity clusters. Lecture Notes in Computer Science, 2327:78– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2327/23270078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270078.
pdf.

Garatti:2002:IAA

[1591] Marco Garatti, Roberto Costa, Stefano Crespi Reghizzi, and Erven Rohou. The impact of alias analysis on VLIW scheduling. Lecture Notes in Computer Science, 2327:93– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270093.
pdf.

Sato:2002:LCV

[1592] Toshinori Sato and Itsujiro Arita.

Low-cost value predictors using frequent value locality. Lecture Notes in Computer Science, 2327:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270106.
pdf.

Tang:2002:ICW

[1593] Weiyu Tang, Alexander Veidenbaum, Alexandru Nicolau, and Rajesh Gupta. Integrated I-cache way predictor and branch target buffer to reduce energy consumption. Lecture Notes in Computer Science, 2327:120–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270120.
                                        pdf.

Santana:2002:CAI

[1594] Oliverio J. Santana, Ayose Falcón, Enrique Fernández, Pedro Medina, Alex Ramírez, and Mateo Valero. A comprehensive analysis of indirect branch prediction. Lecture Notes in Computer Science, 2327:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270133.
                                        pdf.

Reinman:2002:HPE

[1595] Glenn Reinman, Brad Calder, and Todd Austin. High performance and energy efficient serial prefetch architecture. Lecture Notes in Computer Science, 2327:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2327/23270146.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2327/23270146.pdf.

Yang:2002:PMH

[1596] Chia-Lin Yang and Alvin Lebeck. A programmable memory hierarchy for prefetching linked data structures. Lecture Notes in Computer Science, 2327: 160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270160.
pdf.

Iwashita:2002:BRB

[1597] Takeshi Iwashita and Masaaki Shimasaki. Block red-black ordering method for parallel processing of ICCG solver. Lecture Notes in Computer Science, 2327:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270175.
pdf.

deStGermain:2002:IPA

[1598] J. Davison de St. Germain, Alan Morris, Steven G. Parker, Allen D. Malony, and Sameer Shende. Integrating performance analysis in the Uintah software development cycle. Lecture Notes in Computer Science, 2327: 190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270190.
pdf.

Noro:2002:PAM

[1599] Ayato Noro, Tomoya Ogawa, Takuma Ohta, Kazuyuki Yamashita, Shigeki Miyaji, and Mitue Den. Performance of adaptive mesh refinement scheme for hydrodynamics on simulations of expanding supernova envelope. Lecture Notes in Computer Science, 2327:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270207.
                                        pdf.

Uehara:2002:MBP

[1600] Hitoshi Uehara, Masanori Tamura, and Mitsuo Yokokawa. An MPI benchmark program library and its application to the Earth simulator. Lecture Notes in Computer Science, 2327:219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270219.
                                        pdf.

Furumura:2002:PSS

[1601] Takashi Furumura. Parallel simulation of seismic wave propagation. Lecture Notes in Computer Science, 2327:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270231.
                                        pdf.

Tsuboki:2002:LSP

[1602] Kazuhisa Tsuboki and Atsushi Sakakibara. Large-scale parallel computing of cloud resolving storm simulator. Lecture Notes in Computer Science, 2327: 243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270243.
pdf.

Oi:2002:RMS

[1603] Hitoshi Oi and Bing rung Tsai. Routing mechanism for static load balancing in a partitioned computer system with a fully connected network. Lecture Notes in Computer Science, 2327: 260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270260.
pdf.

Falcon:2002:SNW

[1604] Ayose Falcón, Oliverio J. Santana, Pedro Medina, Enrique Fernández, Alex Ramírez, and Mateo Valero. Studying new ways for improving adaptive history length branch predictors. Lecture Notes in Computer Science, 2327: 271–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270271.
pdf.

Henry:2002:SCC

[1605] Dana S. Henry, Gabriel H. Loh, and Rahul Sami. Speculative clustered caches for clustered processors. Lecture Notes in Computer Science, 2327: 281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270281.
                                        pdf.

Matsubara:2002:ETD

[1606] Yasuo Matsubara and Takahiro Shakushi. The effects of timing dependence and recursion on parallel program schemata. Lecture Notes in Computer Science, 2327:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2327/23270291.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2327/23270291.pdf.

Kondo:2002:CLI

[1607] Masaaki Kondo, Mitsugu Iwamoto, and Hiroshi Nakamura. Cache line impact on 3D PDE solvers. Lecture Notes in Computer Science, 2327:301– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270301.
                                        pdf.

Carter:2002:EPP

[1608] Lori Carter, Weihaw Chuang, and Brad Calder. An EPIC processor with pending functional units. Lecture Notes in Computer Science, 2327:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270310.
pdf.

Kumar:2002:SEO

[1609] Rakesh Kumar, Tusar Kanti Patra, and Anupam Basu. Software energy optimization of real time preemptive tasks by minimizing cacherelated preemption costs. Lecture Notes in Computer Science, 2327:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270321.
pdf.

Kim:2002:DGA

[1610] Jung-Sook Kim. Distributed genetic algorithm with multiple populations using multi-agent. Lecture Notes in Computer Science, 2327:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270329.
pdf.

Alpert:2002:NWP

[1611] Pinhas Alpert, Alexander Goikhman, Jacob Katzenelson, and Marina Tsidulko. Numerical weather prediction on the supercomputer toolkit. Lecture Notes in Computer Science, 2327: 335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270335.
                                        pdf.

Mello:2002:OPP

[1612] Rodrigo F. de Mello, Maria Stela V. Paiva, Luís Carlos Trevelin, and Adilson Gonzaga. OpenTella: a peerto-peer protocol for the load balancing in a system formed by a cluster from clusters. Lecture Notes in Computer Science, 2327:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270346.
                                        pdf.

Julien:2002:PEC

[1613] Nathalie Julien, Johann Laurent, Eric Senn, and Eric Martin. Power estimation of a C algorithm based on the functional-level power analysis of a digital signal processor. Lecture Notes in Computer Science, 2327:354– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2327/23270354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270354.
pdf.

Arenaz:2002:IAC

[1614] Manuel Arenaz, Juan Touriño, and Ramón Doallo. Irregular assignment computations on cc-NUMA multiprocessors. Lecture Notes in Computer Science, 2327:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270361.
pdf.

Talk:2002:LSP

[1615] (WOMPEI Invited Talk). Large system performance of SPEC OMP2001 benchmarks. Lecture Notes in Computer Science, 2327:370–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2327/23270370.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2327/23270370.pdf.

Muller:2002:SMB

[1616] Matthias S. Müller. A shared memory benchmark in OpenMP. Lecture Notes in Computer Science, 2327:380– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270380.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2327/23270380.
                                        pdf.

Takahashi:2002:PEH

[1617] Daisuke Takahashi, Mitsuhisa Sato, and Taisuke Boku. Performance evaluation of the hitachi SR8000 using OpenMP benchmarks. Lecture Notes in Computer Science, 2327:390– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270390.
                                        pdf.

Rabenseifner:2002:CBP

[1618] Rolf Rabenseifner. Communication bandwidth of parallel programming models on hybrid architectures. Lecture Notes in Computer Science, 2327: 401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270401.
                                        pdf.

Prabhakar:2002:PCB

[1619] Achal Prabhakar, Vladimir Getov, and Barbara Chapman. Performance comparisons of basic OpenMP constructs. Lecture Notes in Computer Science, 2327:413–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2327/23270413.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2327/23270413.
pdf.

Krawezik:2002:SOV

[1620] Géraud Krawezik, Guillaume Alléon, and Franck Cappello. SPMD OpenMP versus MPI on a IBM SMP for 3 kernels of the NAS benchmarks. Lecture Notes in Computer Science, 2327:425– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270425.
pdf.

Nakajima:2002:PIS

[1621] Kengo Nakajima and Hiroshi Okuda.

Parallel iterative solvers for unstructured grids using an OpenMP/MPI hybrid programming model for the GeoFEM platform on SMP cluster architectures. Lecture Notes in Computer Science, 2327:437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270437.
pdf.

Montleau:2002:PCM

[1622] Pierre de Montleau, Jose Maria Cela, Serge Moto Mpong, and André Godinass. A parallel computing model for the acceleration of a finite element software. Lecture Notes in Computer Science, 2327:449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270449.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270449.
                                        pdf.

Basumallik:2002:TOE

[1623] Ayon Basumallik, Seung-Jai Min, and Rudolf Eigenmann. Towards OpenMP execution on software distributed shared memory systems. Lecture Notes in Computer Science, 2327: 457–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270457.
                                        pdf.

Gonzalez:2002:DLP

[1624] Marc González, Eduard Ayguadé, Xavier Martorell, Jesús Labarta, and Phu V. Luong. Dual-level parallelism exploitation with OpenMP in coastal ocean circulation modeling. Lecture Notes in Computer Science, 2327:469– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2327/23270469.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2327/23270469.
                                        pdf.

Nakano:2002:SCG

[1625] Hirofumi Nakano, Kazuhisa Ishizaka, Motoki Obata, Keiji Kimura, and Hironori Kasahara. Static coarse grain task scheduling with cache optimization using OpenMP. Lecture Notes in Computer Science, 2327:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270479.
pdf.

Zima:2002:HPF

[1626] Hans P. Zima. High Performance Fortran — history, status and future (HiWEP invited talk). Lecture Notes in Computer Science, 2327:490– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270490.
pdf.

Sakagami:2002:PEJ

[1627] Hitoshi Sakagami and Shingo Furubayashi. Performance evaluation for Japanese HPF compilers with special benchmark suite. Lecture Notes in Computer Science, 2327:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270491.
pdf.

Asaoka:2002:EHJ

[1628] Kae Asaoka, Akio Hirano, Yasuo Okabe, and Masanori Kanazawa. Evaluation of the HPF/JA extensions on Fujitsu VPP using the NAS parallel benchmarks. Lecture Notes in Computer Science, 2327:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270503.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270503.
pdf.

Cai:2002:TDE

[1629] DongSheng Cai, Yaoting Li, Ken ichi Nishikawa, Chiejie Xiao, and Xiaoyan Yan. Three-dimensional electromagnetic particle-in-cell code using High Performance Fortran on PC cluster. Lecture Notes in Computer Science, 2327:515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270515.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270515.
pdf.

Iwashita:2002:TLH

[1630] Hidetoshi Iwashita, Kohichiro Hotta, Sachio Kamiya, and Matthijs van Waveren. Towards a lightweight HPF compiler. Lecture Notes in Computer Science, 2327:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270526.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2327/23270526.
pdf.

Brezany:2002:PSH

[1631] Peter Brezany, Jonghyun Lee, and Marianne Winslett. Parallel I/O support for HPF on computational grids. Lecture Notes in Computer Science, 2327:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2327/23270539.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2327/23270539.
pdf.

Araki:2002:OHP

[1632] Takuya Araki, Hitoshi Murai, Tsunehiko Kamachi, and Yoshiki Seo. Optimization of HPF programs with dynamic recompilation technique. Lecture Notes in Computer Science, 2327: 551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2327/23270551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2327/23270551.
pdf.

Anonymous:2002:AIaw

[1633] Anonymous. Author index. Lecture Notes in Computer Science, 2327:563– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2327/2327auth.pdf.

Borkowski:2002:ICS

[1634] Janusz Borkowski. Interrupt and cancellation as synchronization methods. Lecture Notes in Computer Science, 2328:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2328/23280003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2328/23280003.
pdf.

Decyk:2002:SMP

[1635] Viktor K. Decyk and Dean E. Dauger. Supercomputing for the masses: a parallel Macintosh cluster. Lecture Notes in Computer Science, 2328:10– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280010.
pdf.

Deelman:2002:AVD

[1636] Ewa Deelman, Carl Kesselman, Roy Williams, Kent Blackburn, Albert Lazzarini, and Scott Koranda. Applications of virtual data in the LIGO experiment. Lecture Notes in Computer Science, 2328:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280023.
pdf.

Sampels:2002:VAV

[1637] Michael Sampels. Visualization of automorphisms and vertex-symmetry. Lecture Notes in Computer Science, 2328:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2328/23280035.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2328/23280035.
pdf.

Schmollinger:2002:MCS

[1638] Martin Schmollinger and Michael Kaufmann. κNUMA: a model for clusters of SMP-machines. Lecture Notes in Computer Science, 2328:42– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280042.
pdf.

Masko:2002:PSA

[1639] Marek Tudruj and Łukasz Masko.

A parallel system architecture based on dynamically configurable shared memory clusters. Lecture Notes in Computer Science, 2328:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280051.
pdf.

Fernandez:2002:SSA

[1640] C. Fernández, F. Torres, and S. T. Puente. SASEPA: Simultaneous allocation and scheduling with exclusion and precedence relations algorithm. Lecture Notes in Computer Science, 2328: 65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280065.
pdf.

Fujimoto:2002:OTS

[1641] Noriyuki Fujimoto and Kenichi Hagihara. Optimal task scheduling of a complete K-ary tree with communication delays. Lecture Notes in Computer Science, 2328:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280071.
pdf.

Gawiejnowicz:2002:GAT

[1642] Stanisław Gawiejnowicz, Wiesław Kurc, and Lidia Pankowska. A greedy approach for a time-dependent scheduling problem. Lecture Notes in Computer Science, 2328:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280079.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2328/23280079.
pdf.

Giaro:2002:DSB

[1643] Krzysztof Giaro, Marek Kubale, Michał Małafiejski, and Konrad Piwakowski. Dedicated scheduling of biprocessor tasks to minimize mean flow time. Lecture Notes in Computer Science, 2328:87–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280087.
pdf.

Laskowski:2002:FSP

[1644] Eryk Laskowski. Fast scheduling and partitioning algorithm in the multi-processor system with redundant communication resources. Lecture Notes in Computer Science, 2328:97– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280097.
pdf.

Rotaru:2002:HDLb

[1645] Tiberiu Rotaru and Hans-Heinrich Nägeli. Heterogeneous dynamic load balancing with a scheme based on the Laplacian polynomial. Lecture Notes in Computer Science, 2328:107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280107.
                                        pdf.

Masko:2002:TSD

[1646] Marek Tudruj and Łukasz Masko. Task scheduling for dynamically configurable multiple SMP clusters based on extended DSC approach. Lecture Notes in Computer Science, 2328:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280115.
                                        pdf.

Wolniewicz:2002:PTM

[1647] Paweł Wolniewicz and Maciej Drozdowski. Processing time and memory requirements for multi-installment divisible job processing. Lecture Notes in Computer Science, 2328:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280125.
                                        pdf.

Drozdowski:2002:EET

[1648] Maciej Drozdowski. Estimating execution time of distributed applications. Lecture Notes in Computer Science, 2328:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2328/23280137.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2328/23280137.
pdf.

Kwiatkowski:2002:EPP

[1649] Jan Kwiatkowski. Evaluation of parallel programs by measurement of its granularity. Lecture Notes in Computer Science, 2328:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280145.
pdf.

Metkowski:2002:PDC

[1650] Rafał Metkowski, Piotr Bała, and Terry Clark. The performance of different communication mechanisms and algorithms used for parallelization of molecular dynamics code. Lecture Notes in Computer Science, 2328:154– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280154.
pdf.

Nikolow:2002:BTS

[1651] Darin Nikolow, Renata Słota, and Jacek Kitowski. Benchmarking tertiary storage systems with file fragmentation. Lecture Notes in Computer Science, 2328:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2328/23280162.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2328/23280162.pdf.

Olas:2002:FCC

[1652] Tomasz Olas, Konrad Karczewski, Adam Tomas, and Roman Wyrzykowski. FEM computations on clusters using different models of parallel programming. Lecture Notes in Computer Science, 2328:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2328/23280170.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2328/23280170.pdf.

Blesa:2002:PST

[1653] Maria J. Blesa, Lluis Hernàndez, and Fatos Xhafa. Parallel skeletons for tabu search method based on search strategies and neighborhood partition. Lecture Notes in Computer Science, 2328: 185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280185.
                                        pdf.

Blazewicz:2002:NPA

[1654] Jacek Błazewicz and Rafał Walkowiak. A new parallel approach for multidimensional packing problem. Lecture Notes in Computer Science, 2328:194– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280194.
pdf.

Brzezinski:2002:CRP

[1655] Jerzy Brzeziński and Dariusz Wawrzyniak. Consistency requirements of Peterson’s algorithm for mutual exclusion of n processes in a distributed shared memory system. Lecture Notes in Computer Science, 2328:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280202.
pdf.

Czech:2002:TPA

[1656] Zbigniew J. Czech. Three parallel algorithms for simulated annealing. Lecture Notes in Computer Science, 2328: 210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280210.
pdf.

Guinand:2002:CPT

[1657] Frédéric Guinand, Gilles Parmentier, and Denis Trystram. Construction of phylogenetic trees on parallel clusters. Lecture Notes in Computer Science, 2328:218–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280218.
                                        pdf.

Kokosinski:2002:PGA

[1658] Zbigniew Kokosiński. On parallel generation of t-ary trees in an associative model. Lecture Notes in Computer Science, 2328:228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280228.
                                        pdf.

Wodecki:2002:SFS

[1659] Mieczysław Wodecki and Wojciech Bożejko. Solving the flow shop problem by parallel simulated annealing. Lecture Notes in Computer Science, 2328: 236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280236.
                                        pdf.

Dembinski:2002:AVI

[1660] Piotr Dembiński, Wojciech Penczek, and Agata Półrola. Automated verification of infinite state concurrent systems. Lecture Notes in Computer Science, 2328:247–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280247.
pdf.

Gonzalez:2002:LCA

[1661] J. A. González, C. León, M. Pristinta, J. L. Roda, C. Rodríguez, J. M. Rodríguez, and F. Sande. A language for the complexity analysis of parallel programs. Lecture Notes in Computer Science, 2328:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280256.
pdf.

Krivoi:2002:CSH

[1662] Sergey Krivoi. Criteria of satisfiability for homogeneous systems of linear Diophantine constraints. Lecture Notes in Computer Science, 2328:264– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280264.
pdf.

Maslennikov:2002:SGE

[1663] Oleg Maslennikov. Systematic generation of executing programs for processor elements in parallel ASIC or FPGA-based systems and their transformation into VHDL-descriptions of processor element control units. Lecture Notes in Computer Science, 2328: 272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280272.
                                        pdf.

Sener:2002:DDP

[1664] Cevat Sener, Yakup Paker, and Ayşe Kiper. Developing a data-parallel application with DaParT. Lecture Notes in Computer Science, 2328:280– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280280.
                                        pdf.

Smyk:2002:AMM

[1665] Adam Smyk and Marek Tudruj. Application of mixed MPI OpenMP programming in a multi SMP cluster computer. Lecture Notes in Computer Science, 2328:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2328/23280288.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2328/23280288.pdf.

Brezany:2002:ICP

[1666] Peter Brezany, Marian Bubak, Maciej Malawski, and Katarzyna Zaja̧c. Irregular and out-of-core parallel computing on clusters. Lecture Notes in Computer Science, 2328:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280299.
pdf.

Bubak:2002:CGA

[1667] Marian Bubak, Włodzimierz Funika, Bartosz Balis, and Roland Wismüller. A concept of grid application monitoring. Lecture Notes in Computer Science, 2328:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280307.
pdf.

Bubak:2002:TMI

[1668] Marian Bubak, Włodzimierz Funika, Piotr Mȩtel, Rafał Orłowski, and Roland Wismüller. Towards a monitoring interface specification for distributed Java applications. Lecture Notes in Computer Science, 2328:315– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280315.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280315.
pdf.

Garstecki:2002:TCP

[1669] Łukasz Garstecki, Paweł Kaczmarek, Jacques Chassin de Kergommeaux, Henryk Krawczyk, and Bogdan Wiszniewski. Testing for conformance of parallel programming pattern languages. Lecture Notes in Computer Science, 2328:323–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280323.
                                        pdf.

Gepner:2002:OIE

[1670] Pawel Gepner. Overview of IA-64 explicitly parallel instruction computing architecture. Lecture Notes in Computer Science, 2328:331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280331.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280331.
                                        pdf.

Goscinski:2002:TOS

[1671] A. Gosciński, M. Hobbs, and J. Silcock. Toward an operating system that supports parallel processing on nondedicated clusters. Lecture Notes in Computer Science, 2328:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280340.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2328/23280340.
pdf.

Kim:2002:LDJ

[1672] Joong-Han Kim, Seong-Soo Yae, R. S.

Ramakrishna, and Yoo-Sung Kim. Load distribution in Jini using JINT. Lecture Notes in Computer Science, 2328:354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2328/23280354.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2328/23280354.
pdf.

Lepiarz:2002:ASL

[1673] Marcin Lepiarz and Zdzislaw Onderka.

Agent system for load monitoring of the heterogeneous computer network. Lecture Notes in Computer Science, 2328: 364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280364.
pdf.

Nguyen:2002:DTR

[1674] G. T. Nguyen, L. Hluchy, V. D. Tran, and M. Kotocova. DDG task recovery for cluster computing. Lecture Notes in Computer Science, 2328:369– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280369.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2328/23280369.
pdf.

Alonso:2002:CBS

[1675] Pedro Alonso, Raquel Cortina, Irene Díaz, Vicente Hernández, and José Ranilla. A columnwise block striping in Neville elimination. Lecture Notes in Computer Science, 2328:379– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280379.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280379.
pdf.

Bjorstad:2002:FLN

[1676] Petter E. Bjørstad and Piotr Krzyżanowski. A flexible 2-level Neumann–Neumann method for structural analysis problems. Lecture Notes in Computer Science, 2328:387–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2328/23280387.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2328/23280387.pdf.

Blaheta:2002:PDD

[1677] Radim Blaheta, Ondřej Jakl, and Jiří Starý. Parallel displacement decomposition solvers for elasticity problems. Lecture Notes in Computer Science, 2328:395–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2328/23280395.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2328/23280395.
pdf.

Ciegis:2002:SPR

[1678] R. Čiegis and G. Šilko. A scheme for partitioning regular graphs. Lecture Notes in Computer Science, 2328:404– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280404.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280404.
pdf.

Cooper:2002:ALE

[1679] A. Cooper, M. Szularz, and J. Weston.

Analysis of the Lanczos error bounds and its application to the explicitly restarted Lanczos algorithm. Lecture Notes in Computer Science, 2328:410– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280410.
pdf.

Gustavson:2002:NGD

[1680] Fred G. Gustavson. New generalized data structures for matrices lead to a variety of high performance algorithms. Lecture Notes in Computer Science, 2328:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280418.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2328/23280418.
                                        pdf.

Petcu:2002:SLS

[1681] Dana Petcu. Solving large systems of differential equations with PaViS. Lecture Notes in Computer Science, 2328: 437–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280437.
                                        pdf.

Saad:2002:PPS

[1682] Y. Saad and M. Sosonkina. pARMS: a package for solving general sparse linear systems on parallel computers. Lecture Notes in Computer Science, 2328: 446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280446.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280446.
                                        pdf.

Sergyienko:2002:IGQ

[1683] Anatoli Sergyienko and Oleg Maslennikov. Implementation of Givens QRdecomposition in FPGA. Lecture Notes in Computer Science, 2328:458– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280458.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280458.
                                        pdf.

Stpiczynski:2002:NMP

[1684] Przemysław Stpiczyński. A new message passing algorithm for solving linear recurrence systems. Lecture Notes in Computer Science, 2328:466– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280466.
pdf.

Burczynski:2002:DEA

[1685] Tadeusz Burczyński and Wacław Kus. Distributed evolutionary algorithms in shape optimization of nonlinear structures. Lecture Notes in Computer Science, 2328:477–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280477.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280477.
pdf.

Hluchy:2002:PNS

[1686] L. Hluchy, D. Froehlich, V. D. Tran, J. Astalos, M. Dobrucky, and G. T. Nguyen. Parallel numerical solution for flood modeling systems. Lecture Notes in Computer Science, 2328:485– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280485.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280485.
pdf.

Jurczyk:2002:ECD

[1687] Tomasz Jurczyk, Barbara Głut, and Jacek Kitowski. An empirical comparison of decomposition algorithms for complex finite element meshes. Lecture Notes in Computer Science, 2328: 493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280493.
                                        pdf.

Kamieniarz:2002:APC

[1688] Grzegorz Kamieniarz, Ryszard Matysiak, Alvaro Caramico D’Auria, Filippo Esposito, and Cristiano Benelli. Application of parallel computing in the transfer — matrix simulations of the supramolecules mn_{6} and ni_{12}. Lecture Notes in Computer Science, 2328:502– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280502.
                                        pdf.

Krawczyk:2002:PEE

[1689] Henryk Krawczyk, Aleksander Neyman, Michal Nowikowski, and Jamil Saif. The parallel environment for endoscopic image analysis. Lecture Notes in Computer Science, 2328:510– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280510.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2328/23280510.
pdf.

Kwiatkowski:2002:UFC

[1690] Jan Kwiatkowski, Wieslawa Kwiatkowska, Krzysztof Kawa, and Piotr Kania. Using fractal coding in medical image magnification. Lecture Notes in Computer Science, 2328:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280517.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280517.
pdf.

Levin:2002:QCS

[1691] Mikhail P. Levin. Quasi-characteristics scheme with parallel facilities for computations of two-phase flows in heterogeneous porous media. Lecture Notes in Computer Science, 2328:526– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280526.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280526.
pdf.

Musial:2002:MCM

[1692] G. Musiał and L. Dȩbski. Monte Carlo method with parallel computation of phase transitions in the threedimensional Ashkin–Teller model. Lecture Notes in Computer Science, 2328: 535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280535.
                                        pdf.

Nilsson:2002:FSO

[1693] Stefan Nilsson. Flow simulations on overlapping grids. Lecture Notes in Computer Science, 2328:544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280544.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280544.
                                        pdf.

Norton:2002:PUA

[1694] Charles D. Norton and Thomas A. Ćwik. Parallel unstructured AMR and gigabit networking for Beowulf-class clusters. Lecture Notes in Computer Science, 2328:552–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2328/23280552.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2328/23280552.pdf.

Sawyer:2002:PGM

[1695] William Sawyer and Peter Messmer. Parallel grid manipulations for general circulation models. Lecture Notes in Computer Science, 2328:564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280564.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2328/23280564.
pdf.

Soloviev:2002:BML

[1696] Alexander Soloviev, Vyacheslav Maksimov, Valerii Rozenberg, and Yurii Ermoliev. Block models of lithosphere dynamics: Approach and algorithms. Lecture Notes in Computer Science, 2328: 572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280572.
pdf.

Szymanski:2002:CMD

[1697] Bolesla K. Szymański and Gilbert Chen. A component model for discrete event simulation. Lecture Notes in Computer Science, 2328:580–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280580.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280580.
pdf.

Kolodziej:2002:MHG

[1698] Joanna Kołodziej. Modelling hierarchical genetic strategy as a family of Markov chains. Lecture Notes in Computer Science, 2328:595–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280595.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280595.
                                        pdf.

Rutkowska:2002:PPI

[1699] Danuta Rutkowska, Robert Nowicki, and Yoichi Hayashi. Parallel processing by implication-based neuro-fuzzy systems. Lecture Notes in Computer Science, 2328:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280599.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280599.
                                        pdf.

Schaefer:2002:CSM

[1700] Robert Schaefer and Zenon J. Jabłoński. On the convergence of sampling measures in the global genetic search. Lecture Notes in Computer Science, 2328: 608–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280608.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280608.
                                        pdf.

Seijas:2002:GAT

[1701] Juan Seijas, Carmen Morató, and José L. Sanz-González. Genetic algorithms: Two different elitism operators for stochastic and deterministic applications. Lecture Notes in Computer Science, 2328:617–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2328/23280617.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280617.
pdf.

Seredynski:2002:ILS

[1702] Franciszek Seredyński and Anna Swiȩcicka. Immune-like system approach to cellular automata-based scheduling. Lecture Notes in Computer Science, 2328:626–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2328/23280626.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2328/23280626.pdf.

Starczewski:2002:CST

[1703] Janusz Starczewski and Leszek Rutkowski.

Connectionist structures of type 2 fuzzy inference systems. Lecture Notes in Computer Science, 2328:634–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280634.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280634.
pdf.

Wojnarski:2002:LCN

[1704] Marcin Wojnarski. LTF-C — neural network for solving classification problems. Lecture Notes in Computer Science, 2328:643–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2328/23280643.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280643.
                                        pdf.

Zola:2002:EJH

[1705] Jarosław Zola and Roman Wyrzykowski. EPL-Julia the high-performance library for evolutionary computations. Lecture Notes in Computer Science, 2328:652–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2328/23280652.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280652.
                                        pdf.

Fialko:2002:AMI

[1706] Sergiy Fialko. Aggregation multilevel iterative solver for analysis of large-scale finite element problems of structural mechanics: Linear statics and natural vibrations. Lecture Notes in Computer Science, 2328:663– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280663.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280663.
                                        pdf.

Grzybowski:2002:CSC

[1707] Andrzej Grzybowski. Computer simulations in constructing a coefficient of uncertainty in regression estimation — methodology and results. Lecture Notes in Computer Science, 2328:671– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280671.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280671.
pdf.

Grzymkowski:2002:MPI

[1708] Radosław Grzymkowski and Damian Słota. Multi-phase inverse Stefan problems solved by approximation method. Lecture Notes in Computer Science, 2328:679–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280679.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280679.
pdf.

Karafiat:2002:EEF

[1709] Andrzej Karafiat, Lech Sławik, and Olga Trzos. Error estimates for BE/ FE method in elastic scattering. Lecture Notes in Computer Science, 2328: 687–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280687.
pdf.

Lesczynski:2002:NMS

[1710] Jacek Lesczyński and Mariusz Ciesielski. A numerical method for solution of ordinary differential equations of fractional order. Lecture Notes in Computer Science, 2328:695– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280695.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280695.
                                        pdf.

Leszczynski:2002:EGU

[1711] Jacek Leszczyński and Sebastian Pluta. The efficient generation of unstructured control volumes in 2D and 3D. Lecture Notes in Computer Science, 2328:703–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280703.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280703.
                                        pdf.

Nagorka:2002:CTM

[1712] Arkadiusz Nagórka, Norbert Sczygiol, and Grzegorz Szwarc. Coupling of thermal and mechanical phenomena by boundary conditions in numerical modelling of solidifying castings. Lecture Notes in Computer Science, 2328:711– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280711.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280711.
                                        pdf.

Paprzycki:2002:SNA

[1713] Marcin Paprzycki, Deborah Dent, and Anna Kucaba-Piȩtal. Solvers for nonlinear algebraic equations; where are we today? Lecture Notes in Computer Science, 2328:719–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280719.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280719.
pdf.

Patan:2002:OLS

[1714] Maciej Patan and Dariusz Uciński.

Optimal location of sensors for parameter estimation of static distributed systems. Lecture Notes in Computer Science, 2328:729–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280729.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280729.
pdf.

Pielorz:2002:AER

[1715] Amalia Pielorz. Application of equations with a retarded argument in physical systems. Lecture Notes in Computer Science, 2328:738–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280738.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280738.
pdf.

Poullikkas:2002:MFS

[1716] Andreas Poullikkas, Andreas Karageorghis, and Georgios Georgiou. The method of fundamental solutions in three-dimensional elastostatics. Lecture Notes in Computer Science, 2328: 747–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280747.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280747.
pdf.

Sevastjanov:2002:CNM

[1717] Pavel V. Sevastjanov, Paweł Róg, and Andrey V. Venberg. A constructive numerical method for the comparison of intervals. Lecture Notes in Computer Science, 2328:756–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280756.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280756.
pdf.

Smyrlis:2002:RSN

[1718] Yiorgos-Sokratis Smyrlis and Andreas Karageorghis. Rotation of the sources and normalization of the fundamental solutions in the MFS. Lecture Notes in Computer Science, 2328:762– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280762.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280762.
pdf.

Vasilyeva:2002:RUP

[1719] Ekaterina Vasilyeva and Valerii Rozenberg. Reconstruction of unknown properties of seismic flows. Lecture Notes in Computer Science, 2328:770– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280770.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280770.
pdf.

Weiner:2002:PTS

[1720] R. Weiner, B. A. Schmitt, and H. Podhaisky. Parallel two-step W-methods on singular perturbation problems. Lecture Notes in Computer Science, 2328:778–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280778.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280778.
pdf.

Habbas:2002:LPC

[1721] Zineb Habbas, Michaël Krajecki, and Daniel Singer. The Langford’s problem: a challenge for parallel resolution of CSP. Lecture Notes in Computer Science, 2328:789–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280789.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280789.
pdf.

Kleymenov:2002:MCS

[1722] A. Kleymenov, D. Petunin, A. Semenov, and I. Vazhev. A model of cooperative solvers for computational problems. Lecture Notes in Computer Science, 2328:797–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280797.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280797.
                                        pdf.

Revol:2002:MPC

[1723] N. Revol, Y. Denneulin, J.-F. Méhaut, and B. Planquelle. A methodology of parallelization for continuous verified global optimization. Lecture Notes in Computer Science, 2328:803– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280803.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280803.
                                        pdf.

Romero:2002:MCC

[1724] Nicolas Romero. Mobile concurrent constraint programming. Lecture Notes in Computer Science, 2328:811–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280811.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280811.
                                        pdf.

Sinz:2002:CPD

[1725] Carsten Sinz, Jörg Denzinger, Jürgen Avenhaus, and Wolfgang Küchlin. Combining parallel and distributed search in automated equational deduction. Lecture Notes in Computer Science, 2328:819–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2328/23280819.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2328/23280819.pdf.

Bujanda:2002:NME

[1726] Blanca Bujanda and Juan Carlos Jorge. Numerical methods for evolutionary convection-diffusion problems with nonlinear reaction terms. Lecture Notes in Computer Science, 2328:833– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280833.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280833.
pdf.

Jankowski:2002:SIN

[1727] Jan Jankowski and Monika Warmowska. Solution of incompressible Navier–Stokes equations using projection methods. Lecture Notes in Computer Science, 2328:841–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280841.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2328/23280841.
                                        pdf.

Auzinger:2002:TST

[1728] Winfried Auzinger, Othmar Koch, and Ewa Weinmüller. Theory and solution techniques for singular boundary value problems in ordinary differential equations. Lecture Notes in Computer Science, 2328:851–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280851.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280851.
                                        pdf.

Bielecki:2002:END

[1729] Andrzej Bielecki and Dariusz Jabłoński. Estimation of numerical dynamics constants of a weakly nonlinear neuron. Lecture Notes in Computer Science, 2328:862–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2328/23280862.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280862.
                                        pdf.

Domoshnitsky:2002:PSD

[1730] Alexander Domoshnitsky, Michael Drakhlin, and Elena Litsyn. On positivity of solutions of delayed differential equation with state dependent impulses. Lecture Notes in Computer Science, 2328:870–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2328/23280870.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280870.
pdf.

Krzhizhanovskaya:2002:DSS

[1731] V. V. Krzhizhanovskaya, M. A. Zatevakhin, A. A. Ignatiev, Y. E. Gorbachev, and P. M. A. Sloot. Distributed simulation of siliconbased film growth. Lecture Notes in Computer Science, 2328:879–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280879.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280879.
pdf.

Magdon-Maksymowicz:2002:BTS

[1732] Maria Stanisława Magdoń-Maksymowicz, Marian Bubak, and Andrzej Zbigniew Maksymowicz. Biological time scale and ageing in the penna model. Lecture Notes in Computer Science, 2328: 888–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2328/23280888.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2328/23280888.
pdf.

Maniatty:2002:SMP

[1733] William A. Maniatty, Thomas Caraco, Niles Lehman, and Boleslaw K. Szymański. Spatial models of persistence in RNA worlds: Exploring the origins of life. Lecture Notes in Computer Science, 2328:896–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2328/23280896.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2328/23280896.
                                        pdf.

Topa:2002:ATN

[1734] Paweł Topa and Mariusz Paszkowski. Anastomosing transportation networks. Lecture Notes in Computer Science, 2328:904–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2328/23280904.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2328/23280904.pdf.

Anonymous:2002:AIax

[1735] Anonymous. Author index. Lecture Notes in Computer Science, 2328:913– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2328/2328auth.pdf.

Hey:2002:USC

[1736] T. Hey and A. E. Trefethen. The UK escience core program and the Grid. Lecture Notes in Computer Science, 2329: 3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290003.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2329/23290003.
pdf.

Fox:2002:CG

[1737] G. Fox, O. Balsoy, S. Pallickara, A. Uyar, D. Gannon, and A. Slominski. Community grids. Lecture Notes in Computer Science, 2329:22– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290022.
pdf.

Marir:2002:CMS

[1738] F. Marir, K. Zerzour, K. Ouazzane, and Y. Xue. A conceptual model for surveillance video content and eventbased indexing and retrieval. Lecture Notes in Computer Science, 2329:41– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290041.
pdf.

Monostori:2002:COD

[1739] K. Monostori, R. Finkel, A. Zaslavsky, G. Hodász, and M. Pataki. Comparison of overlap detection techniques. Lecture Notes in Computer Science, 2329: 51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290051.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2329/23290051.
                                        pdf.

Llopis:2002:UPR

[1740] F. Llopis, J. L. Vicedo, and A. Ferrández. Using a passage retrieval system to support question answering process. Lecture Notes in Computer Science, 2329: 61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290061.
                                        pdf.

Emmen:2002:XDP

[1741] A. Emmen. XML design patterns used in the EnterTheGrid portal. Lecture Notes in Computer Science, 2329:70– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290070.
                                        pdf.

Fernandez-Iglesias:2002:MME

[1742] M. J. Fernández-Iglesias, J. S. Rodríguez, L. Anido, J. Santos, M. Caeiro, and M. Llamas. Modeling metadata-enabled information retrieval. Lecture Notes in Computer Science, 2329:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2329/23290078.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2329/23290078.pdf.

Merks:2002:SBP

[1743] R. Merks, A. Hoekstra, J. Kaandorp, and P. Sloot. Spontaneous branching in a polyp oriented model of stony coral growth. Lecture Notes in Computer Science, 2329:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290088.
pdf.

Topa:2002:LMP

[1744] P. Topa and J. Tyszka. Local minimization paradigm in numerical modeling of foraminiferal shells. Lecture Notes in Computer Science, 2329:97– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290097.
pdf.

Suppi:2002:UPS

[1745] R. Suppi, P. Munt, and E. Luque.

Using PDES to simulate individualoriented models in ecology: a case study. Lecture Notes in Computer Science, 2329:107–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2329/23290107.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2329/23290107.pdf.

Kamerman:2002:SMH

[1746] D. J. Kamerman and M. H. F. Wilkinson. In silico modeling of the human intestinal microflora. Lecture Notes in Computer Science, 2329:117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290117.
                                        pdf.

Liu:2002:MAM

[1747] Y. Liu and H. J. Ruskin. A mesoscopic approach to modeling immunological memory. Lecture Notes in Computer Science, 2329:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290127.
                                        pdf.

Gonzalez:2002:NMO

[1748] L. González. A new method for ordering binary states probabilities in reliability and risk analysis. Lecture Notes in Computer Science, 2329:137– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290137.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2329/23290137.
pdf.

Sanseverino:2002:REU

[1749] C. M. Rocco Sanseverino and J. A.

Moreno. Reliability evaluation using Monte Carlo simulation and support vector machine. Lecture Notes in Computer Science, 2329:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290147.
pdf.

Meriste:2002:MTS

[1750] M. Meriste and L. Motus. On models for time-sensitive interactive computing. Lecture Notes in Computer Science, 2329:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290156.
pdf.

Ferri-Ramirez:2002:IDM

[1751] C. Ferri-Ramírez, J. Hernández-Orallo, and M. J. Ramírez-Quintana. Induction of decision multi-trees using Levin search. Lecture Notes in Computer Science, 2329:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290166.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290166.
                                        pdf.

Spinnato:2002:VSM

[1752] P. F. Spinnato, G. D. van Albada, and P. M. A. Sloot. A versatile simulation model for hierarchical treecodes. Lecture Notes in Computer Science, 2329: 176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290176.
                                        pdf.

Semoushin:2002:CPI

[1753] I. V. Semoushin and O. Yu. Gorokhov. Computational processes in iterative stochastic control design. Lecture Notes in Computer Science, 2329:186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290186.
                                        pdf.

Ratto:2002:EAD

[1754] M. Ratto and A. Saltelli. An efficient approach to deal with the curse of dimensionality in sensitivity analysis computations. Lecture Notes in Computer Science, 2329:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290196.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2329/23290196.
pdf.

Pflug:2002:BQM

[1755] G. Ch. Pflug and L. Halada. Birge and Qi method for three-stage stochastic programs using IPM. Lecture Notes in Computer Science, 2329:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290206.
pdf.

Boukhanovsky:2002:MSM

[1756] A. V. Boukhanovsky. Multivariate stochastic models of metocean fields: Computational aspects and applications. Lecture Notes in Computer Science, 2329:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290216.
pdf.

Frolova:2002:SGA

[1757] J. Frolova and V. Korobitsin. Simulation of gender artificial society: Multi-agent models of subjectobject interactions. Lecture Notes in Computer Science, 2329:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2329/23290226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290226.
                                        pdf.

Wedemann:2002:MFP

[1758] R. S. Wedemann, R. Donangelo, L. A. V. de Carvalho, and I. H. Martins. Memory functioning in psychopathology. Lecture Notes in Computer Science, 2329:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290236.
                                        pdf.

Debenham:2002:IME

[1759] J. Debenham. Investigating emarket evolution. Lecture Notes in Computer Science, 2329:246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290246.
                                        pdf.

Nakai:2002:MGS

[1760] J. Nakai. Markets as global scheduling mechanisms: The current state. Lecture Notes in Computer Science, 2329: 256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290256.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2329/23290256.
pdf.

Kenjeres:2002:NSC

[1761] S. Kenjereš, K. Hanjalić, and G. Krstović.

Numerical simulations of combined effects of terrain orography and thermal stratification on pollutant distribution in a town valley. Lecture Notes in Computer Science, 2329:266– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290266.
pdf.

Kim:2002:DCP

[1762] C. Kim, B. Choi, K. Kim, and S. Han. The differentiated call processing based on the simple priorityscheduling algorithm in SIP6. Lecture Notes in Computer Science, 2329:276– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290276.
pdf.

Fatta:2002:FAN

[1763] G. Di Fatta, G. Lo Re, and A. Urso.

A fuzzy approach for the network congestion problem. Lecture Notes in Computer Science, 2329:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2329/23290286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290286.
                                        pdf.

Lobosco:2002:PEF

[1764] M. Lobosco, V. Santos Costa, and C. L. de Amorim. Performance evaluation of Fast Ethernet, Giganet, and Myrinet on a cluster. Lecture Notes in Computer Science, 2329:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290296.
                                        pdf.

Datta:2002:BOP

[1765] A. Datta and S. Soundaralakshmi. Basic operations on a partitioned optical passive stars network with large group size. Lecture Notes in Computer Science, 2329:306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290306.
                                        pdf.

Zudilova:2002:MGR

[1766] E. V. Zudilova and M. O. Borisov. 3D mesh generation for the results of anisotropic etch simulation. Lecture Notes in Computer Science, 2329:316– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2329/23290316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290316.
pdf.

Daoud:2002:FSA

[1767] D. S. Daoud and D. S. Subasi. A fractional splitting algorithm for non-overlapping domain decomposition. Lecture Notes in Computer Science, 2329:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290324.
pdf.

Montenegro:2002:TMG

[1768] R. Montenegro, G. Montero, J. M.

Escobar, E. Rodríguez, and J. M. González-Yuste. Tetrahedral mesh generation for environmental problems over complex terrains. Lecture Notes in Computer Science, 2329:335– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290335.
pdf.

Tai:2002:DDM

[1769] X.-C. Tai. Domain decomposition and multigrid methods for obstacle problems. Lecture Notes in Computer Science, 2329:345–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290345.
                                        pdf.

Jurczyk:2002:DDC

[1770] T. Jurczyk and B. Głut. Domain decomposition coupled with Delaunay mesh generation. Lecture Notes in Computer Science, 2329:353–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290353.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290353.
                                        pdf.

Artoli:2002:APF

[1771] A. M. Artoli, A. G. Hoekstra, and P. M. A. Sloot. Accuracy of 2D pulsatile flow in the lattice Boltzmann BGK method. Lecture Notes in Computer Science, 2329:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290361.
                                        pdf.

Raney:2002:TMT

[1772] B. Raney, A. Voellmy, N. Cetin, M. Vrtic, and K. Nagel. Towards a microscopic traffic simulation of all of switzerland. Lecture Notes in Computer Science, 2329:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290371.
pdf.

Ruskin:2002:MTF

[1773] H. J. Ruskin and R. Wang. Modeling traffic flow at an urban unsignalized intersection. Lecture Notes in Computer Science, 2329:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290381.
pdf.

Gonzalez-Santos:2002:DMO

[1774] G. González-Santos and C. Vargas-Jarillo. A discrete model of oil recovery. Lecture Notes in Computer Science, 2329:391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2329/23290391.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2329/23290391.
pdf.

Feng:2002:VPD

[1775] Y. Feng, H. J. Ruskin, and B. Zhu.

Virtual phase dynamics for constrained geometries in a soap froth. Lecture Notes in Computer Science, 2329:399– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290399.
                                        pdf.

Beletskyy:2002:CMP

[1776] V. Beletskyy. A correction method for parallel loop execution. Lecture Notes in Computer Science, 2329:409– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290409.
                                        pdf.

Mark:2002:CSA

[1777] P. van der Mark, R. van Engelen, K. Gallivan, and W. Dewar. A case study for automatic code generation on a coupled oceanatmosphere model. Lecture Notes in Computer Science, 2329:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290419.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290419.
                                        pdf.

Maubach:2002:DFO

[1778] J. M. Maubach and W. Drenth. Dataflow oriented visual programming libraries for scientific computing. Lecture Notes in Computer Science, 2329: 429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290429.
pdf.

Ferdinandova:2002:ODD

[1779] I. Ferdinandova. One dilemma — different points of view. Lecture Notes in Computer Science, 2329:439– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290439.
pdf.

Meng:2002:BA

[1780] I.-H. Meng, W.-P. Yang, W.-C. Chen, and L.-P. Chang. Business agent. Lecture Notes in Computer Science, 2329: 449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290449.
pdf.

Benavent:2002:ULD

[1781] X. Benavent, F. Vegara, J. Domingo, and G. Ayala. On the use of longitudinal data techniques for modeling the behavior of a complex system. Lecture Notes in Computer Science, 2329: 458–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2329/23290458.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290458.
                                        pdf.

Kwiesielewicz:2002:PIC

[1782] M. Kwiesielewicz and E. van Uden. Problem of inconsistent and contradictory judgements in pairwise comparison method in sense of AHP. Lecture Notes in Computer Science, 2329:468– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290468.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290468.
                                        pdf.

Nguyen:2002:IPM

[1783] T. Nguyen and C. Plumejeaud. An integration platform for metacomputing applications. Lecture Notes in Computer Science, 2329:474–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290474.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290474.
                                        pdf.

Brezany:2002:LSS

[1784] P. Brezany, M. Bubak, M. Malawski, and K. Zajaac. Large-scale scientific irregular computing on clusters and grids. Lecture Notes in Computer Science, 2329:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2329/23290484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290484.
pdf.

Helstrup:2002:HLT

[1785] H. Helstrup, J. Lien, V. Lindenstruth, D. Röhrich, B. Skaali, T. Steinbeck, K. Ullaland, A. Vestbø, and A. Wiebalck. High level trigger system for the LHC ALICE experiment. Lecture Notes in Computer Science, 2329: 494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290494.
pdf.

Pierce:2002:GCW

[1786] M. Pierce, C. Youn, and G. Fox. The gateway computational Web portal: Developing Web services for high performance computing. Lecture Notes in Computer Science, 2329:503–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290503.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290503.
pdf.

Abdalhaq:2002:EOT

[1787] B. Abdalhaq, A. Cortés, T. Margalef, and E. Luque. Evolutionary optimization techniques on computational grids. Lecture Notes in Computer Science, 2329:513–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290513.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290513.
                                        pdf.

Goodyer:2002:EEP

[1788] C. Goodyer and M. Berzins. Eclipse and Ellipse: PSEs for EHL solutions using IRIS Explorer and SCIRun. Lecture Notes in Computer Science, 2329: 523–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290523.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290523.
                                        pdf.

Murillo:2002:PNK

[1789] M. Murillo and X.-C. Cai. Parallel Newton–Krylov–Schwarz method for solving the anisotropic bidomain equations from the excitation of the heart model. Lecture Notes in Computer Science, 2329:533–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290533.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290533.
                                        pdf.

Hluchy:2002:PFM

[1790] L. Hluchy, V. D. Tran, J. Astalos, M. Dobrucky, G. T. Nguyen, and D. Froehlich. Parallel flood modeling systems. Lecture Notes in Computer Science, 2329:543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290543.
pdf.

Nowinski:2002:WBR

[1791] A. Nowiński, K. Nowiński, and P. Bała.

Web based real time system for wavepacket dynamics. Lecture Notes in Computer Science, 2329:552–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290552.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290552.
pdf.

Gofen:2002:TCP

[1792] A. Gofen. The Taylor center for PCs:

Exploring, graphing and integrating ODEs with the ultimate accuracy. Lecture Notes in Computer Science, 2329: 562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290562.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290562.
pdf.

Cotofrei:2002:CRT

[1793] P. Cotofrei and K. Stoffel. Classification rules + time = temporal rules. Lecture Notes in Computer Science, 2329:572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2329/23290572.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290572.
                                        pdf.

Tam:2002:POD

[1794] L. Tam, D. Taniar, and K. Smith. Parametric optimization in data mining incorporated with GA-based search. Lecture Notes in Computer Science, 2329: 582–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290582.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290582.
                                        pdf.

Ladanyi:2002:ISP

[1795] L. Ladányi, T. K. Ralphs, and M. J. Saltzman. Implementing scalable parallel search algorithms for dataintensive applications. Lecture Notes in Computer Science, 2329:592–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290592.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290592.
                                        pdf.

Krishnaswamy:2002:TEC

[1796] S. Krishnaswamy, A. Zaslavsky, and S. W. Loke. Techniques for estimating the computation and communication costs of distributed data mining. Lecture Notes in Computer Science, 2329: 603–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290603.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290603.
pdf.

Cai:2002:DRA

[1797] Z. Cai and M. Lu. Distributed resource allocation in ad hoc networks. Lecture Notes in Computer Science, 2329:613– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290613.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290613.
pdf.

Karagiorgos:2002:ADM

[1798] G. Karagiorgos and N. M. Missirlis.

The average diffusion method for the load balancing problem. Lecture Notes in Computer Science, 2329:623–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290623.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290623.
pdf.

Tehver:2002:RAS

[1799] M. Tehver, E. Vainikko, K. Skaburskas, and J. Vedru. Remote access and scheduling for parallel applications on distributed systems. Lecture Notes in Computer Science, 2329:633–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290633.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290633.
                                        pdf.

Hong:2002:WSF

[1800] J. Hong, H. Sung, H. Lee, K. Kim, and S. Han. Workload scheduler with fault tolerance for MMSC. Lecture Notes in Computer Science, 2329:643– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290643.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290643.
                                        pdf.

Santoso:2002:SEJ

[1801] J. Santoso, G. D. van Albada, T. Basaruddin, and P. M. A. Sloot. A simulation environment for job scheduling on distributed systems. Lecture Notes in Computer Science, 2329:653– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290653.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290653.
                                        pdf.

Vankan:2002:IEM

[1802] W. J. Vankan, R. Maas, and M. ten Dam. ICT environment for multidisciplinary design and multi-objective optimisation: a case study. Lecture Notes in Computer Science, 2329:663– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290663.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290663.
pdf.

Koulisianis:2002:WBP

[1803] M. D. Koulisianis, G. K. Tsolis, and T. S. Papatheodorou. A Web-based problem solving environment for solution of option pricing problems and comparison of methods. Lecture Notes in Computer Science, 2329:673–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290673.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290673.
pdf.

Nechaev:2002:CCG

[1804] Yu.I. Nechaev, A. B. Degtyarev, and A. V. Boukhanovsky. Cognitive computer graphics for information interpretation in real time intelligence systems. Lecture Notes in Computer Science, 2329:683–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290683.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290683.
pdf.

Zhao:2002:AIA

[1805] Z. Zhao, R. G. Belleman, G. D. van Albada, and P. M. A. Sloot. AG-IVE: An agent based solution to constructing interactive simulation systems. Lecture Notes in Computer Science, 2329: 693–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290693.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290693.
                                        pdf.

Ruiz:2002:CAL

[1806] I. L. Ruiz, E. L. Espinosa, G. C. García, and M.Á. Gómez-Nieto. Computerassisted learning of chemical experiments through a 3D virtual lab. Lecture Notes in Computer Science, 2329: 704–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290704.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290704.
                                        pdf.

Derksen:2002:LBB

[1807] J. Derksen. Lattice-Boltzmann based large-eddy simulations applied to industrial flows. Lecture Notes in Computer Science, 2329:713–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290713.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290713.
                                        pdf.

Niaei:2002:CSP

[1808] A. Niaei, J. Towfighi, M. Sadrameli, and M. E. Masoumi. Computational study of the pyrolysis reactions and coke deposition in industrial naphtha cracking. Lecture Notes in Computer Science, 2329:723–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290723.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290723.
pdf.

Bianco:2002:AEF

[1809] M. Bianco, G. Bilardi, F. Pesavento, G. Pucci, and B. A. Schrefler. An accurate and efficient frontal solver for fully-coupled hygro-thermo-mechanical problems. Lecture Notes in Computer Science, 2329:733–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290733.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290733.
pdf.

Christakis:2002:UCF

[1810] N. Christakis, P. Chapelle, M. K. Patel, M. Cross, I. Bridle, H. Abou-Chakra, and J. Baxter. Utilising computational fluid dynamics (CFD) for the modelling of granular material in largescale engineering processes. Lecture Notes in Computer Science, 2329:743– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290743.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290743.
                                        pdf.

Gloukhov:2002:PII

[1811] V. Gloukhov. Parallel implementation of the INM atmospheric general circulation model on distributed memory multiprocessors. Lecture Notes in Computer Science, 2329:753–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290753.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290753.
                                        pdf.

Campari:2002:RSH

[1812] E. G. Campari and G. Levi. A realistic simulation for highway traffic by the use of cellular automata. Lecture Notes in Computer Science, 2329:763– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290763.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290763.
                                        pdf.

Kroc:2002:ACA

[1813] J. Kroc. Application of cellular automata simulations to modeling of dynamic recrystallization. Lecture Notes in Computer Science, 2329:773– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2329/23290773.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290773.
pdf.

Topa:2002:DCA

[1814] P. Topa. A distributed cellular automata simulation on cluster of PCs. Lecture Notes in Computer Science, 2329:783–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2329/23290783.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2329/23290783.
pdf.

Jimenez-Morales:2002:EOD

[1815] F. Jiménez-Morales, M. Mitchell, and J. P. Crutchfield. Evolving one dimensional cellular automata to perform non-trivial collective behavior task: One case study. Lecture Notes in Computer Science, 2329:793–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290793.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290793.
pdf.

Kole:2002:NUS

[1816] J. S. Kole, M. T. Figge, and H. De Raedt. New unconditionally stable algorithms to solve the time-dependent Maxwell equations. Lecture Notes in Computer Science, 2329:803–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290803.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290803.
                                        pdf.

Dincov:2002:CDF

[1817] D. D. Dinc̆ov, K. A. Parrott, and K. A. Pericleous. Coupled 3-D finite difference time domain and finite volume methods for solving microwave heating in porous media. Lecture Notes in Computer Science, 2329: 813–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290813.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290813.
                                        pdf.

Perminov:2002:NSR

[1818] V. Perminov. Numerical solution of Reynolds equations for forest fire spread. Lecture Notes in Computer Science, 2329:823–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290823.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290823.
                                        pdf.

Heitzer:2002:FBS

[1819] M. Heitzer. FEM-based structural optimization with respect to shakedown constraints. Lecture Notes in Computer Science, 2329:833–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290833.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290833.
pdf.

Leopold:2002:TBC

[1820] C. Leopold. Tight bounds on capacity misses for 3D stencil codes. Lecture Notes in Computer Science, 2329: 843–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290843.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290843.
pdf.

Walkley:2002:DCO

[1821] M. Walkley, J. Wood, and K. Brodlie.

A distributed co-operative problem solving environment. Lecture Notes in Computer Science, 2329:853–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290853.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290853.
pdf.

Gang:2002:SAP

[1822] X. J. Gang, W. H. An, and D. G.

Zhong. The software architecture of a problem solving environment for enterprise computing. Lecture Notes in Computer Science, 2329:862–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290862.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290862.
                                        pdf.

Johansson:2002:SAG

[1823] P. Johansson and D. Kressner. Semi-automatic generation of Webbased computing environments for software libraries. Lecture Notes in Computer Science, 2329:872–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290872.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290872.
                                        pdf.

Scurr:2002:DGB

[1824] A. D. Scurr and A. J. Keane. The development of a grid based engineering design problem solving environment. Lecture Notes in Computer Science, 2329: 881–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290881.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290881.
                                        pdf.

Nguyen:2002:TPP

[1825] G. T. Nguyen, V. D. Tran, and M. Kotocova. TOPAS — parallel programming environment for distributed computing. Lecture Notes in Computer Science, 2329:890–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290890.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290890.
pdf.

Nool:2002:PIL

[1826] M. Nool and M. M. J. Proot. Parallel implementation of a least-squares spectral element solver for incompressible flow problems. Lecture Notes in Computer Science, 2329:900–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290900.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290900.
pdf.

Meng:2002:SIS

[1827] S. Meng, X. K. Li, and G. Evans.

Smooth interfaces for spectral element approximations of Navier– Stokes equations. Lecture Notes in Computer Science, 2329:910–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290910.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290910.
pdf.

Chambarel:2002:SCF

[1828] A. Chambarel and H. Bolvin. Simulation of a compressible flow by the finite element method using a general parallel computing approach. Lecture Notes in Computer Science, 2329:920– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290920.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290920.
                                        pdf.

Banda:2002:CRS

[1829] M. K. Banda and M. Seaïd. A class of the relaxation schemes for two-dimensional Euler systems of gas dynamics. Lecture Notes in Computer Science, 2329:930–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290930.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290930.
                                        pdf.

Keppens:2002:OPM

[1830] R. Keppens and G. Tóth. OpenMP parallelism for multi-dimensional gridadaptive magnetohydrodynamic simulations. Lecture Notes in Computer Science, 2329:940–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2329/23290940.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2329/23290940.pdf.

Rodriguez:2002:PET

[1831] E. Rodríguez, G. Montero, R. Montenegro, J. M. Escobar, and J. M. González-Yuste. Parameter estimation in a three-dimensional wind field model using genetic algorithms. Lecture Notes in Computer Science, 2329: 950–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290950.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290950.
pdf.

Li:2002:MIM

[1832] S. Li, S. C. La, W. H. Yu, and L. Wang. Minimizing interference in mobile communications using genetic algorithms. Lecture Notes in Computer Science, 2329:960–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290960.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290960.
pdf.

Castellano:2002:KMT

[1833] G. Castellano, C. Castiello, and A. M.

Fanelli. KERNEL: a Matlab toolbox for knowledge extraction and refinement by NEural learning. Lecture Notes in Computer Science, 2329:970– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23290970.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23290970.
pdf.

Zacharias:2002:DRC

[1834] J. Zacharias, C. Hartmann, and A. Delgado. Damages recognition on crates of beverages by artificial neural networks trained with data obtained from numerical simulation. Lecture Notes in Computer Science, 2329:980– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23290980.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23290980.
                                        pdf.

Watanabe:2002:SMS

[1835] T. Watanabe, E. Kume, and K. Kato. Simulation monitoring system using AVS. Lecture Notes in Computer Science, 2329:990–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2329/23290990.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2329/23290990.pdf.

Gofen:2002:ORC

[1836] A. Gofen. ODEs and redefining the concept of elementary functions. Lecture Notes in Computer Science, 2329: 1000–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2329/23291000.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23291000.
                                        pdf.

Schoemaker:2002:CDS

[1837] R. M. Schoemaker, P. C. A. de Haas, H. J. H. Clercx, and R. M. M. Mattheij. Contour dynamics simulations with a parallel hierarchicalelement method. Lecture Notes in Computer Science, 2329:1010–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23291010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23291010.
pdf.

Cheng:2002:PAD

[1838] J.-R. C. Cheng and P. E. Plassmann. A parallel algorithm for the dynamic partitioning of particle-mesh computational systems. Lecture Notes in Computer Science, 2329:1020–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23291020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23291020.
pdf.

Okunbor:2002:SSI

[1839] D. Okunbor and E. Akinjide. Stable symplectic integrators for power systems. Lecture Notes in Computer Science, 2329:1030–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2329/23291030.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2329/23291030.pdf.

Prodan:2002:CJC

[1840] A. Prodan and R. Prodan. A collection of Java class libraries for stochastic modeling and simulation. Lecture Notes in Computer Science, 2329:1040– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23291040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23291040.
                                        pdf.

Ochmanska:2002:TOP

[1841] E. Ochmanska. Task-oriented Petri net models for discrete event simulation. Lecture Notes in Computer Science, 2329:1049–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2329/23291049.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23291049.
                                        pdf.

Oliveira:2002:SSP

[1842] S. Oliveira, D. Stewart, and T. Soma. A subspace semidefinite programming for spectral graph partitioning. Lecture Notes in Computer Science, 2329:1058– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2329/23291058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2329/23291058.
                                        pdf.

Yoo:2002:SPE

[1843] H. S. Yoo and J.-H. Jang. A study on the pollution error in r-h methods using singular shape functions. Lecture Notes in Computer Science, 2329:1068– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2329/23291068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2329/23291068.
pdf.

Scotney:2002:DSD

[1844] B. W. Scotney, S. A. Coleman, and M. G. Herron. Device space design for efficient scale-space edge detection. Lecture Notes in Computer Science, 2329:1077–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2329/23291077.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2329/23291077.
pdf.

Anonymous:2002:AIay

[1845] Anonymous. Author index. Lecture Notes in Computer Science, 2329:1087– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2329/2329auth.pdf.